A Complete Bibliography of Publications in Limnology and Oceanography: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

10 April 2024 Version 1.14

Title word cross-reference

13 [HZC^{+}24, LRRP20, WWGP21]. ^{15} [HZC^{+}24, SWS^{+}22, WWGP21]. ^{18} [DLL20]. ^{222} [ACW^{+}22]. ^{223} [LvBS^{+}22]. ^{224} [LvBS^{+}22]. ^{34} [HZC^{+}24]. ^{th} [BGM^{+}22]. _{2} [AAC^{+}21, AFH^{+}21, AFH^{+}22, ACW^{+}22, AKT^{+}21, BLS20, CGR22, CT23, CTM^{+}21, CCB^{+}20, CSC^{+}21, CMCC24, CHS^{+}20, DRA^{+}23, EDC24, EBW^{+}20, EWGP22, GPB^{+}20, HBS^{+}22, HGA24, LLL^{+}23, LSL^{+}20, LTLH22, MGP^{+}21, MDWT^{+}22, MRT^{+}24, MGG^{+}24, NB20, QBJG21, RRHS^{+}21, RGW^{+}23, RSP^{+}24, SCC^{+}21, SDB^{+}23, SPLJA^{+}20, TZF^{+}22, TMS^{+}21, hTRH20, VLP^{+}21, VKB^{+}22, WRHR22, WDG^{+}21, WMRL22, XWL^{+}21, YMSH20, YSHS22, YSHS23, YG23]. _{3} [LL23]. _{4} [AFH^{+}21, DRA^{+}23]. : [KMV^{+}20]. δ [HZC^{+}24, LRRP20, SWS^{+}22, WWGP21]. p [ACW^{+}22, YG23].

10.1002 [Ano21a]. 10.1002/lno.10504 [Ano21a]. 12-yr [BPG^{+}22].

2011 [TNK^{+}22]. 2016 [CMT^{+}20].

30 [RCMA21]. 30^{o} [BSGA21]. 36-month [GWS^{+}21].

59^{o} [BSGA21].

62 [Ano21a]. 64 [Ano21b].

8-m [BGM^{+}22].

ability [KMM^{+}22]. abiotic [KPNJ23, LGF22, MTP^{+}23, TZF^{+}22, WRMP22]. above [BZK^{+}22, GPK^{+}21, NH23]. Abrupt [BGM^{+}22, GBO^{+}20]. absorption [AVJ^{+}24, SSB^{+}22]. abundance [BMR^{+}21b, BBGM20, GFW^{+}20, LBBM21, LPMBR23, MBD^{+}22, MCYR20, SSH^{+}22]. abundances [EPR^{+}24, SML^{+}21a]. acantharian [BBGM20]. accelerate [CCC^{+}21]. Accelerated [AJMG21]. Acclimation [GTKW^{+}20, SMD21]. accounting [KHCG^{+}22, LAB^{+}22]. accounts [SMW^{+}21]. accretion [CEE^{+}22, MEQBV20, MVBG20]. accumulation [BAG^{+}24, HMP^{+}21, LLK^{+}22, LAB^{+}22, MEQBV20, PBP^{+}23, SSWM20, hTRH20, WSLC22, ZWM22]. accurate [KHCG^{+}22]. achieved [KTA^{+}21]. Achilles [ZvBZ^{+}20]. acid [KLG^{+}23, SRM23, VCT^{+}23, WWGP21]. acidic [MNMJ^{+}21, SPLJA^{+}20]. acidification [ARPFGS23, HDS^{+}20, HSV22, JHS^{+}21, RVBP23, RBD22, RLP^{+}20, SMG^{+}22, SSEO21, ZLU^{+}20]. acids [CLG^{+}24, DMS^{+}21, EST^{+}24, KHSP21, VST^{+}22, WDS^{+}23]. acoustically [GNO22]. acquisition [DK20, SdBW^{+}23]. across [ARB^{+}22, BMA^{+}20, BCTH20, BRHS20, CLG^{+}24, CEW^{+}21, CZK^{+}23, CEE^{+}22, CLP^{+}24, DHNH22, DRA^{+}23, DBFC24, ETMD23, FDF^{+}21, GSP^{+}23, GGLH^{+}22, GCW^{+}23, GLAHH^{+}23, GSD^{+}23, HDD^{+}22, HLPT20, HM20, JLS^{+}20, JHBB22, KAP^{+}20, KSF^{+}23, LTQS^{+}21, LHL^{+}21, LTW^{+}24, LHD^{+}20, LCZ^{+}21, MTP^{+}23, MLP^{+}21a, MSB^{+}22, MCH^{+}21, MKP^{+}23, NH22, PYM^{+}20, PGEB22, RWMP^{+}23, RSLW23, RBAC^{+}20, RVS^{+}22, SWOR20, SML21b, SHZ^{+}21, SSR^{+}23, TKA21, VCBB20, VMT^{+}24, VCT^{+}23, VST^{+}22, VMP^{+}21, WRC^{+}21, WCC20, vHHP^{+}22]. act [OYM^{+}20]. active [BGG^{+}20, EBK^{+}20, PC21, RTK23, SSS^{+}24, XYL^{+}23]. activities [BSGA21, LZD^{+}23]. activity [CTP21, CLT^{+}22, GFW^{+}20, KGG^{+}23, PGN^{+}21, SCC^{+}21, SAD^{+}20, ZLJ^{+}24, ZYZ^{+}21, vHLV^{+}20]. acute [EFP^{+}21]. adaptation [LBvE23, LCS^{+}20, MUGL20, RCS^{+}20, ZZCD23]. adaptations [FRRG^{+}21]. adapted [CBD^{+}23, KKP^{+}21, MRC^{+}24, TKA21]. Adaptive [WSB^{+}20, CBT^{+}21, DKW^{+}21]. addition [BVM^{+}20]. additionality [HGA24]. additions [DCF^{+}22, PBC21, VWR^{+}23]. Adélie [NCC^{+}21]. Adirondack [WK21]. adjacent [DKPS23, KSF^{+}23, NCC^{+}21, SORGC^{+}21]. Adriatic [CDA^{+}20]. advances [SHS^{+}22]. advancing [MVBG20]. advected [SLZ^{+}21b]. advection [CJB^{+}21]. advective [LDG^{+}21]. Aerobic [AWB^{+}23, SSM^{+}22b]. aerosol [YAE^{+}23]. aeruginosa [MDP^{+}23]. affect [DGTW21, OJUN23, PSNS22, SVBT20, SOT^{+}20, XSZ^{+}21]. affected [IÁSNCR21, QMG21, TUR^{+}20]. affecting [ZLJ^{+}24]. affects [ARPFGS23, LRNG24, LHS^{+}20, RWL^{+}24, RBL^{+}20, WWB^{+}20, vdPAV^{+}20]. affiliation [HNS21]. Africa [GBB^{+}21]. African [LL23]. after [BWB^{+}22, GPB^{+}20, HBR^{+}22, RVS^{+}22]. aftermath [KPS21]. age [FHMK20]. aged [KSH^{+}23]. aggregate [HVMS^{+}21, KL22, ZLIP20]. aggregate-colonizing [KL22]. aggregates [EWGP22, LCMG^{+}20, ZLIP20]. aggregation [ALG^{+}23]. aggregations [MBLA^{+}21]. aging [SSS^{+}24]. Aglantha [MMKWZ22]. aid [LHH24]. aids [LBJS22]. air [ACW^{+}22, CCB^{+}20, MBA^{+}21, SGM^{+}21]. al [Ano21b, ZR23, Ano21a]. al. [CSH21, PPB^{+}21]. Alaska [CCDM21, MDDL22, VJH^{+}22]. Alaskan [MK21, SSR^{+}23]. alewife [DFJ22, RPA^{+}22]. Alexandrium [BH21, PHM21]. alga [BK20, GR24, WCH^{+}21]. algae [FPBG^{+}22, LLSQ24, THS21]. algal [BMGK22, FPG^{+}24, GTKW^{+}20, HPF^{+}23, HM20, KAC^{+}23, KRC23, KWB21, KNG^{+}23, LZG21, SKB^{+}22, ZXZ^{+}23]. algal-virus [HPF^{+}23]. algorithms [MT23]. alicyclic [LPO^{+}20]. aligns [CGB^{+}21]. alkaline [BWV^{+}20]. Alkalinity [YXC^{+}22, DSY20, IMB^{+}22]. alkenone [THT^{+}20]. alkenone-producing [THT^{+}20]. alleviates [RS20]. allocation [ZLS^{+}21]. allochthonous [CTP21, KBB^{+}22]. allochthony [HHK^{+}22]. allometries [LLMM^{+}23]. allow [RTK23]. allows [HMBB21]. along [ADLW22, AKT^{+}21, BGAD^{+}22, BWV^{+}20, BBL^{+}21, CÖL^{+}20, CFC^{+}20, CCY^{+}21, CCDM21, DWS^{+}22, EJM^{+}23, GEHD20, HHMH22, IMB^{+}22, IYW^{+}22, JSL^{+}23, JWAvD^{+}22, KPS21, LAFV23, LSC^{+}21, MHP^{+}23, NSK^{+}21, PMF^{+}21, RCMA21, SO23, TKW^{+}23, TKS^{+}20, ZZO^{+}21, ZYZ^{+}21]. Alosa [DFJ22]. alpha [GCEM^{+}23]. Alpine [SPR22, BVL^{+}23, PSE^{+}20, PMD^{+}23, SDS^{+}23, XTL^{+}20, ZLJ^{+}24, DLC^{+}22]. Alps [ESW^{+}21]. alter [DKBG21, HEH^{+}20, HZS21, MRT^{+}24, ZZWL22]. alterations [BCFI^{+}22, CMB^{+}21]. altered [BMH^{+}23, LVZ21, RSLW23]. alternative [BSIdG23, SH22, SHBA22]. alters [JHBB22, KTH^{+}21, MMKWZ22, RLP^{+}20, SMD21, ZZT^{+}21, ZWM22]. altitude [DCF^{+}22]. Aluminum [ZLL^{+}21, MCS^{+}23]. Amazon [CBF^{+}23, LvBS^{+}22, PBP^{+}23, PCN22, dMKB^{+}20]. Amazonia [KDZ^{+}24]. American [ETG21, WSB^{+}20, YWZ^{+}21]. Amino [SRM23, WWGP21, DMS^{+}21, GAS^{+}23, KHSP21, VCT^{+}23, WDS^{+}23]. ammonia [FSS^{+}23]. ammonium [ACA^{+}22, AWGV20, CEW^{+}21, HBR^{+}22, MSB^{+}22, PCC^{+}23, RWS^{+}22]. among [BRB^{+}20, BFE^{+}20, BRB^{+}22, CGGC^{+}21, CLW^{+}21, EWF22, HEH^{+}20, HDS^{+}20, LTJ^{+}20, LLBC20, MGG^{+}24, RCH^{+}20, RBR^{+}23, SVMM^{+}22, SRM23]. amphipod [BRB^{+}22]. amplified [BTK^{+}21]. Anaerobic [LPA^{+}23, ACA^{+}22, MWHR^{+}22, SSM^{+}22b, SZY^{+}20]. Analyses [GWS^{+}21, KLG^{+}23, WWGP21, WDS^{+}23]. Analysis [MMLPRHCG22, vHHP^{+}21, CDH^{+}21, DMS^{+}21, FHMK20, HTLP23, LN21, MGWS23, SGM^{+}22, WZCK20]. anammox [CEW^{+}21, CHP^{+}24, ZQB^{+}21]. anchialine [HBY^{+}22]. Andean [BLS^{+}22, PBGGRB21]. anhydrase [SAD^{+}20]. animals [LG21]. Annual [CVF^{+}20, vLWV^{+}20, FR23, SDL22, vdPMF^{+}21]. Anomalous [CJB^{+}21]. anoxia [AJMG21, CCC^{+}21]. anoxic [HBR^{+}22, KdED^{+}20, SSM^{+}22b]. Antarctic [CFC^{+}20, CMT^{+}20, NGSS23, BTT23, BJK^{+}21, BBL^{+}21, CT23, CP22, CSNS24, DAB^{+}21, JHvL^{+}23, JWAvD^{+}22, KTH^{+}21, LDS^{+}20, MSGW^{+}20, MP23, MHP^{+}23, SGSF20, WOMWR20, YAH^{+}21, vLWV^{+}20]. Antarctica [LWvdM^{+}21, NCC^{+}21, SSP^{+}20, DAB^{+}21]. Antecedent [TKM^{+}22, RVS^{+}22]. Anthropocene [BAG^{+}24, KDW^{+}23]. Anthropogenic [MLL^{+}21, CLW^{+}21, FBK^{+}22, LDB^{+}20, WM20, ZYZ^{+}21]. anticyclone [WCC20]. apex [MWV22]. aphotic [LLL^{+}23]. appendicularian [MLS^{+}21]. application [LCMG^{+}20, SHPW22]. Applications [DAN^{+}22]. Applying [WRMP22]. approach [ALG^{+}23, ASL^{+}24, DNA^{+}22, FMT^{+}20, GCEM^{+}23, GBW24, KBSA21, KHCG^{+}22, VMP^{+}21]. approaches [GWS^{+}21, KSF^{+}23, LAB^{+}22, MLB^{+}21, PMJ^{+}21, ST22, WAPC20]. approaching [AR20]. aquatic [BBMD20, CBPNA22, GF21, GMMW20, HEH^{+}20, HTSG21, HGMK22, MLB^{+}21, MGWS23, QS21, SHS^{+}22, SPHH22, YJB^{+}20, ZN20]. aquifer [BKMY21, KPM^{+}20]. archaea [SZY^{+}20]. archaeaplankton [WPY^{+}20]. archetypes [LWH^{+}23]. Archipelago [vHLV^{+}20, SBV^{+}21]. archives [TJW^{+}22]. Arctic [SBM^{+}21, TMS^{+}21, BCA^{+}20, BOC^{+}21, BPG^{+}22, CLD^{+}23, CM20b, CGB^{+}21, GPK^{+}21, HZM^{+}21, HTSG21, HGH^{+}21, IKN^{+}21, JLMW21, KAP^{+}20, KLG^{+}23, KGG^{+}23, LAFV23, LMM^{+}21, LLC21, LWC^{+}24, LXC^{+}21b, LDB^{+}21, MMKWZ22, MKBSK19, MN20, MKD22, MGGRR^{+}21, NB23, PGN^{+}21, PLB^{+}21, RBR^{+}24, RRHS^{+}21, RGW^{+}23, RCS^{+}20, SJR21, SMCS23, SHE^{+}21, SZA^{+}21, SJJ^{+}21, SSR^{+}23, TVB^{+}21, VCT^{+}23, VMP^{+}21, WCH^{+}21, WRHR22, WM20, XKP^{+}22, ZZO^{+}21, ZXE^{+}23, vdPAV^{+}20, vdPMF^{+}21]. area [CHS^{+}20, CBS^{+}21, IJSC20, PPB^{+}21]. areas [DvOH^{+}20]. Argo [WF22]. argus [BFCC^{+}21]. arid [LRM21]. arsenic [PCJ^{+}21]. arthropod [KPNJ23]. artificial [ARB^{+}22, MTB^{+}21, SJH20]. artificially [DMCW23]. ascidians [JSW^{+}21]. ASLO [Ano20a]. assemblage [PATL^{+}23, RLP^{+}20]. assemblages [ASL^{+}24, SKB^{+}22, SSB^{+}22]. assembly [IYW^{+}22, SGS22]. assessed [CDH^{+}21]. Assessing [AVJ^{+}24, HTLP23, KSR20]. assessment [DAN^{+}22, GAS^{+}23, MLP^{+}21a, SRL^{+}20, TRS^{+}20, WMTJW22]. assessments [CBC^{+}22]. Assimilation [ORE20, SHE^{+}21, AHCJ^{+}20]. associated [AAC^{+}21, BSGA21, BRB^{+}20, BDMGH^{+}23, CKH^{+}21, CMB^{+}21, Edm21b, FKN21, HMP^{+}21, HZZ^{+}20, HGF^{+}21, KFHS24, LBBM21, MEQBV20, RVBP23, WPY^{+}20, WRHR22]. association [Edm21a, WO21, XKP^{+}22]. Associations [LEK^{+}22, LVZ21, RCMA21]. asynchronous [WSZ22a, WSZ22b]. asynchrony [VCBB20]. Atacama [SML^{+}21a, FUUG^{+}21]. Atchafalaya [Roy20]. Atlantic [LL23, SBW^{+}20, YAE^{+}24, BMR^{+}21a, BAD^{+}21, BDL^{+}21, BFE^{+}20, BAHH^{+}20, CGR22, CDH^{+}21, CBN21, GLAHH^{+}23, HBG^{+}22, JLS^{+}20, JSL^{+}23, KBvdD^{+}20, MSG^{+}22, SORGC^{+}21, SPKM22, SSM^{+}22a, WF22, YAH^{+}21, YAE^{+}23]. Atlantification [MMKWZ22]. atmosphere [BCL^{+}22]. Atmospheric [JOC^{+}20, SGM^{+}21, SMNT22, VLP^{+}21, ZZHZ23, ZLW21b]. atoll [FHR^{+}21, GLB^{+}22]. atrazine [BCTH20]. attenuate [LPA^{+}23]. Attenuation [PC21, WDS^{+}23, WAPC20, ZXE^{+}23, ZXZ^{+}23]. attraction [SHBA22]. attributes [PFB^{+}22]. atyid [HBY^{+}22]. atypical [EJM^{+}23]. Aufeis [HGH^{+}21]. augments [PFH21]. Aurelia [MBC^{+}21]. Australia [MVBG20, LAB^{+}22]. autochthonous [KBB^{+}22]. automated [FHMK20, KAC^{+}23]. Autonomous [LDG^{+}21, BRCO21, WDG^{+}21]. autotrophic [DWW^{+}21]. autumn [CLP^{+}24, SBM^{+}21, SBV^{+}21]. availability [FBJ^{+}23, FNB^{+}23, HZM^{+}21, HPP^{+}21, JHSG^{+}24, JWAvD^{+}22, KLP23, LTLH22, LBG^{+}22, LHS^{+}20, LSCL21, NCC^{+}21, RBL^{+}20, SMKS22, SSK24, WWFS21]. avoid [GE22].

back [EDC24]. background [GLN^{+}20, GLN^{+}21]. bacteria [BMLV24, CCC^{+}21, GAS^{+}23, HPB^{+}21, JHSG^{+}24, LES^{+}21, MLK^{+}22, PDK^{+}23, PATL^{+}23, RHS^{+}22, SJR21, VKGW22, ZLW^{+}21a]. Bacterial [BLRM24, CCY^{+}21, YPF^{+}23, ZLL^{+}24, BYS^{+}23, CLD^{+}23, CVF^{+}20, KPG^{+}23, MZW^{+}21, PGN^{+}21, PKV20, SPLJA^{+}20, ZQB^{+}21, ZZT^{+}21]. Bacterioplankton [CGRGP^{+}21, CTP21, WCC20, FWvA^{+}21, IYW^{+}22, LPO^{+}20, WCZ^{+}20, WPY^{+}20, WZCK20, YML^{+}21, dMKB^{+}20]. bacterivory [AH20]. Baikal [CYZZ^{+}20, ONH^{+}20]. balance [BMR^{+}21a, ESW^{+}21, KBSA21, LVDM22]. balances [RST^{+}21]. Baltic [TMKB^{+}22, vHLV^{+}20, BDMGH^{+}23, CGRGP^{+}21, CBB^{+}22, HPNU20, KPL^{+}20, LCHB22, RBR^{+}23]. bank [RVS^{+}22]. bank-full [RVS^{+}22]. barcoding [CSM^{+}23]. Barents [DGW^{+}21]. barite [LN21]. barium [FSH^{+}23]. barnacle [FJP^{+}21]. baroclinic [ILPL20]. barriers [DCJB22, SORGC^{+}21]. Basal [MHL^{+}20, VCT^{+}23]. base [SPR22]. based [BMA^{+}20, BCWS23, GWS^{+}21, HRK^{+}22, LWH^{+}23, LEB21, LLBC20, MLP^{+}21a, MLB^{+}21, MK20, RVS^{+}22, SBD^{+}21, VMP^{+}21, WSZ22a, WSZ22b, vHHP^{+}22, CBN21]. Baseline [CLW^{+}21, GBW24, SOO^{+}21]. baselines [DAN^{+}22]. Basin [MGI^{+}22, AJMG21, CWF^{+}22, GGR24, KdED^{+}20, KVKS22, LEK^{+}22, MACC24, Roy20, SORGC^{+}21, SSM^{+}22b, DSY20, VCBB20]. Basin-scale [MGI^{+}22]. basin-wide [SORGC^{+}21]. basins [LLK^{+}22, RBR^{+}23, SHBA22, SHZ^{+}21]. bathypelagic [CYZZ^{+}20, SSS^{+}24]. Bay [HSD^{+}20, LZG21, MN20, AAB^{+}22, LAMG^{+}22, SSM^{+}22a, WPY^{+}20, XYL^{+}23, CCB^{+}20, FHMK20, MLK^{+}22, SCB^{+}20, SCT^{+}21, TTC^{+}22, TRS^{+}20, WZCK20]. bays [JvKvT^{+}20, MZW^{+}21]. be [MCS^{+}23]. beach [vEMF^{+}20]. bearing [BO20, IMB^{+}22]. Beaufort [CCDM21]. bed [ARBvS^{+}21, MDBI20, PMD^{+}23]. bed-load [MDBI20]. bedrocks [MHVC22]. beds [MRH^{+}23, YvdHC^{+}20]. Beggiatoaceae [MLK^{+}22]. behavior [CLT^{+}22, DKBG21, KCT23]. behavioral [GFG^{+}22]. behaviors [CSB22]. behind [VSB^{+}23]. being [MLL^{+}21]. below [BZK^{+}22, QBJG21]. belowground [CEE^{+}22]. bending [SMF20]. benefit [FPG^{+}24]. benefits [JOC^{+}20, OLR^{+}23]. Benner [LD22]. Benthic [HZM^{+}21, KPL^{+}20, ABS^{+}21, BSB^{+}20, BMR^{+}21b, BWV^{+}20, BZK^{+}22, BOC^{+}21, CEW^{+}21, EFG^{+}24, FKN21, HANW21, HPCH24, HCS^{+}21, HBR^{+}22, KVKS22, LHS^{+}21, LBG^{+}22, MV22a, MV22b, MBLA^{+}21, MHdS^{+}21, PSY^{+}22, SLZ^{+}21a, TKS^{+}20, VCT^{+}23, WHZ^{+}21]. Benthic-pelagic [KPL^{+}20, MBLA^{+}21]. bentho [NvBE^{+}21]. bentho-pelagic [NvBE^{+}21]. benthos [CP22, SSK24]. Bering [LNC^{+}23]. Bermuda [CBN21]. beta [GCEM^{+}23, LHL^{+}21]. betaine [MAD^{+}22]. better [SBD^{+}21]. between [BHSG^{+}20, BMA^{+}20, BM22, BMLV24, BH21, CBS^{+}21, DWW^{+}21, DKPS23, FW21, GWS^{+}21, GBB^{+}21, HZM^{+}21, HMB^{+}20, LEK^{+}22, LABM20, MCH^{+}22, PPB^{+}21, QS21, SDL22, SJJ^{+}21, SWS^{+}22, SDSGR21, SGSD^{+}22, VVMJ23, WCZ^{+}20, XKP^{+}22, XTJW21, YAE^{+}24, ZQB^{+}21, ZLU^{+}20, ZFQ^{+}20]. Beyond [KHCG^{+}22]. bias [MMLPRHCG22]. biased [KSI24]. Bidirectional [EST^{+}24]. Bight [LMC^{+}21]. bio [BHG^{+}24, BC20]. bio-optical [BHG^{+}24]. bioactive [BBA^{+}23]. bioavailability [DVL^{+}23, EJRRC^{+}23, RBSB22]. biochemical [Ano21c, WRWPG19]. bioconvection [SBW21]. bioconvection-induced [SBW21]. biodegradability [SEK20]. biodegradation [BKC20, CWF^{+}22]. biodiversity [GCMJ23, KPNJ23]. bioerosion [BFW^{+}22]. Biofilm [CTF^{+}21, ARBvS^{+}21]. biofilms [MKP^{+}23, RBAC^{+}20]. biofouled [GLM^{+}23]. biofouling [KWB21]. biogenic [BBCA^{+}21, IKN^{+}21, JHS^{+}21, LWC^{+}24, MZXY22, MY23, RNW^{+}23, RWL^{+}24, ZBNH^{+}21]. Biogeochemical [BYB^{+}23, DGTW21, RFB^{+}22, SBM^{+}21, TBH^{+}22, VCBB20, WF22, WAPC20, BVM^{+}20, CBD^{+}23, EFC^{+}23, FDF^{+}21, GMMW20, KSF^{+}23, LLMM^{+}23, LGI^{+}20, LZG21, LBG^{+}22, MHD^{+}22, MBD^{+}23, MPW^{+}23, SSP^{+}20, TKH21, TCP^{+}22, VMT^{+}24, WOK^{+}21, ZSD^{+}20]. Biogeochemical-Argo [WF22]. Biogeochemistry [JLMW21, FBJ^{+}23, JSB^{+}20, KDZ^{+}24, LWvdM^{+}21]. Biogeography [MCH^{+}21, MZW^{+}21, BLRM24, MBD^{+}22, PYM^{+}20]. Biogeomorphic [QMG21, CZH^{+}21]. Biological [CGD22, GMNG^{+}24, MGL22, NTJ^{+}20, ARBvS^{+}21, BGM^{+}22, DMS^{+}21, DLL20, GBB^{+}21, KTH^{+}21, KBL^{+}20, LTQS^{+}21, MK21, ORRMD^{+}21, PFH21, PVSP^{+}22, PJRV23, QRWT^{+}22, TKS^{+}20]. biomarkers [PMF^{+}21]. biomass [BRCO21, FBV23, HNS21, KTWGT20, LCS^{+}20, LDB^{+}20, LCZ^{+}21, MTP^{+}23, MSJ21, PBC21, QS21, SBB20, YAP^{+}22]. biomasses [EG20]. biophysical [ASH^{+}20]. biospheric [BJK^{+}21]. Biotic [BC20, LGF22, MTP^{+}23, MOW^{+}22, TZF^{+}22, WWB^{+}20, WRMP22]. biotoxins [THS21]. bioturbation [BSB^{+}20]. bioturbators [ZSF^{+}23]. biovolume [MNM^{+}21]. bivalve [GPB^{+}20, MIR^{+}22, WTvD^{+}22]. Biwa [LDB^{+}20, LDB^{+}21]. black [BND^{+}23, RCH^{+}20, SO23, SGSD^{+}22]. black-box [RCH^{+}20]. blackwater [TUR^{+}20]. bleaching [Ano21c, CBC^{+}22, RFB^{+}22, WRWPG19]. bloom [BMH^{+}23, CGRGP^{+}21, CMT^{+}20, GMNG^{+}24, GAS^{+}21, HPP^{+}21, KRC23, LZG21, MBB^{+}20, MLS^{+}21, PHM21, WOT^{+}23, XMP^{+}21]. bloom-forming [BMH^{+}23]. blooming [MBC^{+}21]. blooms [BAHH^{+}20, HM20, HMB^{+}22, KAC^{+}23, Roy20, SVC^{+}22, SLJ^{+}22, SKB^{+}22, SRL^{+}20, THT^{+}20, THS21, ZXZ^{+}23]. Blue [HMP^{+}21, HGA24, FGP22, KBB^{+}22, KHCG^{+}22, LGQ^{+}21, LAB^{+}22, XWL^{+}21, OGD20]. bluefin [OFM^{+}21]. Body [GNO22, CHAA21, LDB^{+}20, PMRG20, dJGCS23]. Boeckella [MSGW^{+}20]. bog [SGL^{+}23]. boreal [AVJ^{+}24, AKT^{+}21, CHS^{+}20, FSM^{+}21, KSH^{+}23, MBA^{+}21, MHSL22, MNMJ^{+}21, RSLW23, RBSB22, VST^{+}22]. Borealization [MN20]. Both [MTP^{+}23]. Bothnian [RvHF^{+}20]. Bottom [BD22, NH22, BMA^{+}20, DHNH22, GÖA^{+}23, LAFV23, LLBC20, PMD^{+}23]. Bottom-up [BD22, BMA^{+}20, GÖA^{+}23, LLBC20]. bound [SdBW^{+}23]. boundaries [KHCG^{+}22, WSZ22a, WSZ22b]. boundary [DHNH22, EFG^{+}24, NH22, NH23]. box [RCH^{+}20, SO23]. brackish [GGS^{+}21, HPB^{+}21]. branching [RBD22]. Brazil [PLLM21]. Brazilian [BSB^{+}20, HMP^{+}21, LvBS^{+}22, LHD^{+}20]. breakpoints [PKKS20]. breakup [PW22]. Bringing [SHPW22]. brown [LLSQ24, SGGB21]. Brownification [SSM^{+}22a]. browning [MHSL22, PC21, MSB^{+}20]. budget [DdEM^{+}23, GB23, VJH^{+}22]. budgets [BBCA^{+}21, CBC^{+}22]. buffering [HSV22]. builder [GTKW^{+}20]. bulk [BSGA21, CZS^{+}22]. buoy [FIH^{+}20]. buoyant [FGO^{+}20, KWB21]. burial [AMO20, BJK^{+}21, BSB^{+}20, SGA^{+}22, SSEO21, ZWM22]. burrows [XWL^{+}21]. business [AAC^{+}21]. business-as-usual [AAC^{+}21].

C [LRNG24, HZC^{+}24, LRRP20, WWGP21]. C-storage [LRNG24]. C. [PPB^{+}21]. cable [HPB^{+}21, MLK^{+}22]. CaCO [LL23]. caerulea [LMP^{+}20]. calanoid [TCS21]. Calanus [CSH21, CBS^{+}21, PPB^{+}21]. calcification [CGR22, CBC^{+}22, GR24, KNG^{+}23, MCH^{+}22, RVBP23, dOKC^{+}21]. calcifiers [MCH^{+}22]. calcifying [JHS^{+}21]. Calcite [DSY20]. Calcium [GAS^{+}21, HXL^{+}21, SSCE20, SKA^{+}24, SCT^{+}21]. calculation [KTWGT20]. Calibrating [LHGJ21]. California [LMC^{+}21, PMRG20, ACE^{+}22, CMB^{+}21, CM20a, KBL^{+}20, LTW^{+}24, NTJ^{+}20, SWB^{+}20]. californianus [NTJ^{+}20]. Can [DMCW23, ETG21, SYMF21, VMT^{+}20, AR20, BVGQN22, KSN^{+}21, MCS^{+}23, MKM^{+}23, THS21, ZLG^{+}23]. Canada [KSH^{+}23]. Canadian [BAG^{+}24, GB21, GB23, PGEB22, RCMA21]. Candidatus [MLK^{+}22]. cannot [ZLG^{+}23]. canopies [TCP^{+}22]. canopy [CCFF21]. Canyon [CFC^{+}20]. canyons [LNR^{+}20]. capabilities [BHSG^{+}20]. Cape [SCC^{+}21]. capture [JSW^{+}21, ZPK^{+}22]. Capturing [LDGB^{+}23]. Carbon [BMCS23, CM20a, FXD^{+}22, GBMKJ22, JHvL^{+}23, KRM^{+}20, MNM^{+}21, SHS^{+}22, ZSK^{+}21, ACW^{+}22, BND^{+}23, BMR^{+}21a, BCWS23, BRB^{+}20, BSF^{+}21, BJK^{+}21, BCL^{+}22, BSB^{+}20, BDL^{+}21, BPL22, CGC^{+}21, CKCB^{+}23, CEE^{+}22, DGTW21, DWS^{+}22, DCF^{+}22, DdEM^{+}23, DBM^{+}23, DK20, DLL20, EJRRC^{+}23, ESW^{+}21, FGP22, FUUG^{+}21, GBB^{+}21, GLM^{+}23, GAS^{+}23, HCB^{+}22, HMP^{+}21, HANW21, HSBC20, HZM^{+}21, HNS21, HGA24, HLS^{+}22, IFU^{+}20, JHSG^{+}24, KHK^{+}22, KBB^{+}22, KHCG^{+}22, LVDM22, LDGB^{+}23, LDGB^{+}24, LZD^{+}23, LCHB22, LAB^{+}22, MRC^{+}24, MRH^{+}23, MGWS23, MMG^{+}21, MRT^{+}24, MCYR20, MKM^{+}23, MNMJ^{+}21, NKH^{+}23, OBL21, PBP^{+}23, PYY^{+}20, PFH21, PC21, PNT^{+}21, PVSP^{+}22, PJRV23, QRWT^{+}22, RVBP23, RMW^{+}21, RHA^{+}23, RSLW23, RVvB^{+}23, RCH^{+}20, SBD^{+}21, SVMM^{+}22, SSS^{+}24, SZA^{+}21, SML21b, SSEO21, SBL^{+}21, SSS^{+}20, SWK^{+}23, SOT^{+}20, SRM23, SKA^{+}24, SGS22, TEK^{+}21, TKS^{+}20, VCT^{+}23, VST^{+}22]. carbon [WF22, WSZ22a, WSZ22b, WM20, XWL^{+}21, XDQ^{+}22, YAP^{+}22, YXC^{+}22, YSHS23, ZPK^{+}22, ZLU^{+}20, ZJW^{+}22, ZLL^{+}21, ZBNH^{+}21, ZWM22, vGDH^{+}21, OGD20]. carbon-uptake [RCH^{+}20]. Carbonate [EWGP22, GBB^{+}21, BBCA^{+}21, CSD^{+}24, CBC^{+}22, LLK^{+}22, SSCE20, SKA^{+}24, SCT^{+}21]. carbonic [SAD^{+}20]. carboxyl [LPO^{+}20]. carboxyl-rich [LPO^{+}20]. carcasses [HCB^{+}22]. Caribbean [ABC^{+}21, GBB^{+}21, MSL21, MHMML^{+}22]. Carpathian [BRB^{+}22]. cascade [CCY^{+}21]. cascades [CAV^{+}22]. cascading [EG20]. Case [JLMW21, LTJ^{+}20, SWOR20, TMKB^{+}22]. catchment [BCWS23, CZK^{+}23, KSF^{+}23, RNW^{+}23, RS21]. catchments [TKW^{+}23]. catenella [PHM21]. causal [MGWS23]. cause [LDGB^{+}24, MWV22, NKH^{+}23]. caused [GBO^{+}20, RSLW23, SLZ^{+}21b, SKB^{+}22]. causes [AFH^{+}22, Wei24]. Cd [LSL^{+}20, BYB^{+}23, LSL^{+}20]. Ce [BYB^{+}23]. cell [HDD^{+}22, KTWGT20, LSL^{+}20, MRC^{+}24, OBL21]. cellular [LSL^{+}20]. center [CTH^{+}20]. central [AJMG21, PGN^{+}21, HABL^{+}21, SBM^{+}21, TMKB^{+}22, TMS^{+}21]. century [BGM^{+}22, NEB^{+}22, TMKB^{+}22]. Ceriodaphnia [HXL^{+}21, ZLG^{+}23]. Cetacean [WKRL^{+}23]. Cetacean-mediated [WKRL^{+}23]. CH [AFH^{+}21, DRA^{+}23]. Chaetoceros [KMM^{+}22]. chain [MFGF^{+}22, SGM^{+}22]. Challenges [MBD^{+}23, BDE^{+}22, SHS^{+}22]. chance [RS20]. change [AAC^{+}21, ABC^{+}21, BSGA21, BMB22, BDE^{+}22, BPG^{+}22, BFW^{+}22, ETG21, FM21, FMT^{+}20, GGLH^{+}22, KDV^{+}23, LAD^{+}22, LZM^{+}21, LMSN23, MHD^{+}22, MDDL22, MPW^{+}23, NEB^{+}22, RFW^{+}24, SRV^{+}20, SEK20, WHAR^{+}20, WBB^{+}21, WCH^{+}21, ZXE^{+}23, dSNIB22]. changepoint [PW22]. Changes [DNB^{+}20, HZS21, MSB^{+}20, BHG^{+}24, BGM^{+}22, DMC^{+}22, DAD^{+}21, EG20, EPR^{+}24, FLH^{+}23, GLAHH^{+}23, KLO23, KHHZ20, LSC^{+}21, LGQ^{+}21, LXC^{+}21b, MY23, MFM^{+}21, MSD23, RSLW23, RGW^{+}23, SSS^{+}20, SSM^{+}22a, SLZ^{+}21b, SGM^{+}21, XYL^{+}23, ZLS^{+}21, vEMF^{+}20]. Changing [YAH^{+}21, CBPNA22, DKBG21, DK20, GTKW^{+}20, HANW21, HTSG21, JLR^{+}22, KBSA21, LGQ^{+}21, MLL^{+}21, MRT^{+}24, PC21, WOMWR20]. channel [LGF22]. Character [PSNG^{+}21]. characteristics [CZS^{+}22, DGTW21, GPB^{+}20, HHMH22, LGF22, MSB^{+}20, VMT^{+}24]. Characterization [HRK^{+}22, CBN21, EJRRC^{+}23, TMM^{+}20]. characterize [OAM^{+}22]. Characterizing [AHF^{+}22, KdED^{+}20, KDV^{+}23, NGSS23]. charophytes [SKA^{+}24]. chemical [BC20, MRD^{+}21, WOMWR20]. chemistry [BLS20, CGR22, CSD^{+}24, DAB^{+}21, EWGP22, FKN21, HSP^{+}20, MHVC22, NTJ^{+}20, YJB^{+}20, YMSH20, ZPK^{+}22]. chemodiversity [SGSD^{+}22]. chemostasis [AFH^{+}21]. chemosynthetic [AGC^{+}21]. chemotone [AGC^{+}21]. Chesapeake [CCB^{+}20, LZG21, MLK^{+}22, SCB^{+}20, SCT^{+}21, TTC^{+}22, WZCK20]. China [MLL^{+}21, LCZ^{+}21, PYY^{+}20, SLHL^{+}22, WPY^{+}20, XDQ^{+}22, XMP^{+}21, XYL^{+}23, ZZWL22, ZZB^{+}20]. chironomid [MDDL22]. chloride [RD23]. chlorophyll [DNB^{+}20, DMM^{+}21, GF21, QFM^{+}21, SWOR20, SPLJA^{+}20, TMSL22, YJ20, vdPMF^{+}21]. chlorophyll-a [DMM^{+}21, TMSL22, vdPMF^{+}21]. chlorophytes [AWGV20]. Chloropicophyceae [EHC^{+}22]. choice [VKGW22]. choline [MAD^{+}22]. Choquet [PPB^{+}21]. chromophoric [AVJ^{+}24, ZZT^{+}21]. chronic [EFP^{+}21]. Chronically [SIP21]. chronometers [LvBS^{+}22]. Chukchi [RSH^{+}20]. Chydorus [WNHY21]. ciliates [TION23, Wei24]. Circumpolar [DAB^{+}21, YAH^{+}21, YAP^{+}22]. cladoceran [WNHY21]. clam [WOMWR20]. clarity [DAN^{+}22]. classification [LWH^{+}23]. clear [MLC^{+}21]. clear-water [MLC^{+}21]. cliff [CZH^{+}21]. Climate [LCS^{+}20, MLC^{+}21, AAC^{+}21, BPG^{+}22, CLA^{+}24, CLG^{+}24, CMB^{+}21, DMM^{+}21, DK20, DMCW23, Edm21b, ETG21, FM21, FMT^{+}20, HEH^{+}20, HFL^{+}20, HZM^{+}21, KHB22, LLSQ24, LZM^{+}21, LMSN23, MHD^{+}22, MPW^{+}23, NEB^{+}22, OSR20, RFW^{+}24, SBH23, SRV^{+}20, WHAR^{+}20, dSNIB22]. climate-driven [ETG21, HFL^{+}20]. climate-productivity [CLG^{+}24]. climate-related [CLA^{+}24]. climate-sensitive [HZM^{+}21]. climatic [CEE^{+}22, CVF^{+}20, KdED^{+}20, MLP^{+}21a, MDDL22, MCH^{+}21, PMRG20]. climatology [PBGGRB21]. clines [LCS^{+}20]. Clockwise [MAP^{+}22]. clonal [CZH^{+}21]. close [MKP^{+}23]. Cloudy [RS20]. Clytia [CUS21]. CO [ACW^{+}22, LSL^{+}20, YG23, MHSL22, SdBW^{+}23, MZW^{+}21, MCYR20, VGvdB^{+}21, YSHS22, YAE^{+}24, ZZT^{+}21, AFH^{+}21, ACW^{+}22, AKT^{+}21, BLS20, CT23, CTM^{+}21, CCB^{+}20, CSC^{+}21, CMCC24, CHS^{+}20, DRA^{+}23, HGA24, MGP^{+}21, MRT^{+}24, NB20, RRHS^{+}21, RSP^{+}24, SDB^{+}23, SPLJA^{+}20, VLP^{+}21, XWL^{+}21, YMSH20, YSHS22, YSHS23, KMV^{+}20, KMM^{+}22]. Co-acquisition [SdBW^{+}23]. co-existence [MCYR20]. co-limitation [YSHS22, YAE^{+}24]. Co-occurrence [MHSL22, MZW^{+}21, ZZT^{+}21]. co-occurring [VGvdB^{+}21]. coal [BWV^{+}20]. coast [CCDM21, NSK^{+}21, PBP^{+}23, SPKM22]. Coastal [GÖA^{+}23, LHS^{+}21, LMS^{+}21, MIR^{+}22, NEB^{+}22, SSS^{+}20, ZLW21b, AMO20, BND^{+}23, BPIBN20, BMR^{+}21b, BGAD^{+}22, BJS^{+}23b, BWSW24, BKMY21, CSM^{+}23, CTP21, CGD22, CPBJR^{+}21, CSNS24, CCG^{+}22, DCJB22, EJRRC^{+}23, FBR^{+}23, FBJ^{+}23, FWDY20, HRK^{+}22, HPB^{+}21, HSV22, IÁSNCR21, JLW^{+}23, KLO23, KHK^{+}22, KdED^{+}20, KSF^{+}23, KRM^{+}20, LRNG24, LTH21, LCHB22, LRM21, MHG^{+}23, MBC^{+}21, MAD^{+}22, MCS^{+}23, MKD22, NSK^{+}21, NGSS23, PYY^{+}20, PLLM21, QS21, SOSN23, SMG^{+}22, SCC^{+}21, SRV^{+}20, SLZ^{+}21b, SOT^{+}20, SSR^{+}23, SSB^{+}22, TNK^{+}22, TJW^{+}22, THVJ23, TMKB^{+}22, WHAR^{+}20, WWB^{+}20, WHZ^{+}21, YML^{+}21, YGD^{+}21, ZSD^{+}20, ZJW^{+}22, ZSZ^{+}22, ZWM22, vHLV^{+}20, vHHP^{+}21, vLWV^{+}20]. coastline [BGAD^{+}22]. cobalt [KMV^{+}20, WXK^{+}22]. coccolithophore [GBB^{+}21, JLR^{+}22, ZLS^{+}21]. coccolithophores [GDB20]. coefficient [ZXE^{+}23]. coerulea [MBC^{+}21]. coexistence [LTJ^{+}20]. coherent [FHR^{+}21]. Cold [BFW^{+}22, GR24, KKP^{+}21, MACC24, RCS^{+}20]. cold-adapted [KKP^{+}21]. Cold-water [BFW^{+}22, GR24, MACC24]. colimitation [LSL^{+}20]. colonial [DTPV21]. colonies [EBW^{+}20, SdBW^{+}23]. colonization [FPG^{+}24, MKBSK19]. colonizing [KL22, dOKC^{+}21]. colony [KOR^{+}20, ZSZ^{+}21]. colony-forming [KOR^{+}20]. color [GBHF21]. Colorado [DSY20]. colored [CM21]. column [BRB^{+}20, HMB^{+}22, KRM^{+}20, LHS^{+}21, LDS^{+}20, PSY^{+}22, SGSD^{+}22, SGK^{+}21, TZF^{+}22, TRS^{+}20]. combinations [GRJ20]. Combined [IFU^{+}20, WHAR^{+}20, SVC^{+}22, SN22]. Combining [APB^{+}23]. Comment [LD22, PPB^{+}21, CSH21, SB22, ZR23]. commercial [HSP^{+}20]. common [BMH^{+}23, BWSW24, MLB^{+}21]. commonality [LTQS^{+}21]. communities [BBH^{+}20, BAD^{+}21, CGRGP^{+}21, CCY^{+}21, DVL^{+}23, EDC24, EJRRC^{+}23, FBP20, FR23, GCW^{+}23, GMS21, HBS^{+}22, HBY^{+}22, HPCH24, IMB^{+}22, KPNJ23, KPS21, LRNG24, LNR^{+}20, LBG^{+}22, MXL^{+}21, MDP^{+}23, MHdS^{+}21, MOW^{+}22, PKKS20, SORGC^{+}21, SPHH22, TMS^{+}21, VJH^{+}22, WPY^{+}20, XKP^{+}22, XYL^{+}23, YML^{+}21, ZQB^{+}21, dMKB^{+}20, dOKC^{+}21, vEMF^{+}20]. community [AR20, AFK^{+}22, BHSG^{+}20, BMA^{+}20, BMLV24, BVGQN22, CB21b, CVF^{+}20, DBFC24, FWvA^{+}21, FAB^{+}20, FKN21, GSP^{+}23, GGHBS^{+}20, HZS21, HLS^{+}22, KHHZ20, MDDL22, MVBG20, OYM^{+}20, PKV20, RvHF^{+}20, RPKC22, SH22, SGL^{+}23, SSM^{+}22a, SBV^{+}21, SGSD^{+}22, TPTS23, WBB^{+}21, WCZ^{+}20, WMRL22, YAE^{+}23]. community-environment [GGHBS^{+}20]. Comparative [MK20, MVBG20, HTLP23]. Comparing [LHD^{+}20, SHZ^{+}21, XTJW21, WTvD^{+}22]. Comparison [EWF22]. compensation [BWB^{+}22, dJGCS23]. competition [DWW^{+}21, MKL^{+}24, PCC^{+}23]. competitive [SDSGR21]. Complex [WSB^{+}20, ASL^{+}24, BRB^{+}22, FBRG23, LSC^{+}21, WNHY21]. complexation [WOK^{+}21]. complexes [KLP23, dCLO^{+}23]. Complexities [LCM^{+}22]. complexity [CLC^{+}22]. complicate [MSJ21]. complicated [WPO^{+}22]. component [DMS^{+}21]. components [AMC21, ZZB^{+}20]. composition [Ano21c, BHSG^{+}20, BPIBN20, BYS^{+}23, BCFI^{+}22, CPB^{+}21, CSM^{+}23, DVL^{+}23, DvOH^{+}20, DBM^{+}23, EHC^{+}22, FBRG23, GAS^{+}23, HANW21, HZS21, HCS^{+}21, JLS^{+}20, KHB22, KAP^{+}20, KSR20, LBBM21, LSC^{+}21, LXC^{+}21b, MO23, PGEB22, PATL^{+}23, PGN^{+}21, RvHF^{+}20, RLP^{+}20, RRAO^{+}21, SSM^{+}22a, SEK20, SBV^{+}21, SSB^{+}22, SGSD^{+}22, TPTS23, TATC^{+}24, TUR^{+}20, WRWPG19, WCZ^{+}20, XTJW21, ZLIP20, ZZT^{+}21, ZYZ^{+}21, dMKB^{+}20, vdPAV^{+}20, vdPMF^{+}21]. compositional [BSF^{+}21]. compositions [SVMM^{+}22]. compound [DMS^{+}21]. compound-specific [DMS^{+}21]. compounds [EBK^{+}20, GDB20, LPO^{+}20, MZXY22, MY23, MRD^{+}21]. comprehend [MMLPRHCG22]. computed [IKN^{+}21, MMLPRHCG22]. concealed [DKW^{+}21]. Concentration [TUR^{+}20, AFH^{+}21, ARPFGS23, CHS^{+}20, CZS^{+}22, DNB^{+}20, HXL^{+}21, LD22, RSP^{+}24, SOSN23, SB22, SZA^{+}21, VCBB20]. concentration-discharge [AFH^{+}21]. concentration-driven [LD22, SB22]. Concentrations [DWS^{+}22, BTT23, BCTH20, BAG^{+}24, DRA^{+}23, DKBG21, EBK^{+}20, RCMA21, TKW^{+}23, TMSL22]. concepts [GM22]. Conceptual [RST^{+}21]. Concurrent [HSB^{+}20]. conditions [AH20, CM21, CVF^{+}20, LSW^{+}20b, MLS^{+}21, PHM21, SN22, SBL^{+}21, TKM^{+}22, WOMWR20, ZLS^{+}21, ZXZ^{+}23]. congruence [GWS^{+}21]. conjugated [SJH20]. conjunction [PMRG20]. connecting [MBD^{+}23]. Connectivity [BGAD^{+}22, CPBJR^{+}21, DCJB22, DPB^{+}23, DFJ22, GLB^{+}22, JSB^{+}20, MACC24, VVMJ23]. Consequences [KTWGT20, MBA^{+}21, RGP^{+}22, MLP21b, Wei24]. conserved [LABM20, SSH^{+}22]. consideration [WSZ22a, WSZ22b]. consistent [EFP^{+}21]. consortia [CYPRG^{+}23]. constituents [KSF^{+}23]. constitutes [PPB^{+}21]. constrain [KVKS22, ORRMD^{+}21]. Constraining [MBD^{+}22]. constrains [MMG^{+}21]. constraints [FDHS20, FXD^{+}22, GFW^{+}20, LSDA21, TSX^{+}23, WLK^{+}21, ZPK^{+}22]. consumer [HHK^{+}22]. consumers [MV22a, MV22b, VGSB23]. consumes [LES^{+}21]. consuming [SCB^{+}20]. consumption [CGC^{+}21, HP20, KHSP21, MP23, MGG^{+}24, TZF^{+}22]. contact [XLN21]. containing [ZLW^{+}21a]. contamination [LMS^{+}21]. content [BMCS23, HLKD23, HNS21, HGA24, LLMM^{+}23, MNM^{+}21, OBL21]. contents [XSZ^{+}21]. conterminous [FBK^{+}22, GGR24, LWH^{+}23]. contextually [HMB^{+}20]. contiguous [BCTH20]. continent [LL23]. continental [BMA^{+}20, CHP^{+}24, DMM^{+}21, HTLP23, HM20, KTH^{+}21, KPS21, LvBS^{+}22, MZXY22, MY23, PGEB22, QRWT^{+}22, THT^{+}20, WRC^{+}21, ZLU^{+}20]. contingency [GGLH^{+}22]. continuous [LWH^{+}23, SEG21]. continuum [BMLV24, EJM^{+}23, ZYZ^{+}21]. contrasted [IJSC20]. Contrasting [CGB^{+}21, DFJ22, DKPS23, GFG^{+}22, GBC^{+}20, MLK^{+}22, RBR^{+}23, WTvD^{+}22, WRHR22, BCL^{+}22, CGL^{+}21, LZM^{+}21, VMP^{+}21]. contrasts [BHSG^{+}20, RHP^{+}23]. Contrib [Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano20y]. contribute [CKCB^{+}23, HCB^{+}22, HVMS^{+}21]. Contribution [LLL^{+}23, RSP^{+}24, VGSB23, LLSQ24, LAMG^{+}22, PJRV23, PMJ^{+}21, SBD^{+}21]. Contributions [XMP^{+}21, HLH^{+}23, LLTT^{+}24, OGD20, WF22]. contributor [GKW^{+}20, GKW^{+}21]. Control [CLD^{+}23, HPP^{+}21, AFH^{+}21, GÖA^{+}23, JvKvT^{+}20, LD22, LGF22, LLBC20, PCJ^{+}21, RS20, SB20, VLP^{+}21, YOXC23, vdPMF^{+}21]. controlled [BKMY21, KPM^{+}20, KRM^{+}20, SCT^{+}21, ZZL^{+}21]. controlling [KNG^{+}23, PCC^{+}23]. Controls [HSV22, LD22, TKW^{+}23, ARBvS^{+}21, ASH^{+}20, CFC^{+}20, ESW^{+}21, FBK^{+}22, HBY^{+}22, HPF^{+}23, LHS^{+}21, LCZ^{+}21, MLK^{+}22, MT23, PMD^{+}23, YMSH20, SB22]. conundrum [BC20]. convection [AHF^{+}22]. conversion [SGSF20]. cope [VMT^{+}20, WTvD^{+}22]. Copepod [EWF22, KSI24, CHAA21, LDB^{+}21, MSGW^{+}20, SYMF21, TION23, TVB^{+}21, VMP^{+}21, XSZ^{+}21]. copepods [BJS23a, CHA^{+}24, HMBB21, KL22, LAFV23, LTJ^{+}20, PJRV23, RTK23, SMF20, TCS21, dJGCS23]. Coping [CBT^{+}21]. copper [KP20, WOK^{+}21]. copy [ZR23]. Copyright [Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21-53, Ano21p, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22a, Ano22q, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano24a, Ano24b, Ano24c, Ano24d]. Coral [FKN21, KDW^{+}23, MHMML^{+}22, SSL22, BST^{+}22, BCFI^{+}22, BBCA^{+}21, BFW^{+}22, CÖL^{+}20, CDW^{+}24, CBC^{+}22, Edm21a, EFP^{+}21, FHR^{+}21, GCMJ23, GLB^{+}22, HPCH24, LMS23, LBG^{+}22, LRRP20, MCH^{+}22, MGP^{+}21, MSD23, MACC24, MMLPRHCG22, MDIY23, OLFH21, RMH^{+}22, RBD22, RCH^{+}20, SHBA22, SWZ^{+}23, SSCE20, VVMJ23, dOKC^{+}21]. coralline [GR24, KNG^{+}23, WCH^{+}21]. corallite [MMLPRHCG22]. corals [ABC^{+}21, Ano21c, CTM^{+}21, Edm21b, FPBG^{+}22, LRdGFP23, PMJ^{+}21, RFB^{+}22, WRWPG19]. core [LTQS^{+}21]. cornuta [HXL^{+}21, ZLG^{+}23]. correlate [MSD23]. corridors [DCJB22]. Corrigendum [Ano21a, Ano21b, Ano21c, CLC^{+}21, GSG^{+}21, GKW^{+}21, LEK^{+}22, MV22a, Phi21, WSZ22a]. cosmopolitan [MRT^{+}24]. count [MBD^{+}23]. counter [YAH^{+}21]. counteracts [GGLH^{+}22]. Countergradient [DKW^{+}21]. Coupled [EG20, HPB^{+}21, RHP^{+}23, ACW^{+}22, BRCO21, BVM^{+}20, LZG21, MBLA^{+}21, MK20, SBB20]. coupling [HJWP24, KPL^{+}20]. covariance [BBMD20, VLP^{+}21]. cover [BGM^{+}21, MGI^{+}22]. covered [CVF^{+}20, LDS^{+}20, LXC^{+}21b, PMD^{+}23, SSM^{+}21a, SMCS23]. crab [XWL^{+}21]. crabs [CJB^{+}21]. Crassostrea [MDWT^{+}22, WSB^{+}20]. creek [RWMP^{+}23]. creeks [WPO^{+}22]. criteria [MGWS23, MSJ21]. Crocosphaera [DFW^{+}20, ZFQ^{+}20]. crop [JHSG^{+}24]. Cross [FGO^{+}20, DAN^{+}22, GLN^{+}20, GLN^{+}21, GPDA^{+}20]. cross-scale [DAN^{+}22]. cross-shelf [GPDA^{+}20]. Cross-shore [FGO^{+}20, GLN^{+}20, GLN^{+}21]. crossflow [JYA21]. crowding [ZLG^{+}23]. cruise [ZZO^{+}21]. crustacean [RVBP23, SO23]. crustaceans [BZSF23, WNHY21]. crustose [GR24, WCH^{+}21]. Cruz [TZF^{+}22]. cryosphere [HGH^{+}21]. cryptophyte [CT23]. ctenophores [CCG^{+}22]. cues [BK20, BMGK22, ZSZ^{+}21]. cultured [LHGJ21]. Current [ACE^{+}22, CMB^{+}21, Lem20, LTW^{+}24, SN22, SSL22, SYMF21, VMT^{+}20, BFCC^{+}21, DAB^{+}21, KBL^{+}20, SWB^{+}20]. currents [FBRG23, SLJ^{+}22]. cyanobacteria [AWGV20, DCK^{+}22, JSPS24, LBvE23, LEK^{+}22, MTP^{+}23, PSNG^{+}21, WOT^{+}23, WWFS21, ZJ22a, ZJ22b, ZFQ^{+}20]. cyanobacteria-derived [PSNG^{+}21]. cyanobacterial [EWGP22, FBP20, HMB^{+}22, KRC23, MBB^{+}20, Roy20, SRL^{+}20, XMP^{+}21]. cyanobacterium [BMH^{+}23, DTPV21, GFW^{+}20, SDS^{+}23, YG23].

Cyanothece [hTRH20]. cycle [BPL22, CVF^{+}20, GBO^{+}20, LAMG^{+}22, SSS^{+}20, XYZ^{+}21, ZSD^{+}20]. cycles [GMMW20, LLMM^{+}23, NH22, TKH21, vdPMF^{+}21]. Cyclical [LMSN23]. cycling [ABS^{+}21, BYB^{+}23, BCA^{+}20, CCK^{+}21, CBF^{+}23, CLD^{+}23, DNA^{+}22, EFG^{+}24, FMT^{+}20, HSBC20, HTSG21, HDK^{+}21, JHvL^{+}23, KWS^{+}21, KHCG^{+}22, LBG^{+}22, MPW^{+}23, ORRMD^{+}21, PCJ^{+}21, RLL^{+}21, RGP^{+}22, SLZ^{+}21a, SBW^{+}20, THL^{+}21, WCC20, WRS^{+}22, WOK^{+}21, XMP^{+}21, ZLL^{+}24]. Cyclone [BJS^{+}23b, BPL22, WCC20]. cyclone-anticyclone [WCC20]. Cyclonic [ZBNH^{+}21, ACE^{+}22]. cylindrus [MLG^{+}20]. Cymodocea [JREVB21]. cytometry [FHMK20].

Daily [CGGC^{+}21, LSS^{+}24]. Dakota [THT^{+}20]. damage [vHHP^{+}21]. dampen [dJGCS23]. damping [LHH24, SN22, ZLN21]. Danish [GBMKJ22]. Daphnia [DKW^{+}21, DMCW23, LBvE23, MTB^{+}21, MKL^{+}24, RVBP23, RBL^{+}20, VGvdB^{+}21, YWZ^{+}21, ZSZ^{+}21]. Dardanelles [CÖL^{+}20]. dark [KGG^{+}23, vdPAV^{+}20]. darkening [GÖA^{+}23]. darkness [CBT^{+}21, CGB^{+}21, MLG^{+}20, QBJG21]. data [ASH^{+}20, BMB22, CZK^{+}23, DAD^{+}21, MVP^{+}20, OAM^{+}22, PSY^{+}22, WF22]. data-limited [ASH^{+}20]. dataset [PPB^{+}21]. day [EWGP22, GK22]. de-eutrophication [DNB^{+}20]. dead [BK20, HCB^{+}22, MHMML^{+}22, dOKC^{+}21]. death [DFW^{+}20]. decadal [BBMD20, LDB^{+}20, LDB^{+}21, SMP21]. decades [BMB22, GGLH^{+}22]. decay [HKM^{+}20]. Deciphering [BND^{+}23]. decline [FBK^{+}22, YLT^{+}20]. declines [BWB^{+}22, HDD^{+}22]. declining [PGN^{+}21]. decomposition [BZK^{+}22, ZLL^{+}21]. Deconvolving [WDS^{+}23]. decoupled [DMC^{+}22]. decouples [SCT^{+}21]. decoupling [NKH^{+}23, ZBNH^{+}21]. Decrease [KLO23, FBV23]. Decreased [HXL^{+}21]. decreases [LLTT^{+}24, RVBP23, ZLL^{+}21]. Decreasing [MGP^{+}21]. Deep [CKCB^{+}23, HMM^{+}22, SWOR20, AGC^{+}21, AHF^{+}22, BMR^{+}21a, BZSF23, CYZZ^{+}20, CLW^{+}21, CPBJR^{+}21, FWDY20, GLAHH^{+}23, HJWP24, Lem20, MBLA^{+}21, MKBSK19, MOW^{+}22, MGG^{+}24, OBB^{+}20, PFH21, PCN22, REF^{+}21, SJH20, SSM^{+}21a, SSP^{+}20, SPLJA^{+}20, SSWM20, SDS^{+}23, SHPW22]. deep-pelagic [BZSF23]. Deep-sea [HMM^{+}22, AGC^{+}21, BMR^{+}21a, CLW^{+}21, MBLA^{+}21, MKBSK19, OBB^{+}20, PCN22]. deep-silled [HJWP24]. deeper [Ano21a, GSG^{+}17]. deeply [CYPRG^{+}23]. defense [HXL^{+}21]. defensive [ZSZ^{+}21]. deficient [AHCJ^{+}20, BYB^{+}23, CBB^{+}22, CKCB^{+}23, DNA^{+}22, EBK^{+}20, LVHK20, MLL^{+}21, SMW^{+}21]. define [CSB22]. Defining [PPB^{+}21]. definition [BO20]. Deforestation [GBO^{+}20]. deformation [GLN^{+}20]. degradation [BMG^{+}23, CPB^{+}21, FXD^{+}22, GAS^{+}23, LLC21, PWS^{+}20, RLP^{+}20, YAH^{+}21]. degree [KSN^{+}21]. delays [RVBP23]. delineated [BSF^{+}21]. delineation [BGAD^{+}22]. delivery [GPDA^{+}20]. Delta [PYY^{+}20, CMCC24, CM21, YLT^{+}20, LTH21]. delta-front [YLT^{+}20]. deltaic [LTH21]. Delving [GSG^{+}17, Ano21a]. demand [BMA^{+}20, MHG^{+}23, MGG^{+}24]. demands [BMR^{+}21a]. demersal [TGE^{+}21]. demography [MMLPRHCG22]. demonstration [SVBT20]. Denitrification [CEW^{+}21, CHP^{+}24, ACA^{+}22, RWS^{+}22]. denitrifier [LPMBR23]. denitrifying [ZQB^{+}21]. Density [CHAA21, HZC^{+}24, YJB^{+}20]. Density-dependent [CHAA21]. Dependency [KLG^{+}23, GZX^{+}22]. dependent [BMLV24, CHAA21, GNO22, HGH^{+}21, KP20, LMP^{+}20, MWHR^{+}22, OBL21, SZY^{+}20, WHZ^{+}21, ZLS^{+}21]. depleted [SBW^{+}20]. depletion [SMCS23, SSWM20]. deposition [BFE^{+}20, GRJ^{+}22, SKA^{+}24, YAE^{+}23, ZLW21b, vEMF^{+}20]. depositional [HMB^{+}20]. deposits [KRM^{+}20]. deprivation [CBT^{+}21]. Depth [DWS^{+}22, KVKS22, LMP^{+}20, BSGA21, BMS^{+}22, CÖL^{+}20, CEE^{+}22, DLC^{+}22, FGO^{+}20, GGR24, GLN^{+}21, LAFV23, MTB^{+}21, PMF^{+}21, SO23, SMKS22]. Depth-dependent [LMP^{+}20]. depth-keeping [FGO^{+}20]. depth-related [BSGA21]. depths [LMS23, SSH^{+}22]. derived [CDW^{+}24, DVL^{+}23, DLDF21, PSNG^{+}21, SWZ^{+}23, SHE^{+}21, TNK^{+}22, TKS^{+}20, vGDH^{+}21]. desert [BHR^{+}23, GRJ^{+}22, RGP^{+}22]. design [KFHS24, RWT^{+}21]. despite [KTA^{+}21, LMM^{+}21, MKP^{+}23]. detect [SYMF21]. Detecting [CLA^{+}24]. deteriorating [QMG21]. deterioration [HSB^{+}20]. determinations [ADLW22]. determine [BHG^{+}24, CSM^{+}23, CUS21, GF21, LPMBR23, PGB24, PHM21, TATC^{+}24, WOT^{+}23, dSNIB22]. determines [JSB^{+}20]. determining [CZH^{+}21]. Deterministic [PATL^{+}23]. Detritivore [FPG^{+}24]. Detritus [BVL^{+}23, LMP^{+}20, SPKM22]. Detritus-hosted [BVL^{+}23]. development [MOW^{+}22, MNMJ^{+}21, OYM^{+}20]. Diatom [CMB^{+}21, XSZ^{+}21, AR20, BVGQN22, CGRGP^{+}21, CMT^{+}20, FR23, GBC^{+}20, HDS^{+}20, KSN^{+}21, KMM^{+}22, KKP^{+}21, KP20, LAMG^{+}22, MGC^{+}20, MRT^{+}24, MLG^{+}20, OBL21, QBJG21, STB22, SKB^{+}22, WRS^{+}22, YSHS22, YSHS23, ZLIP20]. diatom-dominated [GBC^{+}20]. diatom-rich [LAMG^{+}22]. Diatoms [DKBG21, ACE^{+}22, AWGV20, BRCO21, CLGH23, CGB^{+}21, FLH^{+}23, HVMS^{+}21, KMV^{+}20, KOR^{+}20, KLP23, LSL^{+}20, LHGJ21, MAP^{+}22, MLT^{+}21, RWL^{+}24, WRHR22, ZLL^{+}21, vdPAV^{+}20]. diazotroph [ALG^{+}23, LTLH22, MBD^{+}22, SCC^{+}21, SWZ^{+}23]. diazotroph-derived [SWZ^{+}23]. diazotrophic [DCK^{+}22, HBS^{+}22, JSPS24]. Diazotrophs [WWFS21, FBP20, GFW^{+}20, RHS^{+}22]. Diazotrophy [WOT^{+}23]. Dichtyochophyceae [LES^{+}21]. die-off [SVBT20]. dieback [HFL^{+}20, YOXC23]. Diego [GNO22]. Diel [KTWGT20, BHG^{+}24, BJS23a, GNO22, LGI^{+}20, MK21]. diet [DvOH^{+}20, TION23, VST^{+}22]. differ [BSIdG23, BMLV24, SWS^{+}22]. differences [HMBB21, VGvdB^{+}21, XSZ^{+}21]. Different [HEH^{+}20, LPO^{+}20, SML21b, BRCO21, BRB^{+}22, CGGC^{+}21, KSN^{+}21, MTB^{+}21, MHMML^{+}22, RBR^{+}23, SSL22, XSZ^{+}21, ZLS^{+}21, ZZB^{+}20]. Differential [KBB^{+}22, RWL^{+}24, VKB^{+}22]. differentiated [LHH24]. differently [BWSW24, WWFS21]. differs [BRB^{+}20, JLS^{+}20]. diffuse [ZXE^{+}23]. diffusion [ZLIP20]. digitale [MMKWZ22]. dimensional [CLT^{+}22]. dimensions [LHL^{+}21]. dimethyl [MT23, XYZ^{+}21]. dimethylated [MZXY22, MY23]. dimethylsulfide [GMNG^{+}24]. dimethylsulfoniopropionate [ORRMD^{+}21]. dimethysulfoxide [HDK^{+}21]. dimictic [ACA^{+}22, PW22]. dinitrogen [SBL^{+}21]. dinoflagellate [CGRGP^{+}21, FHMK20, PHM21]. dinoflagellates [CLGH23, FHMK20]. dioica [MLS^{+}21]. dioxide [BCL^{+}22, CM20a, DGTW21, HNS21, RVBP23, RHA^{+}23, SBL^{+}21, SWK^{+}23, XDQ^{+}22]. dipole [LXC^{+}21a, WCC20]. direct [MEQBV20]. direction [WWB^{+}20]. directly [FWvA^{+}21]. discharge [AFH^{+}21, FKN21, IÁSNCR21, KRB^{+}21, KRM^{+}20, MGI^{+}22, PCN22, RTM^{+}21, SZA^{+}21, WLK^{+}21, vdPMF^{+}21]. discharging [BKMY21]. disconnected [BGAD^{+}22]. discovery [DNA^{+}22]. Disease [ARB^{+}22]. disentangled [GMNG^{+}24]. Disko [MN20]. disparate [SMW^{+}21]. Dispersal [PHM21, DCJB22, GCW^{+}23, MLS^{+}21, ZZL^{+}21, ZSZ^{+}22]. dispersion [CTH^{+}20, RPA^{+}22]. disproportionately [BKC20]. disruption [RS20]. dissimilatory [CEW^{+}21, RWS^{+}22]. dissipation [PLLM21]. dissipative [EWF22]. dissolution [BFW^{+}22, EBW^{+}20, LLK^{+}22, LN21, MSG^{+}22, RWL^{+}24, SSCE20, SCT^{+}21]. dissolvable [MCS^{+}23]. Dissolved [BMR^{+}21a, BCWS23, BSF^{+}21, MRD^{+}21, PNT^{+}21, RMW^{+}21, SGS22, AVJ^{+}24, AKT^{+}21, AMC21, BND^{+}23, BMCS23, BPIBN20, BAD^{+}21, BGG^{+}20, BYS^{+}23, BGM^{+}21, CGGC^{+}21, CPB^{+}21, CTP21, CLC^{+}22, CM21, CCDM21, DVL^{+}23, DPB^{+}23, EJRRC^{+}23, FAB^{+}20, GB23, HYPG^{+}22, HOB^{+}22, HP20, JSL^{+}23, JSB^{+}20, KAP^{+}20, KMV^{+}20, KDZ^{+}24, KHSP21, KDV^{+}23, LBBM21, LRdGFP23, LMM^{+}21, LD22, LPMBR23, LSC^{+}21, LXC^{+}21b, LPO^{+}20, LHMR23, LMS^{+}21, LSS^{+}24, MGL22, MCS^{+}23, MHP^{+}23, PBC21, PCJ^{+}21, PSNG^{+}21, PWS^{+}20, PC21, QRWT^{+}22, RVBP23, RHA^{+}23, RWMP^{+}23, RSLW23, RCMA21, SJH20, SB20, SB22, SZA^{+}21, SOT^{+}20, SEK20, SCT^{+}21, SGSD^{+}22, TSX^{+}23, VST^{+}22, WMRL22, YWLY20, ZZT^{+}21, ZYZ^{+}21, ZZB^{+}20, dMKB^{+}20]. Distinct [AFH^{+}21, CBD^{+}23, LHMR23, WCZ^{+}20, XYL^{+}23, LPO^{+}20, LNC^{+}23, SJJ^{+}21, SKB^{+}22, WWGP21]. distinguish [WRMP22, ZLG^{+}23]. Distinguishing [DMS^{+}21, DAB^{+}21]. distributed [HGF^{+}21]. Distribution [BCA^{+}20, BRB^{+}22, dCLO^{+}23, BFCC^{+}21, BBGM20, CBHB21, DCJB22, EBK^{+}20, FBRG23, GGR24, KBL^{+}20, LBBM21, LL23, MLK^{+}22, OFM^{+}21, RBR^{+}23, SMKS22, SCM22, TRS^{+}20, VSB^{+}23, ZZO^{+}21]. distributional [GBC^{+}20]. Distributions [WOK^{+}21, EPR^{+}24, HBG^{+}22, JSL^{+}23, LAFV23, LWC^{+}24, LLBC20, MT23, PYY^{+}20, YAH^{+}21]. Disturbance [WPO^{+}22, BPG^{+}22, LCM^{+}22, SYMF21, ZSD^{+}20, ZTT23]. disturbances [LDB^{+}20, RPKC22]. disturbed [BCWS23]. diurnal [BWSW24]. Divergent [QRWT^{+}22, BRB^{+}22]. Diverse [FM21, GSP^{+}23, BMCS23, DBFC24, KSN^{+}21, LABM20, MCH^{+}22]. Diversification [Gil22]. diversity [BOC^{+}21, BBL^{+}21, CML23, CLC^{+}22, CKH^{+}21, CLP^{+}24, FBV23, GCEM^{+}23, GZX^{+}22, HMM^{+}22, LHL^{+}21, LWS^{+}22, MHdS^{+}21, MKP^{+}23, PYM^{+}20, PGEB22, RPKC22, SLHL^{+}22, SSK24, VSB^{+}23, WNHY21, XTJW21]. division [KTWGT20]. DNA [CSM^{+}23, CBN21, GWS^{+}21, OGD20, SLHL^{+}22]. DNA-based [CBN21, GWS^{+}21]. DNRA [CHP^{+}24]. Do [WRA^{+}23, KGG^{+}23, RMH^{+}22, SEK20, ZXZ^{+}23]. DOC [BMR^{+}21a, HLKD23, MBH^{+}23, MFGF^{+}22]. does [CHS^{+}20, LPA^{+}23, RBD22, ZZCD23]. doi [Ano21a]. Dolioletta [FLW^{+}21]. doliolid [FLW^{+}21]. DOM [BHSG^{+}20, BKC20]. Dominance [NH22, FHMK20, HEH^{+}20, PHM21, SDS^{+}23]. dominant [BSIdG23, BMR^{+}21a, KSN^{+}21, KRB^{+}21, ONH^{+}20, SYT^{+}21]. dominate [FBP20, MSG^{+}22]. dominated [CMCC24, DKPS23, GBC^{+}20, HBS^{+}22, JvKvT^{+}20, KVKS22, KDV^{+}23, NDP^{+}24, OMM21, QS21, RMW^{+}21, SH22, WCZ^{+}20]. dominates [VVMJ23]. Dongshan [XYL^{+}23]. DOP [BPIBN20]. dormant [ZJ22b]. Dosidicus [PMRG20]. double [SBH23]. down [BMA^{+}20, BD22, GÖA^{+}23, LLBC20, RS20]. downslope [NH22]. downstream [RD23]. drainage [AFH^{+}22, BWV^{+}20]. drift [FGO^{+}20, LLC21, SBM^{+}21]. drifters [CTH^{+}20]. drive [AWGV20, AR20, DPB^{+}23, DCF^{+}22, DLC^{+}22, HMB^{+}20, LCS^{+}20, MV22a, MV22b, PBC21, PATL^{+}23, PSK21, SLJ^{+}22, SORGC^{+}21, SSS^{+}24, SGM^{+}21, XDQ^{+}22]. driven [AHF^{+}22, BVGQN22, CGRGP^{+}21, CSC^{+}21, ETG21, EPR^{+}24, FBRG23, FHR^{+}21, GBC^{+}20, GKW^{+}20, GKW^{+}21, HFL^{+}20, HPP^{+}21, LD22, LMSN23, MCH^{+}22, MKL^{+}24, SB22, SLZ^{+}21b, TEK^{+}21, ZCZ^{+}23, vdPMF^{+}21]. driver [NvBE^{+}21, SGS22, ZZWL22]. Drivers [ACW^{+}22, CCA^{+}21, FHMK20, IOS24, MFM^{+}21, MAG^{+}22, VSB^{+}23, BAG^{+}24, BMS^{+}22, CGGC^{+}21, CBPNA22, DAN^{+}22, DRA^{+}23, GCMJ23, GSD^{+}23, GLB^{+}22, HTLP23, HHK^{+}22, JHBB22, KAP^{+}20, KdED^{+}20, KOR^{+}20, KAC^{+}23, KBvdD^{+}20, KDV^{+}23, LDGB^{+}24, LHL^{+}21, LHMR23, MBB^{+}20, PGEB22, RWMP^{+}23, RKBT^{+}23, SLHL^{+}22, TMKB^{+}22, VCBB20, WRC^{+}21, WRMP22, ZSF^{+}23, vLWV^{+}20, CHS^{+}20]. drives [DBB^{+}21, DBM^{+}23, KDZ^{+}24, KBL^{+}20, LBL^{+}24, MHdS^{+}21, RTM^{+}20, RTM^{+}21]. driving [CZH^{+}21, CLP^{+}24]. dry [YAE^{+}23]. drying [CTF^{+}21, SLRS^{+}24]. Dual [TMM^{+}20, LLC^{+}23]. due [AFH^{+}21, BHSL21, BGM^{+}22, MLP21b, XSZ^{+}21]. duration [ABC^{+}21, CTF^{+}21, PW22]. during [ACA^{+}22, AJMG21, Ano21c, BGM^{+}22, BPL22, CCC^{+}21, CM20b, CLP^{+}24, FPG^{+}24, GB23, GK22, HRK^{+}22, HKM^{+}20, KRC23, KTH^{+}21, KGG^{+}23, MLS^{+}21, MDDL22, MAP^{+}22, PFH21, SBM^{+}21, SGA^{+}22, VJH^{+}22, WRWPG19, WSB^{+}20, WMRL22, XTL^{+}20, YWLY20, ZLW21b, vHHP^{+}22, vdPAV^{+}20]. dust [BFE^{+}20, GRJ^{+}22, LCHB22]. dye [CTH^{+}20]. Dynamic [LXC^{+}21b, RRAO^{+}21, KDV^{+}23, WRMP22]. Dynamics [CMT^{+}20, ZLU^{+}20, AFH^{+}21, ACW^{+}22, ACA^{+}22, AWB^{+}23, AR20, AFM23, AVS^{+}22, ASL^{+}24, BJG^{+}22, CGRGP^{+}21, CZH^{+}21, CÖL^{+}20, CAV^{+}22, CBHB21, CGL^{+}21, CMB^{+}21, CKCB^{+}23, CLD^{+}23, DCF^{+}22, DGW^{+}21, DFW^{+}20, FWvA^{+}21, FKN21, FSH^{+}23, GGLH^{+}22, HJWP24, HPB^{+}21, HPF^{+}23, HMB^{+}20, HS21, HLS^{+}22, JvKvT^{+}20, JHBB22, KRB^{+}21, KdED^{+}20, KDW^{+}23, KWB21, LVDM22, LMC^{+}21, Lem20, LHS^{+}21, LN21, LZM^{+}21, LNC^{+}23, MCH^{+}22, MSD23, MAD^{+}22, MRD^{+}21, MIR^{+}22, MPW^{+}23, MAP^{+}22, NB20, PNT^{+}21, QRWT^{+}22, RvHF^{+}20, RTM^{+}20, RTM^{+}21, RBR^{+}24, RRHS^{+}21, RHM^{+}21, SPHH22, SBB20, SO23, SLRS^{+}24, SLZ^{+}21b, STAK22, VJH^{+}22, WOMWR20, XMP^{+}21, ZSF^{+}23, vHHP^{+}22].

Earlier [PW22]. early [KPG^{+}23, MDWT^{+}22, OFM^{+}21, SBM^{+}21, VCT^{+}23, vdPAV^{+}20]. Earth [CLA^{+}24]. east [LHS^{+}20, MLL^{+}21]. Eastern [HABL^{+}21, AHCJ^{+}20, CDH^{+}21, CJB^{+}21, DNA^{+}22, ESW^{+}21, EBK^{+}20, RCMA21, SORGC^{+}21, WRS^{+}22, BYB^{+}23, EBK^{+}20, ETMD23, FSS^{+}23, LMM^{+}21, MDWT^{+}22, SMW^{+}21]. eastward [DAB^{+}21]. eastward-flowing [DAB^{+}21]. Eavesdropping [THS21]. Ebullition [LSW^{+}20b, LHD^{+}20, PSK21, RWT^{+}21]. ebullitive [KSH^{+}23]. echosounder [PSY^{+}22]. Ecklonia [BRHS20, BS20]. eco [RNW^{+}23]. eco-surroundings [RNW^{+}23]. Ecological [BLS^{+}22, DTPV21, MKBSK19, ACE^{+}22, ADLW22, ASL^{+}24, FDF^{+}21, HLH^{+}23, KAC^{+}23, LZM^{+}21, MWV22, SBM^{+}21, SOO^{+}21, TJW^{+}22, VMT^{+}24, VMP^{+}21]. ecologically [LTJ^{+}20]. ecologists [MLB^{+}21]. ecology [CDH^{+}21, GM22, GMMW20, HDK^{+}21, HMB^{+}20, LLMM^{+}23, LGQ^{+}21, MSL21, MBC^{+}21, PMRG20, SHPW22, ZR23]. Ecophysiological [GK22]. ecophysiology [LES^{+}21]. Ecosystem [MA21, SWB^{+}20, VCBB20, BD22, BJS^{+}23b, BLS20, BRM^{+}23, CLG^{+}24, DLDF21, DMC^{+}22, DFJ22, GGS^{+}21, GMS21, HFL^{+}20, HBY^{+}22, HHK^{+}22, KVKS22, KDV^{+}23, LBBM21, LDG^{+}21, LGQ^{+}21, LSDA21, MHD^{+}22, MHL^{+}20, MGWS23, MSG^{+}22, MIR^{+}22, NTJ^{+}20, OBB^{+}20, PHB^{+}22, PFB^{+}22, PSNS22, PHM21, RVvB^{+}23, RGP^{+}22, SML21b, TKM^{+}22, TJW^{+}22, VJH^{+}22, WWB^{+}20, WRC^{+}21, WMTJW22]. ecosystem-metabolism [PHB^{+}22]. ecosystems [BSB^{+}20, BR22, BGM^{+}21, BFW^{+}22, CBPNA22, CML23, CLW^{+}21, CCDM21, FMT^{+}20, GF21, GSD^{+}23, HGMK22, HGH^{+}21, JLMW21, KHK^{+}22, SHS^{+}22, SPHH22, SOT^{+}20]. ecotone [AGC^{+}21]. ecotypes [TRS^{+}20]. ectotherm [MLP21b]. eddies [ACE^{+}22, EWF22, LXC^{+}21a, SSA^{+}23, ZBNH^{+}21]. Eddy [GPDA^{+}20, BBMD20, LBL^{+}24, VLP^{+}21, WCC20]. edge [BST^{+}22]. edges [CZH^{+}21]. edible [FJP^{+}21]. Editorial [Ham22, Kir20]. Edward [SBV^{+}21]. Eelgrass [MIR^{+}22, ARB^{+}22, GGS^{+}21, LDGB^{+}23, MDBI20, SIP21]. Effect [EJRRC^{+}23, HLKD23, SYT^{+}21, YSHS22, GMS21, LSL^{+}20, MXL^{+}21, MVBG20]. effective [KAP^{+}20]. Effects [CGR22, DWR^{+}23, EDC24, HANW21, HSBC20, HSP^{+}20, LLC^{+}23, LLC21, MCH^{+}21, MOW^{+}22, OSR20, PBGGRB21, SSL22, SSWM20, ZLIP20, AFK^{+}22, BYS^{+}23, BS20, BLS20, BPG^{+}22, CGC^{+}21, FMT^{+}20, FPB^{+}23, GGLH^{+}22, GLN^{+}21, GFG^{+}22, JHSG^{+}24, JHBB22, KSR20, LRdGFP23, LTLH22, MVP^{+}20, MLC^{+}21, MO23, MDWT^{+}22, MSB^{+}20, MGC^{+}20, MGG^{+}24, QRWT^{+}22, STB22, SSEO21, SCM22, VKB^{+}22, VWR^{+}23, WRS^{+}22, YG23, ZTT23, dJGCS23, BZK^{+}22]. efficiency [ALH^{+}21, BHR^{+}23, FDHS20, GZX^{+}22, HPCH24, SGSF20]. Efficient [KMV^{+}20, SWZ^{+}23]. effort [DFJ22]. egg [KSI24]. eggs [OLR^{+}23, ZLG^{+}23]. eight [vHHP^{+}22]. ejection [WSB^{+}20]. Electrona [DAB^{+}21]. electrothrix [MLK^{+}22]. Element [HTSG21]. Elemental [EHC^{+}22, LLMM^{+}23, HMBB21, MV22a, MV22b, STB22, ZLS^{+}21]. Elements [GGHBS^{+}20]. Elevated [DCK^{+}22, QBJG21, RVBP23, BMH^{+}23, BLS20, CGR22, DWR^{+}23, EDC24, GPB^{+}20, MDWT^{+}22, SIP21, VKB^{+}22, WRHR22, YG23]. elevation [CM20a, DGTW21, SLHL^{+}22, VMT^{+}24, ZQB^{+}21]. elongated [LHH24]. Elucidating [SCC^{+}21]. embayment [FWDY20, NSK^{+}21]. emerge [QMG21]. emergence [SDS^{+}23]. emergent [HZZ^{+}20, VJS^{+}20]. Emerging [SKB^{+}22, LTH21]. Emiliania [KTWGT20, MLS^{+}21, ZLS^{+}21]. emission [BSIdG23, GKW^{+}20, GKW^{+}21, ZSK^{+}21]. emissions [AKT^{+}21, CMCC24, HABL^{+}21, KSH^{+}23, LWC^{+}24, LPA^{+}23, MGI^{+}22, PLB^{+}21, SPBK21, YOXC23]. empirical [GMI^{+}21, WRMP22]. enclosed [XYL^{+}23]. encode [RCS^{+}20]. encroaching [SLJ^{+}22]. encrustations [SKA^{+}24]. end [CHS^{+}20]. endemic [HBY^{+}22]. Endogenous [CCC^{+}21]. endorheic [BGM^{+}22, YJB^{+}20]. endosymbionts [RCH^{+}20]. energetic [CGGC^{+}21]. Energy [BMA^{+}20, ALH^{+}21, CBT^{+}21, EST^{+}24, MHL^{+}20, SGSF20, SDB^{+}23, VKGW22]. Energy-based [BMA^{+}20]. engineer [NTJ^{+}20]. engineered [YJB^{+}20]. engineers [MIR^{+}22]. England [HHMH22]. enhance [BOC^{+}21]. Enhanced [HBR^{+}22, AMO20, KFHS24, LWC^{+}24, LBL^{+}24, MDIY23, ZPK^{+}22]. enhances [FWT^{+}20, QBJG21]. enriched [KWS^{+}21, ZLW^{+}21a]. enrichment [BAG^{+}24, BS20, BGM^{+}21, BRM^{+}23, CAV^{+}22, DKPS23, GGS^{+}21, HHMH22, LZM^{+}21, LLTT^{+}24, RLL^{+}21, YvdHC^{+}20]. entrained [SYMF21]. entropy [CGGC^{+}21]. environment [AWB^{+}23, CPB^{+}21, GGHBS^{+}20, KWS^{+}21, LVZ21, OYM^{+}20, RBD22, TRS^{+}20]. Environmental [CLP^{+}24, KOR^{+}20, KAC^{+}23, LTJ^{+}20, OGD20, PGEB22, SORGC^{+}21, CVF^{+}20, GGLH^{+}22, GCW^{+}23, HZC^{+}24, IYW^{+}22, JHBB22, MY23, MDDL22, PSNG^{+}21, PCC^{+}23, SSL22, SLHL^{+}22, WBB^{+}21, WRC^{+}21, WCH^{+}21, ZLS^{+}21, vLWV^{+}20]. environments [AMC21, MP23, RCS^{+}20, WLSD21]. enzymatic [BHSG^{+}20, BSGA21]. enzymes [BSGA21]. Eodiaptomus [LDB^{+}21]. epibiont [RBL^{+}20]. epibionts [JHS^{+}21]. epifaunal [CCFF21]. epilimnetic [HSB^{+}20]. episodic [CJB^{+}21, PFH21, RPKC22, YAE^{+}23]. equatorial [BCA^{+}20]. Erie [AJMG21, CWF^{+}22, LEK^{+}22, MDP^{+}23]. erodibility [HLPT20]. eroding [WPO^{+}22]. erosion [BBCA^{+}21, MHMML^{+}22, SYT^{+}21, WTvD^{+}22, YLT^{+}20]. Erratum [Ano21d]. essential [BMR^{+}21a, VCT^{+}23, VST^{+}22]. established [PSNS22]. establishment [SH22, ZZL^{+}21, vHHP^{+}22]. estimate [LvBS^{+}22]. estimated [GLN^{+}21, RST^{+}21]. Estimates [Wei24, DHNH22, KSI24, MGI^{+}22, WLSD21]. estimating [LAB^{+}22, PMJ^{+}21]. estimation [VGSB23]. estuaries [ACW^{+}22, DKPS23, EFC^{+}23, HTLP23, IOS24, MLP^{+}21a, MHdS^{+}21, RTM^{+}20, RS21, SDL22, YMSH20]. Estuarine [GFG^{+}22, AVS^{+}22, CEW^{+}21, HTLP23, LSC^{+}21, PKKS20, SMNT22, SSEO21, SLZ^{+}21a]. Estuary [LVHK20, BPIBN20, BPL22, BKMY21, FR23, HBS^{+}22, HSV22, KBSA21, KDZ^{+}24, MURK20, MK21, PGB24, RTM^{+}21, RVvB^{+}23, TBH^{+}22, WAPC20, WLK^{+}21, YOXC23, BM22, BMB22, YGD^{+}21]. ethylene [SBW^{+}20]. eukaryote [XKP^{+}22]. eukaryotes [BMLV24, MK20]. Eukaryotic [DLL20, PATL^{+}23, SRM23]. Euphotic [SWB^{+}20]. European [DMM^{+}21, ESW^{+}21, MMKWZ22, ZSK^{+}21]. eutrophic [AWGV20, BBH^{+}20, BHSL21, CGC^{+}21, DHNH22, HMB^{+}22, JLW^{+}23, KRC23, KHHZ20, MLC^{+}21, TBH^{+}22, THVJ23, TATC^{+}24, WHAR^{+}20, WLK^{+}21, vGDH^{+}21]. Eutrophication [LWS^{+}22, XDQ^{+}22, ZZT^{+}21, BAG^{+}24, DNB^{+}20, GÖA^{+}23, LHS^{+}21, MAP^{+}22, SVC^{+}22, SGA^{+}22, WHAR^{+}20, ZLU^{+}20]. eutrophied [QRWT^{+}22]. Evaluating [CFC^{+}20, GWS^{+}21, SHBA22]. evaluation [ASH^{+}20, IKN^{+}21]. evaporation [FBK^{+}22, VLP^{+}21]. evaporation-to-inflow [FBK^{+}22]. evasion [SDB^{+}23]. Evasive [JSW^{+}21]. event [AFH^{+}22, CBC^{+}22]. events [GRJ^{+}22, KdED^{+}20, MAG^{+}22, TUR^{+}20, WSB^{+}20, WTvD^{+}22, dCLO^{+}23]. Evidence [DKPS23, HGF^{+}21, KPG^{+}23, MSGW^{+}20, ZLL^{+}21, BVM^{+}20, CSH21, CBS^{+}21, DP22, PPB^{+}21]. evidenced [SSK24]. evolution [BST^{+}22, LBJS22, OBL21, ZJ22a]. evolutionary [HLH^{+}23, MTB^{+}21, VGvdB^{+}21, ZJ22b]. Evolving [GM22, ZZCD23]. exacerbates [GÖA^{+}23, RLL^{+}21]. examination [WZCK20]. examined [FHMK20]. Examining [ACE^{+}22, AFM23, SPKM22, SMKS22]. exceeds [SDB^{+}23]. exceptionally [JLW^{+}23]. excess [TMM^{+}20]. exchange [CSC^{+}21, EFG^{+}24, MBA^{+}21, NvBE^{+}21, SMNT22, TEK^{+}21]. exchange-driven [TEK^{+}21]. exchangeable [WXK^{+}22]. exchanges [TBH^{+}22]. exhibit [TBH^{+}22]. exhibits [AFH^{+}21]. existence [MCYR20]. exotic [WHZ^{+}21]. expansion [CZH^{+}21, GAS^{+}21, KSH^{+}23, NB23, QMG21]. expansions [CJB^{+}21]. experiment [BPG^{+}22, GMNG^{+}24, KRM^{+}20, MKBSK19, OJUN23, SGL^{+}23, SSM^{+}22a]. Experimental [BRM^{+}23, DP22, CBPNA22, HKM^{+}20, KTA^{+}21, KFHS24, SVC^{+}22, ZJ22a]. experiments [BDE^{+}22, BBA^{+}23]. explain [BVGQN22, DMM^{+}21, MTP^{+}23]. explained [vLWV^{+}20]. explaining [LPMBR23]. Exploring [HM20]. export [BCWS23, BRCO21, DBM^{+}23, GLM^{+}23, KBvdD^{+}20, MMG^{+}21, RMW^{+}21, SDL22, SPKM22, TEK^{+}21, WSZ22a, WSZ22b, YAP^{+}22, YXC^{+}22, ZBNH^{+}21, ZYZ^{+}21]. exports [SOT^{+}20]. exposed [BOC^{+}21, DGTW21]. Exposure [ABC^{+}21, KDV^{+}23, BMH^{+}23, CTM^{+}21, DWR^{+}23, EDC24, GEHD20, GPB^{+}20, GK22, MDWT^{+}22, MIR^{+}22, SSL22, ZFQ^{+}20, dSNIB22]. exposures [EFP^{+}21]. Expression [LOW^{+}20, BMH^{+}23, LTLH22]. extension [GR24]. extent [BKMY21, DAD^{+}21, SMW^{+}21, ZZCD23]. external [XMP^{+}21]. extinction [GRJ20]. extracted [BAD^{+}21]. extrapallial [CGR22]. Extratropical [BJS^{+}23b]. Extreme [BTK^{+}21, SRN^{+}21, TKM^{+}22, TUR^{+}20]. extremes [KDV^{+}23, MK21]. extremities [TPTS23]. exudates [BMG^{+}23]. exudation [RRAO^{+}21].

facilitate [GPDA^{+}20]. facilitated [MBLA^{+}21]. facilitates [CJB^{+}21, MCYR20]. facilitating [MBB^{+}20]. facility [CLGH23]. factor [ZLJ^{+}24]. Factors [JREVB21, SOSN23, SBB^{+}21, BHG^{+}24, CESS21, JSPS24, KBL^{+}20, KPNJ23, LGF22, PCC^{+}23]. failed [SRV^{+}20]. far [BTK^{+}21]. fast [KBB^{+}22]. fat [MN20]. fate [CDW^{+}24, PGB24, QBJG21, SPKM22, SGK^{+}21, ZCZ^{+}23]. fatty [CLG^{+}24, EST^{+}24, KLG^{+}23, VST^{+}22]. fauna [BDMGH^{+}23]. favorable [SSA^{+}23]. Fe [RvHF^{+}20]. Fear [PVSP^{+}22]. features [ZZL^{+}21]. featuring [THL^{+}21]. February [CMT^{+}20]. Fecal [SSM21b, DMS^{+}21, LG21, MLS^{+}21]. feedback [MBLA^{+}21]. feedbacks [MHD^{+}22]. feeding [CP22, CSNS24, CCG^{+}22, CUS21, FRRG^{+}21, FLW^{+}21, KGG^{+}23, RTK23, SYMF21, WSB^{+}20, WSZ22a, WSZ22b]. feeding-current [SYMF21]. female [KSI24]. fermentation [NKH^{+}23]. ferruginous [TZF^{+}22]. fertilization [BPG^{+}22, HSBC20]. few [OMM21]. field [APB^{+}23, BDE^{+}22, GLN^{+}20, SVBT20, WO21]. fields [HGH^{+}21]. filamentous [FWT^{+}20]. filter [AMO20, THVJ23]. final [BFCC^{+}21]. Fine [DFW^{+}20, ASL^{+}24, GBC^{+}20, GSD^{+}23, GLB^{+}22, HCS^{+}21, WO21]. fine-scale [ASL^{+}24, GBC^{+}20, GSD^{+}23, GLB^{+}22, WO21]. fingerprints [SRM23]. finmarchicus [CBS^{+}21, PPB^{+}21]. Fiona [BJS^{+}23b]. first [SRL^{+}20]. fish [APB^{+}23, BMA^{+}20, BMS^{+}22, CESS21, CLT^{+}22, CPBJR^{+}21, DWR^{+}23, EST^{+}24, GCMJ23, GPDA^{+}20, HXL^{+}21, KCT23, LDB^{+}20, OLR^{+}23, SBD^{+}21, SSA^{+}23]. fish-based [SBD^{+}21]. fisher [RMH^{+}22]. fisheries [FGP22]. fishes [BZSF23, LDG^{+}21, NDP^{+}24, WRMP22, WSZ22a, WSZ22b]. fishing [DFJ22]. fitness [TVB^{+}21]. five [BMB22]. fixation [BMCS23, HBS^{+}22, LLL^{+}23, LTH21, LTLH22, SCC^{+}21, SMW^{+}21, SBL^{+}21, hTRH20, YSHS23, ZLL^{+}21]. fixed [FRRG^{+}21]. fixers [AAC^{+}21]. fixing [FBP20, HMB^{+}22, YG23, ZFQ^{+}20]. fjord [BHSG^{+}20, BJK^{+}21, BOC^{+}21, HJWP24, KHB22, vdPMF^{+}21]. fjords [HZM^{+}21, JLMW21]. flagellate [AH20]. flagellates [ELS23, LOW^{+}20, OMM21, vdPAV^{+}20]. flat [BD22, SYT^{+}21]. flats [ZvBZ^{+}20]. Flexible [VKGW22, ZLN21]. floating [RHM^{+}21, SYMF21, VJS^{+}20]. floating-leaved [VJS^{+}20]. floe [SBM^{+}21]. flood [RVS^{+}22]. flooding [TNK^{+}22, TUR^{+}20]. floodplain [LTH21, MGWS23, dMKB^{+}20]. floodplains [MKP^{+}23]. Florenciella [LES^{+}21]. Florida [CLGH23, HSD^{+}20, OLFH21]. Flow [ZLIP20, ZN20, BMR^{+}21b, FHR^{+}21, FHMK20, GLN^{+}21, GBHF21, LDG^{+}21, MDBI20, RBAC^{+}20, SVBT20, VKGW22, WSB^{+}20]. Flow-induced [ZN20]. flowing [DAB^{+}21]. flowpaths [GEHD20]. flows [NH22]. Fluctuating [WBB^{+}21, GK22, SPBK21, ZXZ^{+}23]. fluctuation [BWSW24]. fluctuations [JHS^{+}21, WRMP22]. Fluid [CUS21, CGR22, STAK22]. flume [WTvD^{+}22]. fluorescence [GF21, LLBC20, RKBT^{+}23]. fluorescence-based [LLBC20]. Fluorescent [BMG^{+}23]. flushing [XWL^{+}21]. fluvial [PSE^{+}20, PGB24, RBAC^{+}20, YLT^{+}20]. flux [CCB^{+}20, CHS^{+}20, CBN21, DHNH22, DGTW21, FUUG^{+}21, HCB^{+}22, HZM^{+}21, IKN^{+}21, MGWS23, NB20, RSP^{+}24, SBD^{+}21, SMNT22, VLP^{+}21, WF22, WDS^{+}23, WSZ22a, WSZ22b, YMSH20]. fluxes [ACW^{+}22, GBB^{+}21, LXC^{+}21a, MCYR20, NSK^{+}21, RST^{+}21, SLRS^{+}24, SWK^{+}23]. focusing [SGA^{+}22]. following [CLGH23, CGD22, CBC^{+}22, HFL^{+}20, KHHZ20, RWS^{+}22, SSS^{+}20, TKM^{+}22]. Food [BHR^{+}23, BOC^{+}21, Gil22, MV22a, MV22b, PSNS22, RBL^{+}20, TSX^{+}23, WRC^{+}21, APB^{+}23, BR22, CDH^{+}21, CESS21, CGL^{+}21, EG20, FSM^{+}21, FWT^{+}20, FLW^{+}21, GGS^{+}21, HCH^{+}24, IFU^{+}20, JSPS24, KPL^{+}20, KLG^{+}23, LTQS^{+}21, LCM^{+}22, LCHB22, LEB21, MMF^{+}21, MLC^{+}21, MSB^{+}20, SWC^{+}23, SSK24, TCS21, VCT^{+}23, WHZ^{+}21, YAP^{+}22]. foraging [CFC^{+}20, CSB22, NCC^{+}21, STAK22]. foraminiferal [XLN21]. forces [CLP^{+}24]. forcing [CMB^{+}21, MK21, SRN^{+}21, ZZHZ23]. forcings [YJB^{+}20]. fore [EDC24]. forecast [SJH20]. forecasting [MO23, THS21]. forest [BRM^{+}23, PBP^{+}23, TCP^{+}22, WLSD21]. forested [AFH^{+}22]. forests [CESS21, FW21, SOO^{+}21, SMKS22]. form [VWR^{+}23]. formation [BC20, CZH^{+}21, GAS^{+}21, GBC^{+}20, HVMS^{+}21, LBL^{+}24, PLLM21, YGD^{+}21, ZSZ^{+}21]. forming [BMH^{+}23, KOR^{+}20]. forms [KRC23, SML21b, ZvBZ^{+}20]. Forward [ZJ22a]. foundations [BV21]. four [BMR^{+}21a, SVMM^{+}22]. fraction [BCL^{+}22]. fractionated [KSF^{+}23, KSR20, LXC^{+}21b]. fractionation [HRK^{+}22, WOK^{+}21, WMRL22]. fractionations [ORRMD^{+}21]. fractions [LHMR23, LMS^{+}21]. Fragilariopsis [MLG^{+}20, MGC^{+}20]. framework [BBCA^{+}21, BFW^{+}22, MLB^{+}21, MO23, SHPW22]. Francisco [BM22, BMB22]. free [AHF^{+}22, KHSP21, SYMF21, WPY^{+}20]. free-floating [SYMF21]. free-living [WPY^{+}20]. frequency [CZK^{+}23, DAD^{+}21, FR23, LGI^{+}20, MK21, MKD22, SWK^{+}23, SSM^{+}22a, VVMJ23, WDG^{+}21]. frequent [BRB^{+}22]. Fresh [IÁSNCR21]. Freshening [WM20]. Freshwater [MLP^{+}21a, AMC21, BHSG^{+}20, BWV^{+}20, BR22, BGM^{+}21, BOC^{+}21, BLS20, CML23, CKCB^{+}23, DWR^{+}23, HSBC20, HP20, KWS^{+}21, LTQS^{+}21, LABM20, MSGW^{+}20, MWHR^{+}22, PGEB22, RVBP23, SBH23, VJS^{+}20, WNHY21, XTJW21, vdPMF^{+}21]. freshwaters [BYS^{+}23, CPB^{+}21]. fringed [KDZ^{+}24]. Front [TKA21, SCC^{+}21, YLT^{+}20]. Frontal [SSA^{+}23]. frustules [XSZ^{+}21]. fucans [HVMS^{+}21]. fuels [CBF^{+}23]. Fukushima [TNK^{+}22]. Fukushima-derived [TNK^{+}22]. Full [SO23, OLR^{+}23, RVS^{+}22]. Full-year [SO23]. function [BJG^{+}22, HTSG21, LBBM21, LHH24, MMF^{+}21, MKM^{+}23, OYM^{+}20, THVJ23, VWR^{+}23]. Functional [KL22, MLB^{+}21, MACC24, SSK24, BWB^{+}22, CSB22, DTPV21, FAB^{+}20, HLH^{+}23, HPCH24, MKP^{+}23, MVBG20, OAM^{+}22, PGEB22, RvHF^{+}20, RLL^{+}21, SPR22, WOT^{+}23, WCZ^{+}20]. functionally [MCH^{+}22]. functioning [GMS21, PSNS22]. functions [XYL^{+}23]. Fundamental [KAP^{+}20]. fungal [FWT^{+}20, XYL^{+}23]. fungi [KSN^{+}21]. fungicides [BZK^{+}22]. Future [RGW^{+}23, BFW^{+}22, EWGP22, FM21, GM22, LLSQ24, NB23, RBD22, VMT^{+}20, WHAR^{+}20]. Fuxian [XTL^{+}20].

G [MK20]. gamma [GCEM^{+}23]. gas [BSIdG23, GPK^{+}21, HABL^{+}21, KSH^{+}23, MBA^{+}21, MGI^{+}22, PLB^{+}21, RSP^{+}24, SMNT22, SLRS^{+}24]. gegenbauri [FLW^{+}21]. gelatinous [CDH^{+}21, HCH^{+}24]. gene [BMH^{+}23, DLL20, MBD^{+}22, MBD^{+}23, WCZ^{+}20, WNHY21, ZR23]. generalists [MZW^{+}21]. genes [BCA^{+}20, FAB^{+}20, LOW^{+}20, MVBG20]. Geneva [Lem20, SLZ^{+}21b]. genomes [RCS^{+}20]. Genomic [MUGL20]. gentoo [NCC^{+}21]. Geochemical [SGA^{+}22, TJW^{+}22]. Geographic [MRH^{+}23, LWH^{+}23, LCS^{+}20]. geologic [CEE^{+}22]. geomorphic [LGF22]. geomorphology [BMR^{+}21b]. George [THT^{+}20]. germination [ZZL^{+}21]. germination-features [ZZL^{+}21]. giant [SBB^{+}21, XKP^{+}22]. gigas [PMRG20]. Giling [Ano21a]. Glacial [MCS^{+}23, NB23]. glacialis [CBS^{+}21, PPB^{+}21]. Glacier [FBJ^{+}23, BJK^{+}21, RBR^{+}24]. glacier-marine [RBR^{+}24]. glacierized [BSF^{+}21]. Glaciers [DBFC24, VMT^{+}24]. Global [GGHBS^{+}20, LG21, ABC^{+}21, BDE^{+}22, CLC^{+}20, CLC^{+}21, CBC^{+}22, DCJB22, JHS^{+}21, KDV^{+}23, OYM^{+}20, OJUN23, PJRV23, RTM^{+}20]. globally [GSD^{+}23, HGF^{+}21]. Glomalin [PYY^{+}20, ZJW^{+}22]. Glomalin-related [PYY^{+}20]. glycerol [ZLL^{+}24]. glycine [MAD^{+}22]. Going [FNB^{+}23]. gold [LCHB22]. Gone [MACC24]. govern [BD22, GLAHH^{+}23]. gradient [ADLW22, AKT^{+}21, BWV^{+}20, BRHS20, CÖL^{+}20, CBHB21, HHMH22, IMB^{+}22, IYW^{+}22, KAP^{+}20, LSC^{+}21, MLP^{+}21a, PMF^{+}21, RCMA21, SWOR20, SO23, SGSD^{+}22, TKW^{+}23, VST^{+}22]. gradients [BSGA21, CLG^{+}24, GLAHH^{+}23, LAFV23, MTB^{+}21, SORGC^{+}21, TKA21, TKS^{+}20]. gravel [ARBvS^{+}21]. gravel-bed [ARBvS^{+}21]. grazer [GSD^{+}23]. grazers [BWB^{+}22, SDSGR21, ZJ22b]. Grazing [HBY^{+}22, TATC^{+}24, DvOH^{+}20, GBMKJ22, JHBB22, LES^{+}21, MMG^{+}21, SDSGR21, ST22, XSZ^{+}21, ZSZ^{+}21]. Great [GSP^{+}23, MFM^{+}21, SWOR20, ETG21, GBO^{+}20, MBB^{+}20, VCBB20, YJB^{+}20]. green [JHBB22, LHH24]. greenhouse [BSIdG23, HABL^{+}21, KSH^{+}23, MGI^{+}22, PLB^{+}21, SLRS^{+}24]. Greening [FSM^{+}21]. Greenland [BHSG^{+}20, MN20, XTJW21]. gregaria [CUS21]. gross [SSM^{+}22a]. Groundwater [KPM^{+}20, BKMY21, FKN21, HGH^{+}21, IÁSNCR21, KRM^{+}20, NSK^{+}21, RTM^{+}21, RST^{+}21, WLK^{+}21]. Groundwater-controlled [KPM^{+}20]. groundwater-dependent [HGH^{+}21]. group [BBL^{+}21, EHC^{+}22, SSB^{+}22]. groups [BDL^{+}21, LTQS^{+}21, LLBC20, SRM23]. Growth [JHS^{+}21, ZLS^{+}21, BS20, BK20, CFC^{+}20, DWR^{+}23, FPG^{+}24, FBJ^{+}23, FPB^{+}23, HSP^{+}20, JHBB22, KPG^{+}23, KFHS24, LSL^{+}20, LDB^{+}21, LSDA21, MRC^{+}24, MVP^{+}20, MDWT^{+}22, MMLPRHCG22, PKV20, QBJG21, RGW^{+}23, SMD21, YSHS22, ZWM22]. Growth-dependent [ZLS^{+}21]. guidelines [LCMG^{+}20]. guilds [RBAC^{+}20, WSZ22b, WSZ22a]. Gulf [DLDF21, DP22, KPS21, MLP^{+}21a, PMRG20, SMP21, WSZ22a, WSZ22b, YMSH20]. guttatus [BFCC^{+}21]. Gyre [HYPG^{+}22, RPKC22, ZBNH^{+}21].

H [LTLH22, hTRH20]. Habitat [BBCA^{+}21, WHZ^{+}21, CYZZ^{+}20, CBD^{+}23, DFJ22, GRJ20, KHCG^{+}22, MO23, PSY^{+}22, SSA^{+}23, VKGW22, YAH^{+}21]. Habitat-dependent [WHZ^{+}21]. Habitat-specific [BBCA^{+}21]. habitats [CLW^{+}21, CEW^{+}21, GPDA^{+}20, HLPT20, MHL^{+}20, RD23, VCT^{+}23, WHAR^{+}20]. hadal [FUUG^{+}21]. hadopelagic [LZD^{+}23]. Halisarca [LMP^{+}20]. halocline [CLD^{+}23]. handling [RTK23]. haplotypes [LGI^{+}20]. haptophytes [THT^{+}20]. hard [MKBSK19]. hard-substrate [MKBSK19]. Harmful [BK20, DTPV21, HM20, HMB^{+}22, KAC^{+}23, KRC23, LZG21, PHM21, SKB^{+}22, THS21]. Harvey [KPS21]. Hatteras [SCC^{+}21]. HAUSGARTEN [MKBSK19]. Hawaiian [SWS^{+}22, BST^{+}22, HBY^{+}22, PMJ^{+}21]. Headward [WPO^{+}22]. Headward-eroding [WPO^{+}22]. headwater [BCL^{+}22, RWT^{+}21, RBSB22]. Headwaters [GBHF21]. healthy [YvdHC^{+}20]. heat [EJRRC^{+}23, LRdGFP23, SSL22, SRV^{+}20, VGvdB^{+}21]. heated [DKW^{+}21, DMCW23]. heatwave [DMM^{+}21, JHSG^{+}24, ZZHZ23]. heatwaves [GBW24, ZSF^{+}23]. heel [ZvBZ^{+}20]. hemisphere [CZK^{+}23]. herbicides [YML^{+}21]. herbivory [FWT^{+}20, JREVB21, LCZ^{+}21]. heterogeneity [CLW^{+}21, GCW^{+}23, HMB^{+}20, RWT^{+}21, SCC^{+}21]. Heterogeneous [BSGA21, MOW^{+}22]. Heterotrophic [LTH21, RHS^{+}22, AH20, BMG^{+}23, DWW^{+}21, ELS23, GAS^{+}23, LOW^{+}20, LABM20, OMM21, PGN^{+}21, PMJ^{+}21, SWC^{+}23, STAK22, ZLW^{+}21a]. Hg [CBB^{+}22]. High [CML23, KRB^{+}21, KSH^{+}23, LZD^{+}23, MDIY23, MKD22, PSE^{+}20, PHB^{+}22, SWK^{+}23, VVMJ23, WDG^{+}21, ZZO^{+}21, AAC^{+}21, BCWS23, BOC^{+}21, BBGM20, CGGC^{+}21, CM20a, CZK^{+}23, CGB^{+}21, DGTW21, DCF^{+}22, DAD^{+}21, EFP^{+}21, JLW^{+}23, JSB^{+}20, JLMW21, KCT23, KMV^{+}20, KTA^{+}21, LDG^{+}21, MCYR20, MHVC22, RBR^{+}24, RRAO^{+}21, SIP21, SSM^{+}22a, SEG21, SLHL^{+}22, VMT^{+}24, WZCK20, ZQB^{+}21, vdPMF^{+}21, PLB^{+}21, RCS^{+}20]. high-altitude [DCF^{+}22]. high-Arctic [BOC^{+}21]. high-elevation [CM20a, DGTW21, SLHL^{+}22, VMT^{+}24, ZQB^{+}21]. High-frequency [MKD22, SWK^{+}23, VVMJ23, WDG^{+}21, CZK^{+}23, DAD^{+}21, SSM^{+}22a]. high-latitude [JSB^{+}20, KCT23, LDG^{+}21, SLHL^{+}22]. high-mountain [CGGC^{+}21, MHVC22]. High-resolution [PHB^{+}22, ZZO^{+}21, BCWS23]. high-throughput [BBGM20, WZCK20]. Higher [GCEM^{+}23, MRT^{+}24, TION23, CLC^{+}22]. Highest [KTA^{+}21]. Highly [ZLW^{+}21a, EFC^{+}23, LDG^{+}21, RBD22]. historic [MDDL22]. Historical [GGLH^{+}22, BAG^{+}24]. history [HXL^{+}21, SSL22]. holobionts [MSD23]. Holocene [MPW^{+}23]. holomictic [XTL^{+}20]. Homeostasis [DBB^{+}21].

Homeostatic [HP20]. host [FPBG^{+}22, KSN^{+}21, LEB21, TVB^{+}21]. host-algae [FPBG^{+}22]. hosted [BVL^{+}23]. hosts [CML23]. Hot [ONH^{+}20]. hotspot [KPNJ23, SCC^{+}21]. hotspots [CFC^{+}20, EFC^{+}23]. Hudson [SMNT22]. Human [ZZWL22, JLW^{+}23, LVZ21, NDP^{+}24, ZSD^{+}20, vHLV^{+}20]. human-altered [LVZ21]. human-dominated [NDP^{+}24]. Humboldt [PMRG20]. humic [EFC^{+}23]. Hurricane [KPS21]. huxleyi [KTWGT20, MLS^{+}21, ZLS^{+}21]. hybridization [CSH21, CBS^{+}21, PPB^{+}21]. Hydraulic [ARBvS^{+}21]. hydroclimate [BJG^{+}22]. Hydrodynamic [GLB^{+}22, GCMJ23, RPA^{+}22, SMF20, SYMF21]. hydrodynamics [SCM22, TCP^{+}22]. hydrogen [SGK^{+}21]. hydrogenase [LTLH22]. Hydrologic [JSB^{+}20, PCJ^{+}21, YMSH20, BMR^{+}21b, DPB^{+}23, YJB^{+}20]. hydrologically [BCL^{+}22, DBFC24, FBRG23]. Hydrology [DBM^{+}23, ESW^{+}21, BCTH20, RSLW23, RVS^{+}22, ZYZ^{+}21]. hydromedusa [CUS21]. hydroperiod [DGTW21]. hydrostatic [SEG21]. hydrothermal [ASL^{+}24, LBG^{+}22, dCLO^{+}23]. hyper [AWGV20]. hyper-eutrophic [AWGV20]. hyperborea [SOO^{+}21]. hypereutrophic [KTA^{+}21]. hypersaline [MVBG20, SSM^{+}21a]. Hypolimnetic [SSWM20, SDS^{+}23]. hypolimnia [LM24]. hypolimnion [Lem20, SPLJA^{+}20]. hyporheic [BZK^{+}22]. hypothesis [LD22, SB22, WRA^{+}23, ZLL^{+}21, CFC^{+}20]. Hypoxia [CBHB21, AAB^{+}22, EFG^{+}24, LHS^{+}21, MKL^{+}24, PMD^{+}23, SLZ^{+}21a, YGD^{+}21, ZLU^{+}20, vHLV^{+}20]. hypoxia-driven [MKL^{+}24]. hypoxic [LM24, MLK^{+}22, QRWT^{+}22, RWS^{+}22, SCB^{+}20]. hysteresis [MAP^{+}22].

ice [AHF^{+}22, CCA^{+}21, CVF^{+}20, HS21, JHvL^{+}23, KPG^{+}23, KFHS24, LAFV23, LDS^{+}20, LXC^{+}21b, MLG^{+}20, OBL20, PMD^{+}23, PW22, PGN^{+}21, PFB^{+}22, RFW^{+}24, SSM^{+}21a, SBM^{+}21, SMCS23, SGL^{+}23, SJJ^{+}21, TMS^{+}21, VCT^{+}23, VMP^{+}21, WOMWR20, WRHR22, YWLY20, ZXE^{+}23]. ice-associated [KFHS24, WRHR22]. ice-covered [CVF^{+}20, LDS^{+}20, LXC^{+}21b, PMD^{+}23, SSM^{+}21a, SMCS23]. ice-floe [SBM^{+}21]. ice-free [AHF^{+}22]. ice-retreat [SJJ^{+}21]. Identification [AMC21]. identifies [OGD20, SLHL^{+}22]. Identifying [vHHP^{+}22]. III [KLP23]. illuminate [BBGM20]. illuminates [BBM21]. illuminations [GTKW^{+}20]. image [OAM^{+}22]. imagery [SGM^{+}22]. images [VMP^{+}21]. imaging [BO20, BBGM20, FHMK20, GSD^{+}23]. imbalances [KTA^{+}21, MSJ21]. immediate [SPKM22]. Impact [LBG^{+}22, SMG^{+}22, SBL^{+}21, YG23, ZSD^{+}20, JLW^{+}23, KWS^{+}21, LG21, MIR^{+}22, RLL^{+}21, RWT^{+}21, YOXC23, ZN20]. impacted [LSCL21]. impactful [CCG^{+}22]. Impacts [BJS^{+}23b, OBB^{+}20, BSB^{+}20, BKC20, CLW^{+}21, HM20, LZM^{+}21, PCN22, PVSP^{+}22, SLRS^{+}24, WHAR^{+}20, WWB^{+}20, WBB^{+}21, WSB^{+}20, WHZ^{+}21, WCC20, YML^{+}21, ZZWL22, ONH^{+}20, SRN^{+}21]. impair [BDMGH^{+}23]. impairs [MGP^{+}21]. Implication [JYA21]. Implications [BH21, FBJ^{+}23, FUUG^{+}21, GBB^{+}21, GAS^{+}23, MSJ21, PYY^{+}20, THT^{+}20, XMP^{+}21, YWLY20, YAP^{+}22, BPL22, EAF^{+}22, FNB^{+}23, Gil22, GM22, KDW^{+}23, QBJG21, ZXE^{+}23]. Importance [BBH^{+}20, MGI^{+}22, BCTH20, BBM21, FBR^{+}23, GCW^{+}23, HS21, LLBC20, MHG^{+}23, TION23, TKH21, TRS^{+}20, ZZT^{+}21]. important [GKW^{+}20, GKW^{+}21, MCS^{+}23, MKM^{+}23, RLL^{+}21]. improve [THS21]. improvement [HSB^{+}20]. In-lake [SEK20]. including [ZN20]. incorporation [FSH^{+}23, GE22]. increase [DKW^{+}21, KLO23, ZZCD23]. Increased [BJK^{+}21, CLC^{+}22, BAG^{+}24, KFHS24, MRT^{+}24, YSHS22]. increases [LMS23, VKGW22, ZLL^{+}21, ZZT^{+}21]. Increasing [ALH^{+}21, CT23, FBV23, OJUN23, SBL^{+}21]. incubation [BBA^{+}23]. incubations [TION23]. independent [JSW^{+}21]. Indian [BGAD^{+}22, BLRM24, GRJ^{+}22, LGI^{+}20]. indicate [CTF^{+}21, HFL^{+}20, LDG^{+}21]. indicates [OJUN23]. indicator [MDP^{+}23]. indicators [WHAR^{+}20]. indices [LEK^{+}22]. indirect [MEQBV20, MSB^{+}20]. Indo [HOB^{+}22, YAH^{+}21]. Indo-Pacific [HOB^{+}22]. induce [AH20, MRT^{+}24]. induced [DAD^{+}21, GMNG^{+}24, IYW^{+}22, LLC^{+}23, LXC^{+}21a, SYMF21, SYT^{+}21, SCM22, SBW21, TNK^{+}22, TSX^{+}23, WTvD^{+}22, ZN20, ZSZ^{+}21]. induces [PW22]. infaunal [LNR^{+}20]. Infection [FWT^{+}20]. infer [RBD22]. inferred [BFCC^{+}21, CZK^{+}23, MNMJ^{+}21]. inflow [BM22, CHS^{+}20, FBK^{+}22, LWC^{+}24, MLP^{+}21a]. inflows [PBGGRB21]. Influence [ABS^{+}21, LL23, TCP^{+}22, dCLO^{+}23, AVJ^{+}24, BJS23a, CZS^{+}22, DLL20, HMM^{+}22, HZC^{+}24, HNS21, JSPS24, KHK^{+}22, KLP23, KPNJ23, MGGRR^{+}21, RMH^{+}22, SML21b, SMKS22, VMT^{+}24, YSHS23, ZYZ^{+}21]. influenced [CESS21, FXD^{+}22, HZM^{+}21, LSC^{+}21, REF^{+}21]. influences [BBCA^{+}21, CGD22, DBFC24, FBJ^{+}23, OYM^{+}20]. influencing [EBK^{+}20, SBB^{+}21]. influx [BHSL21]. inform [ASH^{+}20]. Information [Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano20y, Ano20-38, Ano20z, Ano20-39, Ano20-27, Ano20-40, Ano20-28, Ano20-41, Ano20-29, Ano20-42, Ano20-30, Ano20-43, Ano20-31, Ano20-44, Ano20-32, Ano20-45, Ano20-33, Ano20-34, Ano20-46, Ano20-35, Ano20-47, Ano20-36, Ano20-48, Ano20-37, Ano20-49, Ano20-50, Ano20-51, Ano20-52, Ano20-53, Ano20-54, Ano20-55, Ano20-56, Ano20-57, Ano20-58, Ano20-59, Ano20-60, Ano20-61, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21-53, Ano21p, Ano21q, Ano21-29, Ano21r, Ano21-30, Ano21s, Ano21-31, Ano21t, Ano21-32, Ano21u, Ano21-33, Ano21v]. Information [Ano21-34, Ano21w, Ano21-35, Ano21x, Ano21-36, Ano21y, Ano21-37, Ano21z, Ano21-38, Ano21-27, Ano21-39, Ano21-54, Ano21-55, Ano21-28, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano21-49, Ano21-50, Ano21-51, Ano21-56, Ano21-52, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22a, Ano22q, Ano22r, Ano22-31, Ano22-30, Ano22-43, Ano22s, Ano22-32, Ano22t, Ano22-33, Ano22u, Ano22-34, Ano22v, Ano22-35, Ano22w, Ano22-36, Ano22x, Ano22-37, Ano22y, Ano22-38, Ano22z, Ano22-39, Ano22-27, Ano22-40, Ano22-28, Ano22-41, Ano22-29, Ano22-42, Ano22b, Ano22c, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano22-48, Ano22-49, Ano22-50, Ano22-51, Ano22-52, Ano22-53, Ano22-54, Ano22-55, Ano22d, Ano22-56]. Information [Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23y, Ano23n, Ano23z, Ano23o, Ano23-27, Ano23p, Ano23-28, Ano23q, Ano23-29, Ano23r, Ano23-30, Ano23s, Ano23-31, Ano23t, Ano23-32, Ano23u, Ano23-33, Ano23v, Ano23-34, Ano23w, Ano23-35, Ano23x, Ano23-36, Ano23-37, Ano23-38, Ano23-39, Ano23-40, Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano23-45, Ano23-46, Ano23-47, Ano23-48, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24i, Ano24f, Ano24j, Ano24g, Ano24k, Ano24h, Ano24l, Ano24m, Ano24n, Ano24o, Ano24p]. informs [KSF^{+}23]. Ingestion [BWSW24, ELS23, BZSF23, TION23]. inhabiting [AHCJ^{+}20]. inhibited [GBMKJ22]. initiates [BFE^{+}20]. initiation [PHM21]. inland [GMI^{+}21]. inlet [IÁSNCR21]. Inorganic [BDL^{+}21, ACW^{+}22, FAB^{+}20, HSD^{+}20, KHHZ20, MHP^{+}23, MKM^{+}23, QRWT^{+}22, RSLW23, RVvB^{+}23, RCH^{+}20, SKA^{+}24, TEK^{+}21, VWR^{+}23, ZLU^{+}20]. input [YML^{+}21]. inputs [BOC^{+}21, CZK^{+}23, LPMBR23, TSX^{+}23]. insect [BWB^{+}22]. Insight [LLC^{+}23]. Insights [ACW^{+}22, BCWS23, BDE^{+}22, HCH^{+}24, LLMM^{+}23, LVDM22, Lem20, LN21, PYM^{+}20, SEG21, WAPC20, APB^{+}23, HGA24, KLG^{+}23, WOK^{+}21, ZZB^{+}20]. Instr [Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano20y]. instrumented [KTH^{+}21]. Integrating [CSM^{+}23, HHK^{+}22, PSY^{+}22, KSF^{+}23]. integrative [ST22]. intelligence [ARB^{+}22]. intense [AFH^{+}22, CBF^{+}23, CMT^{+}20, DBB^{+}21, LRM21]. intensifies [SVC^{+}22]. intensify [ZXZ^{+}23]. intensity [MTB^{+}21, RGW^{+}23]. inter [KMV^{+}20, vdPMF^{+}21]. inter-annual [vdPMF^{+}21]. inter-replacement [KMV^{+}20]. Interaction [ZLU^{+}20, EWF22, GAS^{+}21, KBL^{+}20, LMSN23]. Interactions [BBA^{+}23, HMB^{+}20, ZFQ^{+}20, CHA^{+}24, GGS^{+}21, MIR^{+}22, PGB24, SDSGR21, TKH21]. Interactive [FPB^{+}23, MGC^{+}20, STB22, AFK^{+}22, WWB^{+}20]. interactome [CLC^{+}20, CLC^{+}21]. interannual [CGRGP^{+}21, CP22, DLDF21, JLW^{+}23, PFB^{+}22, RFW^{+}24]. interface [BKMY21, RWMP^{+}23]. interferes [HXL^{+}21]. Interleaving [HPNU20]. intermediate [RS20]. intermittency [PSE^{+}20]. intermittent [SPR22]. Internal [BMS^{+}22, FWDY20, GLN^{+}20, LHH24, FGO^{+}20, GLN^{+}21, HMB^{+}22, MXL^{+}21, NH22, PRS^{+}20, XMP^{+}21]. Interplay [SGSD^{+}22, CYPRG^{+}23, HZM^{+}21]. Interrelated [YSHS23]. Interspecific [VGvdB^{+}21, BHG^{+}24, MKL^{+}24]. intertidal [DvOH^{+}20, EJM^{+}23, FNB^{+}23, ORE20, PYM^{+}20, RS21, SYT^{+}21, TEK^{+}21]. intertidal-flat [SYT^{+}21]. intraguild [PSNS22]. intraspecific [LLSQ24]. introduced [LCS^{+}20, WTvD^{+}22]. introgression [PPB^{+}21, WNHY21]. intrusion [KTH^{+}21]. intrusions [HPNU20, NH23]. invader [LBJS22]. invasion [YWZ^{+}21]. invasions [SBH23, WHZ^{+}21]. invasive [BDMGH^{+}23, MWV22]. inversions [AAB^{+}22]. invertebrate [DBFC24]. invertebrates [HANW21, MKBSK19, SRV^{+}20]. inverted [LSW^{+}20a]. Investigating [FHR^{+}21]. involved [LOW^{+}20, ZSZ^{+}21]. iodine [SOSN23]. Ireland [SOO^{+}21]. iron [BTT23, EBW^{+}20, EFG^{+}24, HYPG^{+}22, HZM^{+}21, HPB^{+}21, JWAvD^{+}22, KLP23, LG21, LHS^{+}21, LTW^{+}24, LHMR23, LMS^{+}21, MGC^{+}20, MWHR^{+}22, SGA^{+}22, SdBW^{+}23, WRS^{+}22, YSHS22, YSHS23, ZJW^{+}22, ZLL^{+}21]. iron-dependent [MWHR^{+}22]. iron-light-CO [YSHS22]. iron-supported [SZY^{+}20]. Irradiance [KKP^{+}21, CT23, KNG^{+}23, MSD23]. irradiated [AFM23]. irradiation [KRB^{+}21]. Island [LVHK20, MSGW^{+}20, MVBG20, FMT^{+}20]. Islands [MSGW^{+}20, SBV^{+}21, SWS^{+}22]. isolation [BRB^{+}22]. isoscapes [YAH^{+}21]. Isotope [MGG^{+}24, CDH^{+}21, DMS^{+}21, HRK^{+}22, KLG^{+}23, LHGJ21, ORRMD^{+}21, SRM23, TMM^{+}20, WDS^{+}23, WAPC20]. isotopes [CDW^{+}24, DWS^{+}22, EST^{+}24, HFL^{+}20, HCH^{+}24, HZC^{+}24, IFU^{+}20, LVDM22, LLC^{+}23, MLP^{+}21a, SWS^{+}22, VCT^{+}23, WRS^{+}22, ZSD^{+}20]. Isotopic [PMJ^{+}21, TZF^{+}22, Ano21c, BSF^{+}21, FXD^{+}22, GAS^{+}23, HLKD23, SSK24, TUR^{+}20, WRWPG19, WMRL22]. isotopomeric [WLK^{+}21]. Issue [Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano20y, Ano20-38, Ano20z, Ano20-39, Ano20-27, Ano20-40, Ano20-28, Ano20-41, Ano20-29, Ano20-42, Ano20-30, Ano20-43, Ano20-31, Ano20-44, Ano20-32, Ano20-45, Ano20-33, Ano20-34, Ano20-46, Ano20-35, Ano20-47, Ano20-36, Ano20-48, Ano20-37, Ano20-49, Ano20-50, Ano20-51, Ano20-52, Ano20-53, Ano20-54, Ano20-55, Ano20-56, Ano20-57, Ano20-58, Ano20-59, Ano20-60, Ano20-61, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21-53, Ano21p, Ano21q, Ano21-29, Ano21r, Ano21-30, Ano21s, Ano21-31, Ano21t, Ano21-32, Ano21u, Ano21-33, Ano21v]. Issue [Ano21-34, Ano21w, Ano21-35, Ano21x, Ano21-36, Ano21y, Ano21-37, Ano21z, Ano21-38, Ano21-27, Ano21-39, Ano21-54, Ano21-55, Ano21-28, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano21-49, Ano21-50, Ano21-51, Ano21-56, Ano21-52, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22a, Ano22q, Ano22r, Ano22-31, Ano22-30, Ano22-43, Ano22s, Ano22-32, Ano22t, Ano22-33, Ano22u, Ano22-34, Ano22v, Ano22-35, Ano22w, Ano22-36, Ano22x, Ano22-37, Ano22y, Ano22-38, Ano22z, Ano22-39, Ano22-27, Ano22-40, Ano22-28, Ano22-41, Ano22-29, Ano22-42, Ano22b, Ano22c, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano22-48, Ano22-49, Ano22-50, Ano22-51, Ano22-52, Ano22-53, Ano22-54, Ano22-55, Ano22d, Ano22-56]. Issue [Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23y, Ano23n, Ano23z, Ano23o, Ano23-27, Ano23p, Ano23-28, Ano23q, Ano23-29, Ano23r, Ano23-30, Ano23s, Ano23-31, Ano23t, Ano23-32, Ano23u, Ano23-33, Ano23v, Ano23-34, Ano23w, Ano23-35, Ano23x, Ano23-36, Ano23-37, Ano23-38, Ano23-39, Ano23-40, Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano23-45, Ano23-46, Ano23-47, Ano23-48, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24i, Ano24f, Ano24j, Ano24g, Ano24k, Ano24h, Ano24l, Ano24m, Ano24n, Ano24o, Ano24p]. Issues [BR22].

Japan [LDB^{+}20, LDB^{+}21, NSK^{+}21]. japonicus [LDB^{+}21]. jelly [CDH^{+}21]. jellyfish [HGF^{+}21, MMKWZ22, MBC^{+}21, OFM^{+}21, TKH21]. juvenile [CSNS24, FBJ^{+}23, LDG^{+}21].

kairomone [HXL^{+}21]. Kaiser [Ano21b]. karst [BPL22]. keep [ETG21]. keeping [FGO^{+}20]. Kelp [vEMF^{+}20, BRHS20, BS20, NB23, SPKM22, SBB^{+}21, SMKS22, TCP^{+}22, WLSD21]. Kenya [MGI^{+}22]. Kerguelen [DAB^{+}21, IJSC20]. kerguelensis [MGC^{+}20]. Kermadec [SML^{+}21a]. key [BTT23, BBL^{+}21, LHS^{+}21, VST^{+}22, ZLJ^{+}24, ZSF^{+}23]. Keys [OLFH21]. killer [CCC^{+}21]. Kinetics [FSS^{+}23, ELS23, GF21, MSB^{+}22]. km [TKW^{+}23]. known [LSS^{+}24]. Kongsfjorden [vdPMF^{+}21]. Krill [NCC^{+}21, CSNS24, YAH^{+}21]. Kuroshio [LHS^{+}20].

lab [BDE^{+}22, SHZ^{+}21]. labile [BKC20, HP20, MFGF^{+}22, NKH^{+}23, WCC20]. lability [BMG^{+}23, PSNG^{+}21]. Lack [KMM^{+}22, BWB^{+}22]. lagoon [BJS^{+}23b, CCDM21, HHMH22, KX22, MBC^{+}21, RWS^{+}22, ZWM22]. lagoonal [YMSH20]. lagoons [BMR^{+}21b]. Lagrangian [BGAD^{+}22, FHR^{+}21, RPA^{+}22]. lake [ACA^{+}22, ASH^{+}20, AHF^{+}22, BVL^{+}23, BAG^{+}24, BTK^{+}21, BGG^{+}20, BGM^{+}22, BMLV24, BR22, CYPRG^{+}23, CESS21, CHS^{+}20, CTH^{+}20, CKCB^{+}23, CZK^{+}23, DHNH22, DAN^{+}22, DCF^{+}22, DdEM^{+}23, EST^{+}24, ESW^{+}21, EPR^{+}24, FBK^{+}22, FWDY20, FBV23, FPB^{+}23, GGHBS^{+}20, GWS^{+}21, GB21, GB23, GKW^{+}20, GKW^{+}21, HZC^{+}24, HNS21, HSB^{+}20, HMB^{+}22, HHK^{+}22, HP20, HZZ^{+}20, JOC^{+}20, JSPS24, KKA22, KPM^{+}20, KRC23, KPG^{+}23, KHSP21, LVDM22, LDS^{+}20, LZM^{+}21, LLBC20, MBA^{+}21, MAG^{+}22, MOW^{+}22, MWHR^{+}22, MGG^{+}24, MNMJ^{+}21, OYM^{+}20, PBC21, PCJ^{+}21, PATL^{+}23, Phi20, Phi21, PFB^{+}22, PSK21, RNW^{+}23, RLL^{+}21, RFW^{+}24, REF^{+}21, RCS^{+}20, SVC^{+}22, SSM^{+}21a, SPBK21, SGA^{+}22, SO23, SGL^{+}23, SCM22, SEK20, SBW21, SGM^{+}21, SZY^{+}20, SDS^{+}23, SGS22, TKM^{+}22, TKS^{+}20, TMM^{+}20, VMT^{+}20, VWR^{+}23, WCZ^{+}20, XTL^{+}20, YWLY20, YJB^{+}20, vGDH^{+}21, AJMG21, CYZZ^{+}20, CWF^{+}22, CTP21]. Lake [DSY20, GZX^{+}22, LEK^{+}22, Lem20, LDB^{+}20, LDB^{+}21, MWV22, MDP^{+}23, MSB^{+}20, NB20, ONH^{+}20, PUH^{+}23, RPA^{+}22, SSP^{+}20, SLZ^{+}21b, SRL^{+}20, THT^{+}20, TZF^{+}22, VMT^{+}20, XDQ^{+}22, XTL^{+}20, XMP^{+}21]. lake-ice [YWLY20]. lake-level [BGM^{+}22]. lake-water [MNMJ^{+}21]. lakebed [NH23]. Lakes [RD23, VCBB20, AVJ^{+}24, AKT^{+}21, Ano21a, BSIdG23, BMA^{+}20, BLS^{+}22, BGM^{+}22, BMS^{+}22, BV21, BPG^{+}22, CLA^{+}24, CCA^{+}21, CLG^{+}24, CM20a, CM20b, DAD^{+}21, DKW^{+}21, DMCW23, EAF^{+}22, GGR24, GSG^{+}17, HM20, HABL^{+}21, IMB^{+}22, JSPS24, JSB^{+}20, KVKS22, KSH^{+}23, LM24, LAD^{+}22, LWH^{+}23, LHH24, LPA^{+}23, MBH^{+}23, MTP^{+}23, MLC^{+}21, MSGW^{+}20, MVBG20, MSJ21, MHVC22, MNMJ^{+}21, PGEB22, PMD^{+}23, PC21, PW22, PFB^{+}22, PLB^{+}21, QFM^{+}21, RHA^{+}23, RCMA21, RSP^{+}24, SJH20, SPBK21, SGGB21, SMCS23, SPLJA^{+}20, SSWM20, SLHL^{+}22, TKW^{+}23, TATC^{+}24, VMT^{+}24, VKK^{+}21, VST^{+}22, XTJW21, YJ20, ZSK^{+}21, ZLJ^{+}24, ZZWL22, ZZB^{+}20, DLC^{+}22, GSP^{+}23, GBO^{+}20, MFM^{+}21, MBB^{+}20, SWOR20]. lakeside [MOW^{+}22]. Laminaria [SOO^{+}21]. Land [BSB^{+}20, BGM^{+}21, RTM^{+}20, BCTH20, MGI^{+}22, SML21b]. landfast [JHvL^{+}23]. Landscape [ASH^{+}20, CZH^{+}21, DBFC24, LHL^{+}21, LRM21, PFB^{+}22, PLB^{+}21, SZA^{+}21]. landscapes [LVZ21, NDP^{+}24]. Large [XWL^{+}21, ACW^{+}22, BCL^{+}22, BGAD^{+}22, CCY^{+}21, CMCC24, CHS^{+}20, CTH^{+}20, DHNH22, DMC^{+}22, EST^{+}24, EPR^{+}24, FWDY20, FBV23, GEHD20, HZC^{+}24, HLPT20, HMB^{+}22, IÁSNCR21, JOC^{+}20, KRC23, LM24, LAFV23, LHL^{+}21, MGL22, MVP^{+}20, MLS^{+}21, MOW^{+}22, PGB24, RHP^{+}23, SH22, SO23, SGM^{+}21, SDSGR21, TION23, TBH^{+}22, VMT^{+}20, WCZ^{+}20, XDQ^{+}22, YWLY20, ZLU^{+}20, ZYZ^{+}21]. large-scale [BGAD^{+}22, SH22]. Larger [FLH^{+}23, SGGB21, TION23]. largest [HTLP23]. larvae [CPBJR^{+}21, GFG^{+}22, GPDA^{+}20, OLR^{+}23]. Larval [GLN^{+}21, APB^{+}23, GLN^{+}20, MLP21b, MURK20, MDWT^{+}22, RPA^{+}22, SSA^{+}23]. last [TMKB^{+}22]. lasting [SH22]. Late [SSR^{+}23, BAHH^{+}20, SBM^{+}21]. Late-season [SSR^{+}23]. Lateral [CTH^{+}20, NH23]. latitude [JSB^{+}20, KCT23, LDG^{+}21, SLHL^{+}22]. latitudes [ARB^{+}22, RHM^{+}21]. Latitudinal [GFW^{+}20, BSGA21, JLS^{+}20, JSL^{+}23, LCS^{+}20, RCMA21, ZZB^{+}20]. Laurentian [GSP^{+}23, MFM^{+}21, SWOR20]. layer [DHNH22, EFG^{+}24, GBC^{+}20, NH22, NH23]. layers [PLLM21, SBW21]. leads [BMH^{+}23]. leaf [FPG^{+}24, FRRG^{+}21, ZN20]. leaf-litter [FRRG^{+}21]. learning [MT23, OAM^{+}22, RWMP^{+}23, SCC^{+}21]. leaved [VJS^{+}20]. leidyi [CCG^{+}22]. Length [SGM^{+}22]. less [MN20]. leucine [GE22]. level [BGM^{+}22, BPG^{+}22, FBK^{+}22, FNB^{+}23, LVHK20, LRM21, QMG21, YLT^{+}20, ZvBZ^{+}20]. levels [EWGP22, KTA^{+}21, RBAC^{+}20, SPBK21, YWLY20]. Leveraging [HMBB21]. Liaohe [PYY^{+}20]. lidar [CZS^{+}22]. Life [BST^{+}22, DFW^{+}20, HXL^{+}21, LZG21, OFM^{+}21, ZLJ^{+}24]. ligand [KLP23, WXK^{+}22]. Light [DPB^{+}23, GEHD20, HPCH24, MCH^{+}22, PCC^{+}23, SSB^{+}22, AVJ^{+}24, BS20, CFC^{+}20, CBT^{+}21, CGB^{+}21, DMM^{+}21, FNB^{+}23, GR24, GK22, HZS21, HPP^{+}21, HS21, JWAvD^{+}22, KKP^{+}21, KSR20, KLP23, MTB^{+}21, MGC^{+}20, MLG^{+}20, Phi20, Phi21, QBJG21, RGW^{+}23, SMKS22, TKA21, TPTS23, WWFS21, WRHR22, YSHS22, YSHS23, ZXZ^{+}23]. Light-driven [MCH^{+}22]. Light-use [HPCH24]. limitation [AFH^{+}21, APB^{+}23, AWGV20, BYS^{+}23, BVGQN22, DCK^{+}22, GCW^{+}23, KHK^{+}22, LTW^{+}24, MDP^{+}23, MGC^{+}20, TPTS23, hTRH20, VWR^{+}23, WK21, WRS^{+}22, YSHS22, YAE^{+}24, ZFQ^{+}20]. Limited [BAD^{+}21, TMS^{+}21, TCP^{+}22, ASH^{+}20, EJM^{+}23, GVF^{+}23]. limiting [NB23]. limits [AR20, GAS^{+}21]. Limnol [Ano21b]. Limnology [Ano21a, Kir20]. Lineage [WNHY21, MNM^{+}21]. lineages [BRB^{+}22, FM21]. Linear [GR24, FGO^{+}20, TBH^{+}22]. link [TVB^{+}21, VGSB23]. Linkages [ZQB^{+}21, FLW^{+}21, HCS^{+}21]. linked [BCFI^{+}22, LNR^{+}20, PWS^{+}20, PFB^{+}22, RvHF^{+}20]. Linking [FAB^{+}20, GMMW20, dMKB^{+}20, FJP^{+}21]. links [ARB^{+}22, KHB22, RNW^{+}23, SJJ^{+}21]. Lipid [PMF^{+}21, HNS21]. lipidome [BCFI^{+}22]. lithium [FSH^{+}23]. litter [BRB^{+}20, FRRG^{+}21]. litter-associated [BRB^{+}20]. littoral [VMT^{+}20, VKK^{+}21]. Live [BFW^{+}22]. livestock [GBMKJ22]. living [WPY^{+}20]. lno.10504 [Ano21a]. load [FBV23, MDBI20, RS21, TBH^{+}22, WHAR^{+}20]. loading [DSY20, HMB^{+}22, JOC^{+}20, LTH21, SDL22, VMT^{+}20, XMP^{+}21]. loads [Roy20]. loathing [PVSP^{+}22]. lobate [CCG^{+}22]. lobsters [BFCC^{+}21]. Local [FRRG^{+}21, KPNJ23, SLJ^{+}22, OYM^{+}20, SMG^{+}22]. Local-scale [KPNJ23]. location [PHM21]. locomotion [JYA21]. Long [BBMD20, BMH^{+}23, CB21a, CHA^{+}24, FBV23, GBHF21, LAD^{+}22, LGQ^{+}21, MVP^{+}20, MNMJ^{+}21, RSLW23, SJH20, BCWS23, DAD^{+}21, EDC24, FHMK20, HBR^{+}22, LVHK20, Lem20, LZM^{+}21, MKBSK19, SH22, SGM^{+}21, YAH^{+}21, MKBSK19]. long-lasting [SH22]. Long-term [BBMD20, BMH^{+}23, CB21a, CHA^{+}24, FBV23, GBHF21, LAD^{+}22, LGQ^{+}21, MVP^{+}20, MNMJ^{+}21, RSLW23, SJH20, BCWS23, DAD^{+}21, FHMK20, HBR^{+}22, LVHK20, Lem20, LZM^{+}21, MKBSK19, SGM^{+}21, YAH^{+}21, MKBSK19]. Longitudinal [HCS^{+}21]. loop [RNW^{+}23, VGSB23, VKGW22]. loops [GM22]. loss [BH21, CM21, GMS21, SSS^{+}20, ZQB^{+}21]. Low [BBL^{+}21, PBP^{+}23, BTT23, BS20, BGM^{+}21, BPG^{+}22, CTM^{+}21, CBHB21, CBD^{+}23, EBK^{+}20, FRRG^{+}21, KRB^{+}21, KKP^{+}21, LDGB^{+}24, MCYR20, OBL20, PUH^{+}23, SDB^{+}23, TKA21]. low-energy [SDB^{+}23]. low-level [BPG^{+}22]. low-light [KKP^{+}21, TKA21]. low-molecular [BGM^{+}21]. low-oxygen-adapted [CBD^{+}23]. low-quality [FRRG^{+}21]. low-salinity [OBL20]. low-sulfate [PUH^{+}23]. lower [EJM^{+}23, FPBG^{+}22, TMSL22, DSY20]. lowland [MHL^{+}20, PHB^{+}22, RWT^{+}21]. lysates [WCC20, ZLW^{+}21a].

m [BGM^{+}22]. MA [LVHK20]. Machine [OAM^{+}22, MT23, RWMP^{+}23, SCC^{+}21]. macro [PBP^{+}23, WLSD21]. macro-tidal [PBP^{+}23, WLSD21]. macroalgae [BDMGH^{+}23, DK20, LRRP20, RS21, SH22, SHBA22, SWS^{+}22]. macroalgae-dominated [SH22]. macroalgal [BMR^{+}21b, CCFF21, RMH^{+}22, TPTS23, WBB^{+}21]. macrobenthic [SYT^{+}21]. Macrocystis [SBB^{+}21]. Macrofauna [NvBE^{+}21, HLPT20]. macrofaunal [DvOH^{+}20]. macroinvertebrate [LLC21, LHL^{+}21, MHdS^{+}21]. macroinvertebrates [VKK^{+}21, VST^{+}22]. macromolecular [EHC^{+}22, RRAO^{+}21, ZLS^{+}21]. macronutrients [HKM^{+}20, MHP^{+}23]. macrophyte [GGHBS^{+}20, LRNG24, MCH^{+}21, OGD20, SLHL^{+}22, WCZ^{+}20]. macrophyte- [WCZ^{+}20]. macrophytes [HGMK22, HZZ^{+}20]. macrotidal [YXC^{+}22]. magellanicus [CGR22]. magna [LBvE23, RVBP23]. magnitude [SMW^{+}21, WOT^{+}23]. main [BAG^{+}24, SWS^{+}22]. Maine [DP22, SMP21]. mainly [OYM^{+}20]. major [BCFI^{+}22, EHC^{+}22, HSD^{+}20, KDZ^{+}24, MBB^{+}20, NvBE^{+}21, YXC^{+}22]. Malawi [NB20]. management [DFJ22, FGP22, XMP^{+}21]. Manganese [LWvdM^{+}21, SZY^{+}20, BTT23, CLD^{+}23, HPB^{+}21, LHS^{+}21, MCS^{+}23, SGA^{+}22]. Manganese/iron [SZY^{+}20]. Manganese/iron-supported [SZY^{+}20]. mangrove [BSB^{+}20, CSC^{+}21, CMCC24, CEE^{+}22, EJM^{+}23, HFL^{+}20, KHK^{+}22, KDZ^{+}24, LAB^{+}22, PBP^{+}23, RMW^{+}21, SSS^{+}20, YOXC23, vHHP^{+}21, vHHP^{+}22]. mangrove-dominated [CMCC24, RMW^{+}21]. mangrove-fringed [KDZ^{+}24]. mangrove-seagrass [KHK^{+}22]. mangroves [HMP^{+}21]. mapping [PSY^{+}22]. Mara [MGI^{+}22]. marginal [FXD^{+}22, LLK^{+}22, LCZ^{+}21, WXK^{+}22, YWZZ20]. mariculture [XYL^{+}23]. marina [GGS^{+}21, SIP21]. Marine [EFC^{+}23, MHD^{+}22, SRV^{+}20, ARPFGS23, AMC21, BMG^{+}23, BHSG^{+}20, BMCS23, BPIBN20, BFE^{+}20, CML23, CBT^{+}21, CHA^{+}24, DCJB22, DBB^{+}21, DWW^{+}21, DCK^{+}22, EHC^{+}22, EJRRC^{+}23, EWF22, GBW24, GM22, GFW^{+}20, GAS^{+}23, HLKD23, HVMS^{+}21, JHSG^{+}24, JHS^{+}21, JLMW21, KOR^{+}20, LCMG^{+}20, LG21, LD22, LCM^{+}22, LSL^{+}20, LN21, LMSN23, MLK^{+}22, MNM^{+}21, MUGL20, MLP21b, MZW^{+}21, MK20, OBL21, OGD20, PPB^{+}21, PDK^{+}23, RBR^{+}24, SJR21, SB22, SEG21, SGK^{+}21, ST22, SSK24, TCS21, VKB^{+}22, WWGP21, YAP^{+}22, YSHS22, ZR23, ZZHZ23, ZLL^{+}24, ZLL^{+}21, ZFQ^{+}20]. Marked [IJSC20]. marker [DAB^{+}21]. marsh [BVM^{+}20, CZH^{+}21, DKPS23, EJM^{+}23, HSBC20, KBB^{+}22, LVHK20, LSC^{+}21, MEQBV20, OJUN23, PNT^{+}21, RWMP^{+}23, SVBT20, SRN^{+}21, SWK^{+}23, XWL^{+}21, ZLN21]. marsh-influenced [LSC^{+}21]. marshes [GBMKJ22, LGF22, MKM^{+}23, QMG21, TEK^{+}21, YLT^{+}20]. masks [BYS^{+}23, LDB^{+}20]. mass [BAG^{+}24, ESW^{+}21, JYA21, KBSA21, LVDM22, LLC^{+}23, RFB^{+}22, RST^{+}21, SH22, YWZ^{+}21]. masses [EBK^{+}20, FBRG23]. Massive [MBLA^{+}21, YAP^{+}22]. Masthead [Ano20z, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano20-37, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-54, Ano21-28, Ano22r, Ano22-30, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano22b, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano24e, Ano24f, Ano24g, Ano24h]. mat [MVBG20]. matching [LEB21]. material [HPF^{+}23]. matrix [GRJ20]. mats [MVBG20]. matter [AVJ^{+}24, ABS^{+}21, AKT^{+}21, AMC21, BAG^{+}24, BSF^{+}21, BBH^{+}20, BAD^{+}21, BGG^{+}20, BYS^{+}23, BJG^{+}22, BZK^{+}22, BGM^{+}21, CPB^{+}21, CTP21, CLC^{+}22, CM21, CCDM21, DVL^{+}23, DLC^{+}22, FAB^{+}20, FXD^{+}22, GMI^{+}21, GAS^{+}23, HOB^{+}22, HHK^{+}22, HCS^{+}21, JSB^{+}20, KAP^{+}20, KDZ^{+}24, KBB^{+}22, KHCG^{+}22, LBBM21, LRdGFP23, LMM^{+}21, LNR^{+}20, LD22, LSC^{+}21, LTH21, LXC^{+}21b, LPO^{+}20, LLTT^{+}24, MGL22, MEQBV20, PBC21, PSNG^{+}21, PWS^{+}20, PGN^{+}21, RvHF^{+}20, RLP^{+}20, RCMA21, SB20, SB22, SEK20, SSWM20, SCB^{+}20, SGSD^{+}22, TSX^{+}23, WCC20, YGD^{+}21, ZCZ^{+}23, ZZT^{+}21, ZYZ^{+}21, ZZB^{+}20, dMKB^{+}20]. matters [FGP22, TPTS23]. maturation [RVBP23]. Maud [DLDF21]. maxima [SWOR20]. maximum [SMNT22, SPLJA^{+}20]. may [HVMS^{+}21, MHD^{+}22, OJUN23, SSEO21]. Meadow [THVJ23, CB21b, SVMM^{+}22, VLP^{+}21]. meadows [AHCF22, CCFF21, JREVB21, SN22, dSNIB22]. means [MGG^{+}24]. measurable [CSD^{+}24]. measured [DAD^{+}21, WO21]. Measurement [EAF^{+}22, MHMML^{+}22]. measurements [BVM^{+}20, CZS^{+}22, DLDF21, KSR20, Lem20, SSM^{+}22a, SEG21]. mechanical [XSZ^{+}21]. mechanics [CCG^{+}22, CUS21, SMF20]. mechanism [GLN^{+}20, YXC^{+}22]. Mechanisms [BWB^{+}22, PLLM21, STAK22, BPL22, WDS^{+}23]. Mechanistic [ALG^{+}23]. mechanoreceptional [SYMF21]. mediate [LWS^{+}22]. mediated [GLM^{+}23, SRN^{+}21, VJS^{+}20, WKRL^{+}23]. mediates [PCJ^{+}21]. mediating [HLPT20]. mediation [CBN21, WMTJW22, ZTT23]. Mediterranean [GCEM^{+}23, GLM^{+}23, MGP^{+}21, MBC^{+}21, RHS^{+}22, SORGC^{+}21]. meet [HMBB21]. Meiler [ZR23]. meiobenthic [LBG^{+}22, RLP^{+}20]. meiofauna [MHG^{+}23, PCN22]. melt [KPG^{+}23, VMP^{+}21]. meltwater [MCS^{+}23]. Members [Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano20-42, Ano20-43, Ano20-44, Ano20-45, Ano20-46, Ano20-47, Ano20-48, Ano20-49, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-55, Ano21-40, Ano22-31, Ano22-43, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano22c, Ano23y, Ano23z, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano24i, Ano24j, Ano24k, Ano24l]. Mercenaria [MURK20]. Mercury [CESS21, RCMA21, BAG^{+}24, BCA^{+}20, SGGB21, TUR^{+}20]. mercury-cycling [BCA^{+}20]. mercy [RHM^{+}21]. Meretrix [SYT^{+}21]. meromictic [CYPRG^{+}23, TZF^{+}22]. mesocosm [SSM^{+}22a]. mesocosms [HS21, RLL^{+}21, TVB^{+}21]. mesopelagic [BBM21, CLT^{+}22, LPO^{+}20, OLR^{+}23, SWC^{+}23, WSZ22a, WSZ22b]. mesophotic [FPBG^{+}22, LMS23, MSL21, RFB^{+}22, SWS^{+}22]. mesopredator [RS20]. mesoscale [ACE^{+}22]. mesotrophic [TMM^{+}20]. Mesozooplankton [CDA^{+}20, GNO22, SSM21b]. meta [BR22, CSM^{+}23, PHM21]. meta-barcoding [CSM^{+}23]. meta-ecosystem [PHM21]. meta-ecosystems [BR22]. Metabolic [Ano21a, DMC^{+}22, GSG^{+}17, BMR^{+}21a, CYPRG^{+}23, MSG^{+}22, MGGRR^{+}21]. Metabolism [RRHS^{+}21, BBMD20, BYS^{+}23, BCL^{+}22, BHA^{+}22, CB21b, CZK^{+}23, EDC24, EJRRC^{+}23, HHK^{+}22, KVKS22, LAD^{+}22, LRdGFP23, MA21, MGWS23, MRT^{+}24, PHB^{+}22, Phi20, Phi21, RVS^{+}22, SBB20, TCP^{+}22, WCC20]. Metabolite [JLS^{+}20, JSL^{+}23]. metabolites [LSS^{+}24]. metabolome [BMGK22]. metacommunity [GGHBS^{+}20, GGLH^{+}22]. Metagenomics [BVM^{+}20, Kir20, BWV^{+}20, GMMW20, MVBG20]. metal [DBB^{+}21]. metalimnion [Ano21a, GSG^{+}17]. metals [BBA^{+}23, HKM^{+}20, KDZ^{+}24, WOMWR20]. Metamorphosis [BFCC^{+}21]. metapopulation [GRJ20]. metatranscriptomes [LABM20]. Meteorological [AH20]. meter [BKMY21]. Methane [LVDM22, LDS^{+}20, YWZZ20, AHCF22, AWB^{+}23, AGC^{+}21, BVL^{+}23, BBH^{+}20, DHNH22, GMI^{+}21, GPK^{+}21, GKW^{+}20, GKW^{+}21, HGMK22, HNS21, KWS^{+}21, LHD^{+}20, LPA^{+}23, MWHR^{+}22, PSK21, RHA^{+}23, RWT^{+}21, RSH^{+}20, SPBK21, SBW^{+}20, SSM^{+}22b, SZY^{+}20, TKS^{+}20, TMM^{+}20, VJS^{+}20, VJH^{+}22, YOXC23, ZLL^{+}24, vGDH^{+}21]. methane-derived [TKS^{+}20, vGDH^{+}21]. methanogen [BBH^{+}20]. methanogenesis [BVL^{+}23, SSM^{+}21a]. Methanogens [EJM^{+}23]. methanotroph [vGDH^{+}21]. method [LAB^{+}22]. methylating [CBB^{+}22]. methylmercury [BC20, TKW^{+}23, YJB^{+}20]. Methylobacter [vGDH^{+}21]. Methylotenera [vGDH^{+}21]. methylotroph [vGDH^{+}21]. metric [KKA22]. metropolitan [RTM^{+}21]. Mexico [KPS21, MLP^{+}21a, PMRG20, WSZ22a, WSZ22b, YMSH20]. Meyer [Ano21b]. Meyer-Kaiser [Ano21b]. Michigan [CTP21, RPA^{+}22]. micro [MMLPRHCG22]. micro-computed [MMLPRHCG22]. microalga [GAS^{+}21, RGW^{+}23]. microalgae [BLS20, WRHR22]. microalgal [SRM23]. microbe [TKH21]. microbes [BRB^{+}20, BFE^{+}20, LABM20, ST22, TSX^{+}23]. Microbial [BRB^{+}20, MAD^{+}22, RvHF^{+}20, RCS^{+}20, BHSG^{+}20, BSGA21, BGG^{+}20, BWV^{+}20, BMLV24, BZK^{+}22, CYPRG^{+}23, CLC^{+}22, CGL^{+}21, DBB^{+}21, DCF^{+}22, FAB^{+}20, FPG^{+}24, FLW^{+}21, GM22, GMMW20, GPK^{+}21, HBY^{+}22, HMB^{+}20, HLS^{+}22, IMB^{+}22, KHSP21, LD22, LZD^{+}23, MVBG20, MCYR20, OBB^{+}20, OYM^{+}20, PWS^{+}20, RLL^{+}21, RBAC^{+}20, SBB20, SB20, SGSD^{+}22, TMS^{+}21, VGSB23, VKGW22, VJH^{+}22, XKP^{+}22, XTJW21, ZR23, ZTT23, vEMF^{+}20]. microbialite [IMB^{+}22, MVBG20]. microbialite-bearing [IMB^{+}22]. Microbially [ZCZ^{+}23]. microbiology [JLMW21]. Microbiome [CYZZ^{+}20, HMM^{+}22, SSS^{+}24]. microbiomes [HZZ^{+}20, XTL^{+}20]. microcrystals [LN21]. microcystin [CWF^{+}22]. Microcystis [CLC^{+}20, CLC^{+}21, MDP^{+}23]. microenvironment [EWGP22]. microenvironments [EBW^{+}20]. microevolutionary [ZZCD23]. microfocus [IKN^{+}21]. microheterogeneity [KHSP21]. microorganisms [CBB^{+}22, CKH^{+}21, EPR^{+}24, MGG^{+}24]. microphytobenthos [DvOH^{+}20, ORE20]. microplankton [KHHZ20]. Microplastic [BZSF23, DWR^{+}23, EAF^{+}22]. microplastics [GLM^{+}23, KWB21]. Microscopy [CBN21, KAC^{+}23]. microsporidian [CML23]. Microzooplankton [MMG^{+}21, LCZ^{+}21, LNC^{+}23]. Mid [KCT23, RHM^{+}21]. mid-latitudes [RHM^{+}21]. Mid-summer [KCT23]. midlatitude [BAHH^{+}20]. midwater [VSB^{+}23]. Midwest [BJG^{+}22]. migrating [PJRV23]. migration [BJS23a, BBM21, GNO22, TGE^{+}21, WSZ22a, WSZ22b]. millennial [KSH^{+}23, vHLV^{+}20]. millennial-aged [KSH^{+}23]. mimics [GLN^{+}20]. mine [BWV^{+}20]. Mineral [EBW^{+}20, PYY^{+}20, SdBW^{+}23, VJS^{+}20]. mineral-bound [SdBW^{+}23]. mineral-soil [VJS^{+}20]. mineralization [LMM^{+}21, vEMF^{+}20]. Minerogenic [MKM^{+}23, KBB^{+}22]. minima [SCT^{+}21]. minimum [CCK^{+}21, THL^{+}21, WRA^{+}23, LZG21]. mining [CLGH23, LMS^{+}21, OBB^{+}20]. minute [GK22]. missing [VGSB23]. Mississippi [GBHF21, LTH21, Roy20, TMSL22]. mitigate [SSEO21]. mitigates [MDWT^{+}22]. mixed [SBW21, SSB^{+}22]. Mixing [CM20b, EPR^{+}24, YWLY20, AVS^{+}22, GK22, HPP^{+}21, IFU^{+}20, LLC^{+}23, LXC^{+}21a, NH22, PFH21, PW22, ZSF^{+}23]. Mixing-driven [EPR^{+}24]. mixoplankton [GM22]. Mixotrophic [GDB20, AH20, ELS23, LES^{+}21]. Mixotrophy [TCS21, BH21]. mixtures [FDHS20]. Mn [BYB^{+}23, CLD^{+}23, dCLO^{+}23]. Mn-oxidizing [CLD^{+}23]. Mnemiopsis [CCG^{+}22]. Model [OLR^{+}23, BST^{+}22, CYPRG^{+}23, DFJ22, DK20, GMNG^{+}24, GMI^{+}21, HPF^{+}23, HDS^{+}20, IFU^{+}20, LZG21, LTLH22, PSNS22, WRA^{+}23, WSZ22a, WSZ22b, ZLN21]. Modeled [LAB^{+}22, MURK20]. Modeling [CLGH23, GLM^{+}23, LVHK20, PC21, SBB20, YGD^{+}21, APB^{+}23, BSF^{+}21, BR22, KBSA21, MAG^{+}22, PHB^{+}22, WRMP22]. models [KSF^{+}23, LCMG^{+}20, MBD^{+}23, RPA^{+}22, WMTJW22]. moderate [MHD^{+}22]. modern [SLHL^{+}22, vHLV^{+}20]. modification [GÖA^{+}23, NTJ^{+}20, TCP^{+}22]. modified [KFHS24, RCMA21]. modify [MK21]. modular [MMLPRHCG22]. modulate [SMF20, WWFS21, ZBNH^{+}21]. modulated [GR24]. modulates [ABC^{+}21, KKP^{+}21, WOT^{+}23]. modulating [JREVB21, TGE^{+}21]. modulation [CHAA21]. Molecular [LD22, SB20, BGM^{+}21, CLC^{+}22, HDS^{+}20, MGC^{+}20, ZYZ^{+}21]. molecules [LPO^{+}20, ZLW^{+}21a]. molybdenum [FSH^{+}23]. Monitoring [ASL^{+}24, BCWS23, FJP^{+}21, THS21]. monocultures [FDHS20]. monomictic [GWS^{+}21, SGS22, XTL^{+}20]. montane [LHL^{+}21]. Monterey [FHMK20, TRS^{+}20]. month [GWS^{+}21]. monthly [IOS24]. moon [OLR^{+}23]. Morocco [NEB^{+}22]. morphological [GWS^{+}21]. morphologically [BRCO21]. Morphology [BRHS20, MDWT^{+}22, OBL21, SRN^{+}21, SGM^{+}22]. mortality [CCC^{+}21, KSI24, SH22, Wei24]. mosaic [CEW^{+}21, LTW^{+}24]. mostly [MBH^{+}23]. motion [JvKvT^{+}20]. Mountain [VMT^{+}24, CCA^{+}21, CGGC^{+}21, CESS21, ESW^{+}21, JSPS24, MHVC22, PFB^{+}22, WK21]. Mountains [VMT^{+}24]. mouths [EST^{+}24]. Moving [FDF^{+}21]. Mu [CLGH23]. mucus [BH21]. Mud [MEQBV20]. Mud-associated [MEQBV20]. muddy [CGD22, MHG^{+}23]. Multi [FR23, ZSZ^{+}21, GCEM^{+}23, vHLV^{+}20]. multi-millennial [vHLV^{+}20]. Multi-omics [ZSZ^{+}21]. multi-taxon [GCEM^{+}23]. Multi-year [FR23]. multiarm [ILPL20]. multibasin [ILPL20]. multibeam [PSY^{+}22]. Multiclonal [LBvE23]. Multidecadal [KLO23, BJG^{+}22]. multifactorial [WBB^{+}21]. Multigenerational [dJGCS23]. multilake [DMM^{+}21]. Multiple [KBvdD^{+}20, CBPNA22, CESS21, GBW24, HHMH22, HMBB21, KDV^{+}23, LHL^{+}21, MVP^{+}20, WPO^{+}22, ZJ22b]. Multiscale [SCM22]. multistressor [DK20]. multitasking [GM22]. multitrophic [BPG^{+}22]. multivariate [ZJ22a]. multiyear [SGS22]. mussel [NTJ^{+}20]. mussels [LGQ^{+}21]. mutualistic [FPBG^{+}22]. mysid [BWSW24]. Mytilus [NTJ^{+}20].

N [ZLW^{+}21a, AAC^{+}21, AFH^{+}22, BSGA21, DRA^{+}23, HBS^{+}22, HZC^{+}24, HMB^{+}22, LLL^{+}23, LTLH22, MLL^{+}21, SCC^{+}21, SWS^{+}22, TZF^{+}22, hTRH20, WWGP21]. N-containing [ZLW^{+}21a]. nano [KKP^{+}21]. nano-diatom [KKP^{+}21]. nanoflagellate [LES^{+}21]. nanoflagellates [STAK22]. nanoparticle [YPF^{+}23]. nanoplankton [KPS21]. nanoscale [XSZ^{+}21]. narcomedusa [VSB^{+}23]. narrow [SLJ^{+}22]. national [PATL^{+}23]. national-scale [PATL^{+}23]. native [BDMGH^{+}23, HSP^{+}20, LCS^{+}20, MWV22, WTvD^{+}22]. Natural [FBK^{+}22, AFM23, CTM^{+}21, DVL^{+}23, HBR^{+}22, LMSN23, NDP^{+}24, RBL^{+}20, SdBW^{+}23, vHLV^{+}20]. naturally [MNMJ^{+}21, SSB^{+}22]. Nature [LSW^{+}20a]. Near [PMD^{+}23, RBAC^{+}20, KPS21, NCC^{+}21, SBV^{+}21]. Near-bed [PMD^{+}23]. nearly [YG23]. nearshore [MOW^{+}22, MK21, SO23, SOT^{+}20]. negates [WM20]. negative [MDWT^{+}22, SSEO21]. neglected [ZZL^{+}21]. negligible [AHCF22]. neighboring [ZvBZ^{+}20]. Neogoniolithon [GTKW^{+}20]. neogracile [KMM^{+}22]. Neotropical [BWB^{+}22]. neritic [CPBJR^{+}21]. nerka [ASH^{+}20]. Net [CB21b, MSG^{+}22, BBCA^{+}21, CBC^{+}22, LSDA21, ZLL^{+}21]. network [AFH^{+}22, GRJ20, LWS^{+}22, MHSL22, PSE^{+}20, RVS^{+}22, SLRS^{+}24, WMTJW22]. network-scale [SLRS^{+}24]. networks [BGAD^{+}22, DPB^{+}23, LGF22, SJH20, ZZT^{+}21]. neural [SJH20]. neutrally [FGO^{+}20]. Nevada [CM20a]. Ni [LTLH22, hTRH20]. Niche [TKA21, BO20, CYPRG^{+}23, CHA^{+}24, CUS21, CGB^{+}21, EJM^{+}23, HCH^{+}24, KKP^{+}21, LTJ^{+}20, MCYR20, SSS^{+}24]. niches [Gil22, RFB^{+}22, SHZ^{+}21]. nifH [MBD^{+}22]. night [KGG^{+}23, MLG^{+}20, vdPAV^{+}20]. Niño [WRS^{+}22]. Nitrate [AWGV20, LXC^{+}21a, vGDH^{+}21, BVM^{+}20, CEW^{+}21, CDW^{+}24, LTH21, RWS^{+}22, SZA^{+}21, WRS^{+}22, WAPC20, XLN21]. Nitrification [LMC^{+}21, HRK^{+}22, HGF^{+}21, PCC^{+}23, SWB^{+}20]. nitrifiers [BMCS23]. nitrite [ETMD23]. Nitrogen [AHCJ^{+}20, JHSG^{+}24, LLTT^{+}24, SWS^{+}22, WLK^{+}21, ABS^{+}21, ARBvS^{+}21, AMO20, AVS^{+}22, BRB^{+}20, BDL^{+}21, BKMY21, BRM^{+}23, CBF^{+}23, CDW^{+}24, DWW^{+}21, DWS^{+}22, FBP20, FPB^{+}23, GBO^{+}20, HRK^{+}22, HHMH22, HANW21, HSBC20, HSD^{+}20, HMB^{+}22, KTA^{+}21, KRC23, KHHZ20, LDGB^{+}23, LTH21, MRC^{+}24, MNM^{+}21, MLL^{+}21, MCYR20, ORE20, RGP^{+}22, RS21, Roy20, STB22, SMW^{+}21, SWZ^{+}23, SHE^{+}21, SML21b, SLZ^{+}21a, SSR^{+}23, THVJ23, VWR^{+}23, WWFS21, WRS^{+}22, WDS^{+}23, WKRL^{+}23, YG23, YAE^{+}24, ZQB^{+}21, ZBNH^{+}21, ZFQ^{+}20, ZXZ^{+}23]. nitrogen-enrichment [HHMH22]. nitrogen-fixing [FBP20, YG23, ZFQ^{+}20]. nitrogenase [LTLH22]. Nitrous [TTC^{+}22, AHCF22, FSS^{+}23, HBR^{+}22, LMC^{+}21, LPMBR23, RTM^{+}20, RTM^{+}21, WLK^{+}21, ZZO^{+}21]. No [CSH21, CBS^{+}21, PPB^{+}21]. nocturnal [MDIY23]. nodosa [JREVB21]. Non [FBP20, HMB^{+}22, ZXZ^{+}23]. non-algal [ZXZ^{+}23]. Non-cyanobacterial [FBP20]. non-N-fixing [HMB^{+}22]. Nonlinear [BJG^{+}22, RFW^{+}24, SPHH22, HHK^{+}22, LHH24, PRS^{+}20]. nonphotochemical [CGB^{+}21]. Nonrandom [GMS21]. nonvisual [BBM21]. North [BYB^{+}23, EBK^{+}20, ETMD23, FSS^{+}23, THT^{+}20, LAD^{+}22, MLC^{+}21, SML21b, AHCJ^{+}20, ADLW22, BMR^{+}21a, BAD^{+}21, BAHH^{+}20, CKH^{+}21, CJB^{+}21, DNB^{+}20, DWS^{+}22, DNA^{+}22, ETG21, FDF^{+}21, FIH^{+}20, GFW^{+}20, HBG^{+}22, HYPG^{+}22, KBvdD^{+}20, LTJ^{+}20, LNC^{+}23, MWV22, MMG^{+}21, NvBE^{+}21, PDK^{+}23, RPKC22, SBW^{+}20, SSM^{+}22a, SAD^{+}20, TKA21, WF22, WRS^{+}22, YWZZ20, YWZ^{+}21, YAE^{+}23, YAE^{+}24, ZBNH^{+}21]. North-Atlantic [BMR^{+}21a]. north-temperate [MLC^{+}21]. Northeast [PYY^{+}20, KMV^{+}20, SWC^{+}23, ZZB^{+}20, MMF^{+}21, SPKM22, SSM21b]. Northern [MURK20, BTK^{+}21, CLG^{+}24, CDA^{+}20, CZK^{+}23, FSM^{+}21, JSB^{+}20, KHB22, KPL^{+}20, KBL^{+}20, MHP^{+}23, NEB^{+}22, RSH^{+}20, RSP^{+}24, VMT^{+}24, CMT^{+}20]. northwest [LSCL21]. northwestern [BDL^{+}21, KPS21, LSS^{+}24, MZXY22, MY23, MLP^{+}21a, YMSH20, SWS^{+}22]. novel [APB^{+}23, ASL^{+}24, GLN^{+}20, HGH^{+}21]. novo [DNA^{+}22]. novo-discovery [DNA^{+}22]. nuisance [GAS^{+}21, RS21]. null [LD22, SB22]. numbers [ZR23]. numerical [GMNG^{+}24]. nursery [GPDA^{+}20]. Nutrient [BYS^{+}23, BS20, BAHH^{+}20, BVGQN22, LSDA21, NSK^{+}21, PKKS20, VMT^{+}20, VWR^{+}23, AWGV20, ALH^{+}21, BGM^{+}21, BRM^{+}23, CAV^{+}22, CFC^{+}20, CLGH23, CM21, CMB^{+}21, DCF^{+}22, DKBG21, DKPS23, EFC^{+}23, FMT^{+}20, FBV23, GGS^{+}21, HSBC20, HCS^{+}21, JOC^{+}20, JLW^{+}23, JHvL^{+}23, KRB^{+}21, KHK^{+}22, LLC^{+}23, LBG^{+}22, LZM^{+}21, LSCL21, MAP^{+}22, MSJ21, MRT^{+}24, OBL21, PBC21, PHM21, RLL^{+}21, RBL^{+}20, RBSB22, SDL22, TBH^{+}22, VST^{+}22, WHAR^{+}20, XMP^{+}21, YvdHC^{+}20]. Nutrients [MGGRR^{+}21, AFK^{+}22, FAB^{+}20, KFHS24, LNC^{+}23]. Nutritional [JLR^{+}22, VKK^{+}21]. NW [GBB^{+}21].

O [AFH^{+}22, DRA^{+}23, EBW^{+}20, MGG^{+}24, TZF^{+}22, WMRL22]. obliquus [ZSZ^{+}21]. Observation [CLA^{+}24]. Observations [KTH^{+}21, ZYZ^{+}21, APB^{+}23, CSM^{+}23, FIH^{+}20, MBD^{+}23, WDG^{+}21]. observatory [MKBSK19]. Observed [KdED^{+}20, HDD^{+}22, NH23, SLZ^{+}21b]. occupancy [CGB^{+}21]. occupation [CYPRG^{+}23]. Occurrence [XYZ^{+}21, MZW^{+}21, MHSL22, ZZT^{+}21]. occurring [MWHR^{+}22, VGvdB^{+}21]. occurs [MBH^{+}23]. Ocean [ARPFGS23, RLP^{+}20, SSEO21, YAP^{+}22, ARB^{+}22, BFW^{+}22, CSD^{+}24, CSM^{+}23, DKPS23, GM22, GLAHH^{+}23, HDS^{+}20, JHS^{+}21, KLO23, KTH^{+}21, LD22, LXC^{+}21a, LBL^{+}24, MO23, MLP21b, MBD^{+}23, MCS^{+}23, MFGF^{+}22, NSK^{+}21, OSR20, ORRMD^{+}21, PKV20, RBD22, SBD^{+}21, SSH^{+}22, SB20, SHZ^{+}21, TNK^{+}22, TKH21, WWB^{+}20, YXC^{+}22, ZLU^{+}20, dCLO^{+}23, ADLW22, BSGA21, BDL^{+}21, BRCO21, BDE^{+}22, BLRM24, BBA^{+}23, CGL^{+}21, CKH^{+}21, CLD^{+}23, CBC^{+}22, DFW^{+}20, EBK^{+}20, ETMD23, HCB^{+}22, IKN^{+}21, IJSC20, JSL^{+}23, LGI^{+}20, LWvdM^{+}21, LL23, MT23, MSB^{+}22, MPW^{+}23, MLT^{+}21, MGC^{+}20, PDK^{+}23, PGN^{+}21, RKBT^{+}23, SBM^{+}21, SORGC^{+}21, SWC^{+}23, SHE^{+}21, SJJ^{+}21, SBW^{+}20, SSM21b, SAD^{+}20, TMS^{+}21, WRS^{+}22, XKP^{+}22, XYZ^{+}21, YSHS23, YAE^{+}24, ZZO^{+}21, ZXE^{+}23]. ocean- [DKPS23]. Oceanic [CCG^{+}22, FIM^{+}22, OMM21, CCK^{+}21, CDH^{+}21, CBD^{+}23, DLL20, EHC^{+}22, KTH^{+}21, KP20, MCH^{+}21, WKRL^{+}23, ZZHZ23]. oceanica [AAC^{+}21, CB21b]. Oceanogr [Ano21b]. Oceanographic [MT23, CZS^{+}22, FJP^{+}21, GSD^{+}23, SSA^{+}23, TKA21]. Oceanography [Ano21a, Kir20]. oceans [SSB^{+}22, WSLC22, ZCZ^{+}23, BCA^{+}20, GRJ^{+}22]. octocorals [PMF^{+}21]. off [BK20, SVBT20, YGD^{+}21]. offs [BH21]. offsets [BS20, JOC^{+}20]. offshore [CTP21, KLO23, KPS21]. Ohrid [VMT^{+}20]. Oikopleura [MLS^{+}21]. Okhotsk [ZZO^{+}21]. old [MHVC22]. oligomesotrophic [KHSP21]. oligotrophic [CHP^{+}24, CM20a, DdEM^{+}23, EPR^{+}24, HSB^{+}20, LLL^{+}23, LXC^{+}21a, MHD^{+}22, MOW^{+}22, RCS^{+}20, SGGB21, SBL^{+}21, SDB^{+}23, SRL^{+}20, SGM^{+}21, WSLC22]. oligotrophication [MHSL22, SDS^{+}23]. omics [ZSZ^{+}21]. Omnivorous [CSNS24, KGG^{+}23, NDP^{+}24]. omnivory [BOC^{+}21]. Oncorhynchus [ASH^{+}20]. one [ZZCD23]. onset [ZZL^{+}21]. Ontogenic [PWS^{+}20]. open [HJWP24, PKV20, SPKM22]. open-coast [SPKM22]. Opening [SO23]. operationally [LHMR23]. opportunities [BR22, MBD^{+}23]. opportunity [OJUN23]. Optical [ADLW22, BHG^{+}24, EJRRC^{+}23, LSW^{+}20a]. optically [BGG^{+}20]. optimization [TVB^{+}21]. Optimizing [FGP22]. Oregon [EFG^{+}24]. Organic [BAG^{+}24, HLS^{+}22, KHCG^{+}22, PUH^{+}23, PGN^{+}21, VWR^{+}23, AVJ^{+}24, ABS^{+}21, AKT^{+}21, AMC21, BMR^{+}21a, BCWS23, BMCS23, BSF^{+}21, BPIBN20, BBH^{+}20, BAD^{+}21, BJK^{+}21, BGG^{+}20, BYS^{+}23, BDL^{+}21, BJG^{+}22, BZK^{+}22, BGM^{+}21, CPB^{+}21, CTP21, CGC^{+}21, CLC^{+}22, CM21, CCDM21, DVL^{+}23, DdEM^{+}23, DLC^{+}22, EJRRC^{+}23, FAB^{+}20, FUUG^{+}21, FXD^{+}22, GDB20, GMI^{+}21, GAS^{+}23, HCB^{+}22, HZM^{+}21, HOB^{+}22, HHK^{+}22, HCS^{+}21, HP20, JSB^{+}20, KAP^{+}20, KDZ^{+}24, KBB^{+}22, LBBM21, LRdGFP23, LMM^{+}21, LNR^{+}20, LDGB^{+}23, LDGB^{+}24, LD22, LSC^{+}21, LTH21, LZD^{+}23, LXC^{+}21b, LPO^{+}20, LLTT^{+}24, MGL22, MEQBV20, MRH^{+}23, MRD^{+}21, MNMJ^{+}21, PBC21, PSNG^{+}21, PWS^{+}20, PC21, PNT^{+}21, RvHF^{+}20, RLP^{+}20, RCMA21, RBR^{+}23, SVMM^{+}22, SSS^{+}24, SB20, SB22, SZA^{+}21, SOT^{+}20, SEK20, SSWM20, SCB^{+}20, SGS22, SGSD^{+}22, TSX^{+}23, VST^{+}22, WCC20, WOK^{+}21, WXK^{+}22, YGD^{+}21, ZCZ^{+}23, ZJW^{+}22, ZLW^{+}21a]. organic [ZZT^{+}21, ZYZ^{+}21, ZZB^{+}20, dMKB^{+}20]. organisms [FGO^{+}20]. organization [KPNJ23]. origin [MRD^{+}21]. original [LHH24]. origins [LNR^{+}20]. Orkney [MSGW^{+}20]. orthophosphate [RHS^{+}22]. oscillating [JYA21]. Oscillation [LDB^{+}21]. oscillations [SGS22]. oscillatory [MDBI20]. other [GFW^{+}20, LES^{+}21, WWFS21]. Otolith [DAB^{+}21]. our [MVBG20]. outcompete [RHS^{+}22]. Outwelling [KDZ^{+}24]. Overestimation [GE22]. overgrowth [RMH^{+}22]. overlying [XLN21]. overrides [RRHS^{+}21, YG23]. overview [SPHH22]. oxic [CYZZ^{+}20, GKW^{+}20, GKW^{+}21, HGMK22, TMM^{+}20, XLN21, YWZZ20]. oxidation [ACA^{+}22, FSS^{+}23, GPK^{+}21, LVDM22, LPO^{+}20, LPA^{+}23, MSB^{+}22, MWHR^{+}22, RRHS^{+}21, RSH^{+}20, SHE^{+}21, SSM^{+}22b, SZY^{+}20, THL^{+}21]. oxidative [MRT^{+}24]. oxide [AHCF22, FSS^{+}23, HBR^{+}22, LMC^{+}21, LPMBR23, RTM^{+}20, RTM^{+}21, TTC^{+}22, WLK^{+}21, ZZO^{+}21]. oxidizing [CLD^{+}23]. Oxygen [BKMY21, CBB^{+}22, HJWP24, SMCS23, AHCJ^{+}20, BYB^{+}23, CCK^{+}21, CGGC^{+}21, CKCB^{+}23, CBD^{+}23, DPB^{+}23, DNA^{+}22, EBK^{+}20, FIH^{+}20, GB23, KDV^{+}23, Lem20, LHGJ21, LSW^{+}20b, MHG^{+}23, MP23, MDIY23, PCJ^{+}21, QRWT^{+}22, RWMP^{+}23, SJH20, SMW^{+}21, SSM^{+}22a, SSWM20, SEG21, SCB^{+}20, SCT^{+}21, THL^{+}21, WLK^{+}21, WRA^{+}23, YWLY20]. oxygen-consuming [SCB^{+}20]. Oxygen-controlled [BKMY21]. Oxygen-deficient [CBB^{+}22, AHCJ^{+}20, BYB^{+}23, CKCB^{+}23, DNA^{+}22]. oxygenated [LDS^{+}20]. Oxygenation [BPL22, HBR^{+}22]. oxygenized [HPNU20]. oyster [CCC^{+}21, HLS^{+}22, WSB^{+}20]. oysters [HSP^{+}20, MDWT^{+}22].

P [LPMBR23, MLL^{+}21]. Pacific [ADLW22, CKH^{+}21, ETMD23, HYPG^{+}22, PDK^{+}23, RPKC22, SSM21b, SAD^{+}20, WRS^{+}22, ZBNH^{+}21, AHCJ^{+}20, BSGA21, BYB^{+}23, BCA^{+}20, CJB^{+}21, CBC^{+}22, DWS^{+}22, DNA^{+}22, DFW^{+}20, EBK^{+}20, FDF^{+}21, FSS^{+}23, FIH^{+}20, GFW^{+}20, HDK^{+}21, HOB^{+}22, KMV^{+}20, LLK^{+}22, LWC^{+}24, LSCL21, LNC^{+}23, MZXY22, MY23, MMG^{+}21, SMW^{+}21, SWC^{+}23, SOT^{+}20, TKA21, WOK^{+}21, XYZ^{+}21, YAH^{+}21, YWZZ20]. Page [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano22a]. Paired [BBGM20]. paleoceanographic [RWL^{+}24]. paleoclimate [THT^{+}20]. paleolimnological [BLS^{+}22]. Palmer [NCC^{+}21]. Panulirus [BFCC^{+}21]. paradigm [GM22]. Paradox [GVF^{+}23, BVL^{+}23]. parameters [HZC^{+}24, KTWGT20]. parametrization [PC21]. parasite [DMCW23, LEB21, ONH^{+}20]. parasites [CML23, DMCW23, FWT^{+}20]. parasitic [KSN^{+}21]. parasitism [BDMGH^{+}23]. Parent [CSH21]. Parental [MDWT^{+}22]. particle [AVS^{+}22, BSGA21, CZS^{+}22, CBN21, GEHD20, JSW^{+}21, KSR20, KBvdD^{+}20, LBL^{+}24, RPA^{+}22, WPY^{+}20, WDS^{+}23]. particle-associated [BSGA21, WPY^{+}20]. particles [CKH^{+}21, CBN21, DBB^{+}21, DLL20, JLS^{+}20, LG21, MGG^{+}24, SWC^{+}23, SAD^{+}20, WF22]. particular [GMI^{+}21]. Particulate [JSL^{+}23, CLD^{+}23, FUUG^{+}21, HCB^{+}22, HCS^{+}21, RMW^{+}21, SOSN23, SOT^{+}20, ZCZ^{+}23, ZBNH^{+}21]. Partitioning [BHA^{+}22, HLKD23, HCH^{+}24, MGG^{+}24, SSS^{+}24, SCB^{+}20, TKA21]. passing [KX22]. past [MFM^{+}21, SSL22, SLHL^{+}22]. pathway [KP20, SWZ^{+}23]. pathways [BSIdG23, BMGK22, MMG^{+}21, ZSZ^{+}21, vEMF^{+}20]. pattern [ZZO^{+}21]. Patterns [BMB22, KRC23, WPY^{+}20, BJS23a, BMR^{+}21b, BMLV24, BFE^{+}20, BRB^{+}22, CGB^{+}21, GCW^{+}23, GBC^{+}20, HEH^{+}20, HCS^{+}21, HABL^{+}21, JREVB21, LABM20, LSS^{+}24, MHdS^{+}21, MZW^{+}21, MGGRR^{+}21, NGSS23, PATL^{+}23, PLB^{+}21, QMG21, Roy20, RBSB22, SVBT20, SPBK21, VMP^{+}21, WWGP21, vLWV^{+}20]. pCO [CGR22, CCB^{+}20, EDC24, EWGP22, GPB^{+}20, MDWT^{+}22, NB20, QBJG21, RGW^{+}23, TMS^{+}21, VKB^{+}22, WRHR22, WDG^{+}21]. Peak [TMSL22]. peaks [SSH^{+}22]. Pearl [YGD^{+}21]. peat [BCWS23, KRM^{+}20]. peatland [DdEM^{+}23, FSM^{+}21, KSH^{+}23, TKW^{+}23]. peatland-rich [TKW^{+}23]. peatlands [FSM^{+}21]. Pelagic [HBS^{+}22, RSH^{+}20, ST22, WRHR22, BZSF23, CJB^{+}21, CHA^{+}24, FWT^{+}20, HCH^{+}24, HGF^{+}21, KPL^{+}20, KLG^{+}23, MBLA^{+}21, MLT^{+}21, NvBE^{+}21, PVSP^{+}22, VCT^{+}23, VKK^{+}21, VGSB23, VKGW22]. pellet [SSM21b]. pellets [DMS^{+}21, MLS^{+}21]. penguin [CFC^{+}20, NCC^{+}21]. Peninsula [BBL^{+}21, CFC^{+}20, CMT^{+}20, JWAvD^{+}22, MHP^{+}23, NGSS23, SGSF20]. peptidomic [DNA^{+}22]. perennially [LDS^{+}20, SSM^{+}21a]. performance [BDMGH^{+}23, BFW^{+}22, DWR^{+}23, MSB^{+}20]. peri [SDS^{+}23]. peri-alpine [SDS^{+}23]. period [ACA^{+}22, XTL^{+}20]. Periodic [NH23]. periodically [RWS^{+}22]. periodicities [LDB^{+}21]. Periphyton [FSM^{+}21, VST^{+}22, BHSL21, WK21]. permafrost [BMLV24, FBP20, LLC21, RRHS^{+}21, TKW^{+}23]. permanently [CKCB^{+}23]. permeable [NKH^{+}23]. peroxide [SGK^{+}21]. Persistence [SBW21]. perspective [BLS^{+}22, FIM^{+}22]. perturbations [TMS^{+}21]. Peruvian [THL^{+}21]. petrogenic [BJK^{+}21]. pH [CTM^{+}21, EBW^{+}20, GAS^{+}21, HTLP23, JvKvT^{+}20, JHS^{+}21, JLR^{+}22, KDV^{+}23, MGP^{+}21, MK21, MKD22, MNMJ^{+}21, PCC^{+}23, SCT^{+}21, WDG^{+}21, ZLU^{+}20]. Phaeodaria [MNM^{+}21]. Phaeophyceae [SBB^{+}21]. phagocytosis [LOW^{+}20]. phagotrophy [BH21]. phase [HLKD23, MLC^{+}21]. phases [CB21a]. phenology [CCA^{+}21, CHA^{+}24, DNB^{+}20, LAD^{+}22, MLP21b, vLWV^{+}20]. phosphate [MUGL20, SBW^{+}20]. phosphate-depleted [SBW^{+}20]. Phospholipid [PKV20]. phospholipids [GVF^{+}23]. Phosphonate [SBW^{+}20]. Phosphorus [DGW^{+}21, AWGV20, AJMG21, AVS^{+}22, BPIBN20, BRM^{+}23, DCK^{+}22, FPB^{+}23, GVF^{+}23, HSBC20, HPB^{+}21, KPM^{+}20, KTA^{+}21, MDP^{+}23, QFM^{+}21, Roy20, RBR^{+}23, SGA^{+}22, SdBW^{+}23, SML21b, YJ20, YAE^{+}24, ZFQ^{+}20, ZXZ^{+}23]. photic [LLL^{+}23, QBJG21, WCC20]. photo [RRHS^{+}21]. photo-oxidation [RRHS^{+}21]. Photoacclimation [MSD23, PFH21]. photoautotrophic [CKCB^{+}23, KFHS24]. photochemical [MGL22]. Photodegradation [BKC20]. photogrammetry [ASL^{+}24]. photomineralization [AKT^{+}21]. photon [MSD23]. photophysiology [ADLW22]. photoprotection [ARPFGS23]. photoreactive [KLP23]. photostable [KLP23]. Photosynthesis [GKW^{+}20, GKW^{+}21, GF21, RGW^{+}23]. Photosynthesis-driven [GKW^{+}20, GKW^{+}21]. Photosynthetic [SGSF20, BRHS20, EHC^{+}22, FIH^{+}20]. photosynthetically [PC21]. phototrophic [LABM20]. phycobiliproteins [WWFS21]. phycosphere [JHSG^{+}24]. phyllosomata [BFCC^{+}21]. Physical [MBB^{+}20, BGM^{+}22, HMB^{+}20, KTH^{+}21, KBL^{+}20, LZG21, QRWT^{+}22, SORGC^{+}21]. physico [BC20]. physicochemical [FWvA^{+}21]. physicochemistry [BCTH20]. Physiological [BHG^{+}24, CT23, HDS^{+}20, Wei24, BTT23, HDK^{+}21, KTWGT20, WRHR22, dJGCS23]. physiology [FPG^{+}24, MGC^{+}20]. phytodetritus [LMS23]. Phytoplankton [HBG^{+}22, PBC21, YAE^{+}23, AVJ^{+}24, ARPFGS23, AFK^{+}22, BMG^{+}23, BHG^{+}24, BTT23, BMA^{+}20, BHSL21, BDE^{+}22, BK20, BBL^{+}21, BAHH^{+}20, CAV^{+}22, CFC^{+}20, CSM^{+}23, CBD^{+}23, CLP^{+}24, FBR^{+}23, FBRG23, FDHS20, FWT^{+}20, FBV23, FPB^{+}23, GVF^{+}23, GCW^{+}23, GMS21, GK22, GZX^{+}22, GAS^{+}23, HLH^{+}23, HANW21, HDK^{+}21, HDD^{+}22, HNS21, HZS21, HKM^{+}20, HPP^{+}21, HS21, IJSC20, IYW^{+}22, JWAvD^{+}22, KRB^{+}21, KBL^{+}20, LHS^{+}20, LTW^{+}24, LNC^{+}23, LBL^{+}24, LVZ21, LLBC20, LSDA21, MXL^{+}21, MUGL20, MHD^{+}22, MAG^{+}22, MSJ21, NGSS23, PKKS20, PCC^{+}23, QS21, QBJG21, RBR^{+}24, SVC^{+}22, SSH^{+}22, SJJ^{+}21, SGM^{+}22, SSM^{+}22a, SMD21, SGS22, SSB^{+}22, TJW^{+}22, TATC^{+}24, VKB^{+}22, VWR^{+}23, WCZ^{+}20, WCC20, WRA^{+}23, YML^{+}21, ZZHZ23, ZLW21b, vLWV^{+}20, vdPAV^{+}20, vdPMF^{+}21]. phytoplankton-dominated [WCZ^{+}20]. phytoplankton-related [TJW^{+}22]. Phytoplanktonic [GRJ^{+}22]. Pico [KPS21]. Pico- [KPS21]. picocyanobacteria [AHCJ^{+}20, CBT^{+}21]. picoeukaryote [RPKC22]. picoeukaryotes [EHC^{+}22]. picophytoplankton [FM21, GSP^{+}23]. picoplankton [BDL^{+}21, DWW^{+}21]. pigment [ARPFGS23, CSM^{+}23]. pigments [CTF^{+}21]. Pine [LCHB22]. pinnacle [GCMJ23]. Placopecten [CGR22]. plain [LTH21]. Plains [ETG21]. Planktivorous [LDB^{+}20]. Plankton [FUUG^{+}21, SBV^{+}21, EG20, FGO^{+}20, JSW^{+}21, KKA22, OAM^{+}22, PSNS22, SGL^{+}23, SCM22, THS21, YWLY20]. Planktonic [IMB^{+}22, ALH^{+}21, CML23, EPR^{+}24, FIM^{+}22, Gil22, GSD^{+}23, LLMM^{+}23, LEB21, MMF^{+}21, MDP^{+}23, Wei24]. Planktothrix [SDS^{+}23]. planning [YAP^{+}22]. Plant [CZH^{+}21, VJS^{+}20, KHK^{+}22, LCS^{+}20, WHZ^{+}21]. Plant-mediated [VJS^{+}20]. plants [ZN20, ZLN21]. plastic [VGvdB^{+}21].

Plasticity [LES^{+}21, BCFI^{+}22, BOC^{+}21, NDP^{+}24, PMF^{+}21, RFB^{+}22, WWGP21]. plate [JYA21]. Plateau [DLC^{+}22, DAB^{+}21, XTJW21]. Plum [LVHK20]. plume [CGC^{+}21, SLZ^{+}21b]. plumes [dCLO^{+}23]. point [CBN21]. points [CBPNA22]. polar [KMM^{+}22, KGG^{+}23, LMSN23, MLG^{+}20, STB22, vdPAV^{+}20]. polarized [CZS^{+}22]. poleward [CJB^{+}21]. pollen [LCHB22]. Pollicipes [FJP^{+}21]. polygonal [PLB^{+}21]. polymictic [TKS^{+}20]. Pond [VJH^{+}22, DGTW21, QMG21]. ponds [FBP20, FW21, GCEM^{+}23, LMM^{+}21, PWS^{+}20, PLB^{+}21, RHA^{+}23]. pool [CLC^{+}22, KPNJ23]. pools [BHSG^{+}20, LLTT^{+}24]. poppei [MSGW^{+}20]. populated [HPB^{+}21]. Population [CÖL^{+}20, CHA^{+}24, DAB^{+}21, FRRG^{+}21, HPF^{+}23, KOR^{+}20, LNC^{+}23, MVP^{+}20, MSGW^{+}20, WRMP22, ZLG^{+}23]. populations [ASH^{+}20, DMCW23, MUGL20, RBL^{+}20, ZJ22a]. Pore [TEK^{+}21, CSC^{+}21]. pore-water [CSC^{+}21]. porewater [KDZ^{+}24, RST^{+}21]. porewaters [DKPS23]. Posidonia [AAC^{+}21, CB21b]. position [GLN^{+}21, TRS^{+}20, VGSB23]. possess [CCG^{+}22]. Potential [DCJB22, DWW^{+}21, HDK^{+}21, HGMK22, XLN21, BSB^{+}20, CYPRG^{+}23, DAB^{+}21, MMLPRHCG22, PWS^{+}20, RvHF^{+}20, SSEO21, TGE^{+}21, VVMJ23, WM20, XWL^{+}21, ZPK^{+}22, ZLL^{+}24]. potentials [RLL^{+}21]. Powell [DSY20]. powered [MXL^{+}21]. Prairie [RVS^{+}22]. precariousness [SHBA22]. precipitation [AFH^{+}22, DSY20, HM20, KAP^{+}20, LN21, SSCE20]. predation [BBM21, LAFV23, LWS^{+}22, LDB^{+}20, OFM^{+}21, OLR^{+}23, PSNS22, ZLG^{+}23]. Predator [BMGK22, DCJB22, DFW^{+}20, LEB21, LMSN23, MVP^{+}20, MWV22, MPW^{+}23]. predators [HCH^{+}24, RS20]. predatory [DWR^{+}23]. predict [FMT^{+}20, GMI^{+}21]. predictability [KKA22]. Predicting [BCTH20, BMR^{+}21b, LM24, RPA^{+}22, AVJ^{+}24, TRS^{+}20]. prediction [MSJ21]. predictors [MTP^{+}23]. Preferential [CM21]. Presence [LBJS22, FPG^{+}24, LM24, MK20]. present [EWGP22]. present-day [EWGP22]. Preservation [ZJW^{+}22]. pressure [BBM21, DvOH^{+}20, MSD23, SEG21, WBB^{+}21]. Prevalence [KGG^{+}23, DMCW23, RBL^{+}20]. previous [CTM^{+}21]. previously [KKP^{+}21]. prey [DFW^{+}20, HMBB21, LEB21, LMSN23, RTK23, SYMF21, WO21]. primacy [HS21]. primary [FNB^{+}23, FSH^{+}23, HEH^{+}20, KTA^{+}21, KBSA21, KPG^{+}23, KVKS22, LD22, LSDA21, OLFH21, RHP^{+}23, SB20, SSM^{+}22a]. Prince [SBV^{+}21]. priority [GGLH^{+}22]. pristine [BPG^{+}22, DBM^{+}23, EJM^{+}23, YvdHC^{+}20]. process [ZZL^{+}21]. processes [Ano21a, BD22, CM20b, GSG^{+}17, HP20, MZXY22, NH22, PATL^{+}23, QRWT^{+}22, SSP^{+}20, TJW^{+}22, WPY^{+}20]. processing [AVS^{+}22, BND^{+}23, DBB^{+}21, HHK^{+}22, LMM^{+}21, ORE20]. Prochlorococcus [LGI^{+}20, LES^{+}21, LSCL21, RRAO^{+}21, SHZ^{+}21, TKA21]. produce [WPO^{+}22]. producers [HEH^{+}20, KVKS22]. producing [THT^{+}20]. Production [CKH^{+}21, LMS23, BMA^{+}20, BBH^{+}20, BGG^{+}20, BD22, BH21, BV21, BBCA^{+}21, CWF^{+}22, CCFF21, CDA^{+}20, DBFC24, FSM^{+}21, FNB^{+}23, FSS^{+}23, FSH^{+}23, FIH^{+}20, GE22, GMI^{+}21, GKW^{+}20, GKW^{+}21, HGMK22, HNS21, HBR^{+}22, KTA^{+}21, KBSA21, KSI24, KPG^{+}23, KBvdD^{+}20, KTWGT20, LLL^{+}23, LEK^{+}22, LMP^{+}20, LDS^{+}20, LTLH22, LDB^{+}21, LSDA21, MCH^{+}22, MLT^{+}21, OLFH21, PDK^{+}23, RHP^{+}23, SPKM22, SJJ^{+}21, SPLJA^{+}20, SSM^{+}22a, SSM21b, SMD21, TTC^{+}22, TZF^{+}22, WOT^{+}23, YWZZ20]. productive [FBRG23, SGGB21]. productivity [ASH^{+}20, CLG^{+}24, CGL^{+}21, GLAHH^{+}23, MCH^{+}21, MGGRR^{+}21, SBV^{+}21, WLSD21]. products [LSW^{+}20a]. profiles [DWS^{+}22, SJH20]. profiling [FIH^{+}20]. proglacial [MKP^{+}23]. progressive [SVBT20]. projections [CCA^{+}21]. prokaryotes [BMG^{+}23, CKCB^{+}23, DWW^{+}21, GLAHH^{+}23, LBBM21]. prokaryotic [BAD^{+}21, CDA^{+}20, DVL^{+}23, GE22, SML^{+}21a, SORGC^{+}21, SSS^{+}24]. Prolific [ETMD23]. prolonged [LLC21, ZZHZ23]. prominent [YG23]. promote [CBB^{+}22, DFJ22, EFC^{+}23, KRB^{+}21]. promoted [ZJW^{+}22]. promotes [LTJ^{+}20, vGDH^{+}21]. promoting [GGS^{+}21]. prone [FLH^{+}23]. pronounced [MKD22]. properties [BMG^{+}23, BHG^{+}24, CBD^{+}23, LD22, MRD^{+}21, PSK21, SJR21, SB20, XSZ^{+}21]. Prorocentrum [LZG21]. protect [RD23]. Protein [DNA^{+}22, MK20, PYY^{+}20]. protein-coupled [MK20]. protist [MNM^{+}21, PYM^{+}20]. protistan [SDSGR21]. protists [BO20, SWC^{+}23]. provide [BVM^{+}20, SSA^{+}23]. provinces [BGAD^{+}22, SSH^{+}22]. proxies [FSH^{+}23, RWL^{+}24]. proximity [MKP^{+}23]. proxy [LPO^{+}20]. pseudofecal [WSB^{+}20]. pseudogonyaulax [BH21]. pseudoharengus [DFJ22]. Psychrophilic [SJR21]. pueruli [BFCC^{+}21]. pulex [YWZ^{+}21]. pulsatile [WSB^{+}20]. pulse [KHHZ20]. Pulsed [WBB^{+}21]. Pulses [NKH^{+}23, BRCO21]. pump [CSC^{+}21, DMS^{+}21, DLL20, FWDY20, GBB^{+}21, LG21, PFH21, PVSP^{+}22, PJRV23]. pumps [GM22]. purely [LD22, SB22]. putative [CLD^{+}23]. pyrifera [SBB^{+}21].

quahog [MURK20]. qualitative [WMTJW22]. quality [ABS^{+}21, DLC^{+}22, DKPS23, FRRG^{+}21, HSB^{+}20, HZS21, LZM^{+}21, MFM^{+}21, MHL^{+}20, PHB^{+}22, PWS^{+}20, RvHF^{+}20, SSS^{+}24, SSK24, TSX^{+}23, TPTS23, TCS21, VKK^{+}21]. Quantification [CWF^{+}22, ZR23]. Quantifying [CCFF21, LVZ21]. Quantitative [LN21]. quantity [ABS^{+}21, BBH^{+}20, DLC^{+}22]. quantum [RKBT^{+}23]. Quasi [LDB^{+}21]. Quasi-decadal [LDB^{+}21]. Queen [DLDF21]. quenching [CGB^{+}21].

r} [CLGH23]. Ra [LvBS^{+}22]. radiata [BRHS20, BS20]. radiation [PC21, YG23, ZFQ^{+}20, vdPMF^{+}21]. radiatively [AHF^{+}22]. radiocarbon [HGA24, IFU^{+}20]. radiocesium [TNK^{+}22]. Radiolaria [MNM^{+}21]. radiolarians [IKN^{+}21]. radium [RST^{+}21]. radon [ACW^{+}22, RST^{+}21]. rain [DBFC24]. rainforest [FBJ^{+}23, SOT^{+}20]. range [CJB^{+}21, CEE^{+}22, PYM^{+}20]. ranges [LCS^{+}20]. Rapid [CBC^{+}22, PFH21, SLZ^{+}21b, THL^{+}21, BVGQN22, YWZ^{+}21, ZLW^{+}21a]. Rapidly [ZZCD23, DKBG21, LGQ^{+}21]. rate [CGR22, HZM^{+}21, LSL^{+}20, LCZ^{+}21, QBJG21]. rates [BAG^{+}24, BWSW24, BVM^{+}20, CWF^{+}22, DGTW21, Edm21b, GR24, HMP^{+}21, LLMM^{+}23, LDGB^{+}24, MSG^{+}22, OLFH21, PKV20, RLP^{+}20, RRAO^{+}21, SBL^{+}21, SSWM20, SMD21, TION23, TBH^{+}22, ZQB^{+}21, ZWM22]. ratio [FBK^{+}22]. ratios [BV21, DWS^{+}22, HMBB21, KMV^{+}20, MDP^{+}23, MRT^{+}24]. ray [IKN^{+}21]. re [SDS^{+}23, WZCK20]. re-examination [WZCK20]. re-oligotrophication [SDS^{+}23]. reaches [PGB24]. reactivity [GMI^{+}21, MGL22]. reactor [CBF^{+}23]. reallocating [MRC^{+}24]. really [FPBG^{+}22]. realm [FUUG^{+}21, WKRL^{+}23]. receptor [MK20]. receptor-based [MK20]. Reciprocal [FW21]. recirculating [BKMY21]. Reconciling [MHG^{+}23]. reconfiguration [ZN20]. Reconstructing [KBSA21]. reconstructions [THT^{+}20]. record [WRS^{+}22, WOMWR20]. recovering [RFB^{+}22]. Recovery [RWS^{+}22, vHLV^{+}20, BPG^{+}22, DFJ22, GRJ20, GPB^{+}20, KDW^{+}23, MAP^{+}22, RVS^{+}22, SHBA22, WRWPG19, WPO^{+}22, ZvBZ^{+}20, Ano21c]. Recruitment [MKBSK19, Edm21a, FJP^{+}21, OSR20]. recurrence [FR23]. Recurrent [SH22]. Recycling [HYPG^{+}22, SKA^{+}24]. red [CJB^{+}21, EFP^{+}21]. redox [CKCB^{+}23, EBK^{+}20, EFG^{+}24, SGSD^{+}22, SGK^{+}21]. redox-stratified [CKCB^{+}23]. redoxcline [HPNU20]. reduce [XWL^{+}21]. Reduced [KFHS24, FNB^{+}23, JOC^{+}20, KDZ^{+}24, ZXZ^{+}23, DMCW23]. reduces [ALH^{+}21, BRM^{+}23, DSY20, OFM^{+}21, QBJG21, SSM^{+}22a]. reducing [SJR21]. reduction [CEW^{+}21, DFJ22, KP20, RWS^{+}22, SSM^{+}21a, THL^{+}21, dJGCS23]. reduction-dependent [KP20]. reductions [FBV23, TBH^{+}22]. reef [BCFI^{+}22, BBCA^{+}21, CDW^{+}24, CBC^{+}22, Edm21b, EDC24, FKN21, GCMJ23, GTKW^{+}20, GLB^{+}22, HPCH24, HLS^{+}22, KDW^{+}23, LMS23, LBG^{+}22, MCH^{+}22, MSG^{+}22, MHMML^{+}22, MDIY23, SWZ^{+}23, SSCE20]. reefs [FPBG^{+}22, GCMJ23, OLFH21, RMH^{+}22, SPKM22, VVMJ23, dOKC^{+}21]. reestablishes [AMO20]. Reference [TMKB^{+}22]. Refining [FSH^{+}23]. reflect [HHMH22, SSK24]. reflectance [LSW^{+}20a]. reflects [CVF^{+}20]. refractory [LLTT^{+}24]. refuges [YAH^{+}21]. regardless [CTM^{+}21]. Regeneration [HKM^{+}20]. regime [BTK^{+}21, DMC^{+}22, MWV22, RNW^{+}23, SPHH22, TMKB^{+}22]. regimes [BHA^{+}22, CGGC^{+}21, CGL^{+}21, LGI^{+}20, MCH^{+}21, SSL22, VCBB20, WCZ^{+}20]. Region [LDGB^{+}24, CLP^{+}24, FUUG^{+}21, KLO23, MFM^{+}21]. Region-specific [LDGB^{+}24]. Regional [LRRP20, GRJ20, HDK^{+}21, KSF^{+}23, RTM^{+}20, SMKS22]. regional-scale [SMKS22]. regionally [SJJ^{+}21]. regions [ASH^{+}20, NCC^{+}21, SML^{+}21a]. regularity [SGM^{+}22]. regulates [BGM^{+}21, JHSG^{+}24, MUGL20, THVJ23, YPF^{+}23]. regulating [SOSN23]. regulation [BAHH^{+}20, DK20, GLN^{+}21, HP20]. regulatory [TMKB^{+}22]. related [BSGA21, CLA^{+}24, LDB^{+}21, PYY^{+}20, TJW^{+}22]. relation [DNB^{+}20, WRC^{+}21]. relationship [BCA^{+}20, FPBG^{+}22, KMV^{+}20]. Relationships [QFM^{+}21, BMA^{+}20, BM22, GGHBS^{+}20, KPL^{+}20, LEB21, MNM^{+}21, QS21, SDL22, SZA^{+}21, TRS^{+}20, YJ20, ZZWL22]. Relative [LLBC20, SSH^{+}22, GCW^{+}23, QMG21]. relatively [GVF^{+}23]. relaxation [REF^{+}21]. release [AJMG21, BMCS23, CLGH23, HP20, KPM^{+}20, KRM^{+}20, LHS^{+}21, RS20, XWL^{+}21]. Relevance [SPBK21, CPB^{+}21]. reliable [PPB^{+}21]. relieves [DCK^{+}22]. Remarkably [EFP^{+}21]. remote [LSW^{+}20a, VVMJ23]. Removal [MLS^{+}21, SGL^{+}23]. reoxidation [ETMD23]. reoxygenation [RWS^{+}22]. repeated [CBT^{+}21]. replacement [KMV^{+}20]. replete [CM21]. Reply [CSH21, MBD^{+}23, SB22]. represent [HSD^{+}20]. Reproducing [TVB^{+}21]. reproduction [MMKWZ22, MGP^{+}21]. reproductive [BRB^{+}22]. require [WRA^{+}23]. requires [KHCG^{+}22]. Research [MKBSK19]. reservoir [AWGV20, BBH^{+}20, CB21a, ILPL20, LHD^{+}20, PBGGRB21]. reservoirs [BJG^{+}22, CCY^{+}21, CM20a, LPMBR23]. residence [EAF^{+}22, LvBS^{+}22, RCMA21, SDL22]. resident [LBJS22]. Resilience [CAV^{+}22, BLS^{+}22, BBMD20, DP22, GGS^{+}21, MHD^{+}22, RMH^{+}22, RBD22, RPKC22, SPHH22, TKM^{+}22, ZvBZ^{+}20, dSNIB22]. resilient [WCH^{+}21]. resist [MRC^{+}24]. resistance [TKM^{+}22, YvdHC^{+}20, vHHP^{+}21]. resolution [BCWS23, PHB^{+}22, ZZO^{+}21]. resolve [GSD^{+}23]. resolved [GNO22]. Resolving [BPIBN20, SMW^{+}21]. resonance [LHH24]. Resource [LHS^{+}20, FBJ^{+}23, FDHS20, GZX^{+}22, MHL^{+}20, VCT^{+}23]. resources [MRC^{+}24]. respect [LBvE23]. Respiration [BHA^{+}22, SEG21, BWSW24, BV21, BVM^{+}20, CGR22, FUUG^{+}21, HCS^{+}21, MHG^{+}23, MSG^{+}22, NKH^{+}23, PBC21, RHP^{+}23, SMCS23, SDB^{+}23, SSM^{+}22a, WMRL22]. respired [HGA24]. respond [BWSW24, CTP21]. responding [MSD23]. Response [AAC^{+}21, CZK^{+}23, FBR^{+}23, LWC^{+}24, MLG^{+}20, PRS^{+}20, SLZ^{+}21a, TSX^{+}23, YJB^{+}20, ABC^{+}21, BTT23, BJK^{+}21, BRHS20, BK20, BPG^{+}22, CT23, CMB^{+}21, CBT^{+}21, DCF^{+}22, DKBG21, GRJ^{+}22, GK22, HXL^{+}21, HLS^{+}22, ILPL20, JHS^{+}21, JLR^{+}22, KX22, LVHK20, LCM^{+}22, LTH21, LTW^{+}24, MAP^{+}22, PC21, SGL^{+}23, SDS^{+}23, TNK^{+}22, WWFS21, XYL^{+}23, YAE^{+}23, ZZWL22, ZSZ^{+}21, dSNIB22]. Responses [MXL^{+}21, MY23, CLG^{+}24, CHAA21, DKW^{+}21, FM21, GFG^{+}22, HDS^{+}20, JWAvD^{+}22, KL22, LLSQ24, MLP^{+}21a, MHD^{+}22, MAG^{+}22, NEB^{+}22, Phi20, Phi21, PW22, RMH^{+}22, RFW^{+}24, TBH^{+}22, TMS^{+}21, VGvdB^{+}21, WPO^{+}22, ZZHZ23, ZJ22a, ZJ22b, ZLW21b]. Resting [ZLG^{+}23]. restoration [AMO20, HLS^{+}22, LAB^{+}22]. restructuring [HFL^{+}20]. result [CLC^{+}22]. Results [MKBSK19]. Resurrection [ZJ22b]. resuspension [CGD22, MDBI20, MCS^{+}23]. retain [LDG^{+}21]. retention [GFG^{+}22, GPDA^{+}20, SDL22, VMT^{+}20]. Rethinking [YJ20]. retreat [BJK^{+}21, NB23, SJJ^{+}21]. retrieval [SSB^{+}22]. retrieved [RCS^{+}20]. Return [FHMK20, MLG^{+}20]. reveal [BRB^{+}22, CLW^{+}21, DHNH22, HS21, LABM20, MVP^{+}20, MSB^{+}20, PMF^{+}21, RPKC22, WWGP21, ZSZ^{+}21]. revealed [BO20, DLL20, EST^{+}24, GLN^{+}20, KAC^{+}23, LVDM22, LGI^{+}20, Lem20, MT23, MCYR20, RWMP^{+}23, SIP21, ST22, TION23, VLP^{+}21, ZSD^{+}20]. reveals [ASL^{+}24, BRCO21, BKMY21, CGGC^{+}21, FR23, HPF^{+}23, HHK^{+}22, MKD22, Roy20, TRS^{+}20, VMP^{+}21, ZJ22b]. reversals [LDG^{+}21]. reverse [ZJ22a]. review [BR22, CLA^{+}24, CBPNA22, MA21, SHPW22]. Rhizaria [BO20, IKN^{+}21, LLMM^{+}23, MNM^{+}21, MLT^{+}21]. rhizosphere [EJM^{+}23, HZZ^{+}20]. ria [NSK^{+}21]. rich [LPO^{+}20, LAMG^{+}22, TKW^{+}23]. richness [RBL^{+}20]. Riemann [MBD^{+}23]. riparian [FW21]. Rise [MWV22, FNB^{+}23, LVHK20, LRM21, QMG21, YLT^{+}20, ZvBZ^{+}20]. rising [IOS24, NB23]. risk [LAFV23, Roy20, ZLG^{+}23]. River [DSY20, LTH21, MGI^{+}22, SLRS^{+}24, YGD^{+}21, ZPK^{+}22, AWB^{+}23, BMLV24, CBHB21, CCY^{+}21, CGC^{+}21, CHS^{+}20, CVF^{+}20, DMC^{+}22, DBM^{+}23, DKPS23, EST^{+}24, FXD^{+}22, IOS24, KHHZ20, LHMR23, MGL22, MHL^{+}20, QS21, RHP^{+}23, SML21b, SLZ^{+}21b, TUR^{+}20, WPY^{+}20, ZYZ^{+}21, CM21, GBHF21, LXC^{+}21b, PCN22, Roy20, SMNT22, TMSL22]. river-bay [WPY^{+}20]. river-dominated [DKPS23, QS21]. river-influenced [FXD^{+}22]. Riverine [IYW^{+}22, HLKD23, KRB^{+}21, PBGGRB21, RMW^{+}21, RD23, YGD^{+}21]. rivers [AFH^{+}21, BHA^{+}22, GEHD20, MBH^{+}23, MA21, NSK^{+}21, PHB^{+}22, RHP^{+}23, SMG^{+}22, ZQB^{+}21]. rK}y [CLGH23]. Rn [ACW^{+}22]. robotic [GLN^{+}20]. rock [KPNJ23, ZPK^{+}22]. rockweeds [DP22]. rocky [DP22, GTKW^{+}20, VMT^{+}24]. Role [AVS^{+}22, GGS^{+}21, GRJ20, LNC^{+}23, MDBI20, MLT^{+}21, YLT^{+}20, AKT^{+}21, DCJB22, EBW^{+}20, EBK^{+}20, EFG^{+}24, HLPT20, HGMK22, HMB^{+}22, KRB^{+}21, KdED^{+}20, KNG^{+}23, LLMM^{+}23, LZG21, MEQBV20, SPLJA^{+}20, TGE^{+}21, WLK^{+}21, YGD^{+}21, ZSZ^{+}22, ZXZ^{+}23, dOKC^{+}21]. roles [CSB22, HDK^{+}21]. rolling [ALG^{+}23]. Roseg [PSE^{+}20]. Ross [CLP^{+}24]. rotating [SEG21]. rotationally [REF^{+}21]. rotifers [Gil22]. Rottnest [MVBG20]. roughness [YPF^{+}23]. rRNA [DLL20]. RS19 [KMM^{+}22]. rubble [KDW^{+}23]. rubescens [SDS^{+}23]. Rubisco [RBR^{+}24]. runoff [FBJ^{+}23, HSP^{+}20, LLC^{+}23].

S [DLL20, BSGA21, HZC^{+}24, RvHF^{+}20]. Sahara [NEB^{+}22]. Saharan [BFE^{+}20]. saline [IÁSNCR21, SSP^{+}20]. Salinity [ZLJ^{+}24, ETG21, GR24, IMB^{+}22, LSC^{+}21, OBL20]. salinity-alkalinity [IMB^{+}22]. salinizing [ZZCD23]. salmon [ASH^{+}20, FBJ^{+}23, KHB22]. salps [SDSGR21]. Salt [YJB^{+}20, BVM^{+}20, DSY20, EJM^{+}23, GBMKJ22, KBB^{+}22, LVHK20, LGF22, MKM^{+}23, OJUN23, SWK^{+}23, TEK^{+}21, WLK^{+}21, XWL^{+}21, YLT^{+}20, ZZCD23]. salt-wedge [WLK^{+}21]. saltmarsh [HGA24, LDGB^{+}24, LCS^{+}20, WPO^{+}22, YXC^{+}22, ZZL^{+}21, ZvBZ^{+}20]. sampled [IMB^{+}22, PYM^{+}20]. samples [GWS^{+}21, KSR20]. sampling [BRCO21, RWT^{+}21]. San [BM22, BMB22, GNO22]. sandy [KPM^{+}20, WAPC20, vEMF^{+}20]. Sanriku [NSK^{+}21]. Sargasso [LPO^{+}20, LSS^{+}24]. satellite [CLA^{+}24]. save [DMCW23]. scale [ASL^{+}24, BMA^{+}20, BGAD^{+}22, DAN^{+}22, DMM^{+}21, DFW^{+}20, EWF22, GBC^{+}20, GSD^{+}23, GLB^{+}22, KPNJ23, MACC24, MSGW^{+}20, MHdS^{+}21, MGI^{+}22, OYM^{+}20, PATL^{+}23, SH22, SLRS^{+}24, SMKS22, WO21]. scales [AH20, AHF^{+}22, BRB^{+}22, CESS21, FBR^{+}23, LM24, MCH^{+}21, PYM^{+}20, RTM^{+}20]. Scaling [CGC^{+}21]. Scallop [TJW^{+}22, CGR22, FSH^{+}23]. Scandinavia [LDGB^{+}23, LDGB^{+}24]. scattering [KSR20]. scavenging [YPF^{+}23]. scenario [AAC^{+}21]. scenarios [HEH^{+}20, JHS^{+}21]. sclerochronology [MMLPRHCG22]. Scotia [DLL20]. Sea [TMS^{+}21, AGC^{+}21, BSGA21, BMR^{+}21a, BMLV24, CLW^{+}21, CHAA21, DNB^{+}20, FNB^{+}23, FXD^{+}22, GVF^{+}23, HMM^{+}22, JHvL^{+}23, LLL^{+}23, LAFV23, LVHK20, LRM21, MZXY22, MY23, MBLA^{+}21, MKBSK19, MLG^{+}20, OSR20, OBB^{+}20, PCN22, PGN^{+}21, QMG21, SH22, SIP21, TMKB^{+}22, VCT^{+}23, WHAR^{+}20, WOMWR20, WXK^{+}22, WDG^{+}21, YLT^{+}20, ZSD^{+}20, ZXE^{+}23, ZvBZ^{+}20, CGRGP^{+}21, CGR22, CBB^{+}22, CDA^{+}20, CCDM21, CLP^{+}24, DNB^{+}20, DGW^{+}21, DLL20, EFP^{+}21, GLM^{+}23, HDD^{+}22, HPNU20, KPL^{+}20, LCHB22, LTJ^{+}20, LPO^{+}20, LNC^{+}23, LSS^{+}24, NvBE^{+}21, RHS^{+}22, RvHF^{+}20, RSH^{+}20, RBR^{+}23, SORGC^{+}21, SGSD^{+}22, TMKB^{+}22, ZZO^{+}21, vHLV^{+}20]. Sea-ice [TMS^{+}21, MLG^{+}20, VCT^{+}23]. sea-level [FNB^{+}23, YLT^{+}20]. seabed [CBF^{+}23, SHPW22, TPTS23]. seabird [CDW^{+}24]. seabird-derived [CDW^{+}24]. seafloor [GPDA^{+}20]. Seagrass [AMO20, AAC^{+}21, AHCF22, BBMD20, HHMH22, JvKvT^{+}20, JHBB22, KHK^{+}22, LLTT^{+}24, MRH^{+}23, RLP^{+}20, RS21, SVMM^{+}22, SN22, VLP^{+}21, YvdHC^{+}20, ZWM22, dSNIB22]. seagrass-dominated [JvKvT^{+}20]. seagrasses [LSW^{+}20b, THVJ23]. seals [KTH^{+}21]. seas [BND^{+}23, LCZ^{+}21, YWZZ20, MLL^{+}21]. seascape [PVSP^{+}22, SHPW22]. Season [WWB^{+}20, GNO22, SZA^{+}21, SSR^{+}23]. season-dependent [GNO22]. Seasonal [AAB^{+}22, ACA^{+}22, CCB^{+}20, CMCC24, CGL^{+}21, CP22, DLDF21, GPK^{+}21, HABL^{+}21, LZG21, LSS^{+}24, MMF^{+}21, PLB^{+}21, RWMP^{+}23, RBSB22, SPR22, SSCE20, ZWM22, AWGV20, AJMG21, EFG^{+}24, HBG^{+}22, HSP^{+}20, IÁSNCR21, JLW^{+}23, KWS^{+}21, LVZ21, MLK^{+}22, MGGRR^{+}21, NGSS23, RHM^{+}21, SKB^{+}22, SCT^{+}21, SGS22, TKS^{+}20, VCBB20, XYL^{+}23, vdPMF^{+}21]. Seasonality [CCFF21, CCDM21, LRNG24, MK21, RPKC22, MKD22]. Seasonally [BM22, CVF^{+}20, LXC^{+}21b, MLK^{+}22, PJRV23, SSM^{+}22b, THVJ23]. Seasons [HCNO^{+}20, GCW^{+}23, JHBB22, SML21b]. seawater [Edm21a, KSR20, MAD^{+}22, NTJ^{+}20, SOSN23, YPF^{+}23, YWZZ20]. seaweed [WWB^{+}20]. seaweeds [RHM^{+}21]. secondary [BD22, CCFF21, DBFC24, FSM^{+}21, LEK^{+}22, ZZL^{+}21]. Secretion [HVMS^{+}21]. sector [YAH^{+}21]. Sediment [CEE^{+}22, HZZ^{+}20, MP23, PGB24, SVMM^{+}22, AJMG21, BBH^{+}20, BBCA^{+}21, CGD22, DLDF21, GWS^{+}21, GMI^{+}21, HBS^{+}22, HMM^{+}22, HLKD23, HHMH22, HLPT20, HNS21, HMB^{+}20, JLMW21, LVHK20, LNR^{+}20, LTH21, MDBI20, MIR^{+}22, MCS^{+}23, MWHR^{+}22, MGG^{+}24, OYM^{+}20, PATL^{+}23, PSK21, RWL^{+}24, SMCS23, YLT^{+}20, ZTT23, ZSF^{+}23, ZWM22, dSNIB22, vHHP^{+}22]. sediment-deficient [LVHK20]. Sedimentary [SLHL^{+}22, HLPT20, LLK^{+}22, LL23, RBR^{+}23, SGA^{+}22, SGGB21, TKS^{+}20]. sedimentation [HZM^{+}21, LLC21, RNW^{+}23, SVBT20]. sediments [AGC^{+}21, BAG^{+}24, BJK^{+}21, BVM^{+}20, CBB^{+}22, CHP^{+}24, CBF^{+}23, DGTW21, DLC^{+}22, FXD^{+}22, HPB^{+}21, HBR^{+}22, KBB^{+}22, LDGB^{+}23, MHG^{+}23, NKH^{+}23, ORE20, OGD20, PYM^{+}20, RvHF^{+}20, RLP^{+}20, RWS^{+}22, RCS^{+}20, SSM^{+}21a, SJR21, SSEO21, SSCE20, SZY^{+}20, THL^{+}21, XLN21]. seed [ZZL^{+}21, ZSZ^{+}22]. seedling [vHHP^{+}22]. seedlings [OJUN23, SRN^{+}21]. seepage [BKMY21, LBG^{+}22]. seeps [AGC^{+}21, GPK^{+}21]. seiche [NH22]. seiches [BMS^{+}22]. select [LPO^{+}20]. selection [MTB^{+}21, RTK23]. Selective [FLW^{+}21, HMBB21, KHSP21]. semi [BKC20, WCC20, XYL^{+}23]. semi-enclosed [XYL^{+}23]. semi-labile [BKC20, WCC20]. semianalytically [LSW^{+}20a]. senescence [ZWM22]. sense [ZLG^{+}23]. sensing [LSW^{+}20a, SMF20]. sensitive [BJS^{+}23b, FLH^{+}23, HZM^{+}21]. sensitivities [LSCL21]. sensitivity [CLW^{+}21, CCFF21]. sensor [BCWS23, SSM^{+}22a]. sensor-based [BCWS23]. sensory [RS20]. separation [LTJ^{+}20, MCYR20, NH23]. September [TNK^{+}22]. sequencing [BBGM20, DLL20, WZCK20]. sequestration [FXD^{+}22, GBMKJ22, JHSG^{+}24, LDGB^{+}24, PYY^{+}20, SHS^{+}22, XWL^{+}21, YXC^{+}22]. series [BHA^{+}22, CBN21, FHMK20, FR23, FIH^{+}20, GWS^{+}21, HRK^{+}22, LABM20, MKD22, SGM^{+}22, TRS^{+}20, WDG^{+}21]. services [OBB^{+}20]. setae [SMF20, SYMF21]. settings [BDE^{+}22, CEE^{+}22]. settlement [MURK20, OSR20]. settling [ABS^{+}21, EAF^{+}22]. setup [REF^{+}21]. severity [CTF^{+}21]. sexual [MGP^{+}21]. Shallow [BSIdG23, BMR^{+}21b, BOC^{+}21, CB21a, FBV23, HBS^{+}22, HZZ^{+}20, JOC^{+}20, LMS23, LBG^{+}22, LZM^{+}21, LAMG^{+}22, PSK21, RFB^{+}22, RHA^{+}23, RLL^{+}21, RS21, SPBK21, SSCE20, SWS^{+}22, TBH^{+}22, TKM^{+}22, WCZ^{+}20]. shallow-lake [LZM^{+}21]. shallow-water [LBG^{+}22, SWS^{+}22]. shape [CEE^{+}22, FWvA^{+}21, FBRG23, GGR24, KVKS22, SGM^{+}22, TKM^{+}22, ZLIP20]. shapes [RBAC^{+}20]. shaping [EBK^{+}20, LNC^{+}23, ZSZ^{+}22]. share [MRD^{+}21]. sharks [PRS^{+}20]. Sharp [BHSG^{+}20]. shelf [CHP^{+}24, CBF^{+}23, CHAA21, EFG^{+}24, FBRG23, GPDA^{+}20, GPK^{+}21, HJWP24, KTH^{+}21, KPS21, LvBS^{+}22, QRWT^{+}22, SLJ^{+}22, WRC^{+}21, ZLU^{+}20, MMF^{+}21]. shell [GPB^{+}20, MDWT^{+}22]. shells [FSH^{+}23, TJW^{+}22, WOMWR20]. sheltering [ZN20]. Shen [LD22]. Shield [GB21, GB23]. Shift [MLP21b, AFH^{+}22, BTK^{+}21, BHSL21, GBO^{+}20, RGP^{+}22]. shifted [MPW^{+}23]. Shifting [DAN^{+}22, ZZHZ23]. shifts [AH20, CLA^{+}24, CHA^{+}24, DMC^{+}22, GPK^{+}21, HLH^{+}23, MWV22, MKL^{+}24, RNW^{+}23, RVvB^{+}23, RS21, SPHH22, TMKB^{+}22, WCZ^{+}20, BWV^{+}20]. shipboard [BBA^{+}23]. shipwrecks [HMM^{+}22]. shore [FGO^{+}20, GLN^{+}20, GLN^{+}21, GTKW^{+}20]. shores [DP22]. Shoreward [KTH^{+}21]. Shoring [BV21]. Short [RTK23, LRNG24, ORE20, TMS^{+}21]. short-term [LRNG24, ORE20, TMS^{+}21]. shortages [LMSN23]. show [CLG^{+}24, KFHS24]. shows [JLW^{+}23, OLR^{+}23]. shredders [FRRG^{+}21]. shrimp [HBY^{+}22]. shunts [GM22]. Siberia [HABL^{+}21, SLHL^{+}22]. sick [ONH^{+}20]. side [MWHR^{+}22]. side-by-side [MWHR^{+}22]. Siderophore [PDK^{+}23]. Sierra [CM20a]. signaling [BMGK22, MK20]. signals [ZLG^{+}23]. signature [LHGJ21]. signatures [TZF^{+}22]. Significant [RVvB^{+}23, GSD^{+}23, MTP^{+}23, RNW^{+}23, WF22, YML^{+}21, ZXE^{+}23]. significantly [RWL^{+}24]. Signy [MSGW^{+}20]. silica [IKN^{+}21, MLT^{+}21, RWL^{+}24, Roy20, XSZ^{+}21, ZBNH^{+}21]. silicon [LLMM^{+}23, LAMG^{+}22, MBLA^{+}21, WSLC22, ZSD^{+}20]. silled [HJWP24]. similar [CTP21, CCG^{+}22, HMB^{+}20, LTJ^{+}20, SVMM^{+}22, VWR^{+}23]. Simple [SDL22, ZLN21]. simple-wave [ZLN21]. Simulated [JHBB22, GRJ^{+}22, JHSG^{+}24, MLG^{+}20, SBL^{+}21]. simulations [BGAD^{+}22]. simultaneous [THL^{+}21]. Single [KSN^{+}21]. sink [AFH^{+}22, PUH^{+}23, RTM^{+}21]. sinking [ALG^{+}23, CKH^{+}21, CBN21, DKBG21, DLL20, JLS^{+}20, LCMG^{+}20, LG21, LBL^{+}24, SAD^{+}20, HCB^{+}22]. sinks [GMNG^{+}24, SGGB21]. siphonophores [HCH^{+}24]. site [OFM^{+}21, TCP^{+}22]. site-specific [TCP^{+}22]. sites [MSG^{+}22]. situ [BO20, BBGM20, DHNH22, GPB^{+}20, GSD^{+}23, MHMML^{+}22, MCYR20, SSM^{+}22a, VMP^{+}21, WLSD21]. Size [BJS23a, BMLV24, JSW^{+}21, KSF^{+}23, SDSGR21, AMC21, ALH^{+}21, CLG^{+}24, CHAA21, CHA^{+}24, FGP22, GNO22, HBG^{+}22, KSR20, LSL^{+}20, LHS^{+}20, LXC^{+}21b, LDB^{+}20, LHMR23, MGI^{+}22, PMRG20, TATC^{+}24, WOK^{+}21, dJGCS23]. size- [GNO22]. Size-dependent [BMLV24]. size-fractionated [KSR20, LXC^{+}21b]. Size-independent [JSW^{+}21]. Size-specific [SDSGR21]. skeleton [CDW^{+}24]. Slope [SMP21]. sloping [NH23]. Slow [KBB^{+}22]. Small [SWC^{+}23, AH20, BCWS23, CM20b, Edm21b, EWF22, GB21, GB23, IJSC20, KSF^{+}23, LBL^{+}24, MBA^{+}21, MSGW^{+}20, MLT^{+}21, OYM^{+}20, PCJ^{+}21, PRS^{+}20, PSK21, WF22]. small-scale [EWF22, MSGW^{+}20, OYM^{+}20]. smaller [MN20, RHP^{+}23]. snow [DBFC24, KFHS24, SGL^{+}23, SEG21]. snowpack [PFB^{+}22]. sockeye [ASH^{+}20]. soft [FPBG^{+}22]. soil [PYY^{+}20, VJS^{+}20, ZJW^{+}22]. soils [HGA24, LDGB^{+}24, PNT^{+}21]. Solar [vdPMF^{+}21]. soluble [LMS^{+}21]. somatic [MRC^{+}24]. sorption [PNT^{+}21]. soundscapes [CLW^{+}21]. Source [SCB^{+}20, AFH^{+}21, AFH^{+}22, BBH^{+}20, BHA^{+}22, BOC^{+}21, DBFC24, DKPS23, HSD^{+}20, IFU^{+}20, LCHB22, PUH^{+}23, RTM^{+}21, STB22, VST^{+}22]. source-specific [BHA^{+}22]. source/sink [RTM^{+}21]. sources [AHCF22, BND^{+}23, BSF^{+}21, BCL^{+}22, CSD^{+}24, GMNG^{+}24, KRC23, KLG^{+}23, LZD^{+}23, MV22a, MV22b, MTB^{+}21, MHL^{+}20, MCS^{+}23, PMJ^{+}21, SSS^{+}24, WLK^{+}21]. South [MSGW^{+}20, SMW^{+}21, AHCJ^{+}20, GRJ^{+}22, JLS^{+}20, WPY^{+}20]. Southeast [LL23]. Southern [XYL^{+}23, YAP^{+}22, DAB^{+}21, LDGB^{+}23, LDGB^{+}24, NvBE^{+}21, BRCO21, BDE^{+}22, BBA^{+}23, CGL^{+}21, GRJ^{+}22, HCB^{+}22, IJSC20, LMC^{+}21, LWvdM^{+}21, MT23, MSB^{+}22, MPW^{+}23, MLT^{+}21, MGC^{+}20, RKBT^{+}23, YSHS23]. Southwest [ZZWL22]. sp [DFW^{+}20, GTKW^{+}20, vGDH^{+}21]. space [FWvA^{+}21, FMT^{+}20, LHD^{+}20, MO23]. space-for-time [FMT^{+}20, MO23]. Spain [TZF^{+}22]. spanwise [JYA21]. Spatial [BBH^{+}20, DvOH^{+}20, KHSP21, LBBM21, MHdS^{+}21, NB20, RHA^{+}23, RWT^{+}21, RKBT^{+}23, SML^{+}21a, SMNT22, WRWPG19, WZCK20, AHF^{+}22, BR22, BRB^{+}22, CGRGP^{+}21, CBHB21, CCB^{+}20, CMCC24, FBR^{+}23, GCW^{+}23, HMP^{+}21, HZC^{+}24, KBL^{+}20, LM24, LAFV23, LGI^{+}20, MLK^{+}22, MKP^{+}23, PYM^{+}20, Roy20, SORGC^{+}21, SCC^{+}21, YAP^{+}22, ZBNH^{+}21, Ano21c]. spatially [MOW^{+}22]. Spatio [LSC^{+}21, AH20]. Spatio-temporal [LSC^{+}21, AH20]. Spatiotemporal [Edm21a, GZX^{+}22, MHP^{+}23, CHS^{+}20, FKN21, GCEM^{+}23, IJSC20, MCH^{+}21, OSR20, PSK21, SPBK21, SLRS^{+}24]. Spawning [OFM^{+}21, OLR^{+}23, SRV^{+}20]. specialists [MZW^{+}21]. speciation [BCA^{+}20, RBR^{+}23]. Species [CSB22, BTT23, BRB^{+}20, BRB^{+}22, CCG^{+}22, EWF22, FRRG^{+}21, GMS21, GBB^{+}21, HANW21, HDD^{+}22, IKN^{+}21, LBJS22, LTJ^{+}20, MVP^{+}20, MN20, NB23, SVMM^{+}22, SRN^{+}21, SYT^{+}21, VGvdB^{+}21, VJS^{+}20, WWGP21, WNHY21, vHHP^{+}21, vHHP^{+}22]. Species-specific [CSB22, IKN^{+}21, WWGP21]. specific [BHA^{+}22, BBCA^{+}21, CLG^{+}24, CSB22, DMS^{+}21, IKN^{+}21, LDGB^{+}24, OLFH21, SDSGR21, TCP^{+}22, WWGP21, vdPAV^{+}20]. specificity [KSN^{+}21]. spectra [ALH^{+}21, SSB^{+}22]. Spectral [TPTS23, HZS21, LLBC20]. sphaericus [WNHY21]. spiny [BFCC^{+}21]. Spitsbergen [vdPMF^{+}21]. Sponge [LAMG^{+}22, LMP^{+}20, LMS23, MBLA^{+}21]. Sponges [HSD^{+}20, BMR^{+}21a, HOB^{+}22, MSL21, MRC^{+}24, MCYR20, MDIY23]. Spread [TNK^{+}22]. spring [ACA^{+}22, AFK^{+}22, BAHH^{+}20, CGRGP^{+}21, CM20b, FBRG23, KPG^{+}23, MLC^{+}21, PW22, REF^{+}21, SIP21, VCT^{+}23, YWLY20, vdPAV^{+}20]. Springtime [JWAvD^{+}22]. sprinkling [LCHB22]. squid [PMRG20]. Stability [MDDL22, BOC^{+}21, LMS^{+}21, SSP^{+}20, VKGW22, WMTJW22, ZTT23, dCLO^{+}23]. Stable [GAS^{+}23, HFL^{+}20, WRS^{+}22, CDH^{+}21, EST^{+}24, HCH^{+}24, HZC^{+}24, IFU^{+}20, KLG^{+}23, LVDM22, MLP^{+}21a, SWS^{+}22, TMM^{+}20, WAPC20]. stage [BFCC^{+}21]. stages [OFM^{+}21]. stalked [FJP^{+}21]. state [DMC^{+}22, RNW^{+}23, RVvB^{+}23, RGP^{+}22, SH22, SWOR20, SSWM20, TMKB^{+}22]. states [BSIdG23, TBH^{+}22, BCTH20, CCA^{+}21, DRA^{+}23, FBK^{+}22, GGR24]. Station [NCC^{+}21, CBN21, SBM^{+}21]. status [THVJ23]. stilling [JOC^{+}20, SGM^{+}21]. stimulates [BVM^{+}20, RGW^{+}23]. Stimulation [LBL^{+}24]. stingrays [CSB22]. Stockholm [vHLV^{+}20]. stocks [CEE^{+}22, HMP^{+}21, LDGB^{+}24, SVMM^{+}22]. Stoichiometric [FDHS20, MSJ21, KTA^{+}21]. Stoichiometry [Roy20, DTPV21, FBV23, FPB^{+}23, LG21, MV22a, MV22b, STB22, SML21b, VKB^{+}22, WOT^{+}23, WWFS21, ZLS^{+}21]. Stokes [FGO^{+}20, LCMG^{+}20]. stony [FPBG^{+}22]. storage [BRM^{+}23, LRNG24, MRH^{+}23]. stores [MKM^{+}23]. storm [DAD^{+}21, KX22, WTvD^{+}22, vHHP^{+}21]. storm-induced [DAD^{+}21, WTvD^{+}22]. storms [LRM21, TKM^{+}22]. strains [HDS^{+}20]. stranded [RHM^{+}21]. Strategies [FRRG^{+}21, BDL^{+}21, CZH^{+}21, DFJ22, HXL^{+}21, RCH^{+}20, TATC^{+}24, WTvD^{+}22, WRHR22, ZLJ^{+}24]. strategists [CLGH23]. strategy [LZG21]. Stratification [DMM^{+}21, KX22, XTL^{+}20, MO23, PMD^{+}23, PW22, SSP^{+}20, SGM^{+}21, YWLY20, YJB^{+}20]. stratified [ACA^{+}22, Ano21a, CYPRG^{+}23, CKCB^{+}23, EAF^{+}22, GSG^{+}17, LVDM22, LSW^{+}20a, MAG^{+}22, SBW21]. Stream [BCL^{+}22, SDB^{+}23, BWB^{+}22, CGGC^{+}21, CZK^{+}23, DPB^{+}23, FAB^{+}20, FMT^{+}20, IFU^{+}20, LHL^{+}21, MACC24, MKP^{+}23, MHSL22, MGI^{+}22, PSE^{+}20, RGP^{+}22, RVS^{+}22, SBB20]. streambed [RBAC^{+}20]. streams [AFH^{+}21, ARBvS^{+}21, BHR^{+}23, BHSL21, BCL^{+}22, BKC20, BRM^{+}23, BRB^{+}22, CTF^{+}21, DRA^{+}23, FBJ^{+}23, LLC21, MA21, MGGRR^{+}21, RSLW23, RWT^{+}21, RRHS^{+}21, RHP^{+}23, RBSB22, SPR22, TKW^{+}23, WK21]. strength [DMM^{+}21, WWB^{+}20]. stress [ALH^{+}21, EFP^{+}21, FLH^{+}23, LRdGFP23, MRC^{+}24, MRT^{+}24, WWB^{+}20]. stressed [BWV^{+}20]. stressors [ABC^{+}21, HHMH22]. Strong [BLS20, GCMJ23, SJJ^{+}21, AH20, KTA^{+}21, LMM^{+}21, LHH24, MSGW^{+}20, WXK^{+}22]. stronger [BZK^{+}22]. strongly [LRNG24, RGW^{+}23]. Structural [KHHZ20]. Structure [LNR^{+}20, BGM^{+}22, BR22, CCFF21, CGD22, DBFC24, FWvA^{+}21, FKN21, GSP^{+}23, GGHBS^{+}20, HCH^{+}24, IJSC20, LWS^{+}22, LHS^{+}20, LDB^{+}20, MMF^{+}21, MSGW^{+}20, MDDL22, PFB^{+}22, PSNS22, RPKC22, SGM^{+}22, SSK24, TMKB^{+}22, WRC^{+}21, WCZ^{+}20, YvdHC^{+}20, YAP^{+}22]. structures [ASL^{+}24, FHR^{+}21, KTH^{+}21]. structuring [SORGC^{+}21]. studies [CBPNA22, KSF^{+}23]. study [BBMD20, CZH^{+}21, CB21a, HKM^{+}20, JLMW21, KRM^{+}20, LBvE23, LTJ^{+}20, MAG^{+}22, MCYR20, PHM21, SWOR20, WTvD^{+}22, CBN21]. sub [BJK^{+}21]. sub-Antarctic [BJK^{+}21]. subalpine [SEK20]. subantarctic [HCB^{+}22, SBV^{+}21]. Subarctic [HDK^{+}21, BGG^{+}20, CBS^{+}21, DBFC24, MTP^{+}23, MDDL22, MMG^{+}21, MK21, PCJ^{+}21, PPB^{+}21, SWC^{+}23, SSM21b, TION23, VKK^{+}21, WOK^{+}21]. subecotypes [TKA21]. subglacial [MP23]. subject [KWB21]. Submarine [RTM^{+}21, FKN21, IÁSNCR21, KRM^{+}20, LNR^{+}20, WLK^{+}21]. submerged [GCMJ23, HGMK22]. Submesoscale [FWvA^{+}21]. subpolar [WF22]. subpopulation [DCJB22]. subsidies [EST^{+}24, EFC^{+}23, FW21]. Subsidized [BWV^{+}20]. substantially [CKCB^{+}23]. substitution [FMT^{+}20, KMM^{+}22, MDP^{+}23]. substrate [CLC^{+}22, MKBSK19, dOKC^{+}21]. substrates [DWW^{+}21, HP20, MHMML^{+}22]. Subsurface [KWB21, GLN^{+}20, SSH^{+}22, SLJ^{+}22]. subterranean [BPL22, WAPC20]. subtidal [SSEO21, SPKM22, SMKS22]. Subtle [LGI^{+}20]. Subtropical [HYPG^{+}22, RPKC22, TKA21, ZBNH^{+}21, BAD^{+}21, CKH^{+}21, FIH^{+}20, GLAHH^{+}23, HZZ^{+}20, LSCL21, LBL^{+}24, MZW^{+}21, SGS22, VLP^{+}21, WPY^{+}20, YMSH20, YOXC23, YAE^{+}23, YAE^{+}24, ZLU^{+}20]. Successful [SBH23]. succession [ACE^{+}22, CLGH23, LVZ21, NGSS23, PWS^{+}20]. Successional [THT^{+}20, BFE^{+}20, vHHP^{+}21]. successive [MWV22]. succumb [FLH^{+}23]. sudden [ZZL^{+}21]. sugars [GAS^{+}23]. suggest [PKV20, WF22]. suggests [DAB^{+}21]. suitable [GPDA^{+}20]. Sulfate [MWHR^{+}22, SSM^{+}21a, KWS^{+}21, PUH^{+}23, SJR21, SZY^{+}20, THL^{+}21]. Sulfate- [MWHR^{+}22]. sulfate-dependent [SZY^{+}20]. sulfate-enriched [KWS^{+}21]. sulfate-reducing [SJR21]. sulfated [HVMS^{+}21]. sulfide [MT23, THL^{+}21, XYZ^{+}21]. Sulfolipid [MDP^{+}23]. Sulfur [CCK^{+}21, ORRMD^{+}21, HDK^{+}21, LWC^{+}24, MZXY22, MY23, PUH^{+}23, THL^{+}21]. Summer [DCF^{+}22, PFB^{+}22, ACA^{+}22, CSNS24, CLP^{+}24, DMM^{+}21, JHvL^{+}23, KCT23, LWC^{+}24, LNC^{+}23, MBH^{+}23, MAG^{+}22, PW22, PGN^{+}21, SBM^{+}21, SHE^{+}21, VJH^{+}22, ZLW21b]. summertime [RHA^{+}23, SLJ^{+}22]. Superior [PUH^{+}23, SRL^{+}20]. superoxide [AFM23]. supersaturated [LSW^{+}20b]. supersaturation [CM20a, SBW^{+}20]. supervised [SCC^{+}21]. supplements [LRdGFP23]. Supply [SCT^{+}21, GMI^{+}21, LLC^{+}23, LL23, MFGF^{+}22]. Supply-controlled [SCT^{+}21]. support [LAB^{+}22, SWC^{+}23]. supported [SZY^{+}20]. supporting [HMB^{+}22]. supports [FSM^{+}21, MK20, SPLJA^{+}20, SBW^{+}20]. Surface [GLAHH^{+}23, BM22, CCB^{+}20, DNB^{+}20, JLS^{+}20, KMV^{+}20, KTH^{+}21, MCS^{+}23, PATL^{+}23, SIP21, TCP^{+}22, WDG^{+}21, YPF^{+}23, ZZO^{+}21]. surficial [ZPK^{+}22]. surges [LHH24]. surrounding [AGC^{+}21]. surroundings [RNW^{+}23]. surveillance [ARB^{+}22]. survey [DMM^{+}21, MVP^{+}20]. surveys [ACW^{+}22, LDG^{+}21]. Survival [SRN^{+}21, CGR22, HSP^{+}20, KHB22, RVBP23, vdPAV^{+}20]. survive [WRA^{+}23]. susceptibility [SIP21]. suspended [DWS^{+}22, DLL20, JLS^{+}20, SAD^{+}20]. suspension [RTK23]. sustained [CCC^{+}21]. sustaining [YLT^{+}20]. sustains [BVL^{+}23, MBLA^{+}21]. Svalbard [JLMW21]. swimming [CLT^{+}22]. Switches [YAE^{+}24]. Switzerland [PSE^{+}20]. Symbiodiniaceae [RCH^{+}20]. symbiont [MSD23]. symbionts [BCFI^{+}22]. symbiosis [WWGP21]. Symbiotic [FPBG^{+}22]. sympatric [CSB22, WRMP22]. sympatry [CBS^{+}21, PPB^{+}21]. symptoms [GÖA^{+}23]. Syn [HCNO^{+}20]. Synchronized [BRCO21]. synchronous [FR23, MRD^{+}21]. synchrony [BVGQN22, DPB^{+}23]. Synechococcus [LSCL21, WSLC22, ZLW^{+}21a]. Synergistic [LZM^{+}21]. synthesis [AWB^{+}23, MA21, MGWS23, WRA^{+}23]. syntopy [BRB^{+}22]. system [FIH^{+}20, KX22, KHCG^{+}22, LTW^{+}24, RMW^{+}21, SWZ^{+}23, VKGW22, WPY^{+}20, ZLU^{+}20, ZJW^{+}22, dMKB^{+}20, ACE^{+}22, CMB^{+}21]. system-wide [KHCG^{+}22]. Systematic [BGG^{+}20]. systematically [PKV20]. systems [HEH^{+}20, HMB^{+}20, KTA^{+}21, KSF^{+}23, LABM20, QS21, RBR^{+}24, SBH23].

Tackling [CDH^{+}21]. tadpole [BWB^{+}22]. Taihu [XMP^{+}21, GZX^{+}22, XDQ^{+}22]. tailings [LMS^{+}21]. Taiwan [LHS^{+}20]. tank [ALG^{+}23, SEG21]. target [BMGK22]. Tasmania [LWvdM^{+}21]. taurine [CDA^{+}20]. taxa [OMM21]. Taxon [OLFH21, vdPAV^{+}20, CLG^{+}24, GCEM^{+}23]. Taxon-specific [OLFH21, vdPAV^{+}20, CLG^{+}24]. taxonomic [GWS^{+}21, HNS21, PGEB22, SRM23, vdPMF^{+}21]. techniques [OAM^{+}22]. Temperate [MWV22, AFH^{+}21, AFH^{+}22, AHCF22, CTM^{+}21, CÖL^{+}20, DCJB22, DdEM^{+}23, FBJ^{+}23, FR23, GWS^{+}21, HBS^{+}22, HEH^{+}20, HSV22, KHB22, LAD^{+}22, LHMR23, MTP^{+}23, MGL22, MLC^{+}21, MURK20, NB23, PSK21, RHA^{+}23, RVvB^{+}23, RWS^{+}22, SPBK21, SML21b, SWK^{+}23, SOT^{+}20, TJW^{+}22, VJS^{+}20, WBB^{+}21, WLK^{+}21]. Temperature [MKL^{+}24, OBL21, PSK21, SMD21, AAB^{+}22, AFH^{+}21, AFK^{+}22, BM22, BMB22, BMH^{+}23, BWSW24, CGR22, CT23, CHAA21, DWR^{+}23, DCK^{+}22, DAD^{+}21, DKW^{+}21, Edm21a, FLH^{+}23, GR24, HM20, IOS24, JLR^{+}22, KDV^{+}23, Lem20, LSCL21, LNC^{+}23, LSDA21, MVP^{+}20, MRT^{+}24, PBC21, Phi20, Phi21, SJH20, STB22, SHZ^{+}21, SMKS22, VLP^{+}21, XDQ^{+}22, ZXE^{+}23]. Temperature-dependent [OBL21]. temperatures [NB23, OJUN23, SIP21, SSL22, SGM^{+}21]. Temporal [HLH^{+}23, HDD^{+}22, ZSF^{+}23, AH20, AHF^{+}22, FBR^{+}23, LSC^{+}21, LHS^{+}20, MV22a, MV22b, NB20, RHA^{+}23, RWT^{+}21, Roy20, RKBT^{+}23, SMNT22, WZCK20, ZBNH^{+}21]. temporally [FR23]. temporary [CTF^{+}21, FW21, GCEM^{+}23, GPB^{+}20]. Ten [FJP^{+}21]. term [BCWS23, BBMD20, BMH^{+}23, CB21a, CHA^{+}24, DAD^{+}21, FHMK20, FBV23, GBHF21, HBR^{+}22, LAD^{+}22, LRNG24, LVHK20, Lem20, LGQ^{+}21, LZM^{+}21, MVP^{+}20, MKBSK19, MNMJ^{+}21, ORE20, RSLW23, SJH20, SGM^{+}21, TMS^{+}21, YAH^{+}21, MKBSK19]. terminated [vHLV^{+}20]. ternary [ZJW^{+}22]. terrain [GPDA^{+}20]. Terrestrial [SOT^{+}20, YML^{+}21, AVS^{+}22, TSX^{+}23]. terrigenous [LL23, ZCZ^{+}23]. test [BO20, PPB^{+}21]. test-bearing [BO20]. Testing [CFC^{+}20, LD22, SB22]. Tetradesmus [ZSZ^{+}21]. tetraethers [ZLL^{+}24]. Thaumarchaeota [TRS^{+}20]. thaw [FBP20, PWS^{+}20]. their [BDMGH^{+}23, BCFI^{+}22, BCA^{+}20, CHA^{+}24, FPBG^{+}22, HBG^{+}22, HMBB21, HGA24, LLMM^{+}23, MLS^{+}21, MIR^{+}22, ZLW^{+}21a, ZLL^{+}21]. Thermal [BCFI^{+}22, AR20, CHA^{+}24, EFP^{+}21, KKP^{+}21, MRC^{+}24, MSD23]. thermally [EAF^{+}22]. thermokarst [HABL^{+}21, KSH^{+}23, LPA^{+}23, PWS^{+}20, ZSK^{+}21]. Thiamine [WK21]. thickest [SRN^{+}21]. Thin [PLLM21, GBC^{+}20]. Think [OYM^{+}20]. Three [CLT^{+}22, EWF22, MHL^{+}20, MZW^{+}21, VGvdB^{+}21]. Three-dimensional [CLT^{+}22]. threshold [HMBB21]. Thresholds [RS21]. throughout [SGSD^{+}22]. throughput [BBGM20, WZCK20]. Tibetan [DLC^{+}22, XTJW21]. Tidal [YOXC23, BD22, HSBC20, PBP^{+}23, PNT^{+}21, RWMP^{+}23, SVBT20, SWK^{+}23, WLSD21, XWL^{+}21, ZvBZ^{+}20]. tidal-flat [BD22]. Tidally [CSC^{+}21, FHR^{+}21]. tides [ZSZ^{+}22]. tidewater [RBR^{+}24]. Tight [HJWP24, XKP^{+}22]. Time [CBN21, FIH^{+}20, LABM20, Phi20, Phi21, TRS^{+}20, BHA^{+}22, FWvA^{+}21, FHMK20, FR23, FMT^{+}20, GWS^{+}21, HRK^{+}22, LvBS^{+}22, LHD^{+}20, MO23, MKD22, QS21, RCMA21, SDL22, SGM^{+}22, WDG^{+}21]. Time-series [CBN21, FIH^{+}20, LABM20, WDG^{+}21]. Time-varying [Phi20, Phi21]. times [EAF^{+}22, RTK23]. timing [BBM21, SJJ^{+}21]. Tipping [CBPNA22]. tissue [SWS^{+}22]. TN [FBV23]. TOC [Ano20-50, Ano20-51, Ano20-52, Ano20-53, Ano20-54, Ano20-55, Ano20-56, Ano20-57, Ano20-58, Ano20-59, Ano20-60, Ano20-61, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano21-49, Ano21-50, Ano21-51, Ano21-56, Ano21-52, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano22-48, Ano22-49, Ano22-50, Ano22-51, Ano22-52, Ano22-53, Ano22-54, Ano22-55, Ano22d, Ano22-56, Ano23-37, Ano23-38, Ano23-39, Ano23-40, Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano23-45, Ano23-46, Ano23-47, Ano23-48, Ano24m, Ano24n, Ano24o, Ano24p]. Tolerance [RBD22, EFP^{+}21, SSL22, ZZCD23, vHHP^{+}22]. tolerant [CTM^{+}21]. tomography [IKN^{+}21, MMLPRHCG22]. tool [ZLL^{+}24]. top [BMA^{+}20, BD22, GÖA^{+}23, LLBC20, RS20]. top-down [BMA^{+}20, BD22, GÖA^{+}23, LLBC20, RS20]. topology [GRJ20]. total [DSY20, MCS^{+}23, MGGRR^{+}21, QFM^{+}21, XYL^{+}23]. toxic [BMGK22, LBvE23, SKB^{+}22, ZJ22b]. toxicity [BH21, BK20, RD23]. toxigenic [MBB^{+}20]. toxin [WOT^{+}23]. TP [FBV23]. Trace [WOMWR20, BBA^{+}23, DBB^{+}21, HKM^{+}20, KDZ^{+}24, ZLL^{+}24]. Tracing [CDW^{+}24, VCT^{+}23, LLC^{+}23]. track [ZZO^{+}21]. trade [BH21]. trade-offs [BH21]. trades [BK20]. Trait [VMP^{+}21, GRJ20, LEB21, LVZ21, MLB^{+}21, MO23, WSZ22a, WSZ22b, vHHP^{+}22]. Trait-based [VMP^{+}21, LEB21, MLB^{+}21, WSZ22a, WSZ22b, vHHP^{+}22]. trait-environment [LVZ21]. traits [CZH^{+}21, DTPV21, EG20, FRRG^{+}21, HMBB21, KHK^{+}22, LEB21, OAM^{+}22, WOT^{+}23]. trajectories [MNMJ^{+}21, WPO^{+}22]. transcriptional [CTP21]. transcriptomics [MK20]. transect [DWS^{+}22, JLS^{+}20, JSL^{+}23, KPS21]. transfer [ALH^{+}21, HPF^{+}23, JYA21, RSP^{+}24, SWZ^{+}23, vGDH^{+}21]. transform [SJH20]. Transformation [HOB^{+}22, MZXY22, BKMY21, KRM^{+}20, ZLW^{+}21a]. transformations [SEK20]. transient [NKH^{+}23]. transition [BVGQN22, CLP^{+}24, KWS^{+}21, SBM^{+}21, YGD^{+}21]. Transitioning [BDE^{+}22]. transitions [FDF^{+}21]. transparency [BJS23a, DLC^{+}22, KLO23]. transplanted [MGP^{+}21]. transport [BGAD^{+}22, FHR^{+}21, FGO^{+}20, GLN^{+}20, GLN^{+}21, GPDA^{+}20, KPM^{+}20, LZG21, MDBI20, MLP21b, MURK20, OBL20, QS21, RPA^{+}22, VJS^{+}20, WKRL^{+}23, XLN21]. trap [BH21, DLDF21, GWS^{+}21, MTB^{+}21]. Trench [FUUG^{+}21, SML^{+}21a]. trends [BBMD20, BJG^{+}22, DKPS23, GBHF21, LDB^{+}20]. triad [SSA^{+}23]. tributary [MBB^{+}20]. Trichodesmium [EBW^{+}20, SdBW^{+}23, ZFQ^{+}20]. triggers [LMS^{+}21]. triple [LHGJ21]. tritrophic [EG20]. Trophic [CDH^{+}21, MSL21, MBC^{+}21, MCYR20, NDP^{+}24, PMRG20, WMTJW22, BOC^{+}21, CAV^{+}22, CUS21, HCH^{+}24, KPL^{+}20, LGQ^{+}21, MPW^{+}23, PMF^{+}21, RFB^{+}22, RBAC^{+}20, SWOR20, SPR22, SSWM20, THVJ23, VGSB23, WWGP21, YvdHC^{+}20]. Tropical [BYB^{+}23, EBK^{+}20, ETMD23, FSS^{+}23, SMW^{+}21, AAB^{+}22, ACW^{+}22, AHCJ^{+}20, BLS^{+}22, BPL22, CCFF21, CDH^{+}21, CMCC24, CBD^{+}23, DBM^{+}23, DNA^{+}22, FMT^{+}20, GCEM^{+}23, GLAHH^{+}23, KBvdD^{+}20, LBG^{+}22, LSDA21, MA21, MHdS^{+}21, MHVC22, NB20, PBGGRB21, RMW^{+}21, SVMM^{+}22, SDB^{+}23, THVJ23]. Trough [GNO22]. tuna [OFM^{+}21, OLR^{+}23]. tundra [LMM^{+}21, ZSK^{+}21]. tunicate [ST22]. Turbidity [RS20, BRHS20]. Turbulence [MBA^{+}21, EWF22, FIM^{+}22, GFG^{+}22, TGE^{+}21]. turbulent [LXC^{+}21a]. Turnover [GB21, BGM^{+}21, KBB^{+}22, LD22, MRC^{+}24, PKV20, SB22]. turtle [JHBB22]. twilight [BO20, KCT23]. Two [AHCF22, CTM^{+}21, ACW^{+}22, BCL^{+}22, CB21a, HTLP23, HOB^{+}22, LHMR23, LHH24, MLC^{+}21, MSG^{+}22, SPBK21, WRHR22, WAPC20]. type [dSNIB22]. types [CGGC^{+}21, HPCH24, KKA22, ZLG^{+}23]. typhoon [SYT^{+}21, TNK^{+}22]. typhoon-induced [SYT^{+}21, TNK^{+}22]. typhoons [LLC^{+}23]. Typical [QS21].

U.S [MMF^{+}21]. U.S. [HM20, MSJ21]. U.S.A [HSD^{+}20]. U.S.A. [HTLP23, YJB^{+}20]. Ubatuba [PLLM21]. UCYN [GFW^{+}20]. UCYN-A [GFW^{+}20]. ultraviolet [PC21, YG23, ZFQ^{+}20]. uncertain [FM21]. uncertainties [MBD^{+}22, RST^{+}21]. Uncertainty [CSD^{+}24]. uncharacterized [CBB^{+}22]. uncultured [LOW^{+}20]. Under-ice [HS21, OBL20, SGL^{+}23, KFHS24]. underappreciated [ZXZ^{+}23]. underestimated [PKV20]. underlies [YvdHC^{+}20]. underlying [BWB^{+}22, CBB^{+}22, HPF^{+}23]. Understanding [BPG^{+}22, KHK^{+}22, MMLPRHCG22, ALG^{+}23, MVBG20, MGG^{+}24, SBD^{+}21]. underwater [FIH^{+}20, HZS21, KAC^{+}23]. undescribed [KKP^{+}21]. undisturbed [EFC^{+}23]. Unexpected [BHSL21, MKP^{+}23, dOKC^{+}21]. unicellular [DCK^{+}22, LTLH22, MK20]. unique [CYZZ^{+}20, SRM23]. United [BCTH20, CCA^{+}21, DRA^{+}23, FBK^{+}22, GGR24]. Universal [LEB21]. Unlocking [RCH^{+}20]. unproductive [BYS^{+}23]. Unraveling [JLW^{+}23]. unvegetated [SSEO21]. upgrades [TCS21]. upland [BGM^{+}22]. upper [BM22, BMB22, ORRMD^{+}21, PGB24]. Uptake [FSH^{+}23, ARBvS^{+}21, BRB^{+}20, BDL^{+}21, CDA^{+}20, GDB20, HANW21, HCS^{+}21, KP20, LLMM^{+}23, LD22, MBH^{+}23, MUGL20, MAD^{+}22, MSB^{+}22, MDIY23, RCH^{+}20, SSR^{+}23, WM20, SB22]. upwelling [BVGQN22, CMB^{+}21, FBR^{+}23, HSP^{+}20, IÁSNCR21, KLO23, KDV^{+}23, LLC^{+}23, LTW^{+}24, MXL^{+}21, REF^{+}21, SBL^{+}21, SLZ^{+}21b]. upwelling-driven [BVGQN22]. urban [TBH^{+}22]. urbanization [IYW^{+}22]. urchin [OSR20, SH22]. urea [SHE^{+}21, SBB^{+}21]. urea-derived [SHE^{+}21]. USA [DP22, LZG21, VMT^{+}24]. Use [LvBS^{+}22, BCTH20, BSB^{+}20, BMS^{+}22, CLA^{+}24, FDHS20, GZX^{+}22, HPCH24, IFU^{+}20, MO23, MGI^{+}22, RTM^{+}20, SBB^{+}21, VCT^{+}23]. use/land [MGI^{+}22]. uses [SML21b]. Using [FMT^{+}20, GF21, BGAD^{+}22, CDW^{+}24, CZK^{+}23, DFJ22, DMS^{+}21, FHR^{+}21, FHMK20, IKN^{+}21, KFHS24, LZG21, MGWS23, RPA^{+}22, SJH20, TVB^{+}21, VMP^{+}21, WMTJW22, WDS^{+}23]. usual [AAC^{+}21]. utility [RWL^{+}24]. utilization [BAD^{+}21, LD22, MBLA^{+}21, PDK^{+}23, RCH^{+}20, SB20].

Val [PSE^{+}20]. Validation [LZG21]. valuable [ZR23]. Vanda [SSP^{+}20]. Variability [AR20, DRA^{+}23, KPG^{+}23, LTW^{+}24, TKS^{+}20, ASL^{+}24, BTK^{+}21, BBH^{+}20, CCB^{+}20, CMCC24, CHS^{+}20, CP22, DvOH^{+}20, DLDF21, DAD^{+}21, FJP^{+}21, HBG^{+}22, HMP^{+}21, JLW^{+}23, KX22, KOR^{+}20, KDV^{+}23, LZD^{+}23, LHD^{+}20, LMSN23, MMF^{+}21, MHP^{+}23, MDIY23, PW22, PMRG20, PFB^{+}22, PSK21, RHA^{+}23, RFW^{+}24, RSP^{+}24, RKBT^{+}23, SBM^{+}21, SML^{+}21a, SSL22, SMP21, SMKS22, SWK^{+}23, SSCE20, VVMJ23, XDQ^{+}22]. variable [BM22, BJG^{+}22, GSP^{+}23, RBD22, SSL22]. variables [CSD^{+}24]. Variation [HHMH22, KKA22, Ano21c, BGM^{+}22, CGRGP^{+}21, DMM^{+}21, DKW^{+}21, Edm21a, Edm21b, GGHBS^{+}20, HLH^{+}23, HZC^{+}24, LLSQ24, LHS^{+}20, LRRP20, MTP^{+}23, MV22a, MV22b, MRH^{+}23, MHVC22, OSR20, PGEB22, RPKC22, SPR22, hTRH20, WRWPG19, ZZB^{+}20]. variations [DHNH22, IJSC20, KTWGT20, SMNT22, WZCK20]. varies [GCW^{+}23, LLBC20, RVS^{+}22]. vary [SZA^{+}21]. varying [Phi20, Phi21]. vegetated [EJRRC^{+}23]. Vegetation [ZvBZ^{+}20, JvKvT^{+}20, PGB24, SVBT20, VMT^{+}20]. vehicle [LDG^{+}21]. velocities [EAF^{+}22]. velocity [GLN^{+}20, LCMG^{+}20, RSP^{+}24]. vent [ASL^{+}24, MGP^{+}21]. ventilation [KdED^{+}20]. vents [CTM^{+}21]. Vertical [BBM21, BO20, CYPRG^{+}23, HCH^{+}24, SSP^{+}20, BJS23a, BBGM20, GLN^{+}21, GNO22, GK22, HPP^{+}21, LLBC20, MEQBV20, MO23, WF22, WO21, WSZ22a, WSZ22b, WKRL^{+}23]. Very [WXK^{+}22, CTH^{+}20]. via [BPG^{+}22, GGLH^{+}22, SCC^{+}21]. viability [GLAHH^{+}23]. viral [SML^{+}21a, ZLJ^{+}24]. virginica [MDWT^{+}22, WSB^{+}20]. virus [HPF^{+}23, TVB^{+}21, XKP^{+}22]. virus-to-copepod [TVB^{+}21]. viruses [MLS^{+}21]. Visual [LAFV23, BBM21, LN21]. Vital [Edm21b]. volcanic [GRJ^{+}22, MHVC22]. volume [GGR24, TION23]. volumes [HDD^{+}22]. vs [BFW^{+}22, CPB^{+}21, LLBC20, SSL22, WTvD^{+}22]. Vulnerability [LRM21].

WA [MVBG20]. Wadden [HDD^{+}22]. Warm [MRC^{+}24]. Warm-adapted [MRC^{+}24]. Warming [BMB22, BDMGH^{+}23, OJUN23, RLL^{+}21, SVC^{+}22, ARB^{+}22, BTK^{+}21, DNB^{+}20, DMCW23, FLH^{+}23, MLP21b, ONH^{+}20, RBD22, RGW^{+}23, SBH23, SIP21, SMP21, SSEO21, SGM^{+}21, VKB^{+}22, WWB^{+}20, dJGCS23, vdPMF^{+}21]. wastewater [BHSL21]. wasting [ARB^{+}22]. Water [DLC^{+}22, DBFC24, JvKvT^{+}20, LWC^{+}24, MHVC22, SSS^{+}24, VLP^{+}21, ACW^{+}22, AWB^{+}23, BSGA21, BM22, BMB22, BRB^{+}20, BCTH20, BGG^{+}20, BFW^{+}22, CBB^{+}22, CCB^{+}20, CSC^{+}21, DAN^{+}22, DKPS23, EBK^{+}20, FBK^{+}22, FBRG23, GGS^{+}21, GWS^{+}21, GBHF21, GR24, GMI^{+}21, GKW^{+}20, GKW^{+}21, HSB^{+}20, HMB^{+}22, HSP^{+}20, IOS24, LLC^{+}23, LHS^{+}21, LDS^{+}20, LBG^{+}22, LZM^{+}21, MBA^{+}21, MFM^{+}21, MACC24, MLC^{+}21, MNMJ^{+}21, PHB^{+}22, PSY^{+}22, RCMA21, SJH20, SPBK21, SGGB21, SWK^{+}23, SWS^{+}22, SGSD^{+}22, SGK^{+}21, TEK^{+}21, TZF^{+}22, TRS^{+}20, XLN21, ZZO^{+}21, vGDH^{+}21, SMP21]. water-column [PSY^{+}22]. water-level [FBK^{+}22]. water-quality [PHB^{+}22]. waterbodies [BCTH20, OBL20]. waterflea [ZJ22a]. waters [AHCJ^{+}20, AFM23, CLGH23, CKCB^{+}23, CLD^{+}23, CPBJR^{+}21, CSNS24, FBRG23, FWDY20, GPK^{+}21, IJSC20, JLW^{+}23, KTH^{+}21, LSW^{+}20a, LvBS^{+}22, LZD^{+}23, LMS^{+}21, MCS^{+}23, MKD22, SSM^{+}21a, SSP^{+}20, THVJ23, TCP^{+}22, TMM^{+}20, YML^{+}21, vLWV^{+}20]. Watershed [JSPS24, MBH^{+}23]. watersheds [BSF^{+}21, PGEB22]. Wave [SN22, GLN^{+}20, LHH24, MXL^{+}21, MIR^{+}22, SRN^{+}21, VGvdB^{+}21, ZLN21]. wave-current [SN22]. wave-forcing [SRN^{+}21]. wave-powered [MXL^{+}21]. wavelet [SJH20]. waves [EJRRC^{+}23, FWDY20, FGO^{+}20, GLN^{+}21, GFG^{+}22, PRS^{+}20, SRV^{+}20]. Weak [LMM^{+}21, WMRL22]. weather [dCLO^{+}23]. weathering [PYY^{+}20, ZPK^{+}22]. web [BHR^{+}23, BOC^{+}21, CDH^{+}21, CGL^{+}21, EG20, FSM^{+}21, FLW^{+}21, GGS^{+}21, HCH^{+}24, IFU^{+}20, JSPS24, LCM^{+}22, MMF^{+}21, MLC^{+}21, PSNS22, SWC^{+}23, VCT^{+}23, WRC^{+}21, YAP^{+}22]. webs [BR22, CDH^{+}21, CESS21, FWT^{+}20, GM22, KPL^{+}20, LTQS^{+}21, LCHB22, LEB21, WHZ^{+}21]. wedge [WLK^{+}21]. weight [BGM^{+}21]. well [BD22]. were [SVMM^{+}22]. West [BBL^{+}21, CFC^{+}20, NGSS23, SGSF20]. Western [MN20, CCA^{+}21, CWF^{+}22, FIH^{+}20, HBG^{+}22, IKN^{+}21, JSL^{+}23, JWAvD^{+}22, KBvdD^{+}20, KSH^{+}23, LEK^{+}22, LNC^{+}23, MSG^{+}22, SOO^{+}21, SBW^{+}20, SSR^{+}23, WM20, XYZ^{+}21, YWZZ20, ZZO^{+}21, BFE^{+}20, BRB^{+}22, LLK^{+}22]. wet [YAE^{+}23]. wetland [LRM21, NEB^{+}22, VJS^{+}20, WHZ^{+}21]. wetlands [PYY^{+}20, ZJW^{+}22, ZSZ^{+}22]. Where [CHS^{+}20]. while [AFH^{+}21]. Whole [SPBK21, PBC21]. Whole-lake [SPBK21]. wide [KHCG^{+}22, SORGC^{+}21]. widespread [LCS^{+}20, OMM21, VSB^{+}23]. width [SGM^{+}22]. Wind [FBRG23, dSNIB22, HPP^{+}21, ILPL20, MAG^{+}22, SLZ^{+}21b, SCM22, TKM^{+}22]. Wind-driven [FBRG23, HPP^{+}21, SLZ^{+}21b]. wind-induced [SCM22]. window [RBR^{+}24]. windows [OJUN23]. winds [RHM^{+}21, SLJ^{+}22, ZSZ^{+}22]. Winter [CB21a, AAC^{+}21, GB23, HS21, JHvL^{+}23, KPG^{+}23, KTH^{+}21, MBH^{+}23, PCJ^{+}21, SIP21, SO23, YWLY20]. wintertime [APB^{+}23]. within [BBH^{+}20, BR22, CCFF21, CCC^{+}21, EWGP22, GLB^{+}22, HZC^{+}24, HDD^{+}22, LTH21, LTW^{+}24, MFM^{+}21, MKP^{+}23, RBAC^{+}20, YGD^{+}21, ZLIP20]. within-lake [BR22]. wooden [HMM^{+}22]. world [ZZCD23]. worldwide [QFM^{+}21]. would [SVBT20].

X [IKN^{+}21]. X-ray [IKN^{+}21].

Yakutia [HABL^{+}21]. year [EDC24, FR23, SO23]. year-long [EDC24]. years [FJP^{+}21, HSB^{+}20, MFM^{+}21]. Yellow [MLL^{+}21]. yield [BMCS23, RKBT^{+}23]. yields [APB^{+}23]. Yucatan [BFCC^{+}21]. Yukon [CM21].

Zealand [LNR^{+}20, YWZ^{+}21]. Zehr [MBD^{+}23]. zinc [KMV^{+}20]. zinc/cobalt [KMV^{+}20]. Zn [KMV^{+}20, KMM^{+}22, LSL^{+}20]. Zn/Co [KMM^{+}22]. zone [AVS^{+}22, BBM21, BYB^{+}23, BO20, BZK^{+}22, DNA^{+}22, EBK^{+}20, KCT23, LRM21, MSL21, PLLM21, QBJG21, SMW^{+}21, SWB^{+}20, SCB^{+}20, THL^{+}21, VMP^{+}21, WCC20, YGD^{+}21]. zones [CCK^{+}21, CBB^{+}22, KWS^{+}21, LZM^{+}21, MTP^{+}23, WRA^{+}23]. zoobenthic [BD22, HZC^{+}24]. Zooglider [WO21]. Zooglidermeasured [WO21]. zooplankter [ONH^{+}20]. Zooplankton [CLG^{+}24, DVL^{+}23, KHB22, CDH^{+}21, CBN21, DMS^{+}21, ETG21, GBC^{+}20, HCB^{+}22, HS21, KLG^{+}23, KGG^{+}23, LWS^{+}22, LTJ^{+}20, LDB^{+}20, MVP^{+}20, MWV22, MO23, MSB^{+}20, MN20, PGEB22, PPB^{+}21, SO23, SWC^{+}23, TGE^{+}21, TSX^{+}23, THS21, VKK^{+}21, WO21, YAP^{+}22, ZJ22b, ZZCD23]. Zooplankton-derived [DVL^{+}23]. Zostera [GGS^{+}21, SIP21].

References

Adelson:2022:SHT

[AAB^{+}22] Anne E. Adelson, Andrew H. Altieri, Ximena Boza, Rachel Collin, Kristen A. Davis, Alan Gaul, Sarah N. Giddings, Victoria Reed, and Geno Pawlak. Seasonal hypoxia and temperature inversions in a tropical bay. Limnology and Oceanography, 67(10):2174–2189, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Agawin:2021:RSP

[AAC^{+}21] Nona S. R. Agawin, Ma. Guadalupe Gil Atorrasagasti, Aida Frank Comas, Víctor Fernández-Juárez, Xabier López-Alforja, and Iris E. Hendriks. Response of the seagrass Posidonia oceanica and its associated N_{2} fixers to high businessas-usual climate change scenario in winter. Limnology and Oceanography, 66(6):2346–2359, June 2021. CODEN LI-OCAH. ISSN 0024-3590.

Aichelman:2021:EDM

[ABC^{+}21] Hannah E. Aichelman, Colleen B. Bove, Karl D. Castillo, Jessica M. Boulton, Alyssa C. Knowlton, Olivia C. Nieves, Justin B. Ries, and Sarah W. Davies. Exposure duration modulates the response of Caribbean corals to global change stressors. Limnology and Oceanography, 66(8):3100–3115, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Albert:2021:ISO

[ABS^{+}21] Séréna Albert, Stefano Bonaglia, Nellie Stjärnkvist, Monika Winder, Bo Thamdrup, and Francisco J. A. Nascimento. Influence of settling organic matter quantity and quality on benthic nitrogen cycling. Limnology and Oceanography, 66 (5):1882–1895, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Alejandra:2022:SDA

[ACA^{+}22] Peña Sanchez Alejandra, Duffner Clara, Wunderlich Anja, Mayer Bernhard, Schulz Stefanie, Michael Schloter, and Einsiedl Florian. Seasonal dynamics of anaerobic oxidation of ammonium and denitrification in a dimictic lake during the stratified spring–summer period. Limnology and Oceanography, 67(5):1194–1210, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Abdala:2022:EES

[ACE^{+}22] Zuzanna M. Abdala, Sophie Clayton, Sveinn V. Einarsson, Kimberly Powell, Claire P. Till, Tyler H. Coale, and P. Dreux Chappell. Examining ecological succession of diatoms in California Current System cyclonic mesoscale eddies. Limnology and Oceanography, 67(11):2586–2602, November 2022. CO-DEN LIOCAH. ISSN 0024-3590.

Akhand:2022:DIC

[ACW^{+}22] Anirban Akhand, Abhra Chanda, Kenta Watanabe, Sourav Das, Tatsuki Tokoro, Sugata Hazra, and Tomohiro Kuwae. Drivers of inorganic carbon dynamics and air–water CO_{2} fluxes in two large tropical estuaries: Insights from coupled radon (^{222}Rn) and pCO_{2} surveys. Limnology and Oceanography, 67(S2):S118–S132, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Allen:2022:ODP

[ADLW22] James G. Allen, Mathilde Dugenne, Ricardo M. Letelier, and Angelicque E. White. Optical determinations of photophysiology along an ecological gradient in the North Pacific Ocean. Limnology and Oceanography, 67(3):713–725, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Aho:2021:DCD

[AFH^{+}21] Kelly S. Aho, Jennifer H. Fair, Jacob D. Hosen, Ethan D. Kyzivat, Laura A. Logozzo, Gerard Rocher-Ros, Lisa C. Weber, Byungman Yoon, and Peter A. Raymond. Distinct concentration-discharge dynamics in temperate streams and rivers: CO_{2} exhibits chemostasis while CH_{4} exhibits source limitation due to temperature control. Limnology and Oceanography, 66(10):3656–3668, October 2021. CODEN LI-OCAH. ISSN 0024-3590.

Aho:2022:IPE

[AFH^{+}22] Kelly S. Aho, Jennifer H. Fair, Jake D. Hosen, Ethan D. Kyzivat, Laura A. Logozzo, Lisa C. Weber, Byungman Yoon, Jay P. Zarnetske, and Peter A. Raymond. An intense precipitation event causes a temperate forested drainage network to shift from N_{2}O source to sink. Limnology and Oceanography, 67(S1):S242–S257, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Anderson:2022:IET

[AFK^{+}22] Stephanie I. Anderson, Gayantonia Franzè, Joshua D. Kling, Paul Wilburn, Colin T. Kremer, Susanne Menden-Deuer, Elena Litchman, David A. Hutchins, and Tatiana A. Rynearson. The interactive effects of temperature and nutrients on a spring phytoplankton community. Limnology and Oceanography, 67(3):634–645, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Arlinghaus:2023:ESD

[AFM23] Kandis Arlinghaus, Amanda A. Frossard, and William L. Miller. Examining superoxide dynamics in irradiated natural waters. Limnology and Oceanography, 68(4):878–890, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Ashford:2021:CEC

[AGC^{+}21] Oliver S. Ashford, Shuzhe Guan, Dante Capone, Katherine Rigney, Katelynn Rowley, Victoria Orphan, Sean W. Mullin, Kat S. Dawson, Jorge Cortés, Greg W. Rouse, Guillermo F. Mendoza, Raymond W. Lee, Erik E. Cordes, and Lisa A. Levin. A chemosynthetic ecotone — “chemotone” — in the sediments surrounding deep-sea methane seeps. Limnology and Oceanography, 66(5):1687–1702, May 2021. CODEN LI-OCAH. ISSN 0024-3590.

Anderson:2020:MCI

[AH20] Ruth Anderson and Per Juel Hansen. Meteorological conditions induce strong shifts in mixotrophic and heterotrophic flagellate bacterivory over small spatio-temporal scales. Limnology and Oceanography, 65(6):1189–1199, June 2020. CO-DEN LIOCAH. ISSN 0024-3590.

Al-Haj:2022:TTS

[AHCF22] Alia N. Al-Haj, Tyler Chidsey, and Robinson W. Fulweiler. Two temperate seagrass meadows are negligible sources of methane and nitrous oxide. Limnology and Oceanography, 67 (S2):S193–S207, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Aldunate:2020:NAP

[AHCJ^{+}20] Montserrat Aldunate, Carlos Henríquez-Castillo, Qixing Ji, Jessica Lueders-Dumont, Margaret R. Mulholland, Bess B. Ward, Peter von Dassow, and Osvaldo Ulloa. Nitrogen assimilation in picocyanobacteria inhabiting the oxygen-deficient waters of the eastern tropical North and South Pacific. Limnology and Oceanography, 65(2):437–453, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Austin:2022:CTS

[AHF^{+}22] Jay Austin, Craig Hill, Jacob Fredrickson, Grace Weber, and Kaelan Weiss. Characterizing temporal and spatial scales of radiatively driven convection in a deep, ice-free lake. Limnology and Oceanography, 67(10):2296–2308, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Anderson:2021:ASP

[AJMG21] Hanna S. Anderson, Thomas H. Johengen, Russ Miller, and Casey M. Godwin. Accelerated sediment phosphorus release in Lake Erie’s central basin during seasonal anoxia. Limnology and Oceanography, 66(9):3582–3595, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Allesson:2021:RPC

[AKT^{+}21] Lina Allesson, Birgit Koehler, Jan-Erik Thrane, Tom Andersen, and Dag O. Hessen. The role of photomineralization for CO_{2} emissions in boreal lakes along a gradient of dissolved organic matter. Limnology and Oceanography, 66(1): 158–170, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Ababou:2023:MUD

[ALG^{+}23] Fatima-Ezzahra Ababou, Frédéric A. C. Le Moigne, Olivier Grosso, Catherine Guigue, Sandra Nunige, Mercedes Camps, and Sophie Bonnet. Mechanistic understanding of diazotroph aggregation and sinking: “A rolling tank approach”. Limnology and Oceanography, 68(3):666–677, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Atkinson:2021:INS

[ALH^{+}21] Angus Atkinson, Martin K. S. Lilley, Andrew G. Hirst, Andrea J. McEvoy, Glen A. Tarran, Claire Widdicombe, Elaine S. Fileman, E. Malcolm S. Woodward, Katrin Schmidt, Tim J. Smyth, and Paul J. Somerfield. Increasing nutrient stress reduces the efficiency of energy transfer through planktonic size spectra. Limnology and Oceanography, 66(2):422–437, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Asmala:2021:IDO

[AMC21] Eero Asmala, Philippe Massicotte, and Jacob Carstensen. Identification of dissolved organic matter size components in freshwater and marine environments. Limnology and Oceanography, 65(3):1381–1393, April 2021. CODEN LI-OCAH. ISSN 0024-3590.

Aoki:2020:SRR

[AMO20] Lillian R. Aoki, Karen J. McGlathery, and Matthew P. J. Oreska. Seagrass restoration reestablishes the coastal nitrogen filter through enhanced burial. Limnology and Oceanography, 65(1):1–12, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:AP

[Ano20a] Anonymous. ASLO page. Limnology and Oceanography, 65 (S1):i–iii, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IICa

[Ano20b] Anonymous. Issue information — copyright page. Limnology and Oceanography, 65(1):iii, January 2020. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2020:IICb

[Ano20c] Anonymous. Issue information — copyright page. Limnology and Oceanography, 65(2):iii, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IICc

[Ano20d] Anonymous. Issue information — copyright page. Limnology and Oceanography, 65(3):iii, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IICd

[Ano20e] Anonymous. Issue information — copyright page. Limnology and Oceanography, 65(4):iii, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IICe

[Ano20f] Anonymous. Issue information — copyright page. Limnology and Oceanography, 65(5):iii, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IICf

[Ano20g] Anonymous. Issue information — copyright page. Limnology and Oceanography, 65(6):iii, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IICg

[Ano20h] Anonymous. Issue information — copyright page. Limnology and Oceanography, 65(7):iii, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IICh

[Ano20i] Anonymous. Issue information — copyright page. Limnology and Oceanography, 65(8):iii, August 2020. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2020:IICi

[Ano20j] Anonymous. Issue information — copyright page. Limnology and Oceanography, 65(9):ii, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IICj

[Ano20k] Anonymous. Issue information — copyright page. Limnology and Oceanography, 65(10):ii, October 2020. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2020:IICk

[Ano20l] Anonymous. Issue information — copyright page. Limnology and Oceanography, 65(11):ii, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IICl

[Ano20m] Anonymous. Issue information — copyright page. Limnology and Oceanography, 65(12):ii, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIIa

[Ano20n] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 65(1):vi, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIIb

[Ano20o] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 65(2):v, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIIc

[Ano20p] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 65(3):v, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIId

[Ano20q] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 65(4):v, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIIe

[Ano20r] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 65(5):v–vii, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIIf

[Ano20s] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 65(6):v–vii, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIIg

[Ano20t] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 65(7):v–vii, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIIh

[Ano20u] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 65(8):v–vii, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIIi

[Ano20v] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 65(9):v–vii, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIIj

[Ano20w] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 65(10):v–vii, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIIk

[Ano20x] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 65(11):v–vii, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIIl

[Ano20y] Anonymous. Issue information — instr to contrib. Limnology and Oceanography, 65(12):v–vii, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMb

[Ano20z] Anonymous. Issue information — masthead. Limnology and Oceanography, 65(1):ii, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMd

[Ano20-27] Anonymous. Issue information — masthead. Limnology and Oceanography, 65(2):ii, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMf

[Ano20-28] Anonymous. Issue information — masthead. Limnology and Oceanography, 65(3):ii, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMh

[Ano20-29] Anonymous. Issue information — masthead. Limnology and Oceanography, 65(4):ii, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMj

[Ano20-30] Anonymous. Issue information — masthead. Limnology and Oceanography, 65(5):ii, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMl

[Ano20-31] Anonymous. Issue information — masthead. Limnology and Oceanography, 65(6):ii, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMn

[Ano20-32] Anonymous. Issue information — masthead. Limnology and Oceanography, 65(7):ii, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMp

[Ano20-33] Anonymous. Issue information — masthead. Limnology and Oceanography, 65(8):ii, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMq

[Ano20-34] Anonymous. Issue information — masthead. Limnology and Oceanography, 65(9):i, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMs

[Ano20-35] Anonymous. Issue information — masthead. Limnology and Oceanography, 65(10):i, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMu

[Ano20-36] Anonymous. Issue information — masthead. Limnology and Oceanography, 65(11):i, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMw

[Ano20-37] Anonymous. Issue information — masthead. Limnology and Oceanography, 65(12):i, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMa

[Ano20-38] Anonymous. Issue information — members. Limnology and Oceanography, 65(1):i, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMc

[Ano20-39] Anonymous. Issue information — members. Limnology and Oceanography, 65(2):i, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMe

[Ano20-40] Anonymous. Issue information — members. Limnology and Oceanography, 65(3):i, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMg

[Ano20-41] Anonymous. Issue information — members. Limnology and Oceanography, 65(4):i, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMi

[Ano20-42] Anonymous. Issue information — members. Limnology and Oceanography, 65(5):i, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMk

[Ano20-43] Anonymous. Issue information — members. Limnology and Oceanography, 65(6):i, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMm

[Ano20-44] Anonymous. Issue information — members. Limnology and Oceanography, 65(7):i, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMo

[Ano20-45] Anonymous. Issue information — members. Limnology and Oceanography, 65(8):i, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMr

[Ano20-46] Anonymous. Issue information — members. Limnology and Oceanography, 65(9):iv, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMt

[Ano20-47] Anonymous. Issue information — members. Limnology and Oceanography, 65(10):iv, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMv

[Ano20-48] Anonymous. Issue information — members. Limnology and Oceanography, 65(11):iv, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IIMx

[Ano20-49] Anonymous. Issue information — members. Limnology and Oceanography, 65(12):iv, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IITa

[Ano20-50] Anonymous. Issue information — TOC. Limnology and Oceanography, 65(1):iv–v, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IITb

[Ano20-51] Anonymous. Issue information — TOC. Limnology and Oceanography, 65(2):iv, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IITc

[Ano20-52] Anonymous. Issue information — TOC. Limnology and Oceanography, 65(3):iv, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IITd

[Ano20-53] Anonymous. Issue information — TOC. Limnology and Oceanography, 65(4):iv, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IITe

[Ano20-54] Anonymous. Issue information — TOC. Limnology and Oceanography, 65(5):iv, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IITf

[Ano20-55] Anonymous. Issue information — TOC. Limnology and Oceanography, 65(6):iv, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IITg

[Ano20-56] Anonymous. Issue information — TOC. Limnology and Oceanography, 65(7):iv, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IITh

[Ano20-57] Anonymous. Issue information — TOC. Limnology and Oceanography, 65(8):iv, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IITi

[Ano20-58] Anonymous. Issue information — TOC. Limnology and Oceanography, 65(9):iii, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IITj

[Ano20-59] Anonymous. Issue information — TOC. Limnology and Oceanography, 65(10):iii, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IITk

[Ano20-60] Anonymous. Issue information — TOC. Limnology and Oceanography, 65(11):iii, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2020:IITl

[Ano20-61] Anonymous. Issue information — TOC. Limnology and Oceanography, 65(12):iii, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:CGL

[Ano21a] Anonymous. Corrigendum for Giling et al. (2017). Delving deeper: Metabolic processes in the metalimnion of stratified lakes. Limnology and Oceanography, 62, 1288–1306 (doi: 10.1002/lno.10504). Limnology and Oceanography, 66(5): 2088–2092, May 2021. CODEN LIOCAH. ISSN 0024-3590. See [GSG^{+}17].

Anonymous:2021:CMK

[Ano21b] Anonymous. Corrigendum for Meyer-Kaiser et al. 2019 [Limnol. Oceanogr. 64(5) 1924–1938]. Limnology and Oceanography, 66(8):3253, August 2021. CODEN LIOCAH. ISSN 0024-3590. See [MKBSK19].

Anonymous:2021:CSV

[Ano21c] Anonymous. Corrigendum: Spatial variation in the biochemical and isotopic composition of corals during bleaching and recovery. Limnology and Oceanography, 65(3):1611– 1612, April 2021. CODEN LIOCAH. ISSN 0024-3590. See [WRWPG19].

Anonymous:2021:E

[Ano21d] Anonymous. Erratum. Limnology and Oceanography, 66(10): 3856, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IICa

[Ano21e] Anonymous. Issue information — copyright page. Limnology and Oceanography, 66(1):ii, January 2021. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2021:IICb

[Ano21f] Anonymous. Issue information — copyright page. Limnology and Oceanography, 66(2):ii, February 2021. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2021:IICc

[Ano21g] Anonymous. Issue information — copyright page. Limnology and Oceanography, 66(3):ii, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IICd

[Ano21h] Anonymous. Issue information — copyright page. Limnology and Oceanography, 66(4):ii, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IICe

[Ano21i] Anonymous. Issue information — copyright page. Limnology and Oceanography, 66(5):ii, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IICf

[Ano21j] Anonymous. Issue information — copyright page. Limnology and Oceanography, 66(6):ii, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IICg

[Ano21k] Anonymous. Issue information — copyright page. Limnology and Oceanography, 66(7):ii, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IICh

[Ano21l] Anonymous. Issue information — copyright page. Limnology and Oceanography, 66(8):ii, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IICi

[Ano21m] Anonymous. Issue information — copyright page. Limnology and Oceanography, 66(9):ii, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IICj

[Ano21n] Anonymous. Issue information — copyright page. Limnology and Oceanography, 66(10):ii, October 2021. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2021:IICk

[Ano21o] Anonymous. Issue information — copyright page. Limnology and Oceanography, 66(11):ii, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IICm

[Ano21p] Anonymous. Issue information — copyright page. Limnology and Oceanography, 66(S1):ii, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMa

[Ano21q] Anonymous. Issue information — masthead. Limnology and Oceanography, 66(1):i, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMc

[Ano21r] Anonymous. Issue information — masthead. Limnology and Oceanography, 66(2):i, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMe

[Ano21s] Anonymous. Issue information — masthead. Limnology and Oceanography, 66(3):i, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMg

[Ano21t] Anonymous. Issue information — masthead. Limnology and Oceanography, 66(4):i, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMi

[Ano21u] Anonymous. Issue information — masthead. Limnology and Oceanography, 66(5):i, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMk

[Ano21v] Anonymous. Issue information — masthead. Limnology and Oceanography, 66(6):i, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMm

[Ano21w] Anonymous. Issue information — masthead. Limnology and Oceanography, 66(7):i, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMo

[Ano21x] Anonymous. Issue information — masthead. Limnology and Oceanography, 66(8):i, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMq

[Ano21y] Anonymous. Issue information — masthead. Limnology and Oceanography, 66(9):i, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMs

[Ano21z] Anonymous. Issue information — masthead. Limnology and Oceanography, 66(10):i, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMu

[Ano21-27] Anonymous. Issue information — masthead. Limnology and Oceanography, 66(11):i, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMy

[Ano21-28] Anonymous. Issue information — masthead. Limnology and Oceanography, 66(S1):i, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMb

[Ano21-29] Anonymous. Issue information — members. Limnology and Oceanography, 66(1):iv, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMd

[Ano21-30] Anonymous. Issue information — members. Limnology and Oceanography, 66(2):iv, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMf

[Ano21-31] Anonymous. Issue information — members. Limnology and Oceanography, 66(3):v, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMh

[Ano21-32] Anonymous. Issue information — members. Limnology and Oceanography, 66(4):v, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMj

[Ano21-33] Anonymous. Issue information — members. Limnology and Oceanography, 66(5):v, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMl

[Ano21-34] Anonymous. Issue information — members. Limnology and Oceanography, 66(6):iv, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMn

[Ano21-35] Anonymous. Issue information — members. Limnology and Oceanography, 66(7):iv, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMp

[Ano21-36] Anonymous. Issue information — members. Limnology and Oceanography, 66(8):iv, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMr

[Ano21-37] Anonymous. Issue information — members. Limnology and Oceanography, 66(9):v, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMt

[Ano21-38] Anonymous. Issue information — members. Limnology and Oceanography, 66(10):v, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMv

[Ano21-39] Anonymous. Issue information — members. Limnology and Oceanography, 66(11):iv, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMz

[Ano21-40] Anonymous. Issue information — members. Limnology and Oceanography, 66(S1):iv, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IITa

[Ano21-41] Anonymous. Issue information — TOC. Limnology and Oceanography, 66(1):iii, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IITb

[Ano21-42] Anonymous. Issue information — TOC. Limnology and Oceanography, 66(2):iii, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IITc

[Ano21-43] Anonymous. Issue information — TOC. Limnology and Oceanography, 66(3):iii–iv, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IITd

[Ano21-44] Anonymous. Issue information — TOC. Limnology and Oceanography, 66(4):iii–iv, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IITe

[Ano21-45] Anonymous. Issue information — TOC. Limnology and Oceanography, 66(5):iii–iv, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IITf

[Ano21-46] Anonymous. Issue information — TOC. Limnology and Oceanography, 66(6):iii, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IITg

[Ano21-47] Anonymous. Issue information — TOC. Limnology and Oceanography, 66(7):iii, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IITh

[Ano21-48] Anonymous. Issue information — TOC. Limnology and Oceanography, 66(8):iii, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IITi

[Ano21-49] Anonymous. Issue information — TOC. Limnology and Oceanography, 66(9):iii–iv, September 2021. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2021:IITj

[Ano21-50] Anonymous. Issue information — TOC. Limnology and Oceanography, 66(10):iii–iv, October 2021. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2021:IITk

[Ano21-51] Anonymous. Issue information — TOC. Limnology and Oceanography, 66(11):iii, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IITm

[Ano21-52] Anonymous. Issue information — TOC. Limnology and Oceanography, 66(S1):iii, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IICl

[Ano21-53] Anonymous. Issue information & copyright. Limnology and Oceanography, 66(12):ii, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMw

[Ano21-54] Anonymous. Issue information & masthead. Limnology and Oceanography, 66(12):i, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IIMx

[Ano21-55] Anonymous. Issue information & members. Limnology and Oceanography, 66(12):v, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2021:IITl

[Ano21-56] Anonymous. Issue information & TOC. Limnology and Oceanography, 66(12):iii–iv, December 2021. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2022:IICm

[Ano22a] Anonymous. Issue information — copyright page. Limnology and Oceanography, 67(S1):ii, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMy

[Ano22b] Anonymous. Issue information — masthead. Limnology and Oceanography, 67(S1):i, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMz

[Ano22c] Anonymous. Issue information — members. Limnology and Oceanography, 67(S1):v, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITm

[Ano22d] Anonymous. Issue information — TOC. Limnology and Oceanography, 67(S1):iii–iv, February 2022. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:2022:IICa

[Ano22e] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(1):ii, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IICb

[Ano22f] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(2):ii, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IICc

[Ano22g] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(3):ii, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IICd

[Ano22h] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(4):ii, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IICe

[Ano22i] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(5):ii, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IICf

[Ano22j] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(6):ii, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IICg

[Ano22k] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(7):ii, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IICh

[Ano22l] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(8):ii, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IICi

[Ano22m] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(9):ii, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IICj

[Ano22n] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(10):ii, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IICk

[Ano22o] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(11):ii, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IICl

[Ano22p] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(12):ii, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IICn

[Ano22q] Anonymous. Issue information & copyright. Limnology and Oceanography, 67(S2):ii, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMa

[Ano22r] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(1):i, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMc

[Ano22s] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(2):i, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMe

[Ano22t] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(3):i, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMg

[Ano22u] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(4):i, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMi

[Ano22v] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(5):i, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMk

[Ano22w] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(6):i, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMm

[Ano22x] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(7):i, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMo

[Ano22y] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(8):i, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMq

[Ano22z] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(9):i, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMs

[Ano22-27] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(10):i, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMu

[Ano22-28] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(11):i, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMw

[Ano22-29] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(12):i, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMba

[Ano22-30] Anonymous. Issue information & masthead. Limnology and Oceanography, 67(S2):i, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMb

[Ano22-31] Anonymous. Issue information & members. Limnology and Oceanography, 67(1):iv, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMd

[Ano22-32] Anonymous. Issue information & members. Limnology and Oceanography, 67(2):iv, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMf

[Ano22-33] Anonymous. Issue information & members. Limnology and Oceanography, 67(3):iv, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMh

[Ano22-34] Anonymous. Issue information & members. Limnology and Oceanography, 67(4):iv, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMj

[Ano22-35] Anonymous. Issue information & members. Limnology and Oceanography, 67(5):iv, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMl

[Ano22-36] Anonymous. Issue information & members. Limnology and Oceanography, 67(6):iv, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMn

[Ano22-37] Anonymous. Issue information & members. Limnology and Oceanography, 67(7):iv, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMp

[Ano22-38] Anonymous. Issue information & members. Limnology and Oceanography, 67(8):iv, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMr

[Ano22-39] Anonymous. Issue information & members. Limnology and Oceanography, 67(9):iv, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMt

[Ano22-40] Anonymous. Issue information & members. Limnology and Oceanography, 67(10):iv, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMv

[Ano22-41] Anonymous. Issue information & members. Limnology and Oceanography, 67(11):iv, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMx

[Ano22-42] Anonymous. Issue information & members. Limnology and Oceanography, 67(12):iv, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IIMbb

[Ano22-43] Anonymous. Issue information & members. Limnology and Oceanography, 67(S2):iv, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITa

[Ano22-44] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(1):iii, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITb

[Ano22-45] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(2):iii, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITc

[Ano22-46] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(3):iii, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITd

[Ano22-47] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(4):iii, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITe

[Ano22-48] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(5):iii, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITf

[Ano22-49] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(6):iii, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITg

[Ano22-50] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(7):iii, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITh

[Ano22-51] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(8):iii, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITi

[Ano22-52] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(9):iii, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITj

[Ano22-53] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(10):iii, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITk

[Ano22-54] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(11):iii, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITl

[Ano22-55] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(12):iii, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2022:IITn

[Ano22-56] Anonymous. Issue information & TOC. Limnology and Oceanography, 67(S2):iii, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IICa

[Ano23a] Anonymous. Issue information & copyright. Limnology and Oceanography, 68(1):ii, January 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IICb

[Ano23b] Anonymous. Issue information & copyright. Limnology and Oceanography, 68(2):ii, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IICc

[Ano23c] Anonymous. Issue information & copyright. Limnology and Oceanography, 68(3):ii, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IICd

[Ano23d] Anonymous. Issue information & copyright. Limnology and Oceanography, 68(4):ii, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IICf

[Ano23e] Anonymous. Issue information & copyright. Limnology and Oceanography, 68(5):ii, May 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IICg

[Ano23f] Anonymous. Issue information & copyright. Limnology and Oceanography, 68(6):ii, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IICh

[Ano23g] Anonymous. Issue information & copyright. Limnology and Oceanography, 68(7):ii, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IICi

[Ano23h] Anonymous. Issue information & copyright. Limnology and Oceanography, 68(8):ii, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IICj

[Ano23i] Anonymous. Issue information & copyright. Limnology and Oceanography, 68(9):ii, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IICk

[Ano23j] Anonymous. Issue information & copyright. Limnology and Oceanography, 68(10):ii, October 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IICl

[Ano23k] Anonymous. Issue information & copyright. Limnology and Oceanography, 68(11):ii, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IICm

[Ano23l] Anonymous. Issue information & copyright. Limnology and Oceanography, 68(12):ii, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMa

[Ano23m] Anonymous. Issue information & masthead. Limnology and Oceanography, 68(1):i, January 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMc

[Ano23n] Anonymous. Issue information & masthead. Limnology and Oceanography, 68(2):i, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMe

[Ano23o] Anonymous. Issue information & masthead. Limnology and Oceanography, 68(3):i, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMg

[Ano23p] Anonymous. Issue information & masthead. Limnology and Oceanography, 68(4):i, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMi

[Ano23q] Anonymous. Issue information & masthead. Limnology and Oceanography, 68(5):i, May 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMk

[Ano23r] Anonymous. Issue information & masthead. Limnology and Oceanography, 68(6):i, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMm

[Ano23s] Anonymous. Issue information & masthead. Limnology and Oceanography, 68(7):i, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMo

[Ano23t] Anonymous. Issue information & masthead. Limnology and Oceanography, 68(8):i, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMq

[Ano23u] Anonymous. Issue information & masthead. Limnology and Oceanography, 68(9):i, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMs

[Ano23v] Anonymous. Issue information & masthead. Limnology and Oceanography, 68(10):i, October 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMu

[Ano23w] Anonymous. Issue information & masthead. Limnology and Oceanography, 68(11):i, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMw

[Ano23x] Anonymous. Issue information & masthead. Limnology and Oceanography, 68(12):i, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMb

[Ano23y] Anonymous. Issue information & members. Limnology and Oceanography, 68(1):iv, January 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMd

[Ano23z] Anonymous. Issue information & members. Limnology and Oceanography, 68(2):iv, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMf

[Ano23-27] Anonymous. Issue information & members. Limnology and Oceanography, 68(3):iv, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMh

[Ano23-28] Anonymous. Issue information & members. Limnology and Oceanography, 68(4):iv, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMj

[Ano23-29] Anonymous. Issue information & members. Limnology and Oceanography, 68(5):iv, May 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMl

[Ano23-30] Anonymous. Issue information & members. Limnology and Oceanography, 68(6):iv, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMn

[Ano23-31] Anonymous. Issue information & members. Limnology and Oceanography, 68(7):iv, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMp

[Ano23-32] Anonymous. Issue information & members. Limnology and Oceanography, 68(8):iv, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMr

[Ano23-33] Anonymous. Issue information & members. Limnology and Oceanography, 68(9):iv, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMt

[Ano23-34] Anonymous. Issue information & members. Limnology and Oceanography, 68(10):iv, October 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMv

[Ano23-35] Anonymous. Issue information & members. Limnology and Oceanography, 68(11):iv, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IIMx

[Ano23-36] Anonymous. Issue information & members. Limnology and Oceanography, 68(12):iv, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IITa

[Ano23-37] Anonymous. Issue information & TOC. Limnology and Oceanography, 68(1):iii, January 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IITb

[Ano23-38] Anonymous. Issue information & TOC. Limnology and Oceanography, 68(2):iii, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IITc

[Ano23-39] Anonymous. Issue information & TOC. Limnology and Oceanography, 68(3):iii, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IITd

[Ano23-40] Anonymous. Issue information & TOC. Limnology and Oceanography, 68(4):iii, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IITe

[Ano23-41] Anonymous. Issue information & TOC. Limnology and Oceanography, 68(5):iii, May 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IITf

[Ano23-42] Anonymous. Issue information & TOC. Limnology and Oceanography, 68(6):iii, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IITg

[Ano23-43] Anonymous. Issue information & TOC. Limnology and Oceanography, 68(7):iii, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IITh

[Ano23-44] Anonymous. Issue information & TOC. Limnology and Oceanography, 68(8):iii, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IITi

[Ano23-45] Anonymous. Issue information & TOC. Limnology and Oceanography, 68(9):iii, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IITj

[Ano23-46] Anonymous. Issue information & TOC. Limnology and Oceanography, 68(10):iii, October 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IITk

[Ano23-47] Anonymous. Issue information & TOC. Limnology and Oceanography, 68(11):iii, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2023:IITl

[Ano23-48] Anonymous. Issue information & TOC. Limnology and Oceanography, 68(12):iii, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IICa

[Ano24a] Anonymous. Issue information & copyright. Limnology and Oceanography, 69(1):ii, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IICb

[Ano24b] Anonymous. Issue information & copyright. Limnology and Oceanography, 69(2):ii, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IICc

[Ano24c] Anonymous. Issue information & copyright. Limnology and Oceanography, 69(3):ii, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IICd

[Ano24d] Anonymous. Issue information & copyright. Limnology and Oceanography, 69(4):ii, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IIMa

[Ano24e] Anonymous. Issue information & masthead. Limnology and Oceanography, 69(1):i, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IIMc

[Ano24f] Anonymous. Issue information & masthead. Limnology and Oceanography, 69(2):i, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IIMe

[Ano24g] Anonymous. Issue information & masthead. Limnology and Oceanography, 69(3):i, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IIMg

[Ano24h] Anonymous. Issue information & masthead. Limnology and Oceanography, 69(4):i, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IIMb

[Ano24i] Anonymous. Issue information & members. Limnology and Oceanography, 69(1):iv, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IIMd

[Ano24j] Anonymous. Issue information & members. Limnology and Oceanography, 69(2):iv, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IIMf

[Ano24k] Anonymous. Issue information & members. Limnology and Oceanography, 69(3):iv, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IIMh

[Ano24l] Anonymous. Issue information & members. Limnology and Oceanography, 69(4):v, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IITa

[Ano24m] Anonymous. Issue information & TOC. Limnology and Oceanography, 69(1):iii, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IITb

[Ano24n] Anonymous. Issue information & TOC. Limnology and Oceanography, 69(2):iii, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IITc

[Ano24o] Anonymous. Issue information & TOC. Limnology and Oceanography, 69(3):iii, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:2024:IITd

[Ano24p] Anonymous. Issue information & TOC. Limnology and Oceanography, 69(4):iii–iv, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Akimova:2023:CMN

[APB^{+}23] Anna Akimova, Myron A. Peck, Gregor Börner, Cindy van Damme, and Marta Moyano. Combining modeling with novel field observations yields new insights into wintertime food limitation of larval fish. Limnology and Oceanography, 68 (8):1865–1879, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Anderson:2020:VAT

[AR20] Stephanie I. Anderson and Tatiana A. Rynearson. Variability approaching the thermal limits can drive diatom community dynamics. Limnology and Oceanography, 65(9):1961–1973, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Aoki:2022:DSA

[ARB^{+}22] Lillian R. Aoki, Brendan Rappazzo, Deanna S. Beatty, Lia K. Domke, Ginny L. Eckert, Morgan E. Eisenlord, Olivia J. Graham, Leah Harper, Timothy L. Hawthorne, Margot Hessing-Lewis, Kevin A. Hovel, Zachary L. Monteith, Ryan S. Mueller, Angeleen M. Olson, Carolyn Prentice, John J. Stachowicz, Fiona Tomas, Bo Yang, J. Emmett Duffy, Carla Gomes, and C. Drew Harvell. Disease surveillance by artificial intelligence links eelgrass wasting disease to ocean warming across latitudes. Limnology and Oceanography, 67(7): 1577–1589, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Anlanger:2021:HBC

[ARBvS^{+}21] Christine Anlanger, Ute Risse-Buhl, Daniel von Schiller, Christian Noss, Markus Weitere, and Andreas Lorke. Hydraulic and biological controls of biofilm nitrogen uptake in gravel-bed streams. Limnology and Oceanography, 66(11): 3887–3900, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Alcaraz-Rocha:2023:OAA

[ARPFGS23] Paulo Alcaraz-Rocha, Júlia Puig-Fàbregas, José Luis Garrido, and Cristina Sobrino. Ocean acidification affects pigment concentration and photoprotection of marine phytoplankton. Limnology and Oceanography, 68(4):831–844, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Atlas:2020:LBC

[ASH^{+}20] William I. Atlas, Daniel T. Selbie, Carrie A. Holt, Steve Cox-Rogers, Charmaine Carr-Harris, Kara J. Pitman, and Jonathan W. Moore. Landscape and biophysical controls of lake productivity to inform evaluation of sockeye salmon (Oncorhynchus nerka) populations in data-limited regions. Limnology and Oceanography, 66(5):2205–2219, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Audenhaege:2024:MED

[ASL^{+}24] Loïc Van Audenhaege, Jozée Sarrazin, Pierre Legendre, Garance Perrois, Mathilde Cannat, Aurélien Arnaubec, and Marjolaine Matabos. Monitoring ecological dynamics on complex hydrothermal structures: a novel photogrammetry approach reveals fine-scale variability of vent assemblages. Limnology and Oceanography, 69(2):325–338, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Ahonen:2024:API

[AVJ^{+}24] Salla A. Ahonen, Kristiina M. Vuorio, Roger I. Jones, Heikki Hämäläinen, Krista Rantamo, Marja Tiirola, and Anssi V. Vähätalo. Assessing and predicting the influence of chromophoric dissolved organic matter on light absorption by phytoplankton in boreal lakes. Limnology and Oceanography, 69(2):422–433, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Asmala:2022:RPD

[AVS^{+}22] Eero Asmala, Joonas J. Virtasalo, Matias Scheinin, Sara Newton, and Tom Jilbert. Role of particle dynamics in processing of terrestrial nitrogen and phosphorus in the estuarine mixing zone. Limnology and Oceanography, 67(1):1–12, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Alowaifeer:2023:AMS

[AWB^{+}23] Abdullah M. Alowaifeer, Qian Wang, Brian Bothner, Ryan J. Sibert, Samantha B. Joye, and Timothy R. McDermott. Aerobic methane synthesis and dynamics in a river water environment. Limnology and Oceanography, 68(8):1762–1774, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Andersen:2020:NAP

[AWGV20] Isabelle M. Andersen, Tanner J. Williamson, Maria J. González, and Michael J. Vanni. Nitrate, ammonium, and phosphorus drive seasonal nutrient limitation of chlorophytes, cyanobacteria, and diatoms in a hyper-eutrophic reservoir. Limnology and Oceanography, 65(5):962–978, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Bercovici:2021:LUE

[BAD^{+}21] Sarah K. Bercovici, Mar C. Arroyo, Daniele De Corte, Taichi Yokokawa, and Dennis A. Hansell. Limited utilization of extracted dissolved organic matter by prokaryotic communities from the subtropical North Atlantic. Limnology and Oceanography, 65(6):2509–2520, June 2021. CODEN LI-OCAH. ISSN 0024-3590.

Beaulieu:2024:OME

[BAG^{+}24] Marieke Beaulieu, Marc Amyot, Katherine Griffiths, Ashu Dastoor, Adam Jeziorski, and Irene Gregory-Eaves. Organic matter, eutrophication, and increased mass accumulation rates in the Anthropocene are the main drivers of mercury concentrations and historical enrichment in Canadian lake sediments. Limnology and Oceanography, 69(2):434–448, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Browning:2020:NRL

[BAHH^{+}20] Thomas J. Browning, Ali A. Al-Hashem, Mark J. Hopwood, Anja Engel, Ewan D. Wakefield, and Eric P. Achterberg. Nutrient regulation of late spring phytoplankton blooms in the midlatitude North Atlantic. Limnology and Oceanography, 65(6):1136–1148, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Burns:2023:IBT

[BBA^{+}23] Shannon M. Burns, Randelle M. Bundy, William Abbott, Zuzanna Abdala, Alexa R. Sterling, P. Dreux Chappell, Bethany D. Jenkins, and Kristen N. Buck. Interactions of bioactive trace metals in shipboard Southern Ocean incubation experiments. Limnology and Oceanography, 68(3):525– 543, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Brown:2021:HSB

[BBCA^{+}21] Kristen T. Brown, Dorothea Bender-Champ, Michelle Achlatis, Rene M. van der Zande, Andreas Kubicek, Storm B. Martin, Carolina Castro-Sanguino, Sophie G. Dove, and Ove Hoegh-Guldberg. Habitat-specific biogenic production and erosion influences net framework and sediment coral reef carbonate budgets. Limnology and Oceanography, 66(2):349– 365, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Brisbin:2020:PHT

[BBGM20] Margaret Mars Brisbin, Otis Davey Brunner, Mary Matilda Grossmann, and Satoshi Mitarai. Paired high-throughput, in situ imaging and high-throughput sequencing illuminate acantharian abundance and vertical distribution. Limnology and Oceanography, 66(6):2953–2965, December 2020. CO-DEN LIOCAH. ISSN 0024-3590.

Berberich:2020:SVS

[BBH^{+}20] Megan E. Berberich, Jake J. Beaulieu, Trinity L. Hamilton, Sarah Waldo, and Ishi Buffam. Spatial variability of sediment methane production and methanogen communities within a eutrophic reservoir: Importance of organic matter source and quantity. Limnology and Oceanography, 65(6): 1336–1358, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Brown:2021:LDK

[BBL^{+}21] Michael S. Brown, Jeff S. Bowman, Yajuan Lin, Colette J. Feehan, Carly M. Moreno, Nicolas Cassar, Adrian Marchetti, and Oscar M. Schofield. Low diversity of a key phytoplankton group along the West Antarctic Peninsula. Limnology and Oceanography, 66(6):2470–2480, June 2021. CODEN LI-OCAH. ISSN 0024-3590.

Benoit-Bird:2021:VMT

[BBM21] Kelly J. Benoit-Bird and Mark A. Moline. Vertical migration timing illuminates the importance of visual and nonvisual predation pressure in the mesopelagic zone. Limnology and Oceanography, 66(8):3010–3019, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Berger:2020:LTT

[BBMD20] Amelie C. Berger, Peter Berg, Karen J. McGlathery, and Marie Lise Delgard. Long-term trends and resilience of seagrass metabolism: a decadal aquatic eddy covariance study. Limnology and Oceanography, 65(7):1423–1438, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Bravo:2020:BFM

[BC20] Andrea G. Bravo and Claudia Cosio. Biotic formation of methylmercury: a bio–physico–chemical conundrum. Limnology and Oceanography, 65(5):1010–1027, May 2020. CO-DEN LIOCAH. ISSN 0024-3590.

Bowman:2020:DMC

[BCA^{+}20] Katlin L. Bowman, R. Eric Collins, Alison M. Agather, Carl H. Lamborg, Chad R. Hammerschmidt, Drishti Kaul, Christopher L. Dupont, Geoff A. Christensen, and Dwayne A. Elias. Distribution of mercury-cycling genes in the Arctic and equatorial Pacific Oceans and their relationship to mercury speciation. Limnology and Oceanography, 66(6):S310–S320, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Botana:2022:TPC

[BCFI^{+}22] Marina T. Botana, Adriano B. Chaves-Filho, Alex Inague, Arthur Z. Güth, Flavia Saldanha-Corrêa, Marius N. Müller, Paulo Y. G. Sumida, Sayuri Miyamoto, Matthias Y. Kellermann, Raymond C. Valentine, and Marcos Y. Yoshinaga. Thermal plasticity of coral reef symbionts is linked to major alterations in their lipidome composition. Limnology and Oceanography, 67(7):1456–1469, July 2022. CODEN LI-OCAH. ISSN 0024-3590.

Bernal:2022:SMS

[BCL^{+}22] Susana Bernal, Mathew J. Cohen, José L. J. Ledesma, Lily Kirk, Eugènia Martí, and Anna Lupon. Stream metabolism sources a large fraction of carbon dioxide to the atmosphere in two hydrologically contrasting headwater streams. Limnology and Oceanography, 67(12):2621–2634, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Beaulieu:2020:PAC

[BCTH20] Marieke Beaulieu, Hubert Cabana, Zofia Taranu, and Yannick Huot. Predicting atrazine concentrations in waterbodies across the contiguous United States: The importance of land use, hydrology, and water physicochemistry. Limnology and Oceanography, 66(6):2966–2983, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Bass:2023:DOC

[BCWS23] Adrian M. Bass, Martin Coleman, Susan Waldron, and Marian Scott. Dissolved organic carbon export in a small, disturbed peat catchment: Insights from long-term, high-resolution, sensor-based monitoring. Limnology and Oceanography, 68(8):1750–1761, August 2023. CODEN LI-OCAH. ISSN 0024-3590.

Beukema:2022:BWT

[BD22] Jan J. Beukema and Rob Dekker. Bottom-up as well as topdown processes govern zoobenthic secondary production in a tidal-flat ecosystem. Limnology and Oceanography, 67(11): 2547–2556, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Boyd:2022:TGC

[BDE^{+}22] Philip W. Boyd, Scott C. Doney, Sam Eggins, Michael J. Ellwood, Marion Fourquez, Brook L. Nunn, Robert Strzepek, and Emma Timmins-Schiffman. Transitioning global change experiments on Southern Ocean phytoplankton from lab to field settings: Insights and challenges. Limnology and Oceanography, 67(9):1911–1930, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Berthelot:2021:IOC

[BDL^{+}21] Hugo Berthelot, Solange Duhamel, Stéphane L’Helguen, Jean-François Maguer, and Nicolas Cassar. Inorganic and organic carbon and nitrogen uptake strategies of picoplankton groups in the northwestern Atlantic Ocean. Limnology and Oceanography, 66(10):3682–3696, October 2021. CODEN LI-OCAH. ISSN 0024-3590.

Bommarito:2023:WPI

[BDMGH^{+}23] Claudia Bommarito, Dakeishla M. Díaz-Morales, Tamar Guy-Haim, Simona Noè, Jules Delasalle, Björn Buchholz, Maral Khosravi, Gil Rilov, Bernd Sures, and Martin Wahl. Warming and parasitism impair the performance of Baltic native and invasive macroalgae and their associated fauna. Limnology and Oceanography, 68(8):1852–1864, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Briones-Fourzan:2021:MSL

[BFCC^{+}21] Patricia Briones-Fourzán, Julio Candela, Laura Carrillo, Alí F. Espinosa-Magaña, Fernando Negrete-Soto, Cecilia Barradas-Ortiz, Edgar Escalante-Mancera, Rubén Muñoz de Cote-Hernández, Rogelio Martínez-Calderón, and Enrique Lozano-Álvarez. Metamorphosis of spiny lobsters (Panulirus argus and Panulirus guttatus) in the Yucatan Current as inferred from the distribution of pueruli and final stage phyllosomata. Limnology and Oceanography, 66(9):3421–3438, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Borchardt:2020:SDD

[BFE^{+}20] Trace Borchardt, Kelsey V. Fisher, Alina M. Ebling, Jason R. Westrich, Peng Xian, Christopher D. Holmes, William M. Landing, Erin K. Lipp, Michael S. Wetz, and Elizabeth A. Ottesen. Saharan dust deposition initiates successional patterns among marine microbes in the Western Atlantic. Limnology and Oceanography, 65(1):191–203, January 2020. CO-DEN LIOCAH. ISSN 0024-3590.

Buscher:2022:CWC

[BFW^{+}22] Janina Vanessa Büscher, Armin Uwe Form, Max Wisshak, Rainer Kiko, and Ulf Riebesell. Cold-water coral ecosystems under future ocean change: Live coral performance vs. framework dissolution and bioerosion. Limnology and Oceanography, 67(11):2497–2515, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Bharti:2022:CND

[BGAD^{+}22] D. K. Bharti, Katell Guizien, M. T. Aswathi-Das, P. N. Vinayachandran, and Kartik Shanker. Connectivity networks and delineation of disconnected coastal provinces along the Indian coastline using large-scale Lagrangian transport simulations. Limnology and Oceanography, 67(6):1416–1428, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Berggren:2020:SMP

[BGG^{+}20] Martin Berggren, Cristian Gudasz, Francois Guillemette, Geert Hensgens, Linlin Ye, and Jan Karlsson. Systematic microbial production of optically active dissolved organic matter in subarctic lake water. Limnology and Oceanography, 65 (5):951–961, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Brailsford:2021:LCN

[BGM^{+}21] Francesca L. Brailsford, Helen C. Glanville, Miles R. Marshall, Christopher A. Yates, Alun T. Owen, Peter N. Golyshin, Penny J. Johnes, and Davey L. Jones. Land cover and nutrient enrichment regulates low-molecular weight dissolved organic matter turnover in freshwater ecosystems. Limnology and Oceanography, 66(8):2979–2987, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Bjorndahl:2022:ACP

[BGM^{+}22] Judith A. Bjorndahl, Cale A. C. Gushulak, Stefano Mezzini, Gavin L. Simpson, Heather A. Haig, Peter R. Leavitt, and Kerri Finlay. Abrupt changes in the physical and biological structure of endorheic upland lakes due to 8-m lake-level variation during the 20^{th} century. Limnology and Oceanography, 67(5):1022–1039, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Blossom:2021:LMA

[BH21] Hannah E. Blossom and Per Juel Hansen. The loss of mixotrophy in Alexandrium pseudogonyaulax: Implications for trade-offs between toxicity, mucus trap production, and phagotrophy. Limnology and Oceanography, 66(2):528–542, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Bertuzzo:2022:RRR

[BHA^{+}22] Enrico Bertuzzo, Erin R. Hotchkiss, Alba Argerich, John S. Kominoski, Diana Oviedo-Vargas, Philip Savoy, Rachel Scarlett, Daniel von Schiller, and James B. Heffernan. Respiration regimes in rivers: Partitioning source-specific respiration from metabolism time series. Limnology and Oceanography, 67(11):2374–2388, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Baetge:2024:PIF

[BHG^{+}24] Nicholas Baetge, Kimberly H. Halsey, Jason R. Graff, Brian Ver Wey, Toby K. Westberry, Amanda E. Appel, Guillaume Bourdin, Charlotte Begouen Demeaux, Emmanuel Boss, and Michael J. Behrenfeld. Physiological and interspecific factors determine diel changes in phytoplankton biooptical properties. Limnology and Oceanography, 69(2):390– 407, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Baruch:2023:FWE

[BHR^{+}23] Ethan M. Baruch, Tamara K. Harms, Albert Ruhi, Mengdi Lu, Leah Gaines-Sewell, and John L. Sabo. Food web efficiency in desert streams. Limnology and Oceanography, 68 (3):723–734, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Balmonte:2020:SCB

[BHSG^{+}20] John Paul Balmonte, Harald Hasler-Sheetal, Ronnie N. Glud, Thorbjørn J. Andersen, Mikael K. Sejr, Mathias Middelboe, Andreas Teske, and Carol Arnosti. Sharp contrasts between freshwater and marine microbial enzymatic capabilities, community composition, and DOM pools in a NE Greenland fjord. Limnology and Oceanography, 65(1):77–95, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Bergbusch:2021:USP

[BHSL21] Nathanael T. Bergbusch, Nicole M. Hayes, Gavin L. Simpson, and Peter R. Leavitt. Unexpected shift from phytoplankton to periphyton in eutrophic streams due to wastewater influx. Limnology and Oceanography, 66(7):2745–2761, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Bhattacharya:2022:NMT

[BJG^{+}22] Ruchi Bhattacharya, John R. Jones, Jennifer L. Graham, Daniel V. Obrecht, Anthony P. Thorpe, James D. Harlan, and Rebecca L. North. Nonlinear multidecadal trends in organic matter dynamics in Midwest reservoirs are a function of variable hydroclimate. Limnology and Oceanography, 67 (11):2531–2546, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Berg:2021:IPB

[BJK^{+}21] Sonja Berg, Sandra Jivcov, Stephanie Kusch, Gerhard Kuhn, Duanne White, Gerhard Bohrmann, Martin Melles, and Janet Rethemeyer. Increased petrogenic and biospheric organic carbon burial in sub-Antarctic fjord sediments in response to recent glacier retreat. Limnology and Oceanography, 66(12):4347–4362, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Barth:2023:STI

[BJS23a] Alex Barth, Rod Johnson, and Joshua Stone. Size and transparency influence diel vertical migration patterns in copepods. Limnology and Oceanography, 68(12):2749–2758, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Bonnington:2023:IEC

[BJS^{+}23b] Abigail C. Bonnington, Rob C. Jamieson, Kathryn A. Smith, Allie Oliver, Lindsay H. Johnston, Nicole K. LeRoux, Lauren D. Somers, and Barret L. Kurylyk. Impacts of Extratropical Cyclone Fiona on a sensitive coastal lagoon ecosystem. Limnology and Oceanography, 68(12):2703–2715, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Brown:2020:HAT

[BK20] Emily R. Brown and Julia Kubanek. Harmful alga trades off growth and toxicity in response to cues from dead phytoplankton. Limnology and Oceanography, 65(8):1723–1733, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Bowen:2020:PDI

[BKC20] Jennifer C. Bowen, Louis A. Kaplan, and Rose M. Cory. Photodegradation disproportionately impacts biodegradation of semi-labile DOM in streams. Limnology and Oceanography, 65(1):13–26, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Brooks:2021:OCR

[BKMY21] Thomas W. Brooks, Kevin D. Kroeger, Holly A. Michael, and Joanna K. York. Oxygen-controlled recirculating seepage meter reveals extent of nitrogen transformation in discharging coastal groundwater at the aquifer–estuary interface. Limnology and Oceanography, 66(8):3055–3069, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Brock:2024:BBI

[BLRM24] Melissa L. Brock, Alyse A. Larkin, Eric J. Raes, and Adam C. Martiny. Bacterial biogeography of the Indian Ocean. Limnology and Oceanography, 69(1):67–80, January 2024. CO-DEN LIOCAH. ISSN 0024-3590.

Brown:2020:SEE

[BLS20] Terry-René W. Brown, Marc J. Lajeunesse, and Kathleen M. Scott. Strong effects of elevated CO_{2} on freshwater microalgae and ecosystem chemistry. Limnology and Oceanography, 65(2):304–313, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Benito:2022:ERT

[BLS^{+}22] Xavier Benito, Melina Luethje, Tobias Schneider, Sherilyn C. Fritz, Paul A. Baker, Eric J. Pedersen, Pierre Gaüzère, Majoi de Novaes Nascimento, Mark Bush, and Albert Ruhi. Ecological resilience in tropical Andean lakes: a paleolimnological perspective. Limnology and Oceanography, 67(S1):S23–S37, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Bashevkin:2022:SVR

[BM22] Samuel M. Bashevkin and Brian Mahardja. Seasonally variable relationships between surface water temperature and inflow in the upper San Francisco Estuary. Limnology and Oceanography, 67(3):684–702, March 2022. CODEN LI-OCAH. ISSN 0024-3590.

Bartrons:2020:EBT

[BMA^{+}20] Mireia Bartrons, Thomas Mehner, Christine Argillier, Meryem Beklioglu, Petr Blabolil, Trygve Hesthagen, Kerstin Holmgren, Erik Jeppesen, Teet Krause, Samo Podgornik, Pietro Volta, Ian J. Winfield, and Sandra Brucet. Energybased top-down and bottom-up relationships between fish community energy demand or production and phytoplankton across lakes at a continental scale. Limnology and Oceanography, 65(4):892–902, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Bashevkin:2022:WUS

[BMB22] Samuel M. Bashevkin, Brian Mahardja, and Larry R. Brown. Warming in the upper San Francisco Estuary: Patterns of water temperature change from five decades of data. Limnology and Oceanography, 67(5):1065–1080, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Bayer:2023:CCC

[BMCS23] Barbara Bayer, Kelsey McBeain, Craig A. Carlson, and Alyson E. Santoro. Carbon content, carbon fixation yield and dissolved organic carbon release from diverse marine nitrifiers. Limnology and Oceanography, 68(1):84–96, January 2023. CODEN LIOCAH. ISSN 0024-3590.

Bachi:2023:FPM

[BMG^{+}23] Giancarlo Bachi, Elisabetta Morelli, Margherita Gonnelli, Cecilia Balestra, Raffaella Casotti, Valtere Evangelista, Daniel J. Repeta, and Chiara Santinelli. Fluorescent properties of marine phytoplankton exudates and lability to marine heterotrophic prokaryotes degradation. Limnology and Oceanography, 68(4):982–1000, April 2023. CODEN LI-OCAH. ISSN 0024-3590.

Brown:2022:PCT

[BMGK22] Emily R. Brown, Sam G. Moore, David A. Gaul, and Julia Kubanek. Predator cues target signaling pathways in toxic algal metabolome. Limnology and Oceanography, 67(6):1227– 1237, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Briddon:2023:LTE

[BMH^{+}23] Charlotte L. Briddon, Mihai Miclăuş, Adriana Hegedüs, Maria Nicoară, Maria-Cecilia Chiriac, and Bogdan Drugă. Long-term exposure to elevated temperature leads to altered gene expression in a common bloom-forming cyanobacterium. Limnology and Oceanography, 68(12):2654–2667, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Blais:2024:SDC

[BMLV24] Marie-Amélie Blais, Alex Matveev, Connie Lovejoy, and Warwick F. Vincent. Size-dependent community patterns differ between microbial eukaryotes and bacteria in a permafrost lake–river–sea continuum. Limnology and Oceanography, 69(3):667–680, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Bart:2021:DOC

[BMR^{+}21a] Martijn C. Bart, Benjamin Mueller, Titus Rombouts, Clea van de Ven, Gabrielle J. Tompkins, Ronald Osinga, Corina P. D. Brussaard, Barry MacDonald, Anja Engel, Hans Tore Rapp, and Jasper M. de Goeij. Dissolved organic carbon (DOC) is essential to balance the metabolic demands of four dominant North-Atlantic deep-sea sponges. Limnology and Oceanography, 66(3):925–938, March 2021. CODEN LI-OCAH. ISSN 0024-3590.

Besterman:2021:PBM

[BMR^{+}21b] Alice F. Besterman, Karen J. McGlathery, Matthew A. Reidenbach, Patricia L. Wiberg, and Michael L. Pace. Predicting benthic macroalgal abundance in shallow coastal lagoons from geomorphology and hydrologic flow patterns. Limnology and Oceanography, 66(1):123–140, January 2021. CO-DEN LIOCAH. ISSN 0024-3590.

Brooks:2022:ISD

[BMS^{+}22] Jill L. Brooks, Jonathan D. Midwood, Adam Smith, Steven J. Cooke, Bryan Flood, Christine M. Boston, Patricia Semecsen, Susan E. Doka, and Mathew G. Wells. Internal seiches as drivers of fish depth use in lakes. Limnology and Oceanography, 67(5):1040–1051, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Bao:2023:DSP

[BND^{+}23] Hongyan Bao, Jutta Niggemann, Moge Du, Weiqiang Zhao, Dekun Huang, Yuanbi Yi, Jin-Yu Terence Yang, Thorsten Dittmar, and Shuh-Ji Kao. Deciphering sources and processing of dissolved black carbon in coastal seas. Limnology and Oceanography, 68(11):2562–2575, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Biard:2020:VND

[BO20] Tristan Biard and Mark D. Ohman. Vertical niche definition of test-bearing protists (Rhizaria) into the twilight zone revealed by in situ imaging. Limnology and Oceanography, 65 (11):2583–2602, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Bridier:2021:FSD

[BOC^{+}21] Guillaume Bridier, Frédéric Olivier, Laurent Chauvaud, Mikael K. Sejr, and Jacques Grall. Food source diversity, trophic plasticity, and omnivory enhance the stability of a shallow benthic food web from a high-Arctic fjord exposed to freshwater inputs. Limnology and Oceanography, 66(S1): S259–S272, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Budy:2022:UEC

[BPG^{+}22] Phaedra Budy, Casey A. Pennock, Anne E. Giblin, Chris Luecke, Daniel L. White, and George W. Kling. Understanding the effects of climate change via disturbance on pristine Arctic lakes — multitrophic level response and recovery to a 12-yr low-level fertilization experiment. Limnology and Oceanography, 67(S1):S224–S241, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Bell:2020:RMD

[BPIBN20] Douglas W. Bell, Perry J. Pellechia, Ellery D. Ingall, and Claudia R. Benitez-Nelson. Resolving marine dissolved organic phosphorus (DOP) composition in a coastal estuary. Limnology and Oceanography, 66(10):2787–2799, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Brankovits:2022:OKS

[BPL22] David Brankovits, John W. Pohlman, and Laura L. Lapham. Oxygenation of a karst subterranean estuary during a tropical cyclone: mechanisms and implications for the carbon cycle. Limnology and Oceanography, 67(12):2691–2705, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Borrelli:2022:RSS

[BR22] Jonathan J. Borrelli and Rick A. Relyea. A review of spatial structure of freshwater food webs: Issues and opportunities modeling within-lake meta-ecosystems. Limnology and Oceanography, 67(8):1746–1759, August 2022. CODEN LI-OCAH. ISSN 0024-3590.

Bastias:2020:MUN

[BRB^{+}20] Elliot Bastias, Miquel Ribot, Susana Bernal, Francesc Sabater, and Eugènia Martí. Microbial uptake of nitrogen and carbon from the water column by litter-associated microbes differs among litter species. Limnology and Oceanography, 65(8):1891–1902, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Bystricky:2022:DPD

[BRB^{+}22] Pavel Karel Bystřický, Tereza Rutová, Vojtěch Brož, Magdalena Gajdošová, Petr Jan Juračka, Denis Copilaş-Ciocianu, and Adam Petrusek. Distribution patterns at different spatial scales reveal reproductive isolation and frequent syntopy among divergent lineages of an amphipod species complex in Western Carpathian streams. Limnology and Oceanography, 67(12):2796–2808, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Blain:2021:SAS

[BRCO21] Stéphane Blain, Mathieu Rembauville, Olivier Crispi, and Ingrid Obernosterer. Synchronized autonomous sampling reveals coupled pulses of biomass and export of morphologically different diatoms in the Southern Ocean. Limnology and Oceanography, 66(3):753–764, March 2021. CODEN LI-OCAH. ISSN 0024-3590.

Blain:2020:MPR

[BRHS20] Caitlin O. Blain, T. Alwyn V. Rees, S. Christine Hansen, and Nick T. Shears. Morphology and photosynthetic response of the kelp Ecklonia radiata across a turbidity gradient. Limnology and Oceanography, 65(3):529–544, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Bumpers:2023:ENE

[BRM^{+}23] Phillip M. Bumpers, Amy D. Rosemond, David W. P. Manning, John S. Kominoski, Jonathan P. Benstead, and Lee M. Demi. Experimental nutrient enrichment of forest streams reduces ecosystem nitrogen and phosphorus storage. Limnology and Oceanography, 68(7):1670–1685, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Blain:2020:NEO

[BS20] Caitlin O. Blain and Nick T. Shears. Nutrient enrichment offsets the effects of low light on growth of the kelp Ecklonia radiata. Limnology and Oceanography, 66(5):2220–2235, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Bernardino:2020:LUI

[BSB^{+}20] Angelo F. Bernardino, Christian J. Sanders, Lorena B. Bissoli, Luiz Eduardo de O. Gomes, J. Boone Kauffman, and Tiago O. Ferreira. Land use impacts on benthic bioturbation potential and carbon burial in Brazilian mangrove ecosystems. Limnology and Oceanography, 65(10):2366–2376, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Behnke:2021:DOM

[BSF^{+}21] Megan I. Behnke, Aron Stubbins, Jason B. Fellman, Eran Hood, Thorsten Dittmar, and Robert G. M. Spencer. Dissolved organic matter sources in glacierized watersheds delineated through compositional and carbon isotopic modeling. Limnology and Oceanography, 66(2):438–451, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Balmonte:2021:SCM

[BSGA21] John Paul Balmonte, Meinhard Simon, Helge-Ansgar Giebel, and Carol Arnosti. A sea change in microbial enzymes: Heterogeneous latitudinal and depth-related gradients in bulk water and particle-associated enzymatic activities from 30^{o}S to 59^{o}N in the Pacific Ocean. Limnology and Oceanography, 66(9):3489–3507, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Balina:2023:SLU

[BSIdG23] Sofia Baliña, María Laura Sánchez, Irina Izaguirre, and Paul A. del Giorgio. Shallow lakes under alternative states differ in the dominant greenhouse gas emission pathways. Limnology and Oceanography, 68(1):1–13, January 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Bhattacharya:2022:LEH

[BST^{+}22] Debashish Bhattacharya, Timothy G. Stephens, Amanda I. Tinoco, Robert H. Richmond, and Phillip A. Cleves. Life on the edge: Hawaiian model for coral evolution. Limnology and Oceanography, 67(9):1976–1985, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Begin:2021:EWR

[BTK^{+}21] Paschale N. Bégin, Yukiko Tanabe, Michio Kumagai, Alexander I. Culley, Michel Paquette, Denis Sarrazin, Masaki Uchida, and Warwick F. Vincent. Extreme warming and regime shift toward amplified variability in a far northern lake. Limnology and Oceanography, 66(S1):S17–S29, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Balaguer:2023:PRA

[BTT23] Jenna Balaguer, Silke Thoms, and Scarlett Trimborn. The physiological response of an Antarctic key phytoplankton species to low iron and manganese concentrations. Limnology and Oceanography, 68(9):2153–2166, September 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Brothers:2021:SFP

[BV21] Soren Brothers and Yvonne Vadeboncoeur. Shoring up the foundations of production to respiration ratios in lakes. Limnology and Oceanography, 66(7):2762–2778, July 2021. CO-DEN LIOCAH. ISSN 0024-3590.

Buttay:2022:NLC

[BVGQN22] Lucie Buttay, David A. Vasseur, Rafael González-Quirós, and Enrique Nogueira. Nutrient limitation can explain a rapid transition to synchrony in an upwelling-driven diatom community. Limnology and Oceanography, 67(S1):S298–S311, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Bartosiewicz:2023:DHM

[BVL^{+}23] Maciej Bartosiewicz, Jessica Venetz, Saskia Läubli, Oscar Sepúlveda Steiner, Damien Bouffard, Jakob Zopfi, and Moritz F. Lehmann. Detritus-hosted methanogenesis sustains the methane paradox in an alpine lake. Limnology and Oceanography, 68(1):248–264, January 2023. CODEN LI-OCAH. ISSN 0024-3590.

Bulseco:2020:MCB

[BVM^{+}20] Ashley N. Bulseco, Joseph H. Vineis, Anna E. Murphy, Amanda C. Spivak, Anne E. Giblin, Jane Tucker, and Jennifer L. Bowen. Metagenomics coupled with biogeochemical rates measurements provide evidence that nitrate addition stimulates respiration in salt marsh sediments. Limnology and Oceanography, 66(6):S321–S339, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Barnum:2022:MUL

[BWB^{+}22] Thomas R. Barnum, John Timothy Wootton, Rebecca J. Bixby, John M. Drake, David Murray-Stoker, Checo Colón-Gaud, Amanda T. Rugenski, Therese C. Frauendorf, Scott Connelly, Susan S. Kilham, Matt R. Whiles, Karen R. Lips, and Catherine M. Pringle. Mechanisms underlying lack of functional compensation by insect grazers after tadpole declines in a neotropical stream. Limnology and Oceanography, 67(S1):S198–S210, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Brinkop:2024:IRR

[BWSW24] Konstanze Brinkop, Laura M. Wienhausen, Erik Sperfeld, and Alexander Wacker. Ingestion and respiration rates of a common coastal mysid respond differently to diurnal temperature fluctuation. Limnology and Oceanography, 69(2):380– 389, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Bier:2020:SSS

[BWV^{+}20] Raven L. Bier, Jennifer J. Wernegreen, Rytas J. Vilgalys, Joseph Christopher Ellis, and Emily S. Bernhardt. Subsidized or stressed? Shifts in freshwater benthic microbial metagenomics along a gradient of alkaline coal mine drainage. Limnology and Oceanography, 66(6):S277–S292, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Bian:2023:BCC

[BYB^{+}23] Xiaopeng Bian, Shun-Chung Yang, Kenneth M. Bolster, Rintaro Moriyasu, James W. Moffett, and Seth G. John. Biogeochemical cycling of Cd, Mn, and Ce in the Eastern Tropical North Pacific oxygen-deficient zone. Limnology and Oceanography, 68(2):483–497, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Berggren:2023:NLM

[BYS^{+}23] Martin Berggren, Linlin Ye, Ryan A. Sponseller, Ann-Kristin Bergström, Jan Karlsson, Hendricus Verheijen, and Geert Hensgens. Nutrient limitation masks the dissolved organic matter composition effects on bacterial metabolism in unproductive freshwaters. Limnology and Oceanography, 68(9): 2059–2069, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Bollinger:2022:ABE

[BZK^{+}22] Eric Bollinger, Jochen P. Zubrod, Marco Konschak, Lenz Sulzer, Jacob Schnurr, Verena C. Schreiner, Ralf Schulz, and Mirco Bundschuh. As above, so below? Effects of fungicides on microbial organic matter decomposition are stronger in the hyporheic than in the benthic zone. Limnology and Oceanography, 67(1):39–52, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Bos:2023:MID

[BZSF23] Ryan P. Bos, Shiye Zhao, Tracey T. Sutton, and Tamara M. Frank. Microplastic ingestion by deep-pelagic crustaceans and fishes. Limnology and Oceanography, 68(7):1595–1610, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:2022:RPD

[CAV^{+}22] Stephen R. Carpenter, Babak M. S. Arani, Egbert H. Van Nes, Marten Scheffer, and Michael L. Pace. Resilience of phytoplankton dynamics to trophic cascades and nutrient enrichment. Limnology and Oceanography, 67(S1):S258–S265, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Cavaliere:2021:WTP

[CB21a] Emily Cavaliere and Helen M. Baulch. Winter in two phases: Long-term study of a shallow reservoir in winter. Limnology and Oceanography, 65(3):1335–1352, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Champenois:2021:NCM

[CB21b] Willy Champenois and Alberto V. Borges. Net community metabolism of a Posidonia oceanica meadow. Limnology and Oceanography, 66(6):2126–2140, June 2021. CODEN LI-OCAH. ISSN 0024-3590.

Capo:2022:ODW

[CBB^{+}22] Eric Capo, Elias Broman, Stefano Bonaglia, Andrea G. Bravo, Stefan Bertilsson, Anne L. Soerensen, Jarone Pinhassi, Daniel Lundin, Moritz Buck, Per O. J. Hall, Francisco J. A. Nascimento, and Erik Björn. Oxygen-deficient water zones in the Baltic Sea promote uncharacterized Hg methylating microorganisms in underlying sediments. Limnology and Oceanography, 67(1):135–146, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Courtney:2022:RAP

[CBC^{+}22] Travis A. Courtney, Hannah C. Barkley, Stephen Chan, Courtney S. Couch, Tye L. Kindinger, Thomas A. Oliver, David J. Kriegman, and Andreas J. Andersson. Rapid assessments of Pacific Ocean net coral reef carbonate budgets and net calcification following the 2014–2017 global coral bleaching event. Limnology and Oceanography, 67(8):1687–1700, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Cox:2023:DHB

[CBD^{+}23] Isabelle Cox, Robert J. W. Brewin, Giorgio Dall’Olmo, Katy Sheen, Shubha Sathyendranath, Rafael Rasse, and Osvaldo Ulloa. Distinct habitat and biogeochemical properties of lowoxygen-adapted tropical oceanic phytoplankton. Limnology and Oceanography, 68(9):2022–2039, September 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Choisnard:2023:ASS

[CBF^{+}23] Noémie Choisnard, Emma Burtscher, Stefan Forster, Claudia Frey, Matthias Moros, and Maren Voss. The Amazon shelf sediments, a reactor that fuels intense nitrogen cycling at the seabed. Limnology and Oceanography, 68(10):2211–2226, October 2023. CODEN LIOCAH. ISSN 0024-3590.

Carter:2021:HDS

[CBHB21] Alice M. Carter, Joanna R. Blaszczak, James B. Heffernan, and Emily S. Bernhardt. Hypoxia dynamics and spatial distribution in a low gradient river. Limnology and Oceanography, 66(6):2251–2265, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Cruz:2021:MDB

[CBN21] Bianca N. Cruz, Samantha Brozak, and Susanne Neuer. Microscopy and DNA-based characterization of sinking particles at the Bermuda Atlantic Time-series Study station point to zooplankton mediation of particle flux. Limnology and Oceanography, 66(10):3697–3713, October 2021. CODEN LI-OCAH. ISSN 0024-3590.

Carrier-Belleau:2022:TPM

[CBPNA22] Charlotte Carrier-Belleau, Ludovic Pascal, Christian Nozais, and Philippe Archambault. Tipping points and multiple drivers in changing aquatic ecosystems: a review of experimental studies. Limnology and Oceanography, 67(S1):S312– S330, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Choquet:2021:NEHb

[CBS^{+}21] Marvin Choquet, Gauthier Burckard, Stig Skreslet, Galice Hoarau, and Janne E. Søreide. No evidence for hybridization between Calanus finmarchicus and Calanus glacialis in a subarctic area of sympatry. Limnology and Oceanography, 66(S1):S314–S325, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Coe:2021:CDA

[CBT^{+}21] Allison Coe, Steven J. Biller, Elaina Thomas, Konstantinos Boulias, Christina Bliem, Aldo Arellano, Keven Dooley, Anna N. Rasmussen, Kristen LeGault, Tyler J. O’Keefe, Sarah Stover, Eric L. Greer, and Sallie W. Chisholm. Coping with darkness: The adaptive response of marine picocyanobacteria to repeated light energy deprivation. Limnology and Oceanography, 66(9):3300–3312, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Caldwell:2021:DPI

[CCA^{+}21] Timothy J. Caldwell, Sudeep Chandra, Thomas P. Albright, Adrian A. Harpold, Thomas E. Dilts, Jonathan A. Greenberg, Steve Sadro, and Michael D. Dettinger. Drivers and projections of ice phenology in mountain lakes in the western United States. Limnology and Oceanography, 66(3):995– 1008, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Chen:2020:SSV

[CCB^{+}20] Baoshan Chen, Wei-Jun Cai, Jean R. Brodeur, Najid Hussain, Jeremy M. Testa, Wenfei Ni, and Qian Li. Seasonal and spatial variability in surface pCO_{2} and air–water CO_{2} flux in the Chesapeake Bay. Limnology and Oceanography, 66(6):3046–3065, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Coffin:2021:KWE

[CCC^{+}21] Michael R. S. Coffin, Jeff C. Clements, Luc A. Comeau, Thomas Guyondet, Michelle Maillet, Laura Steeves, Keryn Winterburn, José M. F. Babarro, Martin A. Mallet, Rémy Haché, Luke A. Poirier, Saswati Deb, and Ramon Filgueira. The killer within: Endogenous bacteria accelerate oyster mortality during sustained anoxia. Limnology and Oceanography, 66(7):2885–2900, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Connolly:2021:SDO

[CCDM21] Craig T. Connolly, Byron C. Crump, Kenneth H. Dunton, and James W. McClelland. Seasonality of dissolved organic matter in lagoon ecosystems along the Alaska Beaufort Sea coast. Limnology and Oceanography, 66(12):4299–4313, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Chen:2021:QES

[CCFF21] Yi-Yang Chen, Paul Cooper, Christopher J. Fulton, and Rebecca J. Fox. Quantifying epifaunal secondary production within tropical macroalgal meadows: Seasonality and sensitivity to canopy structure. Limnology and Oceanography, 66 (12):4267–4284, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Cordeiro:2022:OLC

[CCG^{+}22] Malaika Cordeiro, John H. Costello, Brad J. Gemmell, Kelly R. Sutherland, and Sean P. Colin. Oceanic lobate ctenophores possess feeding mechanics similar to the impactful coastal species Mnemiopsis leidyi. Limnology and Oceanography, 67(12):2706–2717, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Callbeck:2021:SCO

[CCK^{+}21] Cameron M. Callbeck, Donald E. Canfield, Marcel M. M. Kuypers, Pelin Yilmaz, Gaute Lavik, Bo Thamdrup, Carsten J. Schubert, and Laura A. Bristow. Sulfur cycling in oceanic oxygen minimum zones. Limnology and Oceanography, 66(6):2360–2392, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Chen:2021:BCC

[CCY^{+}21] Qiuwen Chen, Yuchen Chen, Jun Yang, Stephen C. Maberly, Jianyun Zhang, Jinren Ni, Gangsheng Wang, Daniele Tonina, Lin Xiao, and Honghai Ma. Bacterial communities in cascade reservoirs along a large river. Limnology and Oceanography, 66(12):4363–4374, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Clifford:2020:MTP

[CDA^{+}20] Elisabeth L. Clifford, Daniele De Corte, Chie Amano, Paolo Paliaga, Ingrid Ivančić, Victor Ortiz, Mirjana Najdek, Gerhard J. Herndl, and Eva Sintes. Mesozooplankton taurine production and prokaryotic uptake in the northern Adriatic Sea. Limnology and Oceanography, 66(10):2730–2747, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Chi:2021:TJW

[CDH^{+}21] Xupeng Chi, Jan Dierking, Henk-Jan Hoving, Florian Lüskow, Anneke Denda, Bernd Christiansen, Ulrich Sommer, Thomas Hansen, and Jamileh Javidpour. Tackling the jelly web: Trophic ecology of gelatinous zooplankton in oceanic food webs of the eastern tropical Atlantic assessed by stable isotope analysis. Limnology and Oceanography, 66(2):289– 305, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Choisnard:2024:TFS

[CDW^{+}24] Noémie Choisnard, Nicolas Noel Duprey, Tanja Wald, Martin Thibault, Fanny Houlbrèque, Alan D. Foreman, Pascale Cuet, Mireille M. M. Guillaume, Hubert Vonhof, Daniel M. Sigman, Gerald H. Haug, Jean-François Maguer, Stéphane L’Helguen, Alfredo Martínez-García, and Anne Lorrain. Tracing the fate of seabird-derived nitrogen in a coral reef using nitrate and coral skeleton nitrogen isotopes. Limnology and Oceanography, 69(2):309–324, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Costa:2022:SDA

[CEE^{+}22] Matthew T. Costa, Exequiel Ezcurra, Paula Ezcurra, Pelayo Salinas de León, Benjamin Turner, Joy Kumagai, James Leichter, and Octavio Aburto-Oropeza. Sediment depth and accretion shape belowground mangrove carbon stocks across a range of climatic and geologic settings. Limnology and Oceanography, 67(S2):S104–S117, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Chiapella:2021:FFM

[CESS21] Ariana M. Chiapella, Collin A. Eagles-Smith, and Angela L. Strecker. From forests to fish: Mercury in mountain lake food webs influenced by factors at multiple scales. Limnology and Oceanography, 66(4):1021–1035, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Chen:2021:DAD

[CEW^{+}21] Jian-Jhih Chen, Dirk V. Erler, Naomi S. Wells, Jianyin Huang, David T. Welsh, and Bradley D. Eyre. Denitrification, anammox, and dissimilatory nitrate reduction to ammonium across a mosaic of estuarine benthic habitats. Limnology and Oceanography, 66(4):1281–1297, April 2021. CO-DEN LIOCAH. ISSN 0024-3590.

Carvalho:2020:TCH

[CFC^{+}20] Filipa Carvalho, Jessica N. Fitzsimmons, Nicole Couto, Nicole Waite, Maxim Gorbunov, Josh Kohut, Matthew J. Oliver, Robert M. Sherrell, and Oscar Schofield. Testing the Canyon Hypothesis: Evaluating light and nutrient controls of phytoplankton growth in penguin foraging hotspots along the West Antarctic Peninsula. Limnology and Oceanography, 65(3):455–470, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Croteau:2021:CNQ

[CGB^{+}21] Dany Croteau, Sébastien Guérin, Flavienne Bruyant, Joannie Ferland, Douglas A. Campbell, Marcel Babin, and Johann Lavaud. Contrasting nonphotochemical quenching patterns under high light and darkness aligns with light niche occupancy in Arctic diatoms. Limnology and Oceanography, 66(S1):S231–S245, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Chen:2021:SEE

[CGC^{+}21] Chung-Chi Chen, Gwo-Ching Gong, Kuo-Ping Chiang, Fuh-Kwo Shiah, Chih-Ching Chung, and Chin-Chang Hung. Scaling effects of a eutrophic river plume on organic carbon consumption. Limnology and Oceanography, 66(5):1867–1881, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Clemo:2022:BIC

[CGD22] William Cyrus Clemo, Katelyn D. Giles, and Kelly M. Dorgan. Biological influences on coastal muddy sediment structure following resuspension. Limnology and Oceanography, 67(11):2466–2482, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Canadell:2021:DED

[CGGC^{+}21] Marta Boix Canadell, Lluís Gómez-Gener, Mélanie Clémençon, Stuart N. Lane, and Tom J. Battin. Daily entropy of dissolved oxygen reveals different energetic regimes and drivers among high-mountain stream types. Limnology and Oceanography, 65(3):1594–1610, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Christaki:2021:SMF

[CGL^{+}21] Urania Christaki, Audrey Gueneugues, Yan Liu, Stéphane Blain, Philippe Catala, Jonathan Colombet, Pavla Debeljak, Ludwig Jardillier, Solène Irion, Fred Planchon, Ingrid Sassenhagen, Telesphore Sime-Ngando, and Ingrid Obernosterer. Seasonal microbial food web dynamics in contrasting Southern Ocean productivity regimes. Limnology and Oceanography, 66(1):108–122, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Cameron:2022:EEP

[CGR22] Louise P. Cameron, Jonathan H. Grabowski, and Justin B. Ries. Effects of elevated pCO_{2} and temperature on the calcification rate, survival, extrapallial fluid chemistry, and respiration of the Atlantic Sea scallop Placopecten magellanicus. Limnology and Oceanography, 67(8):1670–1686, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Camarena-Gomez:2021:BDD

[CGRGP^{+}21] María Teresa Camarena-Gómez, Clara Ruiz-González, Jonna Piiparinen, Tobias Lipsewers, Cristina Sobrino, Ramiro Logares, and Kristian Spilling. Bacterioplankton dynamics driven by interannual and spatial variation in diatom and dinoflagellate spring bloom communities in the Baltic Sea. Limnology and Oceanography, 66(1):255–271, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Corona:2024:LTS

[CHA^{+}24] Stefano Corona, Andrew G. Hirst, David Atkinson, Jasmin Renz, Maarten Boersma, and Angus Atkinson. Longterm shifts in phenology, thermal niche, population size, and their interactions in marine pelagic copepods. Limnology and Oceanography, 69(3):482–497, March 2024. CODEN LI-OCAH. ISSN 0024-3590.

Corona:2021:DDM

[CHAA21] Stefano Corona, Andrew Hirst, David Atkinson, and Angus Atkinson. Density-dependent modulation of copepod body size and temperature–size responses in a shelf sea. Limnology and Oceanography, 66(11):3916–3927, November 2021. CO-DEN LIOCAH. ISSN 0024-3590.

Cheung:2024:DAD

[CHP^{+}24] Henry L. S. Cheung, Jenny R. Hillman, Conrad A. Pilditch, Candida Savage, Isaac R. Santos, Ronnie N. Glud, Francisco J. A. Nascimento, Simon F. Thrush, and Stefano Bonaglia. Denitrification, anammox, and DNRA in oligotrophic continental shelf sediments. Limnology and Oceanography, 69(3): 621–637, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Chmiel:2020:WDR

[CHS^{+}20] Hannah E. Chmiel, Hilmar Hofmann, Sebastian Sobek, Tatyana Efremova, and Natacha Pasche. Where does the river end? Drivers of spatiotemporal variability in CO_{2} concentration and flux in the inflow area of a large boreal lake. Limnology and Oceanography, 65(6):1161–1174, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Cimino:2021:APA

[CJB^{+}21] Megan A. Cimino, Michael G. Jacox, Steven J. Bograd, Stephanie Brodie, Gemma Carroll, Elliott L. Hazen, Bertha E. Lavaniegos, Mark M. Morales, Erin Satterthwaite, and Ryan R. Rykaczewski. Anomalous poleward advection facilitates episodic range expansions of pelagic red crabs in the eastern North Pacific. Limnology and Oceanography, 66 (8):3176–3189, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Cohen:2023:DPP

[CKCB^{+}23] Ashley B. Cohen, Vanja Klepac-Ceraj, Kristen Bidas, Felix Weber, Arkadiy I. Garber, Lisa N. Christensen, Jacob A. Cram, Michael L. McCormick, and Gordon T. Taylor. Deep photoautotrophic prokaryotes contribute substantially to carbon dynamics in oxygen-deficient waters in a permanently redox-stratified freshwater lake. Limnology and Oceanography, 68(1):232–247, January 2023. CODEN LIOCAH. ISSN 0024-3590.

Church:2021:PDM

[CKH^{+}21] Matthew J. Church, Eint Kyi, Robert O. Hall,, Jr., David M. Karl, Markus Lindh, Alexa Nelson, and Emma K. Wear. Production and diversity of microorganisms associated with sinking particles in the subtropical North Pacific Ocean. Limnology and Oceanography, 66(9):3255–3270, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Calamita:2024:DCR

[CLA^{+}24] Elisa Calamita, J. Jelle Lever, Clement Albergel, R. Iestyn Woolway, and Daniel Odermatt. Detecting climate-related shifts in lakes: a review of the use of satellite Earth Observation. Limnology and Oceanography, 69(4):723–741, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Cook:2020:GMI

[CLC^{+}20] Katherine V. Cook, Chuang Li, Haiyuan Cai, Lee R. Krumholz, K. David Hambright, Hans W. Paerl, Morgan M. Steffen, Alan E. Wilson, Michele A. Burford, Hans-Peter Grossart, David P. Hamilton, Helong Jiang, Assaf Sukenik, Delphine Latour, Elisabeth I. Meyer, Judit Padisák, Boqiang Qin, Richard M. Zamor, and Guangwei Zhu. The global Microcystis interactome. Limnology and Oceanography, 65(S1): S194–S207, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Cook:2021:CGM

[CLC^{+}21] K. V. Cook, C. Li, H. Cai, L. R. Krumholz, K. D. Hambright, H. W. Paerl, M. M. Steffen, A. E. Wilson, M. A. Burford, H.-P. Grossart, D. P. Hamilton, H. Jiang, A. Sukenik, D. Latour, E. I. Meyer, J. Padisák, B. Qin, R. M. Zamor, and G. Zhu. Corrigendum to: The global Microcystis interactome. Limnology and Oceanography, 66(6):2496–2497, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Chen:2022:IMS

[CLC^{+}22] Qi Chen, Christian Lønborg, Feng Chen, Michael Gonsior, Yunyun Li, Ruanhong Cai, Chen He, Jiaxin Chen, Yu Wang, Quan Shi, Nianzhi Jiao, and Qiang Zheng. Increased microbial and substrate complexity result in higher molecular diversity of the dissolved organic matter pool. Limnology and Oceanography, 67(11):2360–2373, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Colombo:2023:CPM

[CLD^{+}23] Manuel Colombo, Julie LaRoche, Dhwani Desai, Jingxuan Li, and Maria T. Maldonado. Control of particulate manganese (Mn) cycling in halocline Arctic Ocean waters by putative Mn-oxidizing bacterial dynamics. Limnology and Oceanography, 68(9):2070–2087, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Chaguaceda:2024:ZNL

[CLG^{+}24] Fernando Chaguaceda, Danny C. P. Lau, Willem Goedkoop, Mariem Fadhlaoui, Isabelle Lavoie, and Tobias Vrede. Zooplankton in northern lakes show taxon-specific responses in fatty acids across climate-productivity and ecosystem size gradients. Limnology and Oceanography, 69(4):947–960, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Chen:2023:MRK

[CLGH23] Yuren Chen, Ming Li, Patricia M. Glibert, and Cynthia Heil. MurK y waters: Modeling the succession from r to K strategists (diatoms to dinoflagellates) following a nutrient release from a mining facility in Florida. Limnology and Oceanography, 68(10):2288–2304, October 2023. CODEN LIOCAH. ISSN 0024-3590.

Cristi:2024:EDF

[CLP^{+}24] Antonia Cristi, Cliff S. Law, Matt Pinkerton, Adriana Lopes dos Santos, Karl Safi, and Andres Gutiérrez-Rodríguez. Environmental driving forces and phytoplankton diversity across the Ross Sea region during a summer–autumn transition. Limnology and Oceanography, 69(4):772–788, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Christiansen:2022:TDS

[CLT^{+}22] Svenja Christiansen, Øystein Langangen, Josefin Titelman, Leif Asbjørn Vøllestad, and Stein Kaartvedt. Threedimensional swimming behavior and activity of a mesopelagic fish. Limnology and Oceanography, 67(12):2677–2690, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Chen:2021:BSD

[CLW^{+}21] Chong Chen, Tzu-Hao Lin, Hiromi Kayama Watanabe, Tomonari Akamatsu, and Shinsuke Kawagucci. Baseline soundscapes of deep-sea habitats reveal heterogeneity among ecosystems and sensitivity to anthropogenic impacts. Limnology and Oceanography, 66(10):3714–3727, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Cohen:2020:CDS

[CM20a] Adam P. Cohen and John M. Melack. Carbon dioxide supersaturation in high-elevation oligotrophic lakes and reservoirs in the Sierra Nevada, California. Limnology and Oceanography, 65(3):612–626, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Cortes:2020:MPS

[CM20b] Alicia Cortés and Sally MacIntyre. Mixing processes in small Arctic lakes during spring. Limnology and Oceanography, 65 (2):260–288, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Clark:2021:PLY

[CM21] J. Blake Clark and Antonio Mannino. Preferential loss of Yukon River delta colored dissolved organic matter under nutrient replete conditions. Limnology and Oceanography, 66 (5):1613–1626, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Closset:2021:DRA

[CMB^{+}21] Ivia Closset, Heather M. McNair, Mark A. Brzezinski, Jeffrey W. Krause, Kimberlee Thamatrakoln, and Janice L. Jones. Diatom response to alterations in upwelling and nutrient dynamics associated with climate forcing in the California Current System. Limnology and Oceanography, 65(3): 1578–1593, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Chielle:2024:SSV

[CMCC24] Raisa S. A. Chielle, Rozane V. Marins, Mariany S. Cavalcante, and Luiz C. Cotovicz, Jr. Seasonal and spatial variability of CO_{2} emissions in a large tropical mangrove-dominated delta. Limnology and Oceanography, 69(2):246–261, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Chauvet:2023:HDM

[CML23] Marina Chauvet, Arthur Monjot, and Cécile Lepère. High diversity of microsporidian parasites and new planktonic hosts in freshwater and marine ecosystems. Limnology and Oceanography, 68(4):928–941, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Costa:2020:DID

[CMT^{+}20] Raul Rodrigo Costa, Carlos Rafael Borges Mendes, Virginia Maria Tavano, Tiago Segabinazzi Dotto, Rodrigo Kerr, Thiago Monteiro, Clarisse Odebrecht, and Eduardo Resende Secchi. Dynamics of an intense diatom bloom in the Northern Antarctic Peninsula, February 2016. Limnology and Oceanography, 65(9):2056–2075, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Caroselli:2020:PDT

[CÖL^{+}20] Erik Caroselli, Hasan Baris Özalp, Maila Lavia, Francesca De Witt, Francesco Raimondi, and Stefano Goffredo. Population dynamics of a temperate coral along a depth gradient in the Dardanelles. Limnology and Oceanography, 66(10): 2676–2687, November 2020. CODEN LIOCAH. ISSN 0024-

Clarke:2022:SIV

[CP22] Andrew Clarke and Helen J. Peat. Seasonal and interannual variability of feeding in Antarctic benthos. Limnology and Oceanography, 67(4):962–972, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Catalan:2021:REV

[CPB^{+}21] Núria Catalán, Ada Pastor, Carles M. Borrego, Joan Pere Casas-Ruiz, Jeffrey A. Hawkes, Carmen Gutiérrez, Daniel von Schiller, and Rafael Marcé. The relevance of environment vs. composition on dissolved organic matter degradation in freshwaters. Limnology and Oceanography, 66(2):306– 320, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Compaire:2021:CCN

[CPBJR^{+}21] Jesus C. Compaire, Paula Pérez-Brunius, Sylvia Patricia Adelheid Jiménez-Rosenberg, Javier Rodríguez Outerelo, Laura del Pilar Echeverri García, and Sharon Z. Herzka. Connectivity of coastal and neritic fish larvae to the deep waters. Limnology and Oceanography, 66(6):2423–2441, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Crook:2022:SSF

[CSB22] Kevin A. Crook, Marcus Sheaves, and Adam Barnett. Species-specific foraging behaviors define the functional roles of sympatric stingrays. Limnology and Oceanography, 67(1): 219–230, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Chen:2021:MCP

[CSC^{+}21] Xiaogang Chen, Isaac R. Santos, Mitchell Call, Gloria M. S. Reithmaier, Damien Maher, Ceylena Holloway, Praktan D. Wadnerkar, Paula Gómez-Álvarez, Christian J. Sanders, and Ling Li. The mangrove CO_{2} pump: Tidally driven porewater exchange. Limnology and Oceanography, 65(3):1563– 1577, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Carter:2024:USM

[CSD^{+}24] Brendan R. Carter, Jonathan D. Sharp, Andrew G. Dickson, Marta Álvarez, Michael B. Fong, Maribel I. García-Ibáñez, Ryan J. Woosley, Yuichiro Takeshita, Leticia Barbero, Robert H. Byrne, Wei-Jun Cai, Melissa Chierici, Simon L. Clegg, Regina A. Easley, Andrea J. Fassbender, Kalla L. Fleger, Xinyu Li, Macarena Martín-Mayor, Katelyn M. Schockman, and Zhaohui Aleck Wang. Uncertainty sources for measurable ocean carbonate chemistry variables. Limnology and Oceanography, 69(1):1–21, January 2024. CO-DEN LIOCAH. ISSN 0024-3590.

Choquet:2021:NEHa

[CSH21] Marvin Choquet, Irina Smolina, and Galice Hoarau. No evidence for hybridization in Calanus: Reply to the comment by Parent et al. Limnology and Oceanography, 66(10):3603– 3606, October 2021. CODEN LIOCAH. ISSN 0024-3590. See [PPB^{+}21].

Catlett:2023:IPP

[CSM^{+}23] Dylan Catlett, David A. Siegel, Paul G. Matson, Emma K. Wear, Craig A. Carlson, Thomas S. Lankiewicz, and M. Debora Iglesias-Rodriguez. Integrating phytoplankton pigment and DNA meta-barcoding observations to determine phytoplankton composition in the coastal ocean. Limnology and Oceanography, 68(2):361–376, February 2023. CODEN LI-OCAH. ISSN 0024-3590.

Conroy:2024:OSF

[CSNS24] John A. Conroy, Deborah K. Steinberg, Schuyler C. Nardelli, and Oscar Schofield. Omnivorous summer feeding by juvenile Antarctic krill in coastal waters. Limnology and Oceanography, 69(4):874–887, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Camoying:2023:PRA

[CT23] Marianne G. Camoying and Scarlett Trimborn. Physiological response of an Antarctic cryptophyte to increasing temperature, CO_{2}, and irradiance. Limnology and Oceanography, 68 (8):1880–1894, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Colls:2021:BPT

[CTF^{+}21] Miriam Colls, Xisca Timoner, Carme Font, Vicenç Acuña, and Sergi Sabater. Biofilm pigments in temporary streams indicate duration and severity of drying. Limnology and Oceanography, 66(9):3313–3326, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Choi:2020:LDD

[CTH^{+}20] Jun Choi, Cary Troy, Nathan Hawley, Michael McCormick, and Mathew Wells. Lateral dispersion of dye and drifters in the center of a very large lake. Limnology and Oceanography, 65(2):336–348, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Carbonne:2021:TTC

[CTM^{+}21] Chloe Carbonne, Núria Teixidó, Billy Moore, Alice Mirasole, Thomas Guttierez, Jean-Pierre Gattuso, and Steeve Comeau. Two temperate corals are tolerant to low pH regardless of previous exposure to natural CO_{2} vents. Limnology and Oceanography, 66(11):4046–4061, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Chaudhary:2021:BRS

[CTP21] Adit Chaudhary, Sarah Turner, and Rachel Poretsky. Bacterioplankton respond with similar transcriptional activity to allochthonous dissolved organic matter in coastal and offshore Lake Michigan. Limnology and Oceanography, 66(8): 3162–3175, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Corrales-Ugalde:2021:FMF

[CUS21] Marco Corrales-Ugalde and Kelly R. Sutherland. Fluid mechanics of feeding determine the trophic niche of the hydromedusa Clytia gregaria. Limnology and Oceanography, 66(3):939–953, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Cruaud:2020:ABC

[CVF^{+}20] Perrine Cruaud, Adrien Vigneron, Marie-Stéphanie Fradette, Caetano C. Dorea, Alexander I. Culley, Manuel J. Rodriguez, and Steve J. Charette. Annual bacterial community cycle in a seasonally ice-covered river reflects environmental and climatic conditions. Limnology and Oceanography, 65(S1): S21–S37, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Chaffin:2022:QMP

[CWF^{+}22] Justin D. Chaffin, Judy A. Westrick, Elliot Furr, Johnna A. Birbeck, Laura A. Reitz, Keara Stanislawczyk, Wei Li, Peter K. Weber, Thomas B. Bridgeman, Timothy W. Davis, and Xavier Mayali. Quantification of microcystin production and biodegradation rates in the western basin of Lake Erie. Limnology and Oceanography, 67(7):1470–1483, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Cabello-Yeves:2023:VNO

[CYPRG^{+}23] Pedro J. Cabello-Yeves, Antonio Picazo, Juan J. Roda-Garcia, Francisco Rodriguez-Valera, and Antonio Camacho. Vertical niche occupation and potential metabolic interplay of microbial consortia in a deeply stratified meromictic model lake. Limnology and Oceanography, 68(11):2492–2511, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Cabello-Yeves:2020:MDL

[CYZZ^{+}20] Pedro J. Cabello-Yeves, Tamara I. Zemskaya, Alexandra S. Zakharenko, Mariya V. Sakirko, Vyacheslav G. Ivanov, Rohit Ghai, and Francisco Rodriguez-Valera. Microbiome of the deep Lake Baikal, a unique oxic bathypelagic habitat. Limnology and Oceanography, 65(7):1471–1488, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Cao:2021:PTD

[CZH^{+}21] Haobing Cao, Zhenchang Zhu, Peter M. J. Herman, Stijn Temmerman, Jaco de Smit, Liquan Zhang, Lin Yuan, and Tjeerd J. Bouma. Plant traits determining biogeomorphic landscape dynamics: a study on clonal expansion strategies driving cliff formation at marsh edges. Limnology and Oceanography, 66(10):3754–3767, October 2021. CODEN LI-OCAH. ISSN 0024-3590.

Corman:2023:RLM

[CZK^{+}23] Jessica R. Corman, Jacob A. Zwart, Jennifer Klug, Denise A. Bruesewitz, Elvira de Eyto, Marcus Klaus, Lesley B. Knoll, James A. Rusak, Michael J. Vanni, María Belén Alfonso, Rocio Luz Fernandez, Huaxia Yao, Kari Austnes, Raoul-Marie Couture, Heleen A. de Wit, Jan Karlsson, and Alo Laas. Response of lake metabolism to catchment inputs inferred using high-frequency lake and stream data from across the northern hemisphere. Limnology and Oceanography, 68(12): 2617–2631, December 2023. CODEN LIOCAH. ISSN 0024-

Collister:2022:IPC

[CZS^{+}22] Brian L. Collister, Richard C. Zimmerman, Charles I. Sukenik, William M. Balch, and Victoria J. Hill. The influence of particle concentration and bulk characteristics on polarized oceanographic lidar measurements. Limnology and Oceanography, 67(6):1374–1387, June 2022. CODEN LI-OCAH. ISSN 0024-3590.

Duan:2021:OCE

[DAB^{+}21] Mi Duan, Julian R. Ashford, Sophie Bestley, Xiaoying Wei, Andrea Walters, and Guoping Zhu. Otolith chemistry of Electrona antarctica suggests a potential population marker distinguishing the southern Kerguelen Plateau from the eastward-flowing Antarctic Circumpolar Current. Limnology and Oceanography, 66(2):405–421, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Doubek:2021:EVS

[DAD^{+}21] Jonathan P. Doubek, Orlane Anneville, Gaël Dur, Aleksandra M. Lewandowska, Vijay P. Patil, James A. Rusak, Nico Salmaso, Christian Torsten Seltmann, Dietmar Straile, Pablo Urrutia-Cordero, Patrick Venail, Rita Adrian, María B. Alfonso, Curtis L. DeGasperi, Elvira de Eyto, Heidrun Feuchtmayr, Evelyn E. Gaiser, Scott F. Girdner, Jennifer L. Graham, Hans-Peter Grossart, Josef Hejzlar, Stéphan Jacquet, Georgiy Kirillin, María E. Llames, Shin-Ichiro S. Matsuzaki, Emily R. Nodine, Maria Cintia Piccolo, Don C. Pierson, Alon Rimmer, Lars G. Rudstam, Steven Sadro, Hilary M. Swain, Stephen J. Thackeray, Wim Thiery, Piet Verburg, Tamar Zohary, and Jason D. Stockwell. The extent and variability of storm-induced temperature changes in lakes measured with long-term and high-frequency data. Limnology and Oceanography, 66(5):1979–1992, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Deeds:2022:SBC

[DAN^{+}22] Jeremy Deeds, Aria Amirbahman, Stephen A. Norton, Linda C. Bacon, and Rachel A. Hovel. Shifting baselines and cross-scale drivers of lake water clarity: Applications for lake assessment. Limnology and Oceanography, 67(S1):S184– S197, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Debeljak:2021:HDI

[DBB^{+}21] Pavla Debeljak, Stéphane Blain, Andrew Bowie, Pier van der Merwe, Barbara Bayer, and Ingrid Obernosterer. Homeostasis drives intense microbial trace metal processing on marine particles. Limnology and Oceanography, 66(10):3842–3855, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Dunkle:2024:GSR

[DBFC24] Matthew R. Dunkle, J. Ryan Bellmore, Jason B. Fellman, and Christopher C. Caudill. Glaciers, snow, and rain: Water source influences invertebrate community structure and secondary production across a hydrologically diverse subarctic landscape. Limnology and Oceanography, 69(2):232–245, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Drake:2023:HDE

[DBM^{+}23] Travis W. Drake, Matti Barthel, Christian Ekemba Mbongo, Davin Mata Mpambi, Simon Baumgartner, Clement Ikene Botefa, Marijn Bauters, Martin R. Kurek, Robert G. M. Spencer, Amy M. McKenna, Negar Haghipour, Godé Lompoko Ekamba, Jose N. Wabakanghanzi, Timothy I. Eglinton, Kristof Van Oost, and Johan Six. Hydrology drives export and composition of carbon in a pristine tropical river. Limnology and Oceanography, 68(11):2476–2491, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Dory:2022:SDD

[DCF^{+}22] Flavia Dory, Laurent Cavalli, Evelyne Franquet, Magalie Claeys-Bruno, Benjamin Misson, Thierry Tatoni, and Céline Bertrand. Summer dynamics drive the microbial response to carbon and nutrient additions in a high-altitude lake. Limnology and Oceanography, 67(5):1142–1156, May 2022. CODEN LIOCAH. ISSN 0024-3590.

DeWysiecki:2022:PGD

[DCJB22] Agustín M. De Wysiecki, Federico Cortés, Andrés J. Jaureguizar, and Adam Barnett. Potential global distribution of a temperate marine coastal predator: The role of barriers and dispersal corridors on subpopulation connectivity. Limnology and Oceanography, 67(8):1805–1819, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Deng:2022:ETR

[DCK^{+}22] Lixia Deng, Shunyan Cheung, Chang-Keun Kang, Kailin Liu, Xiaomin Xia, and Hongbin Liu. Elevated temperature relieves phosphorus limitation of marine unicellular diazotrophic cyanobacteria. Limnology and Oceanography, 67 (1):122–134, January 2022. CODEN LIOCAH. ISSN 0024-3590.

deChanvalon:2023:DSM

[dCLO^{+}23] Aubin Thibault de Chanvalon, George W. Luther III, Véronique E. Oldham, Bradley M. Tebo, Nicole R. Coffey, and Timothy F. Shaw. Distribution and stability of Mn complexes in the ocean: Influence of hydrothermal plumes and weather events. Limnology and Oceanography, 68(2):455–466, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Doyle:2023:OCB

[DdEM^{+}23] Brian C. Doyle, Elvira de Eyto, Valerie McCarthy, Mary Dillane, and Eleanor Jennings. The organic carbon budget of an oligotrophic temperate peatland lake. Limnology and Oceanography, 68(3):544–556, March 2023. CODEN LI-OCAH. ISSN 0024-3590.

Dias:2022:CFE

[DFJ22] Beatriz S. Dias, Michael G. Frisk, and Adrian Jordaan. Contrasting fishing effort reduction and habitat connectivity as management strategies to promote alewife (Alosa pseudoharengus) recovery using an ecosystem model. Limnology and Oceanography, 67(S1):S5–S22, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Dugenne:2020:LDC

[DFW^{+}20] Mathilde Dugenne, Fernanda Henderikx Freitas, Samuel T. Wilson, David M. Karl, and Angelicque E. White. Life and death of Crocosphaera sp. in the Pacific Ocean: Fine scale predator–prey dynamics. Limnology and Oceanography, 65 (11):2603–2617, November 2020. CODEN LIOCAH. ISSN 0024-3590.

DelVecchia:2021:BCH

[DGTW21] Amanda Gay DelVecchia, Steven Gougherty, Brad W. Taylor, and Scott A. Wissinger. Biogeochemical characteristics and hydroperiod affect carbon dioxide flux rates from exposed high-elevation pond sediments. Limnology and Oceanography, 66(4):1050–1067, April 2021. CODEN LI-OCAH. ISSN 0024-3590.

Downes:2021:PDB

[DGW^{+}21] Patrick P. Downes, Stephen J. Goult, E. Malcolm S. Woodward, Claire E. Widdicombe, Karen Tait, and Joanna L. Dixon. Phosphorus dynamics in the Barents Sea. Limnology and Oceanography, 66(S1):S326–S342, February 2021. CO-DEN LIOCAH. ISSN 0024-3590.

DAmbrosio:2022:SFE

[DHNH22] Sofia L. D’Ambrosio, Stephen M. Henderson, Jeffrey R. Nielson, and John A. Harrison. In situ flux estimates reveal large variations in methane flux across the bottom boundary layer of a eutrophic lake. Limnology and Oceanography, 67(10): 2119–2139, October 2022. CODEN LIOCAH. ISSN 0024-3590.

deJuan:2023:MPC

[dJGCS23] Carlos de Juan, Kaiene Griffell, Albert Calbet, and Enric Saiz. Multigenerational physiological compensation and body size reduction dampen the effects of warming on copepods. Limnology and Oceanography, 68(5):1037–1047, May 2023. CODEN LIOCAH. ISSN 0024-3590.

Dudgeon:2020:MMC

[DK20] Steve Dudgeon and Janet E. Kübler. A multistressor model of carbon acquisition regulation for macroalgae in a changing climate. Limnology and Oceanography, 65(10):2541–2555, October 2020. CODEN LIOCAH. ISSN 0024-3590.

DuClos:2021:DRA

[DKBG21] Kevin T. Du Clos, Lee Karp-Boss, and Brad J. Gemmell. Diatoms rapidly alter sinking behavior in response to changing nutrient concentrations. Limnology and Oceanography, 66(3): 892–900, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Dunn:2023:CTW

[DKPS23] Robert P. Dunn, Julie L. Krask, James L. Pinckney, and Erik M. Smith. Contrasting trends in water quality between adjacent ocean- and river-dominated estuaries: Evidence for marsh porewaters as a source of nutrient enrichment? Limnology and Oceanography, 68(9):2040–2058, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Dziuba:2021:CVC

[DKW^{+}21] Marcin Krzysztof Dziuba, Lechosław Kuczyński, Łukasz Wejnerowski, Slawek Cerbin, and Justyna Wolinska. Countergradient variation concealed adaptive responses to temperature increase in Daphnia from heated lakes. Limnology and Oceanography, 66(4):1268–1280, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Du:2022:WDT

[DLC^{+}22] YingXun Du, ChunYan Luo, FeiZhou Chen, QiaoYing Zhang, YongQiang Zhou, Kyoung-Soon Jang, YiBo Zhang, ChunQiao Song, YongDong Zhang, YunLin Zhang, and Yue-Han Lu. Water depth and transparency drive the quantity and quality of organic matter in sediments of Alpine Lakes on the Tibetan Plateau. Limnology and Oceanography, 67 (9):1959–1975, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Dezutter:2021:SIV

[DLDF21] Thibaud Dezutter, Catherine Lalande, Gérald Darnis, and Louis Fortier. Seasonal and interannual variability of the Queen Maud Gulf ecosystem derived from sediment trap measurements. Limnology and Oceanography, 66(S1):S411– S426, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Duret:2020:EIO

[DLL20] Manon T. Duret, Richard S. Lampitt, and Phyllis Lam. Eukaryotic influence on the oceanic biological carbon pump in the Scotia Sea as revealed by ^{18}S rRNA gene sequencing of suspended and sinking particles. Limnology and Oceanography, 65(S1):S49–S70, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Diamond:2022:MRS

[DMC^{+}22] Jacob S. Diamond, Florentina Moatar, Matthew J. Cohen, Alain Poirel, Cécile Martinet, Anthony Maire, and Gilles Pinay. Metabolic regime shifts and ecosystem state changes are decoupled in a large river. Limnology and Oceanography, 67(S1):S54–S70, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Dziuba:2023:CCW

[DMCW23] Marcin Krzysztof Dziuba, Florent Manzi, Slawek Cerbin, and Justyna Wolinska. Can climate warming save Daphnia from parasites? Reduced parasite prevalence in Daphnia populations from artificially heated lakes. Limnology and Oceanography, 68(1):181–191, January 2023. CODEN LI-OCAH. ISSN 0024-3590.

deMelo:2020:LDO

[dMKB^{+}20] Michaela L. de Melo, Dolly N. Kothawala, Stefan Bertilsson, João Henrique Amaral, Bruce Forsberg, and Hugo Sarmento. Linking dissolved organic matter composition and bacterioplankton communities in an Amazon floodplain system. Limnology and Oceanography, 65(1):63–76, January 2020. CO-DEN LIOCAH. ISSN 0024-3590.

Donis:2021:SSL

[DMM^{+}21] Daphne Donis, Evanthia Mantzouki, Daniel F. McGinnis, Dominic Vachon, Irene Gallego, Hans-Peter Grossart, Lisette N. de Senerpont Domis, Sven Teurlincx, Laura Seelen, Miquel Lürling, Yvon Verstijnen, Valentini Maliaka, Jeremy Fonvielle, Petra M. Visser, Jolanda Verspagen, Maria van Herk, Maria G. Antoniou, Nikoletta Tsiarta, Valerie McCarthy, Victor C. Perello, Danielle Machado-Vieira, Alinne Gurjão de Oliveira, Dubravka Špoljarić Maronić, Filip Stević, Tanja Žuna Pfeiffer, Itana Bokan Vucelić, Petar Žutinić, Marija Gligora Udovič, An– delka Plenković-Moraj, Luděk Bláha, Rodan Geriš, Markéta Fránková, Kirsten Seestern Christoffersen, Trine Perlt Warming, Tõnu Feldmann, Alo Laas, Kristel Panksep, Lea Tuvikene, Kersti Kangro, Judita Koreivienė, Jūratė Karosienė, Jūratė Kasperovičienė, Ksenija Savadova-Ratkus, Irma Vitonytė, Kerstin Häggqvist, Pauliina Salmi, Lauri Arvola, Karl Rothhaupt, Christos Avagianos, Triantafyllos Kaloudis, Spyros Gkelis, Manthos Panou, Theodoros Triantis, Sevasti-Kiriaki Zervou, Anastasia Hiskia, Ulrike Obertegger, Adriano Boscaini, Giovanna Flaim, Nico Salmaso, Leonardo Cerasino, Sigrid Haande, Birger Skjelbred, Magdalena Grabowska, Maciej Karpowicz, Damian Chmura, Lidia Nawrocka, Justyna Kobos, Hanna Mazur-Marzec, Pablo Alcaraz-Párraga, Elżbieta Wilk-Woźniak, Wojciech Krztoń, Edward Walusiak, Ilona Gagala-Borowska, Joana Mankiewicz-Boczek, Magdalena Toporowska, Barbara Pawlik-Skowronska, Michał Niedźwiecki, Wojciech Peczuła, Agnieszka Napiórkowska-Krzebietke, Julita Dunalska, Justyna Sieńska, Daniel Szymański, Marek Kruk, Agnieszka Budzyńska, Ryszard Goldyn, Anna Kozak, Joanna Rosińska, Elżbieta Szelag-Wasielewska, Piotr Domek, Natalia Jakubowska-Krepska, Kinga Kwasizur, Beata Messyasz, Aleksandra Pełechata, Mariusz Pełechaty, Mikolaj Kokocinski, Beata Madrecka-Witkowska, Iwona Kostrzewska-Szlakowska, Magdalena Frak, Agnieszka Bańkowska-Sobczak, Michał Wasilewicz, Agnieszka Ochocka, Agnieszka Pasztaleniec, Iwona Jasser, Ana M. Antão-Geraldes, Manel Leira, Vitor Vasconcelos, Joao Morais, Micaela Vale, Pedro M. Raposeiro, Vítor Gonçalves, Boris Aleksovski, Svetislav Krstić, Hana Nemova, Iveta Drastichova, Lucia Chomova, Spela Remec-Rekar, Tina Elersek, Lars-Anders Hansson, Pablo Urrutia-Cordero, Andrea G. Bravo, Moritz Buck, William Colom-Montero, Kristiina Mustonen, Don Pierson, Yang Yang, Jessica Richardson, Christine Edwards, Hannah Cromie, Jordi Delgado-Martín, David García, Jose Luís Cereijo, Joan Gomà, Mari Carmen Trapote, Teresa Vegas-Vilarrúbia, Biel Obrador, Ana García-Murcia, Monserrat Real, Elvira Romans, Jordi Noguero-Ribes, David Parreño Duque, Elísabeth Fernández-Morán, Bárbara Úbeda, José Ángel Gálvez, Núria Catalán, Carmen Pérez-Martínez, Eloísa Ramos-Rodríguez, Carmen Cillero-Castro, Enrique Moreno-Ostos, José María Blanco, Valeriano Rodríguez, Jorge Juan Montes-Pérez, Roberto L. Palomino, Estela Rodríguez-Pérez, Armand Hernández, Rafael Carballeira, Antonio Camacho, Antonio Picazo, Carlos Rochera, Anna C. Santamans, Carmen Ferriol, Susana Romo, Juan Miguel Soria, Arda Özen, Tünay Karan, Nilsun Demir, Meryem Beklioğlu, Nur Filiz, Eti Levi, Uğur Iskin, Gizem Bezirci, Ülkü Nihan Tavşanoğlu, Kemal Çelik, Koray Ozhan, Nusret Karakaya, Mehmet Ali Turan Koçer, Mete Yilmaz, Faruk Maraşlıoğlu, Özden Fakioglu, Elif Neyran Soylu, Meral Apaydın Yağcı, Şakir Çınar, Kadir Çapkın, Abdulkadir Yağcı, Mehmet Cesur, Fuat Bilgin, Cafer Bulut, Rahmi Uysal, Köker Latife, Reyhan Akçaalan, Meriç Albay, Mehmet Tahir Alp, Korhan Özkan, Tuğba Ongun Sevindik, Hatice Tunca, Burçin Önem, Hans Paerl, Cayelan C. Carey, and Bastiaan W. Ibelings. Stratification strength and light climate explain variation in chlorophyll-a at the continental scale in a European multilake survey in a heatwave summer. Limnology and Oceanography, 66(12):4314–4333, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Doherty:2021:DZF

[DMS^{+}21] Shannon C. Doherty, Amy E. Maas, Deborah K. Steinberg, Brian N. Popp, and Hilary G. Close. Distinguishing zooplankton fecal pellets as a component of the biological pump using compound-specific isotope analysis of amino acids. Limnology and Oceanography, 66(7):2827–2841, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Duffy:2022:PCE

[DNA^{+}22] Megan E. Duffy, Jacquelyn A. Neibauer, Jamee Adams, Rachel A. Lundeen, Gabrielle Rocap, Anitra E. Ingalls, Clara A. Fuchsman, and Richard G. Keil. Protein cycling in the eastern tropical North Pacific oxygen-deficient zone: a de novo-discovery peptidomic approach. Limnology and Oceanography, 67(2):498–510, February 2022. CODEN LI-OCAH. ISSN 0024-3590.

Desmit:2020:CCC

[DNB^{+}20] Xavier Desmit, Anja Nohe, Alberto Vieira Borges, Theo Prins, Karien De Cauwer, Ruth Lagring, Dimitry Van der Zande, and Koen Sabbe. Changes in chlorophyll concentration and phenology in the North Sea in relation to deeutrophication and sea surface warming. Limnology and Oceanography, 65(4):828–847, April 2020. CODEN LIOCAH. ISSN 0024-3590.

deOrte:2021:URC

[dOKC^{+}21] Manoela Romanó de Orte, David A. Koweek, Tyler Cyronak, Yuichiro Takeshita, Alyssa Griffin, Kennedy Wolfe, Alina Szmant, Robert Whitehead, Rebecca Albright, and Ken Caldeira. Unexpected role of communities colonizing dead coral substrate in the calcification of coral reefs. Limnology and Oceanography, 66(5):1793–1803, May 2021. CO-DEN LIOCAH. ISSN 0024-3590.

Dudgeon:2022:EER

[DP22] Steve R. Dudgeon and Peter S. Petraitis. Experimental evidence for resilience of rockweeds on rocky shores in the Gulf of Maine, USA. Limnology and Oceanography, 67(S1):S211– S223, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Diamond:2023:LHC

[DPB^{+}23] Jacob S. Diamond, Gilles Pinay, Susana Bernal, Matthew J. Cohen, David Lewis, Anna Lupon, Jay Zarnetske, and Florentina Moatar. Light and hydrologic connectivity drive dissolved oxygen synchrony in stream networks. Limnology and Oceanography, 68(2):322–335, February 2023. CODEN LI-OCAH. ISSN 0024-3590.

DelVecchia:2023:VDC

[DRA^{+}23] Amanda G. DelVecchia, Spencer Rhea, Kelly S. Aho, Emily H. Stanley, Erin R. Hotchkiss, Alice Carter, and Emily S. Bernhardt. Variability and drivers of CO_{2}, CH_{4}, and N_{2}O concentrations in streams across the United States. Limnology and Oceanography, 68(2):394–408, February 2023. CODEN LIOCAH. ISSN 0024-3590.

deSmit:2022:WES

[dSNIB22] Jaco C. de Smit, Muhammad S. Bin Mohd Noor, Eduardo Infantes, and Tjeerd J. Bouma. Wind exposure and sediment type determine the resilience and response of seagrass meadows to climate change. Limnology and Oceanography, 67(S1):S121–S132, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Deemer:2020:CPL

[DSY20] Bridget R. Deemer, Edward G. Stets, and Charles B. Yackulic. Calcite precipitation in Lake Powell reduces alkalinity and total salt loading to the Lower Colorado River Basin. Limnology and Oceanography, 65(7):1439–1455, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Duan:2021:ESF

[DTPV21] Zhipeng Duan, Xiao Tan, Hans W. Paerl, and Dedmer B. Van de Waal. Ecological stoichiometry of functional traits in a colonial harmful cyanobacterium. Limnology and Oceanography, 66(5):2051–2062, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Corte:2023:ZDD

[DVL^{+}23] Daniele De Corte, Marta M. Varela, Angeles M. Louro, Sarah K. Bercovici, Joaquín Valencia-Vila, Eva Sintes, Federico Baltar, Tamara Rodríguez-Ramos, Meinhard Simon, Antonio Bode, Thorsten Dittmar, and Jutta Niggemann.

Zooplankton-derived dissolved organic matter composition and its bioavailability to natural prokaryotic communities. Limnology and Oceanography, 68(2):336–347, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Daggers:2020:SVM

[DvOH^{+}20] Tisja D. Daggers, Dick van Oevelen, Peter M. J. Herman, Henricus T. S. Boschker, and Daphne van der Wal. Spatial variability in macrofaunal diet composition and grazing pressure on microphytobenthos in intertidal areas. Limnology and Oceanography, 66(10):2819–2834, November 2020. CODEN LIOCAH. ISSN 0024-3590.

DAvignon:2023:EET

[DWR^{+}23] Geneviève D’Avignon, Duncan Wang, Heather B. Reid, Irene Gregory-Eaves, and Anthony Ricciardi. Effects of elevated temperature and microplastic exposure on growth and predatory performance of a freshwater fish. Limnology and Oceanography, 68(10):2245–2260, October 2023. CODEN LI-OCAH. ISSN 0024-3590.

Dong:2022:DPS

[DWS^{+}22] Sijia Dong, Xingchen Tony Wang, Adam V. Subhas, Frank J. Pavia, Jess F. Adkins, and William M. Berelson. Depth profiles of suspended carbon and nitrogen along a North Pacific transect: Concentrations, isotopes, and ratios. Limnology and Oceanography, 67(1):247–260, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Deng:2021:PCB

[DWW^{+}21] Wenchao Deng, Shanlin Wang, Xianhui Wan, Zhenzhen Zheng, Nianzhi Jiao, Shuh-Ji Kao, Jefferson Keith Moore, and Yao Zhang. Potential competition between marine heterotrophic prokaryotes and autotrophic picoplankton for nitrogen substrates. Limnology and Oceanography, 66(9):3338– 3355, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Elagami:2022:MMS

[EAF^{+}22] Hassan Elagami, Pouyan Ahmadi, Jan H. Fleckenstein, Sven Frei, Martin Obst, Seema Agarwal, and Benjamin S. Gilfedder. Measurement of microplastic settling velocities and implications for residence times in thermally stratified lakes. Limnology and Oceanography, 67(4):934–945, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Evans:2020:RWM

[EBK^{+}20] Natalya Evans, Elisabeth Boles, Jarek V. Kwiecinski, Susan Mullen, Martin Wolf, Allan H. Devol, Rintaro Moriyasu, SungHyun Nam, Andrew R. Babbin, and James W. Moffett. The role of water masses in shaping the distribution of redox active compounds in the Eastern Tropical North Pacific oxygen deficient zone and influencing low oxygen concentrations in the eastern Pacific Ocean. Limnology and Oceanography, 65(8):1688–1705, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Eichner:2020:MID

[EBW^{+}20] Meri Eichner, Subhajit Basu, Siyuan Wang, Dirk de Beer, and Yeala Shaked. Mineral iron dissolution in Trichodesmium colonies: The role of O_{2} and pH microenvironments. Limnology and Oceanography, 65(6):1149–1160, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Edmunds:2024:EYL

[EDC24] Peter J. Edmunds, Steve S. Doo, and Robert C. Carpenter. Effects of year-long exposure to elevated pCO_{2} on the metabolism of back reef and fore reef communities. Limnology and Oceanography, 69(3):533–547, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Edmunds:2021:SVC

[Edm21a] Peter J. Edmunds. Spatiotemporal variation in coral recruitment and its association with seawater temperature. Limnology and Oceanography, 65(3):1394–1408, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Edmunds:2021:VRS

[Edm21b] Peter J. Edmunds. Vital rates of small reef corals are associated with variation in climate. Limnology and Oceanography, 66(3):901–913, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Evans:2023:MNS

[EFC^{+}23] Chris D. Evans, Stacey L. Felgate, Steffi Carter, Mark Stinchcombe, Edward Mawji, Andrew P. Rees, Inma Lebron, Richard Sanders, Paul Brickle, and Daniel J. Mayor. Marine nutrient subsidies promote biogeochemical hotspots in undisturbed, highly humic estuaries. Limnology and Oceanography, 68(8):1802–1820, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Evans:2024:RSH

[EFG^{+}24] Natalya Evans, Alexis E. Floback, Justin Gaffney, Peter J. Chace, Zachary Luna, Joël Knoery, Clare E. Reimers, and James W. Moffett. The role of seasonal hypoxia and benthic boundary layer exchange on iron redox cycling on the Oregon shelf. Limnology and Oceanography, 69(4):742–756, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Evensen:2021:RHC

[EFP^{+}21] Nicolas R. Evensen, Maoz Fine, Gabriela Perna, Christian R. Voolstra, and Daniel J. Barshis. Remarkably high and consistent tolerance of a Red Sea coral to acute and chronic thermal stress exposures. Limnology and Oceanography, 66(5): 1718–1729, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Ehrlich:2020:CCT

[EG20] Elias Ehrlich and Ursula Gaedke. Coupled changes in traits and biomasses cascading through a tritrophic plankton food web. Limnology and Oceanography, 65(10):2502–2514, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Ebenezer:2022:EMC

[EHC^{+}22] Vinitha Ebenezer, Yingyu Hu, Olga Carnicer, Andrew J. Irwin, Michael J. Follows, and Zoe V. Finkel. Elemental and macromolecular composition of the marine chloropicophyceae, a major group of oceanic photosynthetic picoeukaryotes. Limnology and Oceanography, 67(3):540–551, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Euler:2023:MLL

[EJM^{+}23] Sebastian Euler, Luke C. Jeffrey, Damien T. Maher, Scott G. Johnston, Ryo Sugimoto, and Douglas R. Tait. Methanogens limited to lower rhizosphere and to an atypical salt marsh niche along a pristine intertidal mangrove continuum. Limnology and Oceanography, 68(9):2167–2182, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Egea:2023:EMH

[EJRRC^{+}23] Luis G. Egea, Rocío Jiménez-Ramos, Cristina Romera-Castillo, Isabel Casal-Porras, Paula Bonet-Melià, Alba Yamuza-Magdaleno, Lucía Cerezo-Sepúlveda, José L. Pérez-Lloréns, and Fernando G. Brun. Effect of marine heat waves on carbon metabolism, optical characterization, and bioavailability of dissolved organic carbon in coastal vegetated communities. Limnology and Oceanography, 68(2):467– 482, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Edwards:2023:IKM

[ELS23] Kyle F. Edwards, Qian Li, and Grieg F. Steward. Ingestion kinetics of mixotrophic and heterotrophic flagellates. Limnology and Oceanography, 68(4):917–927, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Evans:2024:MDC

[EPR^{+}24] Kate A. Evans, Logan M. Peoples, John R. Ranieri, Emma K. Wear, and Matthew J. Church. Mixing-driven changes in distributions and abundances of planktonic microorganisms in a large, oligotrophic lake. Limnology and Oceanography, 69(3):604–620, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Eglite:2024:BES

[EST^{+}24] Elvita Eglite, Sarah R. Stein, Benjamin A. Turschak, Gabriel J. Bowen, and Tomas O. Höök. Bidirectional energy subsidies for fish in river mouths of a large lake as revealed by stable isotopes and fatty acids. Limnology and Oceanography, 69(3):498–514, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Ejarque:2021:HCC

[ESW^{+}21] Elisabet Ejarque, Katharina Scholz, Georg Wohlfahrt, Tom J. Battin, Martin J. Kainz, and Jakob Schelker. Hydrology controls the carbon mass balance of a mountain lake in the eastern European Alps. Limnology and Oceanography, 66(6):2110–2125, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Elmarsafy:2021:CZN

[ETG21] Mariam Elmarsafy, Kayla L. Tasky, and Derek K. Gray. Can zooplankton on the North American Great Plains “keep up” with climate-driven salinity change? Limnology and Oceanography, 66(3):865–877, March 2021. CODEN LI-OCAH. ISSN 0024-3590.

Evans:2023:PNR

[ETMD23] Natalya Evans, Juliana Tichota, James W. Moffett, and Allan H. Devol. Prolific nitrite reoxidation across the Eastern Tropical North Pacific Ocean. Limnology and Oceanography, 68(8):1719–1733, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Elmi:2022:CIS

[EWF22] Dorsa Elmi, Donald R. Webster, and David M. Fields. Copepod interaction with small-scale, dissipative eddies in turbulence: Comparison among three marine species. Limnology and Oceanography, 67(8):1820–1835, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Eichner:2022:CCM

[EWGP22] Meri Eichner, Dieter Wolf-Gladrow, and Helle Ploug. Carbonate chemistry in the microenvironment within cyanobacterial aggregates under present-day and future pCO_{2} levels. Limnology and Oceanography, 67(1):203–218, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Fasching:2020:LSM

[FAB^{+}20] Christina Fasching, Christian Akotoye, Mina Bižić, Jeremy Fonvielle, Danny Ionescu, Sabateeshan Mathavarajah, Luca Zoccarato, David A. Walsh, Hans-Peter Grossart, and Marguerite A. Xenopoulos. Linking stream microbial community functional genes to dissolved organic matter and inorganic nutrients. Limnology and Oceanography, 65(S1):S71– S87, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Fellman:2023:GRI

[FBJ^{+}23] Jason B. Fellman, J. Ryan Bellmore, Connor Johnson, Matthew R. Dunkle, and Eran Hood. Glacier runoff influences biogeochemistry and resource availability in coastal temperate rainforest streams: Implications for juvenile salmon growth. Limnology and Oceanography, 68(1):70–83, January 2023. CODEN LIOCAH. ISSN 0024-3590.

Fergus:2022:NAC

[FBK^{+}22] C. Emi Fergus, J. Renée Brooks, Philip R. Kaufmann, Amina I. Pollard, Richard Mitchell, G. John Geldhof, Ryan A. Hill, Steven G. Paulsen, Paul Ringold, and Marc Weber. Natural and anthropogenic controls on lake water-level decline and evaporation-to-inflow ratio in the conterminous United States. Limnology and Oceanography, 67(7):1484–1501, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Fernandez:2020:NCD

[FBP20] Leyden Fernandez, Stefan Bertilsson, and Sari Peura. Noncyanobacterial diazotrophs dominate nitrogen-fixing communities in permafrost thaw ponds. Limnology and Oceanography, 65(S1):S180–S193, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Favareto:2023:RPC

[FBR^{+}23] Luciane Favareto, Vanda Brotas, Natalia Rudorff, Nuno Zacarias, Andreia Tracana, Luisa Lamas, Ângela Nascimento, Afonso Ferreira, Mara Gomes, Carlos Borges, Carla Palma, and Ana C. Brito. Response of phytoplankton to coastal upwelling: The importance of temporal and spatial scales. Limnology and Oceanography, 68(6):1376–1387, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Ferronato:2023:WDC

[FBRG23] Carola Ferronato, Giuliana Berden, Martín Rivarossa, and Valeria A. Guinder. Wind-driven currents and water masses shape spring phytoplankton distribution and composition in hydrologically complex, productive shelf waters. Limnology and Oceanography, 68(10):2195–2210, October 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Frenken:2023:LTN

[FBV23] Thijs Frenken, Karen M. Brandenburg, and Dedmer B. Van de Waal. Long-term nutrient load reductions and increasing lake TN : T P stoichiometrydecreasephytoplanktonbiomassanddiversityinalargeshallowlake.LimnologyandOceanography, 68 (10) : 2389−−2401, October2023.CODENLIOCAH.ISSN0024− 3590.

Follett:2021:MEB

[FDF^{+}21] Christopher L. Follett, Stephanie Dutkiewicz, Gael Forget, B. B. Cael, and Michael J. Follows. Moving ecological and biogeochemical transitions across the North Pacific. Limnology and Oceanography, 66(6):2442–2454, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Frank:2020:SCP

[FDHS20] Franziska Frank, Michael Danger, Helmut Hillebrand, and Maren Striebel. Stoichiometric constraints on phytoplankton resource use efficiency in monocultures and mixtures. Limnology and Oceanography, 65(8):1734–1746, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Franks:2020:SDP

[FGO^{+}20] Peter J. S. Franks, Jessica C. Garwood, Michael Ouimet, Jorge Cortes, Ruth C. Musgrave, and Andrew J. Lucas. Stokes drift of plankton in linear internal waves: Cross-shore transport of neutrally buoyant and depth-keeping organisms. Limnology and Oceanography, 65(6):1286–1296, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Falciani:2022:OFB

[FGP22] Jonathan E. Falciani, Maria Grigoratou, and Andrew J. Pershing. Optimizing fisheries for blue carbon management: Why size matters. Limnology and Oceanography, 67(S2): S171–S179, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Fischer:2020:RAD

[FHMK20] Alexis D. Fischer, Kendra Hayashi, Anna McGaraghan, and Raphael M. Kudela. Return of the “age of dinoflagellates” in Monterey Bay: Drivers of dinoflagellate dominance examined using automated imaging flow cytometry and long-term time series analysis. Limnology and Oceanography, 66(5):2125– 2141, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Filippi:2021:ITT

[FHR^{+}21] Margaux Filippi, Alireza Hadjighasem, Matt Rayson, Irina I. Rypina, Greg Ivey, Ryan Lowe, James Gilmour, and Thomas Peacock. Investigating transport in a tidally driven coral atoll flow using Lagrangian coherent structures. Limnology and Oceanography, 66(11):4017–4027, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Fujiki:2020:TSO

[FIH^{+}20] Tetsuichi Fujiki, Ryuichiro Inoue, Makio C. Honda, Masahide Wakita, Yoshihisa Mino, Chiho Sukigara, and Osamu Abe. Time-series observations of photosynthetic oxygen production in the subtropical western North Pacific by an underwater profiling buoy system. Limnology and Oceanography, 65(5):1072–1084, May 2020. CODEN LIOCAH. ISSN 0024-

Franks:2022:OTP

[FIM^{+}22] Peter J. S. Franks, Bryce G. Inman, Jennifer A. MacKinnon, Matthew H. Alford, and Amy F. Waterhouse. Oceanic turbulence from a planktonic perspective. Limnology and Oceanography, 67(2):348–363, February 2022. CODEN LI-OCAH. ISSN 0024-3590.

Fernandes:2021:TYM

[FJP^{+}21] Joana Nascimento Fernandes, David Jacinto, Nélia Penteado, Alina Sousa, David Mateus, Maria Inês Seabra, Teresa Silva, Paola Castellanos, João José Castro, and Teresa Cruz. Ten years of monitoring recruitment of the edible stalked barnacle Pollicipes pollicipes: linking to oceanographic variability. Limnology and Oceanography, 66(6):2309–2318, June 2021. CODEN LIOCAH. ISSN 0024-3590.

FloresLaValle:2021:CRB

[FKN21] Florybeth Flores La Valle, Michael B. Kantar, and Craig E. Nelson. Coral reef benthic community structure is associated with the spatiotemporal dynamics of submarine groundwater discharge chemistry. Limnology and Oceanography, 66(1): 188–200, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Fan:2023:LDM

[FLH^{+}23] Jiale Fan, Futian Li, Siyu Hu, Kunshan Gao, and Juntian Xu. Larger diatoms are more sensitive to temperature changes and prone to succumb to warming stress. Limnology and Oceanography, 68(11):2512–2528, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Frischer:2021:SFL

[FLW^{+}21] Marc E. Frischer, Lauren M. Lamboley, Tina L. Walters, Jay A. Brandes, Erin Arneson, Lulu E. Lacy, Natalia B. López-Figueroa, Áurea E. Rodríguez-Santiago, and Deidre M. Gibson. Selective feeding and linkages to the microbial food web by the doliolid Dolioletta gegenbauri. Limnology and Oceanography, 66(5):1993–2010, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Flombaum:2021:DUR

[FM21] Pedro Flombaum and Adam C. Martiny. Diverse but uncertain responses of picophytoplankton lineages to future climate change. Limnology and Oceanography, 66(12):4171– 4181, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Frauendorf:2020:UST

[FMT^{+}20] Therese C. Frauendorf, Richard A. MacKenzie, Ralph W. Tingley III, Dana M. Infante, and Rana W. El-Sabaawi. Using a space-for-time substitution approach to predict the effects of climate change on nutrient cycling in tropical island stream ecosystems. Limnology and Oceanography, 66 (6):3114–3127, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Flowers:2023:GUI

[FNB^{+}23] Georgina J. L. Flowers, Hazel R. Needham, Richard H. Bulmer, Andrew M. Lohrer, and Conrad A. Pilditch. Going under: The implications of sea-level rise and reduced light availability on intertidal primary production. Limnology and Oceanography, 68(6):1301–1315, June 2023. CODEN LI-OCAH. ISSN 0024-3590.

Frost:2023:IEN

[FPB^{+}23] Paul C. Frost, Nolan J. T. Pearce, Stella A. Berger, Mark O. Gessner, A. Katharina Makower, Vanessa Marzetz, Jens C. Nejstgaard, Aljona Pralle, Svenja Schälicke, Alexander Wacker, Nicole D. Wagner, and Marguerite A. Xenopoulos. Interactive effects of nitrogen and phosphorus on growth and stoichiometry of lake phytoplankton. Limnology and Oceanography, 68(5):1172–1184, May 2023. CODEN LI-OCAH. ISSN 0024-3590.

Ferrier-Pages:2022:SSS

[FPBG^{+}22] Christine Ferrier-Pagès, Vanessa Bednarz, Renaud Grover, Yehuda Benayahu, Jean-François Maguer, Cecile Rottier, Joerg Wiedenmann, and Maoz Fine. Symbiotic stony and soft corals: Is their host-algae relationship really mutualistic at lower mesophotic reefs? Limnology and Oceanography, 67(1): 261–271, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Feckler:2024:DPG

[FPG^{+}24] Alexander Feckler, Sebastian Pietz, Sara Gonçalves, Verena Gerstle, Ute Risse-Buhl, and Mirco Bundschuh. Detritivore physiology and growth benefit from algal presence during microbial leaf colonization. Limnology and Oceanography, 69(4): 848–860, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Fontaine:2023:MYT

[FR23] Diana N. Fontaine and Tatiana A. Rynearson. Multi-year time series reveals temporally synchronous diatom communities with annual frequency of recurrence in a temperate estuary. Limnology and Oceanography, 68(9):1982–1994, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Fenoy:2021:SSW

[FRRG^{+}21] Encarnación Fenoy, Juan Rubio-Ríos, José M. González, María J. Salinas, Francisco J. Moyano, and J. Jesús Casas. Strategies of shredders when feeding on low-quality leaflitter: Local population adaptations or fixed species traits? Limnology and Oceanography, 66(5):2063–2077, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Frohlich:2023:UBM

[FSH^{+}23] Lukas Fröhlich, Valentin Siebert, Qian Huang, Julien Thébault, Brivaëla Moriceau, Klaus Peter Jochum, and Bernd R. Schöne. Uptake of barium, molybdenum, and lithium and incorporation into scallop shells: Refining proxies for primary production dynamics. Limnology and Oceanography, 68(11):2544–2561, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Ferguson:2021:GBP

[FSM^{+}21] Hannah M. Ferguson, Elizabeth J. Slagle, Ann Ashlea Mc-Cann, Jeremy T. Walls, Kevin H. Wyatt, and Allison R. Rober. Greening of the boreal peatland food web: Periphyton supports secondary production in northern peatlands. Limnology and Oceanography, 66(5):1743–1758, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Frey:2023:KNO

[FSS^{+}23] Claudia Frey, Xin Sun, Laura Szemberski, Karen L. Casciotti, Emilio Garcia-Robledo, Amal Jayakumar, Colette L. Kelly, Moritz F. Lehmann, and Bess B. Ward. Kinetics of nitrous oxide production from ammonia oxidation in the Eastern Tropical North Pacific. Limnology and Oceanography, 68 (2):424–438, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Fernandez-Urruzola:2021:PRA

[FUUG^{+}21] Igor Fernández-Urruzola, Osvaldo Ulloa, Ronnie N. Glud, Matthew H. Pinkerton, Wolfgang Schneider, Frank Wenzhöfer, and Rubén Escribano. Plankton respiration in the Atacama Trench region: Implications for particulate organic carbon flux into the hadal realm. Limnology and Oceanography, 66 (8):3134–3148, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Fritz:2021:RSB

[FW21] Kelley A. Fritz and Matt R. Whiles. Reciprocal subsidies between temporary ponds and riparian forests. Limnology and Oceanography, 66(8):3149–3161, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Flood:2020:IWP

[FWDY20] Bryan Flood, Mathew Wells, Erin Dunlop, and Joelle Young. Internal waves pump waters in and out of a deep coastal embayment of a large lake. Limnology and Oceanography, 65 (2):205–223, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Frenken:2020:IFP

[FWT^{+}20] Thijs Frenken, Justyna Wolinska, Yile Tao, Thomas Rohrlack, and Ramsy Agha. Infection of filamentous phytoplankton by fungal parasites enhances herbivory in pelagic food webs. Limnology and Oceanography, 66(10):2618–2626, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Fadeev:2021:SPD

[FWvA^{+}21] Eduard Fadeev, Matthias Wietz, Wilken-Jon von Appen, Morten H. Iversen, Eva-Maria Nöthig, Anja Engel, Julia Grosse, Martin Graeve, and Antje Boetius. Submesoscale physicochemical dynamics directly shape bacterioplankton community structure in space and time. Limnology and Oceanography, 66(7):2901–2913, July 2021. CODEN LI-OCAH. ISSN 0024-3590.

Fu:2022:CIC

[FXD^{+}22] Wenjing Fu, Xiaomei Xu, Ellen M. Druffel, Xuchen Wang, Shuwen Sun, Chunle Luo, Hongmei Zhang, and Di Fan. Carbon isotopic constraints on the degradation and sequestration of organic matter in river-influenced marginal sea sediments. Limnology and Oceanography, 67(S2):S61–S75, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Gollnisch:2021:CPI

[GAS^{+}21] Raphael Gollnisch, Teodor Alling, Maria Stockenreiter, Dag Ahrén, Magdalena Grabowska, and Karin Rengefors. Calcium and pH interaction limits bloom formation and expansion of a nuisance microalga. Limnology and Oceanography, 66(9):3523–3534, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Guo:2023:SCI

[GAS^{+}23] Jinqiang Guo, Eric P. Achterberg, Yuan Shen, Huamao Yuan, Jinming Song, Jin Liu, Xuegang Li, and Liqin Duan. Stable carbon isotopic composition of amino sugars in heterotrophic bacteria and phytoplankton: Implications for assessment of marine organic matter degradation. Limnology and Oceanography, 68(12):2814–2825, December 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Ghane:2021:TSC

[GB21] Alireza Ghane and Leon Boegman. Turnover in a small Canadian Shield lake. Limnology and Oceanography, 66(9):3356– 3373, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Ghane:2023:DOB

[GB23] Alireza Ghane and Leon Boegman. The dissolved oxygen budget of a small Canadian Shield lake during winter. Limnology and Oceanography, 68(1):265–283, January 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Guerreiro:2021:CFC

[GBB^{+}21] Catarina V. Guerreiro, Karl-Heinz Baumann, Geert-Jan A. Brummer, André Valente, Gerhard Fischer, Patrizia Ziveri, Vanda Brotas, and Jan-Berend W. Stuut. Carbonate fluxes by coccolithophore species between NW Africa and the Caribbean: Implications for the biological carbon pump. Limnology and Oceanography, 66(8):3190–3208, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Greer:2020:CFS

[GBC^{+}20] Adam T. Greer, Adam D. Boyette, Valerie J. Cruz, Mustafa Kemal Cambazoglu, Brian Dzwonkowski, Luciano M. Chiaverano, Steven L. Dykstra, Christian Briseño-Avena, Robert K. Cowen, and Jerry D. Wiggert. Contrasting fine-scale distributional patterns of zooplankton driven by the formation of a diatom-dominated thin layer. Limnology and Oceanography, 66(5):2236–2258, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Germolus:2021:LTW

[GBHF21] Noah P. Germolus, Patrick L. Brezonik, Raymond M. Hozalski, and Jacques C. Finlay. Long-term water color and flow trends in the Mississippi River headwaters, 1944–2010. Limnology and Oceanography, 66(9):3552–3567, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Graversen:2022:CSI

[GBMKJ22] Anna Elizabeth Løvgren Graversen, Gary T. Banta, Pere Masque, and Dorte Krause-Jensen. Carbon sequestration is not inhibited by livestock grazing in Danish salt marshes. Limnology and Oceanography, 67(S2):S19–S35, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Guiry:2020:DCA

[GBO^{+}20] Eric J. Guiry, Michael Buckley, Trevor J. Orchard, Alicia L. Hawkins, Suzanne Needs-Howarth, Erling Holm, and Paul Szpak. Deforestation caused abrupt shift in Great Lakes nitrogen cycle. Limnology and Oceanography, 65(8):1921–1935, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Gimenez:2024:MBA

[GBW24] Luis Giménez, Maarten Boersma, and Karen H. Wiltshire. A multiple baseline approach for marine heatwaves. Limnology and Oceanography, 69(3):638–651, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Galvez:2023:HAG

[GCEM^{+}23] Ángel Gálvez, Andreu Castillo-Escrivà, Anne Magurran, Iván Alambiaga, Fabián Bonilla, Antonio Camacho, Eduardo M. García-Roger, Sanda Iepure, Javier Miralles-Lorenzo, Juan S. Monrós, Carla Olmo, Antonio Picazo, Carmen Rojo, Juan Rueda, Mahmood Sasa, Mati Segura, Xavier Armengol, and Francesc Mesquita-Joanes. Higher alpha and gamma, but not beta diversity in tropical than in Mediterranean temporary ponds: a multi-taxon spatiotemporal approach. Limnology and Oceanography, 68(10):2402–2414, October 2023. CODEN LIOCAH. ISSN 0024-3590.

Galbraith:2023:SHD

[GCMJ23] Gemma F. Galbraith, Benjamin J. Cresswell, Mark I. Mc-Cormick, and Geoffrey P. Jones. Strong hydrodynamic drivers of coral reef fish biodiversity on submerged pinnacle coral reefs. Limnology and Oceanography, 68(11):2415–2430, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Ge:2023:RIE

[GCW^{+}23] Feiyang Ge, Bin Chen, Shaopeng Wang, Qianqian Zhou, Zhiyuan Ma, Guangcheng Chen, Weiwei Yu, Wenxuan Han, Senming Tang, Shangke Su, Dian Zhang, Wanting Zhang, Xiaoqian Lu, and Shenqun An. The relative importance of environmental heterogeneity and dispersal limitation on spatial patterns of phytoplankton communities varies across seasons. Limnology and Oceanography, 68(9):1995–2007, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Godrijan:2020:MUO

[GDB20] Jelena Godrijan, David Drapeau, and William M. Balch. Mixotrophic uptake of organic compounds by coccolithophores. Limnology and Oceanography, 65(6):1410–1421, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Giering:2022:OPP

[GE22] Sarah L. C. Giering and Claire Evans. Overestimation of prokaryotic production by leucine incorporation — and how to avoid it. Limnology and Oceanography, 67(3):726–738, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Gardner:2020:LEA

[GEHD20] John R. Gardner, Scott H. Ensign, Jeffrey N. Houser, and Martin W. Doyle. Light exposure along particle flowpaths in large rivers. Limnology and Oceanography, 65(1):128–142, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Gorbunov:2021:UCF

[GF21] Maxim Y. Gorbunov and Paul G. Falkowski. Using chlorophyll fluorescence kinetics to determine photosynthesis in aquatic ecosystems. Limnology and Oceanography, 66(1):1– 13, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Garwood:2022:ERL

[GFG^{+}22] Jessica C. Garwood, Heidi L. Fuchs, Gregory P. Gerbi, Elias J. Hunter, Robert J. Chant, and John L. Wilkin. Estuarine retention of larvae: Contrasting effects of behavioral responses to turbulence and waves. Limnology and Oceanography, 67(4):992–1005, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Gradoville:2020:LCA

[GFW^{+}20] Mary R. Gradoville, Hanna Farnelid, Angelicque E. White, Kendra A. Turk-Kubo, Brittany Stewart, François Ribalet, Sara Ferrón, Paulina Pinedo-Gonzalez, E. Virginia Armbrust, David M. Karl, Seth John, and Jonathan P. Zehr. Latitudinal constraints on the abundance and activity of the cyanobacterium UCYN-A and other marine diazotrophs in the North Pacific. Limnology and Oceanography, 65(8):1858– 1875, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Garcia-Giron:2020:ELM

[GGHBS^{+}20] Jorge García-Girón, Jani Heino, Lars Baastrup-Spohr, John Clayton, Mary de Winton, Tõnu Feldmann, Camino Fernández-Aláez, Frauke Ecke, Mark V. Hoyer, Agnieszka Kolada, Sarian Kosten, Balázs A. Lukács, Roger P. Mormul, Laila Rhazi, Mouhssine Rhazi, Laura Sass, Jun Xu, and Janne Alahuhta. Elements of lake macrophyte metacommunity structure: Global variation and community-environment relationships. Limnology and Oceanography, 66(6):2883– 2895, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Garcia-Giron:2022:HCP

[GGLH^{+}22] Jorge García-Girón, Marja Lindholm, Jani Heino, Heikki Toivonen, and Janne Alahuhta. Historical contingency via priority effects counteracts environmental change on metacommunity dynamics across decades. Limnology and Oceanography, 67(S1):S38–S53, February 2022. CODEN LI-OCAH. ISSN 0024-3590.

Ganz:2024:DDV

[GGR24] Keenan J. Ganz, Max R. Glines, and Kevin C. Rose. The distribution of depth, volume, and basin shape for lakes in the conterminous United States. Limnology and Oceanography, 69(1):22–36, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Gagnon:2021:RFW

[GGS^{+}21] Karine Gagnon, Camilla Gustafsson, Tiina Salo, Francesca Rossi, Sonja Gunell, J. Paul Richardson, Pamela L.

Reynolds, J. Emmett Duffy, and Christoffer Boström. Role of food web interactions in promoting resilience to nutrient enrichment in a brackish water eelgrass (Zostera marina) ecosystem. Limnology and Oceanography, 66(7):2810–2826, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:2022:FNP

[Gil22] John J. Gilbert. Food niches of planktonic rotifers: Diversification and implications. Limnology and Oceanography, 67(10):2218–2251, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Guislain:2022:MDE

[GK22] Alexis Lucas Norbert Guislain and Jan Köhler. From minute to day: Ecophysiological response of phytoplankton to fluctuating light exposure during vertical mixing. Limnology and Oceanography, 67(12):2809–2820, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Gunthel:2020:PDM

[GKW^{+}20] Marco Günthel, Isabell Klawonn, Jason Woodhouse, Mina Bižić, Danny Ionescu, Lars Ganzert, Steffen Kümmel, Ivonne Nijenhuis, Luca Zoccarato, Hans-Peter Grossart, and Kam W. Tang. Photosynthesis-driven methane production in oxic lake water as an important contributor to methane emission. Limnology and Oceanography, 65(12):2853–2865, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Gunthel:2021:CPD

[GKW^{+}21] Marco Günthel, Isabell Klawonn, Jason Woodhouse, Mina Bižić, Danny Ionescu, Lars Ganzert, Steffen Kümmel, Ivonne Nijenhuis, Luca Zoccarato, Hans-Peter Grossart, and Kam W. Tang. Corrigendum to: Photosynthesis-driven methane production in oxic lake water as an important contributor to methane emission. Limnology and Oceanography, 65(6):2583, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Gomez-Letona:2023:SPG

[GLAHH^{+}23] Markel Gómez-Letona, Javier Arístegui, Nauzet Hernández-Hernández, María Pérez-Lorenzo, Xosé Antón Álvarez-Salgado, Eva Teira, and Marta Sebastián. Surface productivity gradients govern changes in the viability of deep ocean prokaryotes across the tropical and subtropical Atlantic. Limnology and Oceanography, 68(1):56–69, January 2023. CODEN LIOCAH. ISSN 0024-3590.

Grimaldi:2022:HDF

[GLB^{+}22] Camille M. Grimaldi, Ryan J. Lowe, Jessica A. Benthuysen, Michael V. W. Cuttler, Rebecca H. Green, Ben Radford, Nicole Ryan, and James Gilmour. Hydrodynamic drivers of fine-scale connectivity within a coral reef atoll. Limnology and Oceanography, 67(10):2204–2217, October 2022. CO-DEN LIOCAH. ISSN 0024-3590.

Guerrini:2023:MCE

[GLM^{+}23] Federica Guerrini, Delphine Lobelle, Lorenzo Mari, Renato Casagrandi, and Erik van Sebille. Modeling carbon export mediated by biofouled microplastics in the Mediterranean Sea. Limnology and Oceanography, 68(5):1078–1090, May 2023. CODEN LIOCAH. ISSN 0024-3590.

Garwood:2020:NCS

[GLN^{+}20] Jessica C. Garwood, Andrew J. Lucas, Perry Naughton, Matthew H. Alford, Paul L. D. Roberts, Jules S. Jaffe, Laura deGelleke, and Peter J. S. Franks. A novel cross-shore transport mechanism revealed by subsurface, robotic larval mimics: Internal wave deformation of the background velocity field. Limnology and Oceanography, 65(7):1456–1470, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Garwood:2021:LCS

[GLN^{+}21] Jessica C. Garwood, Andrew J. Lucas, Perry Naughton, Paul L. D. Roberts, Jules S. Jaffe, Laura deGelleke, and Peter J. S. Franks. Larval cross-shore transport estimated from internal waves with a background flow: The effects of larval vertical position and depth regulation. Limnology and Oceanography, 66(3):678–693, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Glibert:2022:WLS

[GM22] Patricia M. Glibert and Aditee Mitra. From webs, loops, shunts, and pumps to microbial multitasking: Evolving concepts of marine microbial ecology, the mixoplankton paradigm, and implications for a future ocean. Limnology and Oceanography, 67(3):585–597, March 2022. CODEN LI-OCAH. ISSN 0024-3590.

Grasset:2021:EMP

[GMI^{+}21] Charlotte Grasset, Simone Moras, Anastasija Isidorova, Raoul-Marie Couture, Annika Linkhorst, and Sebastian Sobek. An empirical model to predict methane production in inland water sediment from particular organic matter supply and reactivity. Limnology and Oceanography, 66(10):3643– 3655, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Grossart:2020:LMA

[GMMW20] Hans-Peter Grossart, Ramon Massana, Katherine D. McMahon, and David A. Walsh. Linking metagenomics to aquatic microbial ecology and biogeochemical cycles. Limnology and Oceanography, 65(S1):S2–S20, January 2020. CODEN LI-OCAH. ISSN 0024-3590.

Gland:2024:BSS

[GMNG^{+}24] Guillaume Le Gland, Marta Masdeu-Navarro, Martí Galí, Sergio M. Vallina, Matti Gralka, Flora Vincent, Otto Cordero, Assaf Vardi, and Rafel Simó. Biological sources and sinks of dimethylsulfide disentangled by an induced bloom experiment and a numerical model. Limnology and Oceanography, 69(1):140–157, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Gerhard:2021:NSL

[GMS21] Miriam Gerhard, Corinna Mori, and Maren Striebel. Nonrandom species loss in phytoplankton communities and its effect on ecosystem functioning. Limnology and Oceanography, 66(3):779–792, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Gastauer:2022:BSS

[GNO22] Sven Gastauer, Catherine F. Nickels, and Mark D. Ohman. Body size- and season-dependent diel vertical migration of mesozooplankton resolved acoustically in the San Diego Trough. Limnology and Oceanography, 67(2):300–313, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Garnier:2023:CDE

[GÖA^{+}23] Aurélie Garnier, Örjan Östman, Jenny Ask, Olivia Bell, Martin Berggren, Mayra P. D. Rulli, Hani Younes, and Magnus Huss. Coastal darkening exacerbates eutrophication symptoms through bottom-up and top-down control modification.

Limnology and Oceanography, 68(3):678–691, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Grear:2020:SRB

[GPB^{+}20] Jason Grear, Adam Pimenta, Harriet Booth, Doranne Borsay Horowitz, Wilson Mendoza, and Matthew Liebman. In situ recovery of bivalve shell characteristics after temporary exposure to elevated pCO_{2}. Limnology and Oceanography, 65(10):2337–2351, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Goldstein:2020:ERS

[GPDA^{+}20] Esther D. Goldstein, Jodi L. Pirtle, Janet T. Duffy-Anderson, William T. Stockhausen, Mark Zimmermann, Matthew T. Wilson, and Calvin W. Mordy. Eddy retention and seafloor terrain facilitate cross-shelf transport and delivery of fish larvae to suitable nursery habitats. Limnology and Oceanography, 66(10):2800–2818, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Grundger:2021:SSM

[GPK^{+}21] Friederike Gründger, David Probandt, Katrin Knittel, Vincent Carrier, Dimitri Kalenitchenko, Anna Silyakova, Pavel Serov, Bénédicte Ferré, Mette M. Svenning, and Helge Niemann. Seasonal shifts of microbial methane oxidation in Arctic shelf waters above gas seeps. Limnology and Oceanography, 66(5):1896–1914, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Gould:2024:LEC

[GR24] Jessica Gould and Justin B. Ries. Linear extension and calcification rates in a cold-water, crustose coralline alga are modulated by temperature, light, and salinity. Limnology and Oceanography, 69(1):158–172, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Gimenez:2020:RTC

[GRJ20] Luis Giménez, Peter Robins, and Stuart R. Jenkins. Role of trait combinations, habitat matrix, and network topology in metapopulation recovery from regional extinction. Limnology and Oceanography, 65(4):775–789, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Geisen:2022:PRS

[GRJ^{+}22] Carla Geisen, Céline Ridame, Emilie Journet, Pierre Delmelle, Dominique Marie, Claire Lo Monaco, Nicolas Metzl, Rawaa Ammar, Joelle Kombo, and Damien Cardinal. Phytoplanktonic response to simulated volcanic and desert dust deposition events in the South Indian and Southern Oceans. Limnology and Oceanography, 67(7):1537–1553, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Greer:2023:SIA

[GSD^{+}23] Adam T. Greer, Moritz S. Schmid, Patrick I. Duffy, Kelly L. Robinson, Mark A. Genung, Jessica Y. Luo, Thelma Panaïotis, Christian Briseño-Avena, Marc E. Frischer, Su Sponaugle, and Robert K. Cowen. In situ imaging across ecosystems to resolve the fine-scale oceanographic drivers of a globally significant planktonic grazer. Limnology and Oceanography, 68(1):192–207, January 2023. CODEN LI-OCAH. ISSN 0024-3590.

Giling:2017:DDM

[GSG^{+}17] Darren P. Giling, Peter A. Staehr, Hans Peter Grossart, Mikkel René Andersen, Bertram Boehrer, Carmelo Escot, Fatih Evrendilek, Lluís Gómez-Gener, Mark Honti, Ian D. Jones, Nusret Karakaya, Alo Laas, Enrique Moreno-Ostos, Karsten Rinke, Ulrike Scharfenberger, Silke R. Schmidt, Michael Weber, R. Iestyn Woolway, Jacob A. Zwart, and Biel Obrador. Delving deeper: Metabolic processes in the metalimnion of stratified lakes. Limnology and Oceanography, 63(3):1288–1306, May 2017. CODEN LIOCAH. ISSN 0024-3590. See corrigendum [Ano21a].

Giling:2021:C

[GSG^{+}21] Darren P. Giling, Peter A. Staehr, Hans Peter Grossart, Mikkel René Andersen, Bertram Boehrer, Carmelo Escot, Fatih Evrendilek, Lluís Gómez-Gener, Mark Honti, Ian D. Jones, Nusret Karakaya, Alo Laas, Enrique Moreno-Ostos, Karsten Rinke, Ulrike Scharfenberger, Silke R. Schmidt, Michael Weber, R. Iestyn Woolway, Jacob A. Zwart, and Biel Obrador. Corrigendum. Limnology and Oceanography, 62(3):2093, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Gale:2023:DVC

[GSP^{+}23] John Gale, Carey Sweeney, Sara Paver, Maureen L. Coleman, and Anne W. Thompson. Diverse and variable community structure of picophytoplankton across the Laurentian Great Lakes. Limnology and Oceanography, 68(10):2327–2345, October 2023. CODEN LIOCAH. ISSN 0024-3590.

Gefen-Treves:2020:ARS

[GTKW^{+}20] Shany Gefen-Treves, Isaac Kedem, Gad Weiss, Dirk Wagner, Dan Tchernov, and Aaron Kaplan. Acclimation of a rocky shore algal reef builder Neogoniolithon sp. to changing illuminations. Limnology and Oceanography, 65(1):27–36, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Gasparovic:2023:PRM

[GVF^{+}23] Blaženka Gašparović, Ivna Vrana, Sanja Frka, Daniela Marić Pfannkuchen, Ivan Vlašiček, Tamara Djakovac, Ingrid Ivančić, Mirta Smodlaka Tanković, Andrea Milinković, Lana Flanjak, Frédéric Chaux, Tihana Novak, Nikola Medić, and Jelena Godrijan. Paradox of relatively more phospholipids in phytoplankton in phosphorus limited sea. Limnology and Oceanography, 68(12):2800–2813, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Gauthier:2021:ECB

[GWS^{+}21] Joanna Gauthier, David Walsh, Daniel T. Selbie, Alyssa Bourgeois, Katherine Griffiths, Isabelle Domaizon, and Irene Gregory-Eaves. Evaluating the congruence between DNAbased and morphological taxonomic approaches in water and sediment trap samples: Analyses of a 36-month time series from a temperate monomictic lake. Limnology and Oceanography, 66(8):3020–3039, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Guo:2022:SDR

[GZX^{+}22] Chaoxuan Guo, Mengyuan Zhu, Hai Xu, Yunlin Zhang, Boqiang Qin, Guangwei Zhu, and Jianjun Wang. Spatiotemporal dependency of resource use efficiency on phytoplankton diversity in Lake Taihu. Limnology and Oceanography, 67(4): 830–842, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Hughes-Allen:2021:SPG

[HABL^{+}21] Lara Hughes-Allen, Frédéric Bouchard, Isabelle Laurion, Antoine Séjourné, Christelle Marlin, Christine Hatté, François Costard, Alexander Fedorov, and Alexey Desyatkin. Seasonal patterns in greenhouse gas emissions from thermokarst lakes in Central Yakutia (Eastern Siberia). Limnology and Oceanography, 66(S1):S98–S116, February 2021. CODEN LI-OCAH. ISSN 0024-3590.

Hambright:2022:E

[Ham22] K. David Hambright. Editorial. Limnology and Oceanography, 67(3):523–526, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Hedberg:2021:ECP

[HANW21] Per Hedberg, Séréna Albert, Francisco J. A. Nascimento, and Monika Winder. Effects of changing phytoplankton species composition on carbon and nitrogen uptake in benthic invertebrates. Limnology and Oceanography, 66(2):469–480, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Haentjens:2022:PSD

[HBG^{+}22] Nils Haëntjens, Emmanuel S. Boss, Jason R. Graff, Alison P. Chase, and Lee Karp-Boss. Phytoplankton size distributions in the western North Atlantic and their seasonal variability. Limnology and Oceanography, 67(8):1865–1878, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Hylen:2022:EBN

[HBR^{+}22] Astrid Hylén, Stefano Bonaglia, Elizabeth Robertson, Ugo Marzocchi, Mikhail Kononets, and Per O. J. Hall. Enhanced benthic nitrous oxide and ammonium production after natural oxygenation of long-term anoxic sediments. Limnology and Oceanography, 67(2):419–433, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Hallstrom:2022:PFD

[HBS^{+}22] Søren Hallstrøm, Mar Benavides, Ellen R. Salamon, Clayton W. Evans, Lindsey J. Potts, Julie Granger, Craig R. Tobias, Pia H. Moisander, and Lasse Riemann. Pelagic N_{2} fixation dominated by sediment diazotrophic communities in a shallow temperate estuary. Limnology and Oceanography, 67 (2):364–378, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Havird:2022:GEA

[HBY^{+}22] Justin C. Havird, Pamela M. Brannock, Reyn M. Yoshioka, Rebecca C. Vaught, Kaile’a Carlson, Collin Edwards, Allison Tracy, Cornelia W. Twining, Yun Zheng, David Chai, Alan E. Wilson, Nelson G. Hairston, and Scott R. Santos. Grazing by an endemic atyid shrimp controls microbial communities in the Hawaiian anchialine ecosystem. Limnology and Oceanography, 67(9):2012–2027, September 2022. CO-DEN LIOCAH. ISSN 0024-3590.

Halfter:2022:SDH

[HCB^{+}22] Svenja Halfter, Emma L. Cavan, Philip Butterworth, Kerrie M. Swadling, and Philip W. Boyd. “Sinking dead” — how zooplankton carcasses contribute to particulate organic carbon flux in the subantarctic Southern Ocean. Limnology and Oceanography, 67(1):13–25, January 2022. CODEN LI-OCAH. ISSN 0024-3590.

Hetherington:2024:VTS

[HCH^{+}24] Elizabeth D. Hetherington, Hilary G. Close, Steven H. D. Haddock, Alejandro Damian-Serrano, Casey W. Dunn, Natalie J. Wallsgrove, Shannon C. Doherty, and C. Anela Choy. Vertical trophic structure and niche partitioning of gelatinous predators in a pelagic food web: Insights from stable isotopes of siphonophores. Limnology and Oceanography, 69 (4):902–919, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Hunter-Cevera:2020:SS

[HCNO^{+}20] Kristen R. Hunter-Cevera, Michael G. Neubert, Robert J. Olson, Alexi Shalapyonok, Andrew R. Solow, and Heidi M. Sosik. Seasons of Syn. Limnology and Oceanography, 65(5): 1085–1102, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Hood:2021:LPL

[HCS^{+}21] James M. Hood, Lyndsie M. Collis, John D. Schade, Rebecca A. Stark, and Jacques C. Finlay. Longitudinal patterns and linkages in benthic fine particulate organic matter composition, respiration, and nutrient uptake. Limnology and Oceanography, 66(7):2684–2696, July 2021. CODEN LI-OCAH. ISSN 0024-3590.

Hillebrand:2022:TDW

[HDD^{+}22] Helmut Hillebrand, Josie Antonucci Di Carvalho, Jan-Claas Dajka, Claus-Dieter Dürselen, Onur Kerimoglu, Lucie Kuczynski, Lena Rönn, and Alexey Ryabov. Temporal declines in Wadden Sea phytoplankton cell volumes observed within and across species. Limnology and Oceanography, 67 (2):468–481, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Herr:2021:PRD

[HDK^{+}21] Alysia E. Herr, John W. H. Dacey, Ronald P. Kiene, Ross D. McCulloch, Nina Schuback, and Philippe D. Tortell. Potential roles of dimethysulfoxide in regional sulfur cycling and phytoplankton physiological ecology in the NE Subarctic Pacific. Limnology and Oceanography, 66(1):76–94, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Huang:2020:PMR

[HDS^{+}20] Ruiping Huang, Jiancheng Ding, Jiazhen Sun, Yang Tian, Chris Bowler, Xin Lin, and Kunshan Gao. Physiological and molecular responses to ocean acidification among strains of a model diatom. Limnology and Oceanography, 66(6):2926– 2936, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Hansson:2020:DCS

[HEH^{+}20] Lars-Anders Hansson, Mattias K. Ekvall, Liang He, Zhongqiang Li, Marie Svensson, Pablo Urrutia-Cordero, and Huan Zhang. Different climate scenarios alter dominance patterns among aquatic primary producers in temperate systems. Limnology and Oceanography, 65(10):2328–2336, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Harada:2020:SII

[HFL^{+}20] Yota Harada, Brian Fry, Shing Yip Lee, Damien T. Maher, James Z. Sippo, and Rod M. Connolly. Stable isotopes indicate ecosystem restructuring following climate-driven mangrove dieback. Limnology and Oceanography, 65(6):1251– 1263, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Houston:2024:BCA

[HGA24] Alex Houston, Mark H. Garnett, and William E. N. Austin. Blue carbon additionality: New insights from the radiocarbon content of saltmarsh soils and their respired CO_{2}. Limnology and Oceanography, 69(3):548–561, March 2024. CO-DEN LIOCAH. ISSN 0024-3590.

Hubot:2021:ENA

[HGF^{+}21] Nathan D. Hubot, Sarah L. C. Giering, Jessika Füssel, Julie Robidart, Antony Birchill, Mark Stinchcombe, Cynthia Dumousseaud, and Cathy H. Lucas. Evidence of nitrification associated with globally distributed pelagic jellyfish. Limnology and Oceanography, 66(6):2159–2173, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Huryn:2021:AFN

[HGH^{+}21] Alexander D. Huryn, Michael N. Gooseff, Patrick J. Hendrickson, Martin A. Briggs, Ken D. Tape, and Neil C. Terry. Aufeis fields as novel groundwater-dependent ecosystems in the Arctic cryosphere. Limnology and Oceanography, 66(3): 607–624, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Hilt:2022:PRS

[HGMK22] Sabine Hilt, Hans-Peter Grossart, Daniel F. McGinnis, and Frank Keppler. Potential role of submerged macrophytes for oxic methane production in aquatic ecosystems. Limnology and Oceanography, 67(S2):S76–S88, November 2022. CO-DEN LIOCAH. ISSN 0024-3590.

Holgerson:2022:IEM

[HHK^{+}22] Meredith A. Holgerson, Rachel A. Hovel, Patrick T. Kelly, Lauren E. Bortolotti, Jennifer A. Brentrup, Amber R. Bellamy, Samantha K. Oliver, and Alexander J. Reisinger. Integrating ecosystem metabolism and consumer allochthony reveals nonlinear drivers in lake organic matter processing. Limnology and Oceanography, 67(S1):S71–S85, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Haviland:2022:VSS

[HHMH22] Katherine Ann Haviland, Robert Warren Howarth, Roxanne Marino, and Melanie Hayn. Variation in sediment and seagrass characteristics reflect multiple stressors along a nitrogen-enrichment gradient in a New England lagoon. Limnology and Oceanography, 67(3):660–672, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Hannah:2024:ODD

[HJWP24] Charles G. Hannah, Sophia C. Johannessen, Cynthia A. Wright, and Stephen J. Page. Oxygen dynamics in a deepsilled fjord: Tight coupling to the open shelf. Limnology and Oceanography, 69(3):652–666, March 2024. CODEN LI-OCAH. ISSN 0024-3590.

Hollister:2020:RMT

[HKM^{+}20] Adrienne P. Hollister, Makenzie Kerr, Kema Malki, Eric Muhlbach, Maya Robert, Charles L. Tilney, Mya Breitbart, Katherine A. Hubbard, and Kristen N. Buck. Regeneration of macronutrients and trace metals during phytoplankton decay: an experimental study. Limnology and Oceanography, 65(8):1936–1960, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Hattich:2023:TVE

[HLH^{+}23] Giannina S. I. Hattich, Luisa Listmann, Jonathan Havenhand, Thorsten B. H. Reusch, and Birte Matthiessen. Temporal variation in ecological and evolutionary contributions to phytoplankton functional shifts. Limnology and Oceanography, 68(2):297–306, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Hauksson:2023:EMS

[HLKD23] Niels Hauksson, Christian B. Lewis, Tomoko Komada, and Ellen R. M. Druffel. Effect of marine sediment on the phase partitioning and isotopic content of riverine DOC. Limnology and Oceanography, 68(9):2008–2021, September 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Hillman:2020:RLM

[HLPT20] Jenny R. Hillman, Carolyn J. Lundquist, Conrad A. Pilditch, and Simon F. Thrush. The role of large macrofauna in mediating sediment erodibility across sedimentary habitats. Limnology and Oceanography, 65(4):683–693, April 2020. CO-DEN LIOCAH. ISSN 0024-3590.

Hurst:2022:OCD

[HLS^{+}22] Nia R. Hurst, Bryan Locher, Havalend E. Steinmuller, Linda J. Walters, and Lisa G. Chambers. Organic carbon dynamics and microbial community response to oyster reef restoration. Limnology and Oceanography, 67(5):1157–1168, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Ho:2020:ETP

[HM20] Jeff C. Ho and Anna M. Michalak. Exploring temperature and precipitation impacts on harmful algal blooms across continental U.S. lakes. Limnology and Oceanography, 65(5): 992–1009, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Hope:2020:IBS

[HMB^{+}20] J. A. Hope, J. Malarkey, J. H. Baas, J. Peakall, D. R. Parsons, A. J. Manning, S. J. Bass, I. D. Lichtman, P. D. Thorne, L. Ye, and D. M. Paterson. Interactions between sediment microbial ecology and physical dynamics drive heterogeneity in contextually similar depositional systems. Limnology and Oceanography, 65(10):2403–2419, October 2020. CODEN LI-OCAH. ISSN 0024-3590.

Hoffman:2022:RIN

[HMB^{+}22] Daniel K. Hoffman, Mark J. McCarthy, Ashlynn R. Boedecker, Justin A. Myers, and Silvia E. Newell. The role of internal nitrogen loading in supporting non-N-fixing harmful cyanobacterial blooms in the water column of a large eutrophic lake. Limnology and Oceanography, 67(9):2028–2041, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Herstoff:2021:LDM

[HMBB21] Emily M. Herstoff, Cédric L. Meunier, Maarten Boersma, and Stephen B. Baines. Leveraging differences in multiple prey traits allows selective copepods to meet their threshold elemental ratios. Limnology and Oceanography, 66(7):2914– 2922, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Hampel:2022:DSW

[HMM^{+}22] Justyna J. Hampel, Rachel D. Moseley, Rachel L. Mugge, Anirban Ray, Melanie Damour, Douglas Jones, and Leila J. Hamdan. Deep-sea wooden shipwrecks influence sediment microbiome diversity. Limnology and Oceanography, 67(2): 482–497, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Hatje:2021:BCS

[HMP^{+}21] Vanessa Hatje, Pere Masqué, Vinicus F. Patire, Antonio Dórea, and Francisco Barros. Blue carbon stocks, accumulation rates, and associated spatial variability in Brazilian mangroves. Limnology and Oceanography, 66(2):321–334, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Hiltunen:2021:ILC

[HNS21] Minna Hiltunen, Hannu Nykänen, and Jari Syväranta. The influence of lipid content and taxonomic affiliation on methane and carbon dioxide production from phytoplankton biomass in lake sediment. Limnology and Oceanography, 66 (5):1915–1925, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Hildebrand:2022:TDO

[HOB^{+}22] Tabea Hildebrand, Helena Osterholz, Carina Bunse, Hendrik Grotheer, Thorsten Dittmar, and Peter J. Schupp. Transformation of dissolved organic matter by two Indo-Pacific sponges. Limnology and Oceanography, 67(11):2483–2496, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Hornak:2020:HRD

[HP20] Karel Horñák and Jakob Pernthaler. Homeostatic regulation of dissolved labile organic substrates by consumption and release processes in a freshwater lake. Limnology and Oceanography, 65(5):939–950, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Hermans:2021:CDI

[HPB^{+}21] Martijn Hermans, Marina Astudillo Pascual, Thilo Behrends, Wytze K. Lenstra, Daniel J. Conley, and Caroline P. Slomp. Coupled dynamics of iron, manganese, and phosphorus in brackish coastal sediments populated by cable bacteria. Limnology and Oceanography, 66(7):2611–2631, July 2021. CO-DEN LIOCAH. ISSN 0024-3590.

Hochberg:2024:LUE

[HPCH24] Eric J. Hochberg, Chiara Pisapia, Robert C. Carpenter, and Siarah Hall. Light-use efficiency for coral reef communities and benthic functional types. Limnology and Oceanography, 69(3):712–722, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Hinson:2023:MAV

[HPF^{+}23] Audra Hinson, Spiro Papoulis, Lucas Fiet, Margaret Knight, Priscilla Cho, Brielle Szeltner, Ioannis Sgouralis, and David Talmy. A model of algal-virus population dynamics reveals underlying controls on material transfer. Limnology and Oceanography, 68(1):165–180, January 2023. CODEN LI-OCAH. ISSN 0024-3590.

Holtermann:2020:IOI

[HPNU20] Peter Holtermann, Ralf Prien, Michael Naumann, and Lars Umlauf. Interleaving of oxygenized intrusions into the Baltic Sea redoxcline. Limnology and Oceanography, 65(3):482–503, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Hopkins:2021:CPB

[HPP^{+}21] Joanne E. Hopkins, Matthew R. Palmer, Alex J. Poulton, Anna E. Hickman, and Jonathan Sharples. Control of a phytoplankton bloom by wind-driven vertical mixing and light availability. Limnology and Oceanography, 66(5):1926–1949, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Haas:2022:CNI

[HRK^{+}22] Sebastian Haas, Subhadeep Rakshit, Tim Kalvelage, Carolyn Buchwald, Christopher K. Algar, and Douglas W. R. Wallace. Characterization of nitrogen isotope fractionation during nitrification based on a coastal time series. Limnology and Oceanography, 67(8):1714–1731, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Hrycik:2021:UIM

[HS21] Allison R. Hrycik and Jason D. Stockwell. Under-ice mesocosms reveal the primacy of light but the importance of zooplankton in winter phytoplankton dynamics. Limnology and Oceanography, 66(2):481–495, February 2021. CODEN LI-OCAH. ISSN 0024-3590.

Hintz:2020:CID

[HSB^{+}20] William D. Hintz, Matthew S. Schuler, Jonathan J. Borrelli, Lawrence W. Eichler, Aaron B. Stoler, Vincent W. Moriarty, Laurie E. Ahrens, Charles W. Boylen, Sandra A. Nierzwicki-Bauer, and Rick A. Relyea. Concurrent improvement and deterioration of epilimnetic water quality in an oligotrophic lake over 37 years. Limnology and Oceanography, 65(5):927– 938, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Herbert:2020:ENP

[HSBC20] Ellen R. Herbert, Joseph P. Schubauer-Berigan, and Christopher B. Craft. Effects of 10 yr of nitrogen and phosphorus fertilization on carbon and nutrient cycling in a tidal freshwater marsh. Limnology and Oceanography, 65(8):1669–1687, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Hoer:2020:SRM

[HSD^{+}20] Daniel R. Hoer, William Sharp, Gabriel Delgado, Niels L. Lindquist, and Christopher S. Martens. Sponges represent a major source of inorganic nitrogen in Florida Bay (U.S.A.). Limnology and Oceanography, 65(6):1235–1250, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Hollarsmith:2020:ESU

[HSP^{+}20] Jordan A. Hollarsmith, Jason S. Sadowski, Manon M. M. Picard, Brian Cheng, James Farlin, Ann Russell, and Edwin D. Grosholz. Effects of seasonal upwelling and runoff on water chemistry and growth and survival of native and commercial oysters. Limnology and Oceanography, 65(2):224–235, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Hunt:2022:CBC

[HSV22] Christopher W. Hunt, Joseph E. Salisbury, and Douglas Vandemark. Controls on buffering and coastal acidification in a temperate estuary. Limnology and Oceanography, 67(6): 1328–1342, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Hall:2023:ADE

[HTLP23] Nathan Hall, Jeremy Testa, Ming Li, and Hans Paerl. Assessing drivers of estuarine pH: a comparative analysis of the continental U.S.A.’s two largest estuaries. Limnology and Oceanography, 68(10):2227–2244, October 2023. CODEN LI-OCAH. ISSN 0024-3590.

Tuo:2020:HAF

[hTRH20] Sing how Tuo, Irene B. Rodriguez, and Tung-Yuan Ho. H_{2} accumulation and N_{2} fixation variation by Ni limitation in Cyanothece. Limnology and Oceanography, 65(2):377–386, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Hernes:2021:ECA

[HTSG21] Peter J. Hernes, Suzanne E. Tank, Mikael K. Sejr, and Ronnie N. Glud. Element cycling and aquatic function in a changing Arctic. Limnology and Oceanography, 66(S1):S1– S16, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Huang:2021:SSF

[HVMS^{+}21] Guoyin Huang, Silvia Vidal-Melgosa, Andreas Sichert, Stefan Becker, Yang Fang, Jutta Niggemann, Morten Hvitfeldt Iversen, Yi Cao, and Jan-Hendrik Hehemann. Secretion of sulfated fucans by diatoms may contribute to marine aggregate formation. Limnology and Oceanography, 66(10):3768– 3782, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Huang:2021:DCC

[HXL^{+}21] Jing Huang, Xiaoqing Xu, Da Li, Yunfei Sun, Lei Gu, Lu Zhang, Kai Lyu, and Zhou Yang. Decreased calcium concentration interferes with life history defense strategies of Ceriodaphnia cornuta in response to fish kairomone. Limnology and Oceanography, 66(8):3237–3252, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Hawco:2022:RDI

[HYPG^{+}22] Nicholas J. Hawco, Shun-Chung Yang, Paulina Pinedo-González, Erin E. Black, Jennifer Kenyon, Sara Ferrón, Xiaopeng Bian, and Seth G. John. Recycling of dissolved iron in the North Pacific Subtropical Gyre. Limnology and Oceanography, 67(11):2448–2465, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Heuvel:2024:IEP

[HZC^{+}24] Cecilia E. Heuvel, Yingming Zhao, Jan J. H. Ciborowski, Li Wang, Emilee Mancini, and Aaron T. Fisk. The influence of environmental parameters on spatial variation in zoobenthic density and stable isotopes (δ ^{13}C, δ ^{15}N, and δ ^{34}S) within a large lake. Limnology and Oceanography, 69(1):173– 188, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Herbert:2021:BIF

[HZM^{+}21] Lisa C. Herbert, Qingzhi Zhu, Alexander B. Michaud, Katja Laufer-Meiser, Christopher K. Jones, Natascha Riedinger, Zachery S. Stooksbury, Robert C. Aller, Bo Barker Jørgensen, and Laura M. Wehrmann. Benthic iron flux influenced by climate-sensitive interplay between organic carbon availability and sedimentation rate in Arctic fjords. Limnology and Oceanography, 66(9):3374–3392, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Hintz:2021:CSQ

[HZS21] Nils Hendrik Hintz, Moritz Zeising, and Maren Striebel. Changes in spectral quality of underwater light alter phytoplankton community composition. Limnology and Oceanography, 66(9):3327–3337, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Huang:2020:SMA

[HZZ^{+}20] Rui Huang, Jin Zeng, Dayong Zhao, Katherine V. Cook, K. David Hambright, and Zhongbo Yu. Sediment microbiomes associated with the rhizosphere of emergent macrophytes in a shallow, subtropical lake. Limnology and Oceanography, 65(S1):S38–S48, January 2020. CODEN LI-OCAH. ISSN 0024-3590.

Ibanhez:2021:FSS

[IÁSNCR21] J. Severino P. Ibánhez, Xosé Antón Álvarez-Salgado, Mar Nieto-Cid, and Carlos Rocha. Fresh and saline submarine groundwater discharge in a large coastal inlet affected by seasonal upwelling. Limnology and Oceanography, 66(6):2141– 2158, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Ishikawa:2020:CUR

[IFU^{+}20] Naoto F. Ishikawa, Jacques C. Finlay, Hiromi Uno, Nanako O. Ogawa, Naohiko Ohkouchi, Ichiro Tayasu, and Mary E. Power. Combined use of radiocarbon and stable carbon isotopes for the source mixing model in a stream food web. Limnology and Oceanography, 66(10):2688–2696, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Irion:2020:MSV

[IJSC20] Solène Irion, Ludwig Jardillier, Ingrid Sassenhagen, and Urania Christaki. Marked spatiotemporal variations in small phytoplankton structure in contrasted waters of the Southern Ocean (Kerguelen area). Limnology and Oceanography, 66(10):2835–2852, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Ikenoue:2021:NES

[IKN^{+}21] Takahito Ikenoue, Katsunori Kimoto, Yuriko Nakamura, Kjell R. Bjørklund, Naoki Kuramoto, Masaaki Ueki, Yuichi Ota, Jonaotaro Onodera, Naomi Harada, Makio C. Honda, Miyako Sato, Eiji Watanabe, Motoyo Itoh, Shigeto Nishino, and Takashi Kikuchi. New evaluation of species-specific biogenic silica flux of radiolarians (Rhizaria) in the western Arctic Ocean using microfocus X-ray computed tomography. Limnology and Oceanography, 66(11):3901–3915, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Imam:2020:BRW

[ILPL20] Yehya E. Imam, Bernard Laval, Roger Pieters, and Gregory Lawrence. The baroclinic response to wind in a multiarm multibasin reservoir. Limnology and Oceanography, 65(3): 582–600, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Iniesto:2022:PMC

[IMB^{+}22] Miguel Iniesto, David Moreira, Karim Benzerara, Guillaume Reboul, Paola Bertolino, Rosaluz Tavera, and Purificación López-García. Planktonic microbial communities from microbialite-bearing lakes sampled along a salinity-alkalinity gradient. Limnology and Oceanography, 67(12):2718–2733, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Itsukushima:2024:DRM

[IOS24] Rei Itsukushima, Kazuaki Ohtsuki, and Tatsuro Sato. Drivers of rising monthly water temperature in river estuaries. Limnology and Oceanography, 69(3):589–603, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Isabwe:2022:RBP

[IYW^{+}22] Alain Isabwe, Jun R. Yang, Yongming Wang, David M. Wilkinson, Emily B. Graham, Huihuang Chen, and Jun Yang. Riverine bacterioplankton and phytoplankton assembly along an environmental gradient induced by urbanization. Limnology and Oceanography, 67(9):1943–1958, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2022:SGT

[JHBB22] Robert A. Johnson, Kathleen M. Hanes, Alan B. Bolten, and Karen A. Bjorndal. Simulated green turtle grazing alters effects of environmental drivers on seagrass growth dynamics across seasons. Limnology and Oceanography, 67(12):2635– 2648, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2021:GRC

[JHS^{+}21] Mildred Jessica Johnson, Laura Margarethe Hennigs, Yvonne Sawall, Christian Pansch, and Marlene Wall. Growth response of calcifying marine epibionts to biogenic pH fluctuations and global ocean acidification scenarios. Limnology and Oceanography, 66(4):1125–1138, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Jiang:2024:NAR

[JHSG^{+}24] Meijia Jiang, Jason M. Hall-Spencer, Lin Gao, Zengling Ma, and Guang Gao. Nitrogen availability regulates the effects of a simulated marine heatwave on carbon sequestration and phycosphere bacteria of a marine crop. Limnology and Oceanography, 69(2):339–354, February 2024. CODEN LI-OCAH. ISSN 0024-3590.

Jones:2023:CNC

[JHvL^{+}23] Elizabeth M. Jones, Sian F. Henley, Maria A. van Leeuwe, Jacqueline Stefels, Michael P. Meredith, Mairi Fenton, and Hugh J. Venables. Carbon and nutrient cycling in Antarctic landfast sea ice from winter to summer. Limnology and Oceanography, 68(1):208–231, January 2023. CODEN LI-OCAH. ISSN 0024-3590.

Jorgensen:2021:BMH

[JLMW21] Bo Barker Jørgensen, Katja Laufer, Alexander B. Michaud, and Laura M. Wehrmann. Biogeochemistry and microbiology of high Arctic marine sediment ecosystems — case study of Svalbard fjords. Limnology and Oceanography, 66(S1):S273– S292, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2022:NRC

[JLR^{+}22] Roberta Johnson, Gerald Langer, Sergio Rossi, Ian Probert, Marta Mammone, and Patrizia Ziveri. Nutritional response of a coccolithophore to changing pH and temperature. Limnology and Oceanography, 67(10):2309–2324, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2020:MCS

[JLS^{+}20] Winifred M. Johnson, Krista Longnecker, Melissa C. Kido Soule, William A. Arnold, Maya P. Bhatia, Steven J. Hallam, Benjamin A. S. Van Mooy, and Elizabeth B. Kujawinski. Metabolite composition of sinking particles differs from surface suspended particles across a latitudinal transect in the South Atlantic. Limnology and Oceanography, 65(1):111– 127, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Jiang:2023:USI

[JLW^{+}23] Long Jiang, Xinyu Lu, Guifen Wang, Mo Peng, Aihong Wei, Yonggang Zhao, and Karline Soetaert. Unraveling seasonal and interannual nutrient variability shows exceptionally high human impact in eutrophic coastal waters. Limnology and Oceanography, 68(5):1161–1171, May 2023. CODEN LI-OCAH. ISSN 0024-3590.

Janatian:2020:ASO

[JOC^{+}20] Nasime Janatian, Kalle Olli, Fabien Cremona, Alo Laas, and Peeter Nõges. Atmospheric stilling offsets the benefits from reduced nutrient loading in a large shallow lake. Limnology and Oceanography, 65(4):717–731, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Jimenez-Ramos:2021:FMH

[JREVB21] Rocío Jiménez-Ramos, Luis G. Egea, Juan J. Vergara, and Fernando G. Brun. Factors modulating herbivory patterns in Cymodocea nodosa meadows. Limnology and Oceanography, 66(6):2218–2233, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Johnston:2020:HCD

[JSB^{+}20] Sarah Ellen Johnston, Robert G. Striegl, Matthew J. Bogard, Mark M. Dornblaser, David E. Butman, Anne M. Kellerman, Kimberly P. Wickland, David C. Podgorski, and Robert G. M. Spencer. Hydrologic connectivity determines dissolved organic matter biogeochemistry in northern high-latitude lakes. Limnology and Oceanography, 65(8):1764–1780, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Johnson:2023:PDM

[JSL^{+}23] Winifred M. Johnson, Melissa C. Kido Soule, Krista Longnecker, Maya P. Bhatia, Steven J. Hallam, Michael W. Lomas, and Elizabeth B. Kujawinski. Particulate and dissolved metabolite distributions along a latitudinal transect of the western Atlantic Ocean. Limnology and Oceanography, 68 (2):377–393, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Jansen:2024:WLF

[JSPS24] Lara S. Jansen, Daniel Sobota, Yangdong Pan, and Angela L. Strecker. Watershed, lake, and food web factors influence diazotrophic cyanobacteria in mountain lakes. Limnology and Oceanography, 69(3):681–699, March 2024. CODEN LI-OCAH. ISSN 0024-3590.

Jacobi:2021:EPS

[JSW^{+}21] Yuval Jacobi, Noa Shenkar, J. Evan Ward, Maria Rosa, Guy Z. Ramon, Uri Shavit, and Gitai Yahel. Evasive plankton: Size-independent particle capture by ascidians. Limnology and Oceanography, 66(4):1009–1020, April 2021. CODEN LIOCAH. ISSN 0024-3590.

James:2020:WMV

[JvKvT^{+}20] Rebecca K. James, Marieke M. van Katwijk, Brigitta I. van Tussenbroek, Tjisse van der Heide, Henk A. Dijkstra, René M. van Westen, Julie D. Pietrzak, Adam S. Candy, Roland Klees, Riccardo E. M. Riva, Cornelis D. Slobbe, Caroline A. Katsman, Peter M. J. Herman, and Tjeerd J. Bouma. Water motion and vegetation control the pH dynamics in seagrass-dominated bays. Limnology and Oceanography, 65 (2):349–362, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Joy-Warren:2022:SPR

[JWAvD^{+}22] Hannah L. Joy-Warren, Anne-Carlijn Alderkamp, Gert L. van Dijken, Loay J. Jabre, Erin M. Bertrand, Evan N. Baldonado, Molly W. Glickman, Kate M. Lewis, Rob Middag, Kyyas Seyitmuhammedov, Kate E. Lowry, Willem van de Poll, and Kevin R. Arrigo. Springtime phytoplankton responses to light and iron availability along the western Antarctic Peninsula. Limnology and Oceanography, 67(4): 800–815, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Jabbari:2021:SOP

[JYA21] Aidin Jabbari, Kazutaka Yanase, and Josef D. Ackerman. A spanwise oscillating plate in a crossflow: Implication for mass transfer and locomotion. Limnology and Oceanography, 66(9):3393–3407, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Kenitz:2023:EED

[KAC^{+}23] Kasia M. Kenitz, Clarissa R. Anderson, Melissa L. Carter, Emily Eggleston, Kristi Seech, Rebecca Shipe, Jayme Smith, Eric C. Orenstein, Peter J. S. Franks, Jules S. Jaffe, and Andrew D. Barton. Environmental and ecological drivers of harmful algal blooms revealed by automated underwater microscopy. Limnology and Oceanography, 68(3):598–615, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Kellerman:2020:FDD

[KAP^{+}20] Anne M. Kellerman, Ana Arellano, David C. Podgorski, Ellen E. Martin, Jonathan B. Martin, Kelly M. Deuerling, Thomas S. Bianchi, and Robert G. M. Spencer. Fundamental drivers of dissolved organic matter composition across an Arctic effective precipitation gradient. Limnology and Oceanography, 65(6):1217–1234, June 2020. CODEN LI-OCAH. ISSN 0024-3590.

Komada:2022:SFB

[KBB^{+}22] Tomoko Komada, Anthony Bravo, Mandy-Tanita Brinkmann, Kevin Lu, Lily Wong, and Gavin Shields. “Slow” and “fast” in blue carbon: Differential turnover of allochthonous and autochthonous organic matter in minerogenic salt marsh sediments. Limnology and Oceanography, 67(S2):S133–S147, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Krause:2020:IPB

[KBL^{+}20] Jeffrey W. Krause, Mark A. Brzezinski, John L. Largier, Heather M. McNair, Michael Maniscalco, Kay D. Bidle, Andrew E. Allen, and Kimberlee Thamatrakoln. The interaction of physical and biological factors drives phytoplankton spatial distribution in the northern California Current. Limnology and Oceanography, 65(9):1974–1989, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Kim:2021:RPP

[KBSA21] Jongsun Kim, Mark J. Brush, Bongkeun Song, and Iris C. Anderson. Reconstructing primary production in a changing estuary: a mass balance modeling approach. Limnology and Oceanography, 65(6):2535–2546, June 2021. CODEN LI-OCAH. ISSN 0024-3590.

Korte:2020:MDP

[KBvdD^{+}20] Laura F. Korte, Geert-Jan A. Brummer, Michèlle van der Does, Catarina V. Guerreiro, Furu Mienis, Chris I. Munday, Leandro Ponsoni, Stefan Schouten, and Jan-Berend W. Stuut. Multiple drivers of production and particle export in the western tropical North Atlantic. Limnology and Oceanography, 65(9):2108–2124, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Kaartvedt:2023:MSF

[KCT23] Stein Kaartvedt, Svenja Christiansen, and Josefin Titelman. Mid-summer fish behavior in a high-latitude twilight zone. Limnology and Oceanography, 68(7):1654–1669, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Kelly:2020:CVE

[KdED^{+}20] Seán Kelly, Elvira de Eyto, Mary Dillane, Russell Poole, and Martin White. Characterizing ventilation events in an anoxic coastal basin: Observed dynamics and the role of climatic drivers. Limnology and Oceanography, 65(10):2420–2442, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Kroeker:2023:EEM

[KDV^{+}23] Kristy J. Kroeker, Emily M. Donham, Kate Vylet, Joseph K. Warren, Julia Cheresh, Jerome Fiechter, Jan Freiwald, and Yuichiro Takeshita. Exposure to extremes in multiple global change drivers: Characterizing pH, dissolved oxygen, and temperature variability in a dynamic, upwelling dominated ecosystem. Limnology and Oceanography, 68(7):1611–1623, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Kenyon:2023:CRD

[KDW^{+}23] Tania M. Kenyon, Christopher Doropoulos, Kennedy Wolfe, Gregory E. Webb, Sophie Dove, Daniel Harris, and Peter J. Mumby. Coral rubble dynamics in the Anthropocene and implications for reef recovery. Limnology and Oceanography, 68(1):110–147, January 2023. CODEN LIOCAH. ISSN 0024-3590.

Knoke:2024:ORP

[KDZ^{+}24] Melina Knoke, Thorsten Dittmar, Oliver Zielinski, Morimaru Kida, Nils E. Asp, Carlos Eduardo de Rezende, Bernhard Schnetger, and Michael Seidel. Outwelling of reduced porewater drives the biogeochemistry of dissolved organic matter and trace metals in a major mangrove-fringed estuary in Amazonia. Limnology and Oceanography, 69(2):262–278, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Knoll:2024:RSI

[KFHS24] Lesley B. Knoll, Benton Fry, Nicole M. Hayes, and Hailey M. Sauer. Reduced snow and increased nutrients show enhanced ice-associated photoautotrophic growth using a modified experimental under-ice design. Limnology and Oceanography, 69(1):203–216, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Kunisch:2023:WWD

[KGG^{+}23] Erin H. Kunisch, Martin Graeve, Rolf Gradinger, Hauke Flores, Øystein Varpe, and Bodil A. Bluhm. What we do in the dark: Prevalence of omnivorous feeding activity in Arctic zooplankton during polar night. Limnology and Oceanography, 68(8):1835–1851, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Keister:2022:ZCL

[KHB22] Julie E. Keister, BethElLee Herrmann, and Julia Bos. Zooplankton composition links to climate and salmon survival in a northern temperate fjord. Limnology and Oceanography, 67(11):2389–2404, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Krause:2022:BHB

[KHCG^{+}22] Johannes Renke Krause, Alejandro Hinojosa-Corona, Andrew B. Gray, Juan Carlos Herguera, Julianna McDonnell, Michael V. Schaefer, Samantha C. Ying, and Elizabeth Burke Watson. Beyond habitat boundaries: Organic matter cycling requires a system-wide approach for accurate blue carbon accounting. Limnology and Oceanography, 67(S2):S6–S18, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Kleinteich:2020:SCM

[KHHZ20] Julia Kleinteich, Sabine Hilt, Andreas Hoppe, and Christiane Zarfl. Structural changes of the microplankton community following a pulse of inorganic nitrogen in a eutrophic river. Limnology and Oceanography, 66(6):S264–S276, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Kammann:2022:UHN

[KHK^{+}22] Sandra Kammann, Daniel Arturo Saavedra Hortua, John S. Kominoski, Theresa-Marie Fett, and Lucy Gwen Gillis. Understanding how nutrient limitation and plant traits influence carbon in mangrove-seagrass coastal ecosystems. Limnology and Oceanography, 67(S2):S89–S103, November 2022. CO-DEN LIOCAH. ISSN 0024-3590.

Krempaska:2021:SMS

[KHSP21] Natalia Krempaska, Karel Hornńák, Marisa O. D. Silva, and Jakob Pernthaler. Spatial microheterogeneity and selective microbial consumption of dissolved free amino acids in an oligomesotrophic lake. Limnology and Oceanography, 66(10): 3728–3739, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Kirchman:2020:EML

[Kir20] David L. Kirchman. Editorial: Metagenomics in Limnology and Oceanography. Limnology and Oceanography, 65(S1):S1, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Kakouei:2022:VPL

[KKA22] Karan Kakouei, Benjamin M. Kraemer, and Rita Adrian. Variation in the predictability of lake plankton metric types. Limnology and Oceanography, 67(3):608–620, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Kling:2021:IMT

[KKP^{+}21] Joshua D. Kling, Kyla J. Kelly, Sophia Pei, Tatiana A. Rynearson, and David A. Hutchins. Irradiance modulates thermal niche in a previously undescribed low-light and coldadapted nano-diatom. Limnology and Oceanography, 66(6): 2266–2277, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Koski:2022:FRA

[KL22] Marja Koski and Fabien Lombard. Functional responses of aggregate-colonizing copepods. Limnology and Oceanography, 67(9):2059–2072, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Kohlbach:2023:DAZ

[KLG^{+}23] Doreen Kohlbach, Benoit Lebreton, Gaël Guillou, Anette Wold, Haakon Hop, Martin Graeve, and Philipp Assmy. Dependency of Arctic zooplankton on pelagic food sources: New insights from fatty acid and stable isotope analyses. Limnology and Oceanography, 68(10):2346–2358, October 2023. CODEN LIOCAH. ISSN 0024-3590.

Kahru:2023:MCO

[KLO23] Mati Kahru, Zhongping Lee, and Mark D. Ohman. Multidecadal changes in ocean transparency: Decrease in a coastal upwelling region and increase offshore. Limnology and Oceanography, 68(7):1546–1556, July 2023. CODEN LI-OCAH. ISSN 0024-3590.

Kong:2023:ILA

[KLP23] Liangliang Kong, Meizhen Li, and Neil M. Price. The influence of light on the availability of photoreactive and photostable iron(III)–ligand complexes to diatoms. Limnology and Oceanography, 68(5):1091–1104, May 2023. CODEN LI-OCAH. ISSN 0024-3590.

Kellogg:2022:LZC

[KMM^{+}22] Riss M. Kellogg, Dawn M. Moran, Matthew R. McIlvin, Adam V. Subhas, Andrew E. Allen, and Mak A. Saito. Lack of a Zn/Co substitution ability in the polar diatom Chaetoceros neogracile RS19. Limnology and Oceanography, 67(10): 2265–2280, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Kellogg:2020:EZC

[KMV^{+}20] Riss M. Kellogg, Matthew R. McIlvin, Jagruti Vedamati, Benjamin S. Twining, James W. Moffett, Adrian Marchetti, Dawn M. Moran, and Mak A. Saito. Efficient zinc/cobalt inter-replacement in northeast Pacific diatoms and relationship to high surface dissolved Co : Zn ratios. Limnology and Oceanography, 65(11):2557–2582, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Krieger:2023:RIC

[KNG^{+}23] Erik C. Krieger, Wendy A. Nelson, Johan Grand, Eric C. Le Ru, Sarah J. Bury, Amelie Cossais, Simon K. Davy, and Christopher E. Cornwall. The role of irradiance in controlling coralline algal calcification. Limnology and Oceanography, 68 (6):1269–1284, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Kenitz:2020:EDP

[KOR^{+}20] Kasia M. Kenitz, Eric C. Orenstein, Paul L. D. Roberts, Peter J. S. Franks, Jules S. Jaffe, Melissa L. Carter, and Andrew D. Barton. Environmental drivers of population variability in colony-forming marine diatoms. Limnology and Oceanography, 65(10):2515–2528, October 2020. CODEN LI-OCAH. ISSN 0024-3590.

Kong:2020:RDC

[KP20] Liangliang Kong and Neil M. Price. A reduction-dependent copper uptake pathway in an oceanic diatom. Limnology and Oceanography, 65(3):601–611, March 2020. CODEN LI-OCAH. ISSN 0024-3590.

Kivila:2023:VLB

[KPG^{+}23] E. Henriikka Kivilä, Vilmantas Prėskienis, Noémie Gaudreault, Catherine Girard, and Milla Rautio. Variability in lake bacterial growth and primary production under lake ice: Evidence from early winter to spring melt. Limnology and Oceanography, 68(11):2603–2616, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Kiljunen:2020:BPC

[KPL^{+}20] Mikko Kiljunen, Heikki Peltonen, Maiju Lehtiniemi, Laura Uusitalo, Tuula Sinisalo, Joanna Norkko, Mervi Kunnasranta, Jyrki Torniainen, Antti J. Rissanen, and Juha Karjalainen. Benthic-pelagic coupling and trophic relationships in northern Baltic Sea food webs. Limnology and Oceanography, 65(8):1706–1722, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Kazmierczak:2020:GCP

[KPM^{+}20] Jolanta Kazmierczak, Dieke Postma, Sascha Müller, Søren Jessen, Bertel Nilsson, Joanna Czekaj, and Peter Engesgaard. Groundwater-controlled phosphorus release and transport from sandy aquifer into lake. Limnology and Oceanography, 66(5):2188–2204, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Kulkarni:2023:LSA

[KPNJ23] Mihir R. Kulkarni, Shruti V. Paripatyadar, Rohit Naniwadekar, and Jahnavi Joshi. Local-scale abiotic factors influence the organization of rock pool arthropod communities in a biodiversity hotspot. Limnology and Oceanography, 68(10):2375–2388, October 2023. CODEN LIOCAH. ISSN 0024-3590.

Kurtay:2021:PNC

[KPS21] Gulce Kurtay, Hans J. Prevost, and Beth A. Stauffer. Picoand nanoplankton communities on a near to offshore transect along the continental shelf of the northwestern Gulf of Mexico in the aftermath of Hurricane Harvey. Limnology and Oceanography, 66(7):2779–2796, July 2021. CODEN LI-OCAH. ISSN 0024-3590.

Kamjunke:2021:HIL

[KRB^{+}21] Norbert Kamjunke, Michael Rode, Martina Baborowski, Julia Vanessa Kunz, Jakob Zehner, Dietrich Borchardt, and Markus Weitere. High irradiation and low discharge promote the dominant role of phytoplankton in riverine nutrient dynamics. Limnology and Oceanography, 66(7):2648–2660, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Kharbush:2023:PSF

[KRC23] Jenan J. Kharbush, Rebecca S. Robinson, and Susan J. Carter. Patterns in sources and forms of nitrogen in a large eutrophic lake during a cyanobacterial harmful algal bloom. Limnology and Oceanography, 68(4):803–815, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Kreuzburg:2020:CRT

[KRM^{+}20] Matthias Kreuzburg, Fereidoun Rezanezhad, Tatjana Milojevic, Maren Voss, Lennart Gosch, Susanne Liebner, Philippe Van Cappellen, and Gregor Rehder. Carbon release and transformation from coastal peat deposits controlled by submarine groundwater discharge: a column experiment study. Limnology and Oceanography, 65(5):1116–1135, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Khoo:2023:SFB

[KSF^{+}23] Celyn L. L. Khoo, Rachel E. Sipler, Sally J. M. Faulkner, Sean G. Boyd, Maryam Beheshti Foroutani, Colleen E. McBride, and Susan E. Ziegler. Size fractionated biogeochemical constituents across adjacent coastal systems informs approaches for integrating small catchment studies into regional models. Limnology and Oceanography, 68(6):1285– 1300, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Kuhn:2023:HEM

[KSH^{+}23] McKenzie A. Kuhn, Megan Schmidt, Liam Heffernan, Jördis Stührenberg, Klaus-Holger Knorr, Cristian Estop-Aragonés, Tanja Broder, Regina Gonzalez Moguel, Peter M. J. Douglas, and David Olefeldt. High ebullitive, millennial-aged greenhouse gas emissions from thermokarst expansion of peatland lakes in boreal western Canada. Limnology and Oceanography, 68(2):498–513, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Kimmerer:2024:CEP

[KSI24] Wim J. Kimmerer, Anne M. Slaughter, and Toni R. Ignoffo. Copepod egg production estimates are biased by female mortality. Limnology and Oceanography, 69(2):408–421, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Kagami:2021:SDD

[KSN^{+}21] Maiko Kagami, Kensuke Seto, Daiki Nozaki, Takaki Nakamura, Hirano Wakana, and Christian Wurzbacher. Single dominant diatom can host diverse parasitic fungi with different degree of host specificity. Limnology and Oceanography, 66(3):667–677, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Koestner:2020:AEP

[KSR20] Daniel Koestner, Dariusz Stramski, and Rick A. Reynolds. Assessing the effects of particle size and composition on light scattering through measurements of size-fractionated seawater samples. Limnology and Oceanography, 65(1):173–190, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Kelly:2021:HPP

[KTA^{+}21] Patrick T. Kelly, Jason M. Taylor, Isabelle M. Andersen, Jasmine Stovall, and J. Thad Scott. Highest primary production achieved at high nitrogen levels despite strong stoichiometric imbalances with phosphorus in hypereutrophic experimental systems. Limnology and Oceanography, 66(12):4375–4390, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Kokubun:2021:SIO

[KTH^{+}21] Nobuo Kokubun, Yukiko Tanabe, Daisuke Hirano, Vigan Mensah, Takeshi Tamura, Shigeru Aoki, and Akinori Takahashi. Shoreward intrusion of oceanic surface waters alters physical and biological ocean structures on the Antarctic continental shelf during winter: Observations from instrumented seals. Limnology and Oceanography, 66(10):3740–3753, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Kottmeier:2020:DVC

[KTWGT20] Dorothee M. Kottmeier, Anja Terbrüggen, Dieter A. Wolf-Gladrow, and Silke Thoms. Diel variations in cell division and biomass production of Emiliania huxleyi — consequences for the calculation of physiological cell parameters. Limnology and Oceanography, 65(8):1781–1800, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Klaus:2022:DBS

[KVKS22] Marcus Klaus, Hendricus A. Verheijen, Jan Karlsson, and David A. Seekell. Depth and basin shape constrain ecosystem metabolism in lakes dominated by benthic primary producers. Limnology and Oceanography, 67(12):2763–2778, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Kreczak:2021:SDB

[KWB21] Hannah Kreczak, Andrew J. Willmott, and Andrew W. Baggaley. Subsurface dynamics of buoyant microplastics subject to algal biofouling. Limnology and Oceanography, 66(9): 3287–3299, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Kleint:2021:ISS

[KWS^{+}21] Jan F. Kleint, Yannic Wellach, Moritz Schroll, Frank Keppler, and Margot Isenbeck-Schröter. The impact of seasonal sulfate–methane transition zones on methane cycling in a sulfate-enriched freshwater environment. Limnology and Oceanography, 66(6):2290–2308, June 2021. CODEN LI-OCAH. ISSN 0024-3590.

Kang:2022:SVL

[KX22] Xinyi Kang and Meng Xia. Stratification variability in a lagoon system in response to a passing storm. Limnology and Oceanography, 67(2):511–521, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Lovelock:2022:MAE

[LAB^{+}22] Catherine E. Lovelock, M. Fernanda Adame, Don W. Butler, Jeffrey J. Kelleway, Sabine Dittmann, Benedikt Fest, Karen J. King, Peter I. Macreadie, Katherine Mitchell, Mark Newnham, Anne Ola, Christopher J. Owers, and Nina Welti. Modeled approaches to estimating blue carbon accumulation with mangrove restoration to support a blue carbon accounting method for Australia. Limnology and Oceanography, 67 (S2):S50–S60, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Linz:2020:TSM

[LABM20] Alexandra M. Linz, Frank O. Aylward, Stefan Bertilsson, and Katherine D. McMahon. Time-series metatranscriptomes reveal conserved patterns between phototrophic and heterotrophic microbes in diverse freshwater systems. Limnology and Oceanography, 65(S1):S101–S112, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Ladwig:2022:LTC

[LAD^{+}22] Robert Ladwig, Alison P. Appling, Austin Delany, Hilary A. Dugan, Qiantong Gao, Noah Lottig, Jemma Stachelek, and Paul C. Hanson. Long-term change in metabolism phenology in north temperate lakes. Limnology and Oceanography, 67 (7):1502–1521, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Langbehn:2023:VPR

[LAFV23] Tom J. Langbehn, Johanna M. Aarflot, Jennifer J. Freer, and Øystein Varpe. Visual predation risk and spatial distributions of large Arctic copepods along gradients of sea ice and bottom depth. Limnology and Oceanography, 68(6): 1388–1405, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Lopez-Acosta:2022:SCS

[LAMG^{+}22] María López-Acosta, Manuel Maldonado, Jacques Grall, Axel Ehrhold, Cèlia Sitjà, Cristina Galobart, Fiz F. Pérez, and Aude Leynaert. Sponge contribution to the silicon cycle of a diatom-rich shallow bay. Limnology and Oceanography, 67(11):2431–2447, November 2022. CODEN LIOCAH. ISSN 0024-3590.

LaBrie:2021:SAD

[LBBM21] Richard LaBrie, Simon Bélanger, Ronald Benner, and Roxane Maranger. Spatial abundance distribution of prokaryotes is associated with dissolved organic matter composition and ecosystem function. Limnology and Oceanography, 66(3): 575–587, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Lichtschlag:2022:ISW

[LBG^{+}22] Anna Lichtschlag, Ulrike Braeckman, Katja Guilini, Soeren Ahmerkamp, Rachael H. James, and Dirk de Beer. Impact of shallow-water hydrothermal seepage on benthic biogeochemical cycling, nutrient availability, and meiobenthic communities in a tropical coral reef. Limnology and Oceanography, 67(3):567–584, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Lachapelle:2022:PRS

[LBJS22] Josianne Lachapelle, Elvire Bestion, Eleanor E. Jackson, and C.-Elisa Schaum. Presence of a resident species aids invader evolution. Limnology and Oceanography, 67(10):2252–2264, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Liu:2024:SSP

[LBL^{+}24] Haoran Liu, Thomas J. Browning, Edward A. Laws, Yibin Huang, Lei Wang, Yiwei Shang, Xiaogang Xing, Kuanbo Zhou, Zong-Pei Jiang, Xin Liu, Bangqin Huang, and Minhan Dai. Stimulation of small phytoplankton drives enhanced sinking particle formation in a subtropical ocean eddy. Limnology and Oceanography, 69(4):834–847, April 2024. CO-DEN LIOCAH. ISSN 0024-3590.

Lange:2023:MSD

[LBvE23] Jacqueline Lange, Amelie Cara Berges, and Eric von Elert. Multiclonal study of Daphnia magna with respect to adaptation to toxic cyanobacteria. Limnology and Oceanography, 68 (2):514–524, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Lienart:2022:SGD

[LCHB22] Camilla Liénart, Alyssa R. Cirtwill, Melanie L. Hedgespeth, and Clare Bradshaw. A sprinkling of gold dust: Pine pollen as a carbon source in Baltic Sea coastal food webs. Limnology and Oceanography, 67(1):53–65, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Lewis:2022:CDR

[LCM^{+}22] Kristy A. Lewis, Robert R. Christian, Charles W. Martin, Kira L. Allen, Ashley M. McDonald, Victoria M. Roberts, Michelle N. Shaffer, and John F. Valentine. Complexities of disturbance response in a marine food web. Limnology and Oceanography, 67(S1):S352–S364, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Laurenceau-Cornec:2020:NGA

[LCMG^{+}20] Emmanuel C. Laurenceau-Cornec, Frédéric A. C. Le Moigne, Morgane Gallinari, Brivaëla Moriceau, Jordan Toullec, Morten H. Iversen, Anja Engel, and Christina L. De La Rocha. New guidelines for the application of Stokes’ models to the sinking velocity of marine aggregates. Limnology and Oceanography, 65(6):1264–1285, June 2020. CODEN LI-OCAH. ISSN 0024-3590.

Liu:2020:CGA

[LCS^{+}20] Wenwen Liu, Xincong Chen, Donald R. Strong, Steven C. Pennings, Matthew L. Kirwan, Xiaolin Chen, and Yihui Zhang. Climate and geographic adaptation drive latitudinal clines in biomass of a widespread saltmarsh plant in its native and introduced ranges. Limnology and Oceanography, 65(6):1399–1409, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Liu:2021:WCM

[LCZ^{+}21] Kailin Liu, Bingzhang Chen, Liping Zheng, Suhong Su, Bangqin Huang, Mianrun Chen, and Hongbin Liu. What controls microzooplankton biomass and herbivory rate across marginal seas of China? Limnology and Oceanography, 66(1): 61–75, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Lennartz:2022:CTM

[LD22] Sinikka T. Lennartz and Thorsten Dittmar. Controls on turnover of marine dissolved organic matter — testing the null hypothesis of purely concentration-driven uptake: Comment on Shen and Benner, “Molecular properties are a primary control on the microbial utilization of dissolved organic matter in the ocean”. Limnology and Oceanography, 67(3): 673–679, March 2022. CODEN LIOCAH. ISSN 0024-3590. See [SB20, SB22].

Liu:2020:PFP

[LDB^{+}20] Xin Liu, Gaël Dur, Syuhei Ban, Yoichiro Sakai, Shinsuke Ohmae, and Takashi Morita. Planktivorous fish predation masks anthropogenic disturbances on decadal trends in zooplankton biomass and body size structure in Lake Biwa, Japan. Limnology and Oceanography, 66(4):667–682, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Liu:2021:QDP

[LDB^{+}21] Xin Liu, Gaël Dur, Syuhei Ban, Yoichiro Sakai, Shinsuke Ohmae, and Takashi Morita. Quasi-decadal periodicities in growth and production of the copepod Eodiaptomus japonicus in Lake Biwa, Japan, related to the Arctic Oscillation. Limnology and Oceanography, 66(10):3783–3795, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Levine:2021:AVS

[LDG^{+}21] Robert M. Levine, Alex De Robertis, Daniel Grünbaum, Rebecca Woodgate, Calvin W. Mordy, Franz Mueter, Edward Cokelet, Noah Lawrence-Slavas, and Heather Tabisola. Autonomous vehicle surveys indicate that flow reversals retain juvenile fishes in a highly advective high-latitude ecosystem. Limnology and Oceanography, 66(4):1139–1154, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Leiva-Duenas:2023:COC

[LDGB^{+}23] Carmen Leiva-Dueñas, Anna Elizabeth Løvgren Graversen, Gary T. Banta, Marianne Holmer, Pere Masque, Peter Anton Upadhyay Stæhr, and Dorte Krause-Jensen. Capturing of organic carbon and nitrogen in eelgrass sediments of southern Scandinavia. Limnology and Oceanography, 68(3): 631–648, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Leiva-Duenas:2024:RSD

[LDGB^{+}24] Carmen Leiva-Dueñas, Anna Elizabeth Løvgren Graversen, Gary T. Banta, Jeppe Najbjerg Hansen, Marie Louise Kjærgaard Schrøter, Pere Masqué, Marianne Holmer, and Dorte Krause-Jensen. Region-specific drivers cause low organic carbon stocks and sequestration rates in the saltmarsh soils of southern Scandinavia. Limnology and Oceanography, 69(2): 290–308, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Li:2020:MPO

[LDS^{+}20] Wei Li, John E. Dore, August J. Steigmeyer, Yong-Joon Cho, Ok-Sun Kim, Yongqin Liu, Rachael M. Morgan-Kiss, Mark L. Skidmore, and John C. Priscu. Methane production in the oxygenated water column of a perennially ice-covered Antarctic lake. Limnology and Oceanography, 65(1):143–156, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Litchman:2021:TTB

[LEB21] Elena Litchman, Kyle F. Edwards, and Philip W. Boyd. Toward trait-based food webs: Universal traits and trait matching in planktonic predator–prey and host–parasite relationships. Limnology and Oceanography, 66(11):3857–3872, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Larson:2022:CAB

[LEK^{+}22] James H. Larson, Mary Anne Evans, Robert J. Kennedy, Sean W. Bailey, Keith A. Loftin, Zachary R. Laughrey, Robin A. Femmer, Jeff S. Schaeffer, William B. Richardson, Timothy T. Wynne, J. C. Nelson, and Joseph W. Duris. Corrigendum: Associations between cyanobacteria and indices of secondary production in the western basin of Lake Erie. Limnology and Oceanography, 67(11):2617–2620, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Lemmin:2020:IDD

[Lem20] Ulrich Lemmin. Insights into the dynamics of the deep hypolimnion of Lake Geneva as revealed by long-term temperature, oxygen, and current measurements. Limnology and Oceanography, 65(9):2092–2107, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Li:2021:PGE

[LES^{+}21] Qian Li, Kyle F. Edwards, Christopher R. Schvarcz, Karen E. Selph, and Grieg F. Steward. Plasticity in the grazing ecophysiology of Florenciella (Dichtyochophyceae), a mixotrophic nanoflagellate that consumes Prochlorococcus and other bacteria. Limnology and Oceanography, 66(1):47– 60, January 2021. CODEN LIOCAH. ISSN 0024-3590.

LeMezo:2021:FIP

[LG21] Priscilla K. Le Mézo and Eric D. Galbraith. The fecal iron pump: Global impact of animals on the iron stoichiometry of marine sinking particles. Limnology and Oceanography, 66 (1):201–213, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Liu:2022:BAF

[LGF22] Zezheng Liu, Olivier Gourgue, and Sergio Fagherazzi. Biotic and abiotic factors control the geomorphic characteristics of channel networks in salt marshes. Limnology and Oceanography, 67(1):89–101, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Larkin:2020:SBR

[LGI^{+}20] Alyse A. Larkin, Catherine A. Garcia, Kimberly A. Ingoglia, Nathan S. Garcia, Steven E. Baer, Benjamin S. Twining, Michael W. Lomas, and Adam C. Martiny. Subtle biogeochemical regimes in the Indian Ocean revealed by spatial and diel frequency of Prochlorococcus haplotypes. Limnology and Oceanography, 66(6):S220–S232, January 2020. CODEN LI-OCAH. ISSN 0024-3590.

Lienart:2021:LTC

[LGQ^{+}21] Camilla Liénart, Andrius Garbaras, Susanne Qvarfordt, Anton Öberg Sysoev, Helena Höglander, Jakob Walve, Ellen Schagerström, Johan Eklöf, and Agnes Ml Karlson. Longterm changes in trophic ecology of blue mussels in a rapidly changing ecosystem. Limnology and Oceanography, 66(3): 694–710, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Linkhorst:2020:CME

[LHD^{+}20] Annika Linkhorst, Carolin Hiller, Tonya DelSontro, Guilherme M. Azevedo, Nathan Barros, Raquel Mendonça, and Sebastian Sobek. Comparing methane ebullition variability across space and time in a Brazilian reservoir. Limnology and Oceanography, 65(7):1623–1634, July 2020. CODEN LI-OCAH. ISSN 0024-3590.

Liljestrand:2021:CTO

[LHGJ21] Frasier L. Liljestrand, Anxhela Hania, Mario Giordano, and David T. Johnston. Calibrating the triple oxygen isotope signature of cultured diatoms. Limnology and Oceanography, 66(12):4254–4266, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Long:2024:IWR

[LHH24] James P. Long, John D. Halfman, and Nathan Hawley. Internal wave resonance, surges, and strong nonlinear damping differentiated in two elongated lakes with the aid of an original Green’s function. Limnology and Oceanography, 69(4): 920–932, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Li:2021:DMD

[LHL^{+}21] Zhengfei Li, Jani Heino, Zhenyuan Liu, Xingliang Meng, Xiao Chen, Yihao Ge, and Zhicai Xie. The drivers of multiple dimensions of stream macroinvertebrate beta diversity across a large montane landscape. Limnology and Oceanography, 66(1):226–236, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Logozzo:2023:DDT

[LHMR23] Laura A. Logozzo, Jacob D. Hosen, Johnae McArthur, and Peter A. Raymond. Distinct drivers of two size fractions of operationally dissolved iron in a temperate river. Limnology and Oceanography, 68(6):1185–1200, June 2023. CODEN LI-OCAH. ISSN 0024-3590.

Lin:2020:RAA

[LHS^{+}20] Fan-Sian Lin, Pei-Chi Ho, Akash R. Sastri, Chung-Chi Chen, Gwo-Ching Gong, Sen Jan, and Chih hao Hsieh. Resource availability affects temporal variation of phytoplankton size structure in the Kuroshio east of Taiwan. Limnology and Oceanography, 65(2):236–246, February 2020. CODEN LI-OCAH. ISSN 0024-3590.

Lenstra:2021:CHE

[LHS^{+}21] Wytze K. Lenstra, Martijn Hermans, Marie J. M. Séguret, Rob Witbaard, Silke Severmann, Thilo Behrends, and Caroline P. Slomp. Coastal hypoxia and eutrophication as key controls on benthic release and water column dynamics of iron and manganese. Limnology and Oceanography, 66(3): 807–826, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Liu:2023:ITS

[LL23] Xiaoqing Liu and Yiming Luo. Influence of terrigenous supply from the African continent on sedimentary CaCO_{3} distribution in the Southeast Atlantic Ocean. Limnology and Oceanography, 68(1):14–25, January 2023. CODEN LI-OCAH. ISSN 0024-3590.

Lofton:2020:RIT

[LLBC20] Mary E. Lofton, Taylor H. Leach, Beatrix E. Beisner, and Cayelan C. Carey. Relative importance of top-down vs. bottom-up control of lake phytoplankton vertical distributions varies among fluorescence-based spectral groups. Limnology and Oceanography, 65(10):2485–2501, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Levenstein:2021:EPS

[LLC21] Brianna Levenstein, Jennifer Lento, and Joseph Culp. Effects of prolonged sedimentation from permafrost degradation on macroinvertebrate drift in Arctic streams. Limnology and Oceanography, 66(S1):S157–S168, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Lao:2023:EUR

[LLC^{+}23] Qibin Lao, Xuan Lu, Fajin Chen, Guangzhe Jin, Chunqing Chen, Xin Zhou, and Qingmei Zhu. Effects of upwelling and runoff on water mass mixing and nutrient supply induced by typhoons: Insight from dual water isotopes tracing. Limnology and Oceanography, 68(1):284–295, January 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Li:2022:SCA

[LLK^{+}22] Lan Li, Yiming Luo, Markus Kienast, Di Qi, and Jerry Tjiputra. On the sedimentary carbonate accumulation and dissolution in Western Pacific marginal basins. Limnology and Oceanography, 67(1):26–38, January 2022. CODEN LI-OCAH. ISSN 0024-3590.

Landou:2023:CPA

[LLL^{+}23] Etai Landou, Boaz Lazar, Julie LaRoche, Katja Fennel, and Ilana Berman-Frank. Contribution of photic and aphotic N_{2} fixation to production in an oligotrophic sea. Limnology and Oceanography, 68(3):692–708, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Laget:2023:ECA

[LLMM^{+}23] Manon Laget, Natalia Llopis-Monferrer, Jean-François Maguer, Aude Leynaert, and Tristan Biard. Elemental content allometries and silicon uptake rates of planktonic Rhizaria: Insights into their ecology and role in biogeochemical cycles. Limnology and Oceanography, 68(2):439–454, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Li:2024:CIV

[LLSQ24] Jing-Jing Li, Zheng-Yi Liu, Wang-Hui Song, and Song Qin. The contribution of intraspecific variation to future climate responses of brown algae. Limnology and Oceanography, 69 (1):53–66, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Luo:2024:NED

[LLTT^{+}24] Hongxue Luo, Songlin Liu, Stacey M. Trevathan-Tackett, Yuzheng Ren, Jiening Liang, Jinlong Li, Zhijian Jiang, Yunchao Wu, and Xiaoping Huang. Nitrogen enrichment decreases seagrass contributions to refractory organic matter pools. Limnology and Oceanography, 69(2):367–379, February 2024. CODEN LIOCAH. ISSN 0024-3590.

LaBrie:2024:PPH

[LM24] Richard LaBrie and Roxane Maranger. Predicting the presence of hypoxic hypolimnia in lakes at large spatial scales. Limnology and Oceanography, 69(2):355–366, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Laperriere:2021:NNO

[LMC^{+}21] Sarah M. Laperriere, Michael Morando, Douglas G. Capone, Troy Gunderson, Jason M. Smith, and Alyson E. Santoro. Nitrification and nitrous oxide dynamics in the Southern California Bight. Limnology and Oceanography, 66(4):1099–1112, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Laurion:2021:WMD

[LMM^{+}21] Isabelle Laurion, Philippe Massicotte, Flora Mazoyer, Karita Negandhi, and Natalie Mladenov. Weak mineralization despite strong processing of dissolved organic matter in Eastern Arctic tundra ponds. Limnology and Oceanography, 66(S1): S47–S63, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Lesser:2020:DDD

[LMP^{+}20] Michael P. Lesser, Benjamin Mueller, M. Sabrina Pankey, Keir J. Macartney, Marc Slattery, and Jasper M. de Goeij. Depth-dependent detritus production in the sponge, Halisarca caerulea. Limnology and Oceanography, 65(6):1200– 1216, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Longhini:2021:CWC

[LMS^{+}21] Cybelle Menolli Longhini, Léo Mahieu, Fabian Sá, Constant M. G. van den Berg, Pascal Salaün, and Renato Rodrigues Neto. Coastal waters contamination by mining tailings: What triggers the stability of iron in the dissolved and soluble fractions? Limnology and Oceanography, 66(1):171– 187, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Lesser:2023:PPC

[LMS23] Michael P. Lesser, Keir J. Macartney, and Marc Slattery. Production of phytodetritus by a coral reef sponge increases from shallow to mesophotic depths. Limnology and Oceanography, 68(6):1247–1255, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Lohmann:2023:CPS

[LMSN23] Amanda C. Lohmann, Joseph P. Morton, Oscar M. Schofield, and Douglas P. Nowacek. Cyclical prey shortages for a marine polar predator driven by the interaction of climate change and natural climate variability. Limnology and Oceanography, 68(12):2668–2687, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Light:2021:QVA

[LN21] Tricia Light and Richard Norris. Quantitative visual analysis of marine barite microcrystals: Insights into precipitation and dissolution dynamics. Limnology and Oceanography, 66(10):3619–3629, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Liu:2023:RNT

[LNC^{+}23] Kailin Liu, Jun Nishioka, Bingzhang Chen, Koji Suzuki, Shunyan Cheung, Yanhong Lu, Huijun Wu, and Hongbin Liu. Role of nutrients and temperature in shaping distinct summer phytoplankton and microzooplankton population dynamics in the western North Pacific and Bering Sea. Limnology and Oceanography, 68(3):649–665, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Leduc:2020:SIC

[LNR^{+}20] Daniel Leduc, Scott D. Nodder, Ashley A. Rowden, Max Gibbs, Katrin Berkenbusch, Anna Wood, Fabio De Leo, Craig Smith, Julie Brown, Sarah J. Bury, and Arne Pallentin. Structure of infaunal communities in New Zealand submarine canyons is linked to origins of sediment organic matter. Limnology and Oceanography, 65(10):2303–2327, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Labarre:2020:EGI

[LOW^{+}20] Aurelie Labarre, Aleix Obiol, Susanne Wilken, Irene Forn, and Ramon Massana. Expression of genes involved in phagocytosis in uncultured heterotrophic flagellates. Limnology and Oceanography, 65(S1):S149–S160, January 2020. CO-DEN LIOCAH. ISSN 0024-3590.

Lotem:2023:AOM

[LPA^{+}23] Noam Lotem, André Pellerin, Katey Walter Anthony, Almog Gafni, Valeria Boyko, and Orit Sivan. Anaerobic oxidation of methane does not attenuate methane emissions from thermokarst lakes. Limnology and Oceanography, 68(6):1316– 1330, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Leon-Palmero:2023:PID

[LPMBR23] Elizabeth León-Palmero, Rafael Morales-Baquero, and Isabel Reche. P inputs determine denitrifier abundance explaining dissolved nitrous oxide in reservoirs. Limnology and Oceanography, 68(8):1734–1749, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Liu:2020:DCR

[LPO^{+}20] Shuting Liu, Rachel Parsons, Keri Opalk, Nicholas Baetge, Stephen Giovannoni, Luis M. Bolaños, Elizabeth B. Kujawinski, Krista Longnecker, YueHan Lu, Elisa Halewood, and Craig A. Carlson. Different carboxyl-rich alicyclic molecules proxy compounds select distinct bacterioplankton for oxidation of dissolved organic matter in the mesopelagic Sargasso Sea. Limnology and Oceanography, 65(7):1532–1553, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Lange:2023:EDO

[LRdGFP23] Kiara Lange, Stéphanie Reynaud, Jasper M. de Goeij, and Christine Ferrier-Pagès. The effects of dissolved organic matter supplements on the metabolism of corals under heat stress. Limnology and Oceanography, 68(12):2774–2788, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Lovelock:2021:VAZ

[LRM21] Catherine E. Lovelock, Ruth Reef, and Pere Masqué. Vulnerability of an arid zone coastal wetland landscape to sea level rise and intense storms. Limnology and Oceanography, 66(11):3976–3989, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Lammerant:2024:SSA

[LRNG24] Roel Lammerant, Eva Karin Rohlfer, Alf Norkko, and Camilla Gustafsson. Seasonality strongly affects short-term c-storage in coastal macrophyte communities. Limnology and Oceanography, 69(3):700–711, March 2024. CODEN LI-OCAH. ISSN 0024-3590.

Lovelock:2020:RVC

[LRRP20] Catherine E. Lovelock, Ruth Reef, John A. Raven, and John M. Pandolfi. Regional variation in δ ^{13}C of coral reef macroalgae. Limnology and Oceanography, 65(10):2291– 2302, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Letourneau:2021:STC

[LSC^{+}21] Maria L. Letourneau, Sylvia C. Schaefer, Huan Chen, Amy M. McKenna, Merryl Alber, and Patricia M. Medeiros. Spatio-temporal changes in dissolved organic matter composition along the salinity gradient of a marsh-influenced estuarine complex. Limnology and Oceanography, 66(8):3040– 3054, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Liu:2021:TSP

[LSCL21] Kailin Liu, Koji Suzuki, Bingzhang Chen, and Hongbin Liu. Are temperature sensitivities of Prochlorococcus and Synechococcus impacted by nutrient availability in the subtropical northwest Pacific? Limnology and Oceanography, 66(3): 639–651, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Lopez-Sandoval:2021:NTC

[LSDA21] Daffne C. López-Sandoval, Carlos M. Duarte, and Susana Agustí. Nutrient and temperature constraints on primary production and net phytoplankton growth in a tropical ecosystem. Limnology and Oceanography, 66(7):2923–2935, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Li:2020:ECS

[LSL^{+}20] Weiying Li, William G. Sunda, Wenfang Lin, Haizheng Hong, and Dalin Shi. The effect of cell size on cellular Zn and Cd and Zn–Cd–CO_{2} colimitation of growth rate in marine diatoms. Limnology and Oceanography, 66(6):2896–2911, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Longnecker:2024:SDP

[LSS^{+}24] Krista Longnecker, Melissa C. Kido Soule, Gretchen J. Swarr, Rachel J. Parsons, Shuting Liu, Winifred M. Johnson, Brittany Widner, Ruth Curry, Craig A. Carlson, and Elizabeth B. Kujawinski. Seasonal and daily patterns in known dissolved metabolites in the northwestern Sargasso Sea. Limnology and Oceanography, 69(3):449–466, March 2024. CO-DEN LIOCAH. ISSN 0024-3590.

Lee:2020:NOP

[LSW^{+}20a] Zhongping Lee, Shaoling Shang, Yongchao Wang, Jianwei Wei, and Joji Ishizaka. Nature of optical products inverted semianalytically from remote sensing reflectance of stratified waters. Limnology and Oceanography, 65(2):387–400, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Long:2020:EOS

[LSW^{+}20b] Matthew H. Long, Kevin Sutherland, Scott D. Wankel, David J. Burdige, and Richard C. Zimmerman. Ebullition of oxygen from seagrasses under supersaturated conditions. Limnology and Oceanography, 65(2):314–324, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Li:2021:HNF

[LTH21] Song Li, Robert R. Twilley, and Aixin Hou. Heterotrophic nitrogen fixation in response to nitrate loading and sediment organic matter in an emerging coastal deltaic floodplain within the Mississippi River Delta plain. Limnology and Oceanography, 66(5):1961–1978, May 2021. CODEN LI-OCAH. ISSN 0024-3590.

Lindegren:2020:ENS

[LTJ^{+}20] Martin Lindegren, Mridul K. Thomas, Sigrún H. Jónasdóttir, Torkel G. Nielsen, and Peter Munk. Environmental niche separation promotes coexistence among ecologically similar zooplankton species — North Sea copepods as a case study. Limnology and Oceanography, 65(3):545–556, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Li:2022:ENA

[LTLH22] Hsin-Ting Li, Sing-How Tuo, Mei-Chen Lu, and Tung-Yuan Ho. The effects of Ni availability on H_{2} production and N_{2} fixation in a model unicellular diazotroph: The expression of hydrogenase and nitrogenase. Limnology and Oceanography, 67(7):1566–1576, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Lansac-Toha:2021:CCB

[LTQS^{+}21] Fernando Miranda Lansac-Tôha, Bárbara Angélio Quirino, Yasmin Rodrigues Souza, Fábio Amodêo Lansac-Tôha, Luiz Felipe Machado Velho, and Matheus Tenório Baumgartner. The commonality of core biological groups across freshwater food webs. Limnology and Oceanography, 65(3):1459–1474, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Lin:2024:VPR

[LTW^{+}24] YuanYu Lin, Olivia Torano, Logan Whitehouse, Emily Pierce, Claire P. Till, Matthew Hurst, Robert Freiberger, Travis Mellett, Maria T. Maldonado, Jian Guo, Mariam Sutton, David Zeitz, and Adrian Marchetti. Variability in the phytoplankton response to upwelling across an iron limitation mosaic within the California current system. Limnology and Oceanography, 69(4):888–901, April 2024. CODEN LI-OCAH. ISSN 0024-3590.

Leon:2022:UCE

[LvBS^{+}22] Morgane Léon, Pieter van Beek, Jan Scholten, Willard S. Moore, Marc Souhaut, Joselene De Oliveira, Catherine Jeandel, Patrick Seyler, and Julien Jouanno. Use of ^{223}Ra and 224Ra as chronometers to estimate the residence time of Amazon waters on the Brazilian continental shelf. Limnology and Oceanography, 67(4):753–767, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Langenegger:2022:MOD

[LVDM22] Timon Langenegger, Dominic Vachon, Daphne Donis, and Daniel F. McGinnis. Methane oxidation dynamics in a stratified lake: Insights revealed from a mass balance and carbon stable isotopes. Limnology and Oceanography, 67(10):2157– 2173, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Langston:2020:MLT

[LVHK20] Amy K. Langston, Orencio Durán Vinent, Ellen R. Herbert, and Matthew L. Kirwan. Modeling long-term salt marsh response to sea level rise in the sediment-deficient Plum Island Estuary, MA. Limnology and Oceanography, 66(5):2142– 2157, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Loewen:2021:QSS

[LVZ21] Charlie J. G. Loewen, Rolf D. Vinebrooke, and Ron W. Zurawell. Quantifying seasonal succession of phytoplankton trait-environment associations in human-altered landscapes. Limnology and Oceanography, 65(3):1409–1423, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Li:2024:RDE

[LWC^{+}24] Cheng-Xuan Li, Bao-Dong Wang, Kan Chen, Gui-Peng Yang, Jian-Fang Chen, Li-Na Lin, and Zi-Cheng Wang. Response of distributions and emissions of summer biogenic sulfur in the Pacific Arctic to enhanced Pacific water inflow. Limnology and Oceanography, 69(1):81–103, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Lapierre:2023:CCL

[LWH^{+}23] Jean-Francois Lapierre, Katherine E. Webster, Ephraim M. Hanks, Tyler Wagner, Patricia A. Soranno, Ian M. McCullough, Kaitlin L. Reinl, Marcella Domka, and Noah R. Lotting. A continuous classification of the 476,697 lakes of the conterminous US based on geographic archetypes. Limnology and Oceanography, 68(12):2759–2773, December 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Li:2022:EPM

[LWS^{+}22] Yun Li, Rong Wang, Haojie Su, Jianjun Wang, Ping Xie, and Feizhou Chen. Eutrophication and predation mediate zooplankton diversity and network structure. Limnology and Oceanography, 67(S1):S133–S145, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Latour:2021:MBS

[LWvdM^{+}21] Pauline Latour, Kathrin Wuttig, Pier van der Merwe, Robert F. Strzepek, Melanie Gault-Ringold, Ashley T. Townsend, Thomas M. Holmes, Matthew Corkill, and Andrew R. Bowie. Manganese biogeochemistry in the Southern Ocean, from Tasmania to Antarctica. Limnology and Oceanography, 65(6):2547–2562, June 2021. CODEN LI-OCAH. ISSN 0024-3590.

Li:2021:NFI

[LXC^{+}21a] Ruihuan Li, Jie Xu, Xianrong Cen, Wanxuan Zhong, Jianzu Liao, Zhen Shi, and Shengqi Zhou. Nitrate fluxes induced by turbulent mixing in dipole eddies in an oligotrophic ocean. Limnology and Oceanography, 66(7):2842–2854, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Lin:2021:DCS

[LXC^{+}21b] Hui Lin, Huacheng Xu, Yihua Cai, Claude Belzile, Robie W. Macdonald, and Laodong Guo. Dynamic changes in sizefractionated dissolved organic matter composition in a seasonally ice-covered Arctic River. Limnology and Oceanography, 66(8):3085–3099, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Li:2023:HVO

[LZD^{+}23] Xinxin Li, Xin Zhao, Hongyue Dang, Chuanlun Zhang, Igor Fernández-Urruzola, Zhiqiang Liu, Frank Wenzhöfer, and Ronnie N. Glud. High variability in organic carbon sources and microbial activities in the hadopelagic waters. Limnology and Oceanography, 68(8):1704–1718, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Li:2021:SLS

[LZG21] Ming Li, Fan Zhang, and Patricia M. Glibert. Seasonal life strategy of Prorocentrum minimum in Chesapeake Bay, USA: Validation of the role of physical transport using a coupled physical–biogeochemical–harmful algal bloom model. Limnology and Oceanography, 66(11):3873–3886, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Lin:2021:SIN

[LZM^{+}21] Qi Lin, Ke Zhang, Suzanne McGowan, Eric Capo, and Ji Shen. Synergistic impacts of nutrient enrichment and climate change on long-term water quality and ecological dynamics in contrasting shallow-lake zones. Limnology and Oceanography, 66(9):3271–3286, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Marzolf:2021:EMT

[MA21] Nicholas S. Marzolf and Marcelo Ardón. Ecosystem metabolism in tropical streams and rivers: a review and synthesis. Limnology and Oceanography, 66(5):1627–1638, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Matos:2024:GSF

[MACC24] Fábio L. Matos, Jacopo Aguzzi, Joan B. Company, and Marina R. Cunha. Gone with the stream: Functional connectivity of a cold-water coral at basin scale. Limnology and Oceanography, 69(2):217–231, February 2024. CODEN LI-OCAH. ISSN 0024-3590.

Mausz:2022:MUD

[MAD^{+}22] Michaela A. Mausz, Ruth L. Airs, Joanna L. Dixon, Claire E. Widdicombe, Glen A. Tarran, Luca Polimene, Sarah Dashfield, Rachael Beale, David J. Scanlan, and Yin Chen. Microbial uptake dynamics of choline and glycine betaine in coastal seawater. Limnology and Oceanography, 67(5):1052– 1064, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Mesman:2022:DPR

[MAG^{+}22] Jorrit P. Mesman, Ana I. Ayala, Stéphane Goyette, Jérôme Kasparian, Rafael Marcé, Hampus Markensten, Julio A. A. Stelzer, Michael W. Thayne, Mridul K. Thomas, Don C. Pierson, and Bas W. Ibelings. Drivers of phytoplankton responses to summer wind events in a stratified lake: a modeling study. Limnology and Oceanography, 67(4):856–873, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Milan:2022:CHD

[MAP^{+}22] Manuela Milan, Nina Albrecht, Frank Peeters, Simone Wengrat, Martin Wessels, and Dietmar Straile. Clockwise hysteresis of diatoms in response to nutrient dynamics during eutrophication and recovery. Limnology and Oceanography, 67(9):2088–2100, September 2022. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:2021:TSB

[MBA^{+}21] Sally MacIntyre, David Bastviken, Lars Arneborg, Adam T. Crowe, Jan Karlsson, Andreas Andersson, Magnus Gålfalk, Anna Rutgersson, Eva Podgrajsek, and John M. Melack. Turbulence in a small boreal lake: Consequences for air– water gas exchange. Limnology and Oceanography, 66(3): 827–854, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Matson:2020:PDF

[MBB^{+}20] Paul G. Matson, Gregory L. Boyer, Thomas B. Bridgeman, George S. Bullerjahn, Douglas D. Kane, Robert M. L. McKay, Katelyn M. McKindles, Heather A. Raymond, Brenda K. Snyder, Richard P. Stumpf, and Timothy W.

Davis. Physical drivers facilitating a toxigenic cyanobacterial bloom in a major Great Lakes tributary. Limnology and Oceanography, 66(6):2866–2882, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Marques:2021:TEB

[MBC^{+}21] Raquel Marques, Delphine Bonnet, Claire Carré, Cécile Roques, and Audrey M. Darnaude. Trophic ecology of a blooming jellyfish (Aurelia coerulea) in a Mediterranean coastal lagoon. Limnology and Oceanography, 66(1):141–157, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Meiler:2022:CUD

[MBD^{+}22] Simona Meiler, Gregory L. Britten, Stephanie Dutkiewicz, Mary Rose Gradoville, Pia H. Moisander, Oliver Jahn, and Michael J. Follows. Constraining uncertainties of diazotroph biogeography from nifH gene abundance. Limnology and Oceanography, 67(4):816–829, April 2022. CODEN LIOCAH. ISSN 0024-3590. See comment [ZR23].

Meiler:2023:COC

[MBD^{+}23] Simona Meiler, Gregory L. Britten, Stephanie Dutkiewicz, Pia H. Moisander, and Michael J. Follows. Challenges and opportunities in connecting gene count observations with ocean biogeochemical models: Reply to Zehr and Riemann (2023). Limnology and Oceanography, 68(6):1413–1416, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Maavara:2023:WDU

[MBH^{+}23] Taylor Maavara, Craig Brinkerhoff, Jacob Hosen, Kelly Aho, Laura Logozzo, James Saiers, Aron Stubbins, and Peter Raymond. Watershed DOC uptake occurs mostly in lakes in the summer and in rivers in the winter. Limnology and Oceanography, 68(3):735–751, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Maldonado:2021:MSU

[MBLA^{+}21] Manuel Maldonado, Lindsay Beazley, María López-Acosta, Ellen Kenchington, Benoit Casault, Ulrike Hanz, and Furu Mienis. Massive silicon utilization facilitated by a benthicpelagic coupled feedback sustains deep-sea sponge aggregations. Limnology and Oceanography, 66(2):366–391, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Menge:2021:BMP

[MCH^{+}21] Bruce A. Menge, Sarah L. Close, Sally D. Hacker, Karina J. Nielsen, and Francis Chan. Biogeography of macrophyte productivity: Effects of oceanic and climatic regimes across spatiotemporal scales. Limnology and Oceanography, 66(3):711– 726, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Mallon:2022:LDD

[MCH^{+}22] Jennifer Mallon, Tyler Cyronak, Emily R. Hall, Anastazia T. Banaszak, Dan A. Exton, and Adrian M. Bass. Light-driven dynamics between calcification and production in functionally diverse coral reef calcifiers. Limnology and Oceanography, 67(2):434–449, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Michael:2023:GMS

[MCS^{+}23] Susanna M. Michael, John Crusius, Andrew W. Schroth, Robert Campbell, and Joseph A. Resing. Glacial meltwater and sediment resuspension can be important sources of dissolved and total dissolvable aluminum and manganese to coastal ocean surface waters. Limnology and Oceanography, 68(6):1201–1215, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Morganti:2020:TNS

[MCYR20] T. Morganti, R. Coma, G. Yahel, and M. Ribes. Trophic niche separation that facilitates co-existence of high and low microbial abundance sponges is revealed by in situ study of carbon and nitrogen fluxes. Limnology and Oceanography, 66(5):2259, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Marin-Diaz:2020:REB

[MDBI20] Beatriz Marin-Diaz, Tjeerd J. Bouma, and Eduardo Infantes. Role of eelgrass on bed-load transport and sediment resuspension under oscillatory flow. Limnology and Oceanography, 65(2):426–436, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Mayfield:2022:SCC

[MDDL22] Roseanna J. Mayfield, John A. Dearing, C. Patrick Doncaster, and Peter G. Langdon. Stability of chironomid community structure during historic climatic and environmental change in subarctic Alaska. Limnology and Oceanography, 67(S1):S444–S460, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Moskovich:2023:HVE

[MDIY23] Raz Moskovich, Rei Diga, Micha Ilan, and Gitai Yahel. High variability and enhanced nocturnal oxygen uptake in coral reef sponges. Limnology and Oceanography, 68(7):1517–1529, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Martin:2023:SSR

[MDP^{+}23] Robbie M. Martin, Maddie K. Denney, Helena L. Pound, Justin D. Chaffin, George S. Bullerjahn, R. Michael L. McKay, Arthur Zastepa, Katarina A. Jones, Hector F. Castro, Shawn R. Campagna, and Steven W. Wilhelm. Sulfolipid substitution ratios of Microcystis aeruginosa and planktonic communities as an indicator of phosphorus limitation in Lake Erie. Limnology and Oceanography, 68(5):1117–1131, May 2023. CODEN LIOCAH. ISSN 0024-3590.

McNally:2022:PEE

[MDWT^{+}22] Elise M. McNally, Alan M. Downey-Wall, F. Dylan Titmuss, Camila Cortina, Kathleen Lotterhos, and Justin B. Ries. Parental exposure of Eastern oysters (Crassostrea virginica) to elevated pCO_{2} mitigates its negative effects on early larval shell growth and morphology. Limnology and Oceanography, 67(8):1732–1745, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Mariotti:2020:MAO

[MEQBV20] G. Mariotti, T. Elsey-Quirk, G. Bruno, and K. Valentine. Mud-associated organic matter and its direct and indirect role in marsh organic matter accumulation and vertical accretion. Limnology and Oceanography, 66(10):2627–2641, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Moran:2022:OLD

[MFGF^{+}22] Mary Ann Moran, Frank X. Ferrer-González, He Fu, Brent Nowinski, Malin Olofsson, McKenzie A. Powers, Jeremy E. Schreier, William F. Schroer, Christa B. Smith, and Mario Uchimiya. The ocean’s labile DOC supply chain. Limnology and Oceanography, 67(5):1007–1021, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Mahdiyan:2021:DWQ

[MFM^{+}21] Octavia Mahdiyan, Alessandro Filazzola, Lewis A. Molot, Derek Gray, and Sapna Sharma. Drivers of water quality changes within the Laurentian Great Lakes region over the past 40 years. Limnology and Oceanography, 66(1):237–254, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Moreno:2020:IEI

[MGC^{+}20] Carly M. Moreno, Weida Gong, Natalie R. Cohen, Kimberly DeLong, and Adrian Marchetti. Interactive effects of iron and light limitation on the molecular physiology of the Southern Ocean diatom fragilariopsis kerguelensis. Limnology and Oceanography, 65(7):1511–1531, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Musan:2024:IEC

[MGG^{+}24] Israela Musan, Hezi Gildor, Thomas Gonsiorczyk, Hans-Peter Grossart, and Boaz Luz. Isotope effects of O_{2} consumption in a deep lake as means for understanding partitioning of O_{2} demand among microorganisms, particles, and sediment. Limnology and Oceanography, 69(4):992–1004, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Myrstener:2021:NIS

[MGGRR^{+}21] Maria Myrstener, Lluís Gómez-Gener, Gerard Rocher-Ros, Reiner Giesler, and Ryan A. Sponseller. Nutrients influence seasonal metabolic patterns and total productivity of Arctic streams. Limnology and Oceanography, 66(S1):S182–S196, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Mwanake:2022:BSE

[MGI^{+}22] Ricky M. Mwanake, Gretchen M. Gettel, Clarisse Ishimwe, Elizabeth G. Wangari, Klaus Butterbach-Bahl, and Ralf Kiese. Basin-scale estimates of greenhouse gas emissions from the Mara River, Kenya: Importance of discharge, stream size, and land use/land cover. Limnology and Oceanography, 67 (8):1776–1793, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Maisonneuve:2022:BPR

[MGL22] Philippe Maisonneuve, François Guillemette, and Jean-François Lapierre. Biological and photochemical reactivity of dissolved organic matter in a large temperate river. Limnology and Oceanography, 67(6):1388–1401, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Marchini:2021:DPI

[MGP^{+}21] Chiara Marchini, Francesca Gizzi, Thomas Pondrelli, Lisa Moreddu, Luca Marisaldi, Francesco Montori, Valentina Lazzari, Valentina Airi, Erik Caroselli, Fiorella Prada, Giuseppe Falini, Zvy Dubinsky, and Stefano Goffredo. Decreasing pH impairs sexual reproduction in a Mediterranean coral transplanted at a CO_{2} vent. Limnology and Oceanography, 66 (11):3990–4000, November 2021. CODEN LIOCAH. ISSN 0024-3590.

McInerney:2023:SFA

[MGWS23] Paul J. McInerney, Darren P. Giling, Ben Wolfenden, and Ashmita Sengupta. A synthesis of floodplain aquatic ecosystem metabolism and carbon flux using causal criteria analysis. Limnology and Oceanography, 68(1):97–109, January 2023. CODEN LIOCAH. ISSN 0024-3590.

Martiny:2022:MPR

[MHD^{+}22] Adam C. Martiny, George I. Hagstrom, Tim DeVries, Robert T. Letscher, Gregory L. Britten, Catherine A. Garcia, Eric Galbraith, David Karl, Simon A. Levin, Michael W. Lomas, Allison R. Moreno, David Talmy, Weilei Wang, and Katsumi Matsumoto. Marine phytoplankton resilience may moderate oligotrophic ecosystem responses and biogeochemical feedbacks to climate change. Limnology and Oceanography, 67(S1):S378–S389, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Medeiros:2021:SSD

[MHdS^{+}21] Carlinda Raílly Medeiros, Jani Heino, Paulo Jorge Parreira dos Santos, Joseline Molozzi, and Raphael Ligeiro. Spatial scale drives diversity patterns of benthic macroinvertebrate communities in tropical estuaries. Limnology and Oceanography, 66(3):727–739, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Maciute:2023:RIM

[MHG^{+}23] Adele Maciute, Oleksandr Holovachov, Ronnie N. Glud, Elias Broman, Peter Berg, Francisco J. A. Nascimento, and Stefano Bonaglia. Reconciling the importance of meiofauna respiration for oxygen demand in muddy coastal sediments. Limnology and Oceanography, 68(8):1895–1905, August 2023. CODEN LIOCAH. ISSN 0024-3590.

McInerney:2020:BRQ

[MHL^{+}20] Paul J. McInerney, Galen Holt, Rebecca E. Lester, Ross M. Thompson, Barbara Robson, Darren S. Ryder, Nick R. Bond, Darren S. Baldwin, Ben Gawne, and Rochelle Petrie. Basal resource quality and energy sources in three habitats of a lowland river ecosystem. Limnology and Oceanography, 66 (10):2757–2771, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Molina-Hernandez:2022:CRE

[MHMML^{+}22] Ana Molina-Hernández, Francisco Medellín-Maldonado, Ines D. Lange, Chris T. Perry, and Lorenzo Álvarez-Filip. Coral reef erosion: In situ measurement on different dead coral substrates on a Caribbean reef. Limnology and Oceanography, 67(12):2734–2749, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Monteiro:2023:SVD

[MHP^{+}23] Thiago Monteiro, Sian F. Henley, Ricardo César Gonçalves Pollery, Carlos Rafael Borges Mendes, Mauricio Mata, Virginia Maria Tavano, Carlos Alberto Eiras Garcia, and Rodrigo Kerr. Spatiotemporal variability of dissolved inorganic macronutrients along the northern Antarctic Peninsula (1996–2019). Limnology and Oceanography, 68(10):2305– 2326, October 2023. CODEN LIOCAH. ISSN 0024-3590.

Mosquera:2022:COB

[MHSL22] Virginia Mosquera, Eliza Maher Hasselquist, Ryan A. Sponseller, and Hjalmar Laudon. Co-occurrence of browning and oligotrophication in a boreal stream network. Limnology and Oceanography, 67(10):2325–2339, October 2022. CODEN LI-OCAH. ISSN 0024-3590.

Mosquera:2022:WCV

[MHVC22] Pablo V. Mosquera, Henrietta Hampel, Raúl F. Vázquez, and Jordi Catalan. Water chemistry variation in tropical high-mountain lakes on old volcanic bedrocks. Limnology and Oceanography, 67(7):1522–1536, July 2022. CODEN LI-OCAH. ISSN 0024-3590.

Meysick:2022:CEE

[MIR^{+}22] Lukas Meysick, Eduardo Infantes, Luca Rugiu, Karine Gagnon, and Christoffer Boström. Coastal ecosystem engineers and their impact on sediment dynamics: Eelgrass– bivalve interactions under wave exposure. Limnology and Oceanography, 67(3):621–633, March 2022. CODEN LI-OCAH. ISSN 0024-3590.

Mojib:2020:CTS

[MK20] Nazia Mojib and Julia Kubanek. Comparative transcriptomics supports the presence of G protein-coupled receptorbased signaling in unicellular marine eukaryotes. Limnology and Oceanography, 65(4):762–774, April 2020. CODEN LI-OCAH. ISSN 0024-3590.

Miller:2021:SBF

[MK21] Cale A. Miller and Amanda L. Kelley. Seasonality and biological forcing modify the diel frequency of nearshore pH extremes in a subarctic Alaskan estuary. Limnology and Oceanography, 65(3):1475–1491, April 2021. CODEN LI-OCAH. ISSN 0024-3590.

Meyer-Kaiser:2019:RAD

[MKBSK19] Kirstin Meyer-Kaiser, Melanie Bergmann, Thomas Soltwedel, and Michael Klages. Recruitment of Arctic deep-sea invertebrates: Results from a long-term hard-substrate colonization experiment at the Long-Term Ecological Research observatory HAUSGARTEN. Limnology and Oceanography, 64(5): 1924–1938, September 2019. CODEN LIOCAH. ISSN 0024-3590. See corrigendum [Ano21b].

Muth:2022:HFP

[MKD22] Arley F. Muth, Amanda L. Kelley, and Kenneth H. Dunton. High-frequency pH time series reveals pronounced seasonality in Arctic coastal waters. Limnology and Oceanography, 67(7): 1429–1442, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Maszczyk:2024:THD

[MKL^{+}24] Piotr Maszczyk, Karol Krajewski, Konrad Leniowski, Szymon Pukos, Julia Wawrzeńczak, Wojciech Wilczynski, Marcin Lukasz Zebrowski, Jae-Seong Lee, and Ewa Babkiewicz. Temperature and hypoxia-driven shifts in Daphnia interspecific competition. Limnology and Oceanography, 69(3):576–588, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Mueller:2023:MSM

[MKM^{+}23] Peter Mueller, Lars Kutzbach, Thomas J. Mozdzer, Emil Jespersen, Donald C. Barber, and Franziska Eller. Minerogenic salt marshes can function as important inorganic carbon stores. Limnology and Oceanography, 68(4):942–952, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Michoud:2023:UFD

[MKP^{+}23] Grégoire Michoud, Tyler J. Kohler, Hannes Peter, Jade Brandani, Susheel Banu Busi, and Tom J. Battin. Unexpected functional diversity of stream biofilms within and across proglacial floodplains despite close spatial proximity. Limnology and Oceanography, 68(9):2183–2194, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Martini:2021:FTB

[MLB^{+}21] Séverine Martini, Floriane Larras, Aurélien Boyé, Emile Faure, Nicole Aberle, Philippe Archambault, Lise Bacouillard, Beatrix E. Beisner, Lucie Bittner, Emmanuel Castella, Michael Danger, Olivier Gauthier, Lee Karp-Boss, Fabien Lombard, Frédéric Maps, Lars Stemmann, Eric Thiébaut, Philippe Usseglio-Polatera, Meike Vogt, Martin Laviale, and Sakina-Dorothée Ayata. Functional trait-based approaches as a common framework for aquatic ecologists. Limnology and Oceanography, 66(3):965–994, March 2021. CODEN LI-OCAH. ISSN 0024-3590.

Matsuzaki:2021:CFW

[MLC^{+}21] Shin-Ichiro S. Matsuzaki, Richard C. Lathrop, Stephen R. Carpenter, Jake R. Walsh, M. Jake Vander Zanden, Mark R. Gahler, and Emily H. Stanley. Climate and food web effects on the spring clear-water phase in two north-temperate eutrophic lakes. Limnology and Oceanography, 66(1):30–46, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Morin:2020:RSI

[MLG^{+}20] Philippe-Israël Morin, Thomas Lacour, Pierre-Luc Grondin, Flavienne Bruyant, Joannie Ferland, Marie-Hélène Forget, Philippe Massicotte, Natalie Donaher, Douglas A. Campbell, Johann Lavaud, and Marcel Babin. Response of the seaice diatom Fragilariopsis cylindrus to simulated polar night darkness and return to light. Limnology and Oceanography, 65(5):1041–1060, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Malkin:2022:CCS

[MLK^{+}22] Sairah Y. Malkin, Pinky Liau, Carol Kim, Kalev G. Hantsoo, Maya L. Gomes, and Bongkeun Song. Contrasting controls on seasonal and spatial distribution of marine cable bacteria (Candidatus electrothrix) and beggiatoaceae in seasonally hypoxic Chesapeake Bay. Limnology and Oceanography, 67(6): 1357–1373, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Moon:2021:ANC

[MLL^{+}21] Ji-Young Moon, Kitack Lee, Weol-Ae Lim, Eunil Lee, Minhan Dai, Yang-Ho Choi, In-Seong Han, Kyoungsoon Shin, Ja-Myung Kim, and Jinho Chae. Anthropogenic nitrogen is changing the East China and Yellow Seas from being N deficient to being P deficient. Limnology and Oceanography, 66(3):914–924, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Marshall:2021:FIR

[MLP^{+}21a] Danielle A. Marshall, Megan K. La Peyre, Terence A. Palmer, Gaël Guillou, Blair D. Sterba-Boatwright, Jennifer Beseres Pollack, and Benoit Lebreton. Freshwater inflow and responses from estuaries across a climatic gradient: an assessment of northwestern Gulf of Mexico estuaries based on stable isotopes. Limnology and Oceanography, 66(9):3568–3581, September 2021. CODEN LIOCAH. ISSN 0024-3590.

McGeady:2021:SLP

[MLP21b] Ryan McGeady, Colm Lordan, and Anne Marie Power. Shift in the larval phenology of a marine ectotherm due to ocean warming with consequences for larval transport. Limnology and Oceanography, 66(2):543–557, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Mayers:2021:RLV

[MLS^{+}21] Kyle Michael James Mayers, Janice Lawrence, Katrine Sandnes Skaar, Joachim Paul Töpper, Elzbieta Petelenz, Marius Rydningen Saltvedt, Ruth-Anne Sandaa, Aud Larsen, Gunnar Bratbak, and Jessica Louise Ray. Removal of large viruses and their dispersal through fecal pellets of the appendicularian Oikopleura dioica during Emiliania huxleyi bloom conditions. Limnology and Oceanography, 66(11):3963–3975, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Monferrer:2021:RSR

[MLT^{+}21] Natalia Llopis Monferrer, Aude Leynaert, Paul Tréguer, Andrés Gutiérrez-Rodríguez, Brivaela Moriceau, Morgane Gallinari, Mikel Latasa, Stéphane L’Helguen, Jean-François Maguer, Karl Safi, Matthew H. Pinkerton, and Fabrice Not. Role of small Rhizaria and diatoms in the pelagic silica production of the Southern Ocean. Limnology and Oceanography, 66(6):2187–2202, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Marrec:2021:SVP

[MMF^{+}21] Pierre Marrec, Heather McNair, Gayantonia Franzè, Françoise Morison, Jacob P. Strock, and Susanne Menden-Deuer. Seasonal variability in planktonic food web structure and function of the Northeast U.S. Shelf. Limnology and Oceanography, 65(3):1440–1458, April 2021. CODEN LIOCAH. ISSN 0024-3590.

McNair:2021:MGC

[MMG^{+}21] Heather M. McNair, Françoise Morison, Jason R. Graff, Tatiana A. Rynearson, and Susanne Menden-Deuer. Microzooplankton grazing constrains pathways of carbon export in the subarctic North Pacific. Limnology and Oceanography, 66(7): 2697–2711, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Manko:2022:AAR

[MMKWZ22] Maciej K. Mańko, Małgorzata Merchel, Sławomir Kwaśniewski, and Agata Weydmann-Zwolicka. Atlantification alters the reproduction of jellyfish Aglantha digitale in the European Arctic. Limnology and Oceanography, 67(8):1836–1849, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Medellin-Maldonado:2022:UCD

[MMLPRHCG22] Francisco Medellín-Maldonado, Andrés López-Pérez, Leopoldo Ruiz-Huerta, and Juan P. Carricart-Ganivet. Understanding corallite demography to comprehend potential bias in sclerochronology: Analysis of coral modular growth by microcomputed tomography. Limnology and Oceanography, 67 (12):2665–2676, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Moller:2020:BAZ

[MN20] Eva Friis Møller and Torkel Gissel Nielsen. Borealization of Arctic zooplankton — smaller and less fat zooplankton species in Disko Bay, Western Greenland. Limnology and Oceanography, 65(6):1175–1188, June 2020. CODEN LI-OCAH. ISSN 0024-3590.

Mansour:2021:CNC

[MNM^{+}21] Joost Samir Mansour, Andreas Norlin, Natalia Llopis Monferrer, Stéphane L’Helguen, and Fabrice Not. Carbon and nitrogen content to biovolume relationships for marine protist of the Rhizaria lineage (Radiolaria and Phaeodaria). Limnology and Oceanography, 66(5):1703–1717, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Myrstener:2021:LTD

[MNMJ^{+}21] Erik Myrstener, Sofia Ninnes, Carsten Meyer-Jacob, Tim Mighall, and Richard Bindler. Long-term development and trajectories of inferred lake-water organic carbon and pH in naturally acidic boreal lakes. Limnology and Oceanography, 66(6):2408–2422, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Matthews:2023:STF

[MO23] Stephanie A. Matthews and Mark D. Ohman. A space-for-time framework for forecasting the effects of ocean stratification on zooplankton vertical habitat use and trait composition. Limnology and Oceanography, 68(12):2688–2702, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Meyer:2022:ESH

[MOW^{+}22] Michael F. Meyer, Ted Ozersky, Kara H. Woo, Kirill Shchapov, Aaron W. E. Galloway, Julie B. Schram, Emma J. Rosi, Daniel D. Snow, Maxim A. Timofeyev, Dmitry Yu. Karnaukhov, Matthew R. Brousil, and Stephanie E. Hampton. Effects of spatially heterogeneous lakeside development on nearshore biotic communities in a large, deep, oligotrophic lake. Limnology and Oceanography, 67(12):2649–2664, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Michaud:2023:SOC

[MP23] Alexander B. Michaud and John C. Priscu. Sediment oxygen consumption in Antarctic subglacial environments. Limnology and Oceanography, 68(7):1557–1566, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Michelson:2023:HCC

[MPW^{+}23] Chantel I. Michelson, Michael J. Polito, Michael B. Wunder, Steven D. Emslie, Matthew D. McCarthy, William P. Patterson, and Kelton W. McMahon. Holocene climate change shifted Southern Ocean biogeochemical cycling and predator trophic dynamics. Limnology and Oceanography, 68(12): 2642–2653, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Maggioni:2024:WAS

[MRC^{+}24] Federica Maggioni, Patrick Raimbault, Olivier Chateau, Mireille Pujo-Pay, Yves Letourneur, and Riccardo Rodolfo-Metalpa. Warm-adapted sponges resist thermal stress by reallocating carbon and nitrogen resources from cell turnover to somatic growth. Limnology and Oceanography, 69(4):976– 991, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Merder:2021:DOC

[MRD^{+}21] Julian Merder, Heidelinde Röder, Thorsten Dittmar, Ulrike Feudel, Jan A. Freund, Gunnar Gerdts, Alexandra Kraberg, and Jutta Niggemann. Dissolved organic compounds with synchronous dynamics share chemical properties and origin. Limnology and Oceanography, 66(11):4001–4016, November 2021. CODEN LIOCAH. ISSN 0024-3590.

McHenry:2023:GVO

[MRH^{+}23] Jennifer McHenry, Andrew Rassweiler, Gema Hernan, Alexandra K. Dubel, Carolyn Curtin, Jakob Barzak, Nicholas Varias, and Sarah E. Lester. Geographic variation in organic carbon storage by seagrass beds. Limnology and Oceanography, 68(6):1256–1268, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Moreno:2024:HTI

[MRT^{+}24] Hugo Duarte Moreno, Sebastian Rokitta, Nelly Tremblay, Maarten Boersma, Elisabeth Groß, Helena C. L. Klip, Karen H. Wiltshire, and Cédric L. Meunier. Higher temperature, increased CO_{2}, and changing nutrient ratios alter the carbon metabolism and induce oxidative stress in a cosmopolitan diatom. Limnology and Oceanography, 69(1):121– 139, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Minguez:2020:CFC

[MSB^{+}20] Laetitia Minguez, Erik Sperfeld, Stella A. Berger, Jens C. Nejstgaard, and Mark O. Gessner. Changes in food characteristics reveal indirect effects of Lake Browning on zooplankton performance. Limnology and Oceanography, 65(5): 1028–1040, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Mdutyana:2022:KAU

[MSB^{+}22] Mhlangabezi Mdutyana, Xin Sun, Jessica M. Burger, Raquel F. Flynn, Shantelle Smith, Natasha R. van Horsten, Alakendra N. Roychoudhury, Hélène Planquette, Eva Bucciarelli, Sandy J. Thomalla, Bess B. Ward, and Sarah E. Fawcett. The kinetics of ammonium uptake and oxidation across the Southern Ocean. Limnology and Oceanography, 67 (4):973–991, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Mason:2023:PDC

[MSD23] Robert A. B. Mason, William J. Skirving, and Sophie G. Dove. Photoacclimation dynamics in coral holobionts responding to thermal and irradiance changes correlate with photon pressure per symbiont. Limnology and Oceanography, 68(11):2529–2543, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Melendez:2022:NED

[MSG^{+}22] Melissa Meléndez, Joseph Salisbury, Dwight Gledhill, Chris Langdon, Julio M. Morell, Derek Manzello, and Adrienne Sutton. Net ecosystem dissolution and respiration dominate metabolic rates at two western Atlantic reef sites. Limnology and Oceanography, 67(3):527–539, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Maturana:2020:ESS

[MSGW^{+}20] Claudia S. Maturana, Nicolás I. Segovia, Claudio A. González-Wevar, Angie Díaz, Sebastián Rosenfeld, Elie Poulin, Jennifer A. Jackson, and Peter Convey. Evidence of strong small-scale population structure in the Antarctic freshwater copepod Boeckella poppei in lakes on Signy Island, South Orkney Islands. Limnology and Oceanography, 65(9):2024–2040, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Moon:2021:SIC

[MSJ21] David L. Moon, J. Thad Scott, and Tom R. Johnson. Stoichiometric imbalances complicate prediction of phytoplankton biomass in U.S. lakes: Implications for nutrient criteria. Limnology and Oceanography, 66(8):2967–2978, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Macartney:2021:TEC

[MSL21] Keir J. Macartney, Marc Slattery, and Michael P. Lesser. Trophic ecology of Caribbean sponges in the mesophotic zone. Limnology and Oceanography, 66(4):1113–1124, April 2021. CODEN LIOCAH. ISSN 0024-3590.

McNabb:2023:OCS

[MT23] Brandon J. McNabb and Philippe D. Tortell. Oceanographic controls on Southern Ocean dimethyl sulfide distributions revealed by machine learning algorithms. Limnology and Oceanography, 68(3):616–630, March 2023. CODEN LI-OCAH. ISSN 0024-3590.

Maszczyk:2021:DDS

[MTB^{+}21] Piotr Maszczyk, Joanna Tałanda, Ewa Babkiewicz, Konrad Leniowski, and Paulina Urban. Daphnia depth selection in gradients of light intensity from different artificial sources: an evolutionary trap? Limnology and Oceanography, 65(3): 1367–1380, April 2021. CODEN LIOCAH. ISSN 0024-3590.

MacKeigan:2023:BBA

[MTP^{+}23] Paul W. MacKeigan, Zofia E. Taranu, Frances R. Pick, Beatrix E. Beisner, and Irene Gregory-Eaves. Both biotic and abiotic predictors explain significant variation in cyanobacteria biomass across lakes from temperate to subarctic zones. Limnology and Oceanography, 68(6):1360–1375, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Martiny:2020:GAM

[MUGL20] Adam C. Martiny, Lucas Ustick, Catherine A. Garcia, and Michael W. Lomas. Genomic adaptation of marine phytoplankton populations regulates phosphate uptake. Limnology and Oceanography, 66(6):S340–S350, January 2020. CODEN LIOCAH. ISSN 0024-3590.

McManus:2020:NQM

[MURK20] M. Conor McManus, David S. Ullman, Scott D. Rutherford, and Christopher Kincaid. Northern quahog (Mercenaria mercenaria) larval transport and settlement modeled for a temperate estuary. Limnology and Oceanography, 65(2): 289–303, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Makelin:2022:CFS

[MV22a] Saara Mäkelin and Anna Villnäs. Corrigendum: Food sources drive temporal variation in elemental stoichiometry of benthic consumers. Limnology and Oceanography, 67(8):1891– 1893, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Makelin:2022:FSD

[MV22b] Saara Mäkelin and Anna Villnäs. Food sources drive temporal variation in elemental stoichiometry of benthic consumers. Limnology and Oceanography, 67(4):784–799, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Monteiro:2020:CMM

[MVBG20] Juliana Mendes Monteiro, Ryan Vogwill, Karl Bischoff, and Deirdre B. Gleeson. Comparative metagenomics of microbial mats from hypersaline lakes at Rottnest Island (WA, Australia), advancing our understanding of the effect of mat community and functional genes on microbialite accretion. Limnology and Oceanography, 66(6):S293–S309, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Marino:2020:LTS

[MVP^{+}20] John A. Marino, Jr., Henry A. Vanderploeg, Steven A. Pothoven, Ashley K. Elgin, and Scott D. Peacor. Long-term survey data reveal large predator and temperature effects on population growth of multiple zooplankton species. Limnology and Oceanography, 65(4):694–706, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Mostovaya:2022:SID

[MWHR^{+}22] Alina Mostovaya, Michael Wind-Hansen, Paul Rousteau, Laura A. Bristow, and Bo Thamdrup. Sulfate- and irondependent anaerobic methane oxidation occurring side-by-side in freshwater lake sediment. Limnology and Oceanography, 67(1):231–246, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Martin:2022:RNA

[MWV22] Benjamin E. Martin, Jake R. Walsh, and M. Jake Vander Zanden. Rise of a native apex predator and an invasive zooplankton cause successive ecological regime shifts in a North Temperate Lake. Limnology and Oceanography, 67(S1):S163– S172, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Ma:2021:RPC

[MXL^{+}21] Lingqi Ma, Wupeng Xiao, Edward A. Laws, Xiaolin Bai, Kuo-Ping Chiang, Xin Liu, Jixin Chen, and Bangqin Huang. Responses of phytoplankton communities to the effect of internal wave-powered upwelling. Limnology and Oceanography, 66(4):1083–1098, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Ma:2023:RBD

[MY23] Qian-Yao Ma and Gui-Peng Yang. Responses of biogenic dimethylated sulfur compounds to environmental changes in the northwestern Pacific continental sea. Limnology and Oceanography, 68(7):1452–1469, July 2023. CODEN LI-OCAH. ISSN 0024-3590.

Mo:2021:BCO

[MZW^{+}21] Yuanyuan Mo, Wenjing Zhang, David M. Wilkinson, Zheng Yu, Peng Xiao, and Jun Yang. Biogeography and cooccurrence patterns of bacterial generalists and specialists in three subtropical marine bays. Limnology and Oceanography, 66(3):793–806, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Ma:2022:TPB

[MZXY22] Qian-Yao Ma, Hong-Hai Zhang, Feng Xu, and Gui-Peng Yang. Transformation processes of biogenic dimethylated sulfur compounds in the northwestern Pacific continental sea. Limnology and Oceanography, 67(4):903–917, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Ngochera:2020:STD

[NB20] Maxon J. Ngochera and Harvey A. Bootsma. Spatial and temporal dynamics of pCO_{2} and CO_{2} flux in tropical Lake Malawi. Limnology and Oceanography, 65(7):1594–1607, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Niedzwiedz:2023:GRR

[NB23] Sarina Niedzwiedz and Kai Bischof. Glacial retreat and rising temperatures are limiting the expansion of temperate kelp species in the future Arctic. Limnology and Oceanography, 68(4):816–830, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Nardelli:2021:KAA

[NCC^{+}21] Schuyler C. Nardelli, Megan A. Cimino, John A. Conroy, William R. Fraser, Deborah K. Steinberg, and Oscar Schofield. Krill availability in adjacent Adélie and gentoo penguin foraging regions near Palmer Station, Antarctica. Limnology and Oceanography, 66(6):2234–2250, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Neves:2024:TPO

[NDP^{+}24] Mayara Pereira Neves, Rosilene Luciana Delariva, Daniel M. Perkins, Clarice Bernhardt Fialho, and Pavel Kratina. Trophic plasticity of omnivorous fishes in natural and humandominated landscapes. Limnology and Oceanography, 69(1): 189–202, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Nogueira:2022:CWR

[NEB^{+}22] Juliana Nogueira, Heitor Evangelista, Lhoussaine Bouchaou, Luciane Moreira, Abdelfettah Sifeddine, Ahmed ElMouden, Fouad Msanda, Sandrine Caquineau, Francisco Javier Briceño-Zuluaga, Marcus Vinicius Licínio, Magloire Mandeng-Yogo, Mercedes Mendez-Millan, Renato C. Cordeiro, Bastiaan Knoppers, Manuel Moreira-Ramírez, and Renato Martins. Coastal wetland responses to a century of climate change in northern Sahara, Morocco. Limnology and Oceanography, 67(2):285–299, February 2022. CODEN LI-OCAH. ISSN 0024-3590.

Nardelli:2023:CCP

[NGSS23] Schuyler C. Nardelli, Patrick C. Gray, Sharon E. Stammerjohn, and Oscar Schofield. Characterizing coastal phytoplankton seasonal succession patterns on the West Antarctic Peninsula. Limnology and Oceanography, 68(4):845–861, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Nielson:2022:BBL

[NH22] Jeffrey R. Nielson and Stephen M. Henderson. Bottom boundary layer mixing processes across internal seiche cycles: Dominance of downslope flows. Limnology and Oceanography, 67(5):1111–1125, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Nielson:2023:PBL

[NH23] Jeffrey R. Nielson and Stephen M. Henderson. Periodic boundary layer separation and lateral intrusions observed above a sloping lakebed. Limnology and Oceanography, 68 (1):26–39, January 2023. CODEN LIOCAH. ISSN 0024-3590.

Nauer:2023:PLC

[NKH^{+}23] Philipp A. Nauer, Adam J. Kessler, Puspitaningsih Hall, Maria Elena Popa, Sophie ten Hietbrink, Tess Hutchinson, Wei Wen Wong, Karl Attard, Ronnie N. Glud, Chris Greening, and Perran L. M. Cook. Pulses of labile carbon cause transient decoupling of fermentation and respiration in permeable sediments. Limnology and Oceanography, 68(9):2141– 2152, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Nakajima:2021:NFR

[NSK^{+}21] Toshimi Nakajima, Ryo Sugimoto, Takahiro Kusunoki, Katsuhide Yokoyama, and Makoto Taniguchi. Nutrient fluxes from rivers, groundwater, and the ocean into the coastal embayment along the Sanriku ria coast, Japan. Limnology and Oceanography, 66(7):2728–2744, July 2021. CODEN LI-OCAH. ISSN 0024-3590.

Ninokawa:2020:BMS

[NTJ^{+}20] Aaron Ninokawa, Yuichiro Takeshita, Brittany M. Jellison, Laura J. Jurgens, and Brian Gaylord. Biological modification of seawater chemistry by an ecosystem engineer, the California mussel, Mytilus californianus. Limnology and Oceanography, 65(1):157–172, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Neumann:2021:MMD

[NvBE^{+}21] Andreas Neumann, Justus E. E. van Beusekom, Annika Eisele, Kay-Christian Emeis, Jana Friedrich, Ingrid Kröncke, Ella Lu Logemann, Julia Meyer, Céline Naderipour, Ulrike Schückel, Alexa Wrede, and Michael L. Zettler. Macrofauna as a major driver of bentho-pelagic exchange in the southern North Sea. Limnology and Oceanography, 66(6):2203–2217, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Orenstein:2022:MLT

[OAM^{+}22] Eric C. Orenstein, Sakina-Dorothée Ayata, Frédéric Maps, Érica C. Becker, Fabio Benedetti, Tristan Biard, Thibault de Garidel-Thoron, Jeffrey S. Ellen, Filippo Ferrario, Sarah L. C. Giering, Tamar Guy-Haim, Laura Hoebeke, Morten Hvitfeldt Iversen, Thomas Kiørboe, Jean-François Lalonde, Arancha Lana, Martin Laviale, Fabien Lombard, Tom Lorimer, Séverine Martini, Albin Meyer, Klas Ove Möller, Barbara Niehoff, Mark D. Ohman, Cédric Pradalier, Jean-Baptiste Romagnan, Simon-Martin Schröder, Virginie Sonnet, Heidi M. Sosik, Lars S. Stemmann, Michiel Stock, Tuba Terbiyik-Kurt, Nerea Valcárcel-Pérez, Laure Vilgrain, Guillaume Wacquet, Anya M. Waite, and Jean-Olivier Irisson. Machine learning techniques to characterize functional traits of plankton from image data. Limnology and Oceanography, 67(8):1647–1669, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Orcutt:2020:IDS

[OBB^{+}20] Beth N. Orcutt, James A. Bradley, William J. Brazelton, Emily R. Estes, Jacqueline M. Goordial, Julie A. Huber, Rose M. Jones, Nagissa Mahmoudi, Jeffrey J. Marlow, Sheryl Murdock, and Maria Pachiadaki. Impacts of deep-sea mining on microbial ecosystem services. Limnology and Oceanography, 65(7):1489–1510, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Olsthoorn:2020:UIS

[OBL20] Jason Olsthoorn, Cynthia E. Bluteau, and Gregory A. Lawrence. Under-ice salinity transport in low-salinity waterbodies. Limnology and Oceanography, 65(2):247–259, February 2020. CODEN LIOCAH. ISSN 0024-3590.

ODonnell:2021:TDE

[OBL21] Daniel R. O’Donnell, Sophia M. Beery, and Elena Litchman. Temperature-dependent evolution of cell morphology and carbon and nutrient content in a marine diatom. Limnology and Oceanography, 66(12):4334–4346, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Ottmann:2021:SSD

[OFM^{+}21] Daniel Ottmann, Øyvind Fiksen, Melissa Martín, Francisco Alemany, Laura Prieto, Diego Álvarez-Berastegui, and Patricia Reglero. Spawning site distribution of a bluefin tuna reduces jellyfish predation on early life stages. Limnology and Oceanography, 66(10):3669–3681, October 2021. CODEN LI-OCAH. ISSN 0024-3590.

Ortega:2020:EDI

[OGD20] Alejandra Ortega, Nathan R. Geraldi, and Carlos M. Duarte. Environmental DNA identifies marine macrophyte contributions to Blue Carbon sediments. Limnology and Oceanography, 66(6):3139–3149, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Ostertag:2023:WEI

[OJUN23] Eva Julia Marie Ostertag, Kai Jensen, Viktoria Unger, and Stefanie Nolte. Warming experiment indicates that increasing global temperatures may not affect windows of opportunity for salt marsh seedlings. Limnology and Oceanography, 68(10):2261–2273, October 2023. CODEN LIOCAH. ISSN 0024-3590.

Owen:2021:TSP

[OLFH21] Daniel P. Owen, Matthew H. Long, William K. Fitt, and Brian M. Hopkinson. Taxon-specific primary production rates on coral reefs in the Florida Keys. Limnology and Oceanography, 66(3):625–638, March 2021. CODEN LI-OCAH. ISSN 0024-3590.

Ottmann:2023:MMF

[OLR^{+}23] Daniel Ottmann, Tom J. Langbehn, Patricia Reglero, Diego Alvarez-Berastegui, and Øyvind Fiksen. Model of mesopelagic fish predation on eggs and larvae shows benefits of tuna spawning under full moon. Limnology and Oceanography, 68(12):2632–2641, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Obiol:2021:OHF

[OMM21] Aleix Obiol, Imer Muhovic, and Ramon Massana. Oceanic heterotrophic flagellates are dominated by a few widespread taxa. Limnology and Oceanography, 66(12):4240–4253, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Ozersky:2020:HSI

[ONH^{+}20] Ted Ozersky, Teofil Nakov, Stephanie E. Hampton, Nicholas L. Rodenhouse, Kara H. Woo, Kirill Shchapov, Katie Wright, Helena V. Pislegina, Lyubov R. Izmest’eva, Eugene A. Silow, Maxim A. Timofeev, and Marianne V. Moore. Hot and sick? Impacts of warming and a parasite on the dominant zooplankter of Lake Baikal. Limnology and Oceanography, 66 (10):2772–2786, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Oakes:2020:AST

[ORE20] Joanne M. Oakes, Philip M. Riekenberg, and Bradley D. Eyre. Assimilation and short-term processing of microphytobenthos nitrogen in intertidal sediments. Limnology and Oceanography, 65(10):2377–2389, October 2020. CODEN LI-OCAH. ISSN 0024-3590.

Osorio-Rodriguez:2021:SIF

[ORRMD^{+}21] Daniela Osorio-Rodriguez, Manuel Razo-Mejia, Nathan F. Dalleska, Alex L. Sessions, Victoria J. Orphan, and Jess F. Adkins. Sulfur isotope fractionations constrain the biological cycling of dimethylsulfoniopropionate in the upper ocean. Limnology and Oceanography, 66(10):3607–3618, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Okamoto:2020:EOC

[OSR20] Daniel K. Okamoto, Stephen C. Schroeter, and Daniel C. Reed. Effects of ocean climate on spatiotemporal variation in sea urchin settlement and recruitment. Limnology and Oceanography, 65(9):2076–2091, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Orland:2020:TGA

[OYM^{+}20] Chloé Orland, Kurt M. Yakimovich, Nadia C. S. Mykytczuk, Nathan Basiliko, and Andrew J. Tanentzap. Think global, act local: The small-scale environment mainly influences microbial community development and function in lake sediment. Limnology and Oceanography, 65(S1):S88–S100, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Pearman:2023:DPD

[PATL^{+}23] John K. Pearman, Janet Adamson, Georgia Thomson-Laing, Lucy Thompson, Sean Waters, Marcus J. Vandergoes, Jamie D. Howarth, and Susanna A. Wood. Deterministic processes drive national-scale patterns in lake surface sediment bacteria and eukaryotic assemblage composition. Limnology and Oceanography, 68(1):40–55, January 2023. CODEN LI-OCAH. ISSN 0024-3590.

Pace:2021:PBD

[PBC21] Michael L. Pace, Cal D. Buelo, and Stephen R. Carpenter. Phytoplankton biomass, dissolved organic matter, and temperature drive respiration in whole lake nutrient additions. Limnology and Oceanography, 66(6):2174–2186, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Posada-Bedoya:2021:ERI

[PBGGRB21] Andrés Posada-Bedoya, Andrés Gómez-Giraldo, and Ricardo Román-Botero. Effects of riverine inflows on the climatology of a tropical Andean reservoir. Limnology and Oceanography, 66(9):3535–3551, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Passos:2023:LCA

[PBP^{+}23] Tiago Passos, Angelo F. Bernardino, Dan Penny, Roberto Barcellos, Francisco U. Passos, Gabriel N. Nobrega, Tiago O. Ferreira, J. Boone Kauffman, and Christian J. Sanders. Low carbon accumulation in a macro-tidal mangrove forest on the Amazon coast. Limnology and Oceanography, 68(8):1936– 1948, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Pilla:2021:APA

[PC21] Rachel M. Pilla and Raoul-Marie Couture. Attenuation of photosynthetically active radiation and ultraviolet radiation in response to changing dissolved organic carbon in browning lakes: Modeling and parametrization. Limnology and Oceanography, 66(6):2278–2289, June 2021. CODEN LI-OCAH. ISSN 0024-3590.

Proctor:2023:LAP

[PCC^{+}23] Claudette Proctor, Pierre Coupel, Karen Casciotti, Jean-Eric Tremblay, Emily Zakem, Kevin R. Arrigo, and Matthew M. Mills. Light, ammonium, pH, and phytoplankton competition as environmental factors controlling nitrification. Limnology and Oceanography, 68(7):1490–1503, July 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Palmer:2021:HCW

[PCJ^{+}21] Michael J. Palmer, John Chételat, Heather E. Jamieson, Murray Richardson, and Marc Amyot. Hydrologic control on winter dissolved oxygen mediates arsenic cycling in a small subarctic lake. Limnology and Oceanography, 66(S1):S30– S46, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Pereira:2022:ARD

[PCN22] Tiago J. Pereira, Paula Vieira Castellões, and Sérgio A. Netto. Amazon River discharge impacts deep-sea meiofauna. Limnology and Oceanography, 67(10):2190–2203, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Park:2023:SPU

[PDK^{+}23] Jiwoon Park, Bryndan P. Durham, Rebecca S. Key, Ryan D. Groussman, Zinka Bartolek, Paulina Pinedo-Gonzalez, Nicholas J. Hawco, Seth G. John, Michael C. G. Carlson, Debbie Lindell, Lauren W. Juranek, Sara Ferrón, Francois Ribalet, E. Virginia Armbrust, Anitra E. Ingalls, and Randelle M. Bundy. Siderophore production and utilization by marine bacteria in the North Pacific Ocean. Limnology and Oceanography, 68(7):1636–1653, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Powers:2022:SES

[PFB^{+}22] Stephen M. Powers, Steven C. Fradkin, William Baccus, Carmen Archambault, John R. Boetsch, Matthew R. Brousil, Rebecca Lofgren, Ashley Rawhouser, and Stephanie E. Hampton. Summer ecosystem structure in mountain lakes linked to interannual variability of lake ice, snowpack, and landscape attributes. Limnology and Oceanography, 67(9): 2073–2087, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Penta:2021:RPD

[PFH21] W. Bryce Penta, James Fox, and Kimberly H. Halsey. Rapid photoacclimation during episodic deep mixing augments the biological carbon pump. Limnology and Oceanography, 66(5): 1850–1866, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Palinkas:2024:SVI

[PGB24] Cindy M. Palinkas, Cassie Gurbisz, and Miles C. Bolton. Sediment–vegetation interactions determine the fate of fluvial sediment in the upper reaches of a large estuary. Limnology and Oceanography, 69(3):515–523, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Paquette:2022:EDT

[PGEB22] Cindy Paquette, Irene Gregory-Eaves, and Beatrix E. Beisner. Environmental drivers of taxonomic and functional variation in zooplankton diversity and composition in freshwater lakes across Canadian continental watersheds. Limnology and Oceanography, 67(5):1081–1097, May 2022. CODEN LI-OCAH. ISSN 0024-3590.

Piontek:2021:OMC

[PGN^{+}21] Judith Piontek, Luisa Galgani, Eva-Maria Nöthig, Ilka Peeken, and Anja Engel. Organic matter composition and heterotrophic bacterial activity at declining summer sea ice in the central Arctic Ocean. Limnology and Oceanography, 66(S1):S343–S362, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Pathak:2022:HRW

[PHB^{+}22] Devanshi Pathak, Michael Hutchins, Lee E. Brown, Matthew Loewenthal, Peter Scarlett, Linda Armstrong, David Nicholls, Mike Bowes, François Edwards, and Gareth Old. High-resolution water-quality and ecosystem-metabolism modeling in lowland rivers. Limnology and Oceanography, 67(6):1313–1327, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Phillips:2020:TVR

[Phi20]  Joseph S. Phillips.  Time-varying responses of lake

metabolism to light and temperature. Limnology and Oceanography, 65(3):652–666, March 2020. CODEN LI-OCAH. ISSN 0024-3590.

Phillips:2021:CTV

[Phi21] Joseph S. Phillips. Corrigendum: Time-varying responses of lake metabolism to light and temperature. Limnology and Oceanography, 66(4):1333–1334, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Purz:2021:DLB

[PHM21] Anneke Kristin Purz, Dorothee Hodapp, and Stefanie Devi Moorthi. Dispersal, location of bloom initiation, and nutrient conditions determine the dominance of the harmful dinoflagellate Alexandrium catenella: a meta-ecosystem study. Limnology and Oceanography, 66(11):3928–3943, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Pinti:2023:GCS

[PJRV23] Jérôme Pinti, Sigrún H. Jónasdóttir, Nicholas R. Record, and André W. Visser. The global contribution of seasonally migrating copepods to the biological carbon pump. Limnology and Oceanography, 68(5):1147–1160, May 2023. CODEN LI-OCAH. ISSN 0024-3590.

Pinckney:2020:NBE

[PKKS20] James L. Pinckney, Eilea R. Knotts, Krystyn J. Kibler, and Erik M. Smith. Nutrient breakpoints for estuarine phytoplankton communities. Limnology and Oceanography, 66(6): 2999–3016, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Popendorf:2020:PTR

[PKV20] Kimberly J. Popendorf, Michal Koblížek, and Benjamin A. S. Van Mooy. Phospholipid turnover rates suggest that bacterial community growth rates in the open ocean are systematically underestimated. Limnology and Oceanography, 65 (8):1876–1890, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Preskienis:2021:SPG

[PLB^{+}21] Vilmantas Prėskienis, Isabelle Laurion, Frédéric Bouchard, Peter M. J. Douglas, Michael F. Billett, Daniel Fortier, and Xiaomei Xu. Seasonal patterns in greenhouse gas emissions from lakes and ponds in a High Arctic polygonal landscape. Limnology and Oceanography, 66(S1):S117–S141, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Penninck:2021:TLC

[PLLM21] Silvana B. Penninck, Rubens M. Lopes, Josiane F. Lima, and Margaret A. McManus. Thin layers in the coastal zone of Ubatuba, Brazil: Mechanisms of formation and dissipation. Limnology and Oceanography, 66(2):558–574, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Perga:2023:NBS

[PMD^{+}23] Marie-Elodie Perga, Camille Minaudo, Tomy Doda, Florent Arthaud, Harsh Beria, Hannah E. Chmiel, Nicolas Escoffier, Thibault Lambert, Raphaelle Napolleoni, Biel Obrador, Pascal Perolo, Janine Rüegg, Hugo Ulloa, and Damien Bouffard. Near-bed stratification controls bottom hypoxia in icecovered alpine lakes. Limnology and Oceanography, 68(6): 1232–1246, June 2023. CODEN LIOCAH. ISSN 0024-3590.

Pupier:2021:LBR

[PMF^{+}21] Chloé Alexandra Pupier, Miguel Mies, Maoz Fine, Ronaldo Bastos Francini-Filho, Frederico Pereira Brandini, Leonardo Zambotti-Villela, Pio Colepicolo, and Christine Ferrier-Pagès. Lipid biomarkers reveal the trophic plasticity of octocorals along a depth gradient. Limnology and Oceanography, 66(5):2078–2087, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Price:2021:IAE

[PMJ^{+}21] James T. Price, Rowan H. McLachlan, Christopher P. Jury, Robert J. Toonen, and Andréa G. Grottoli. Isotopic approaches to estimating the contribution of heterotrophic sources to Hawaiian corals. Limnology and Oceanography, 66(6):2393–2407, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Portner:2020:TEH

[PMRG20] Elan J. Portner, Unai Markaida, Carlos J. Robinson, and William F. Gilly. Trophic ecology of Humboldt squid, Dosidicus gigas, in conjunction with body size and climatic variability in the Gulf of California, Mexico. Limnology and Oceanography, 65(4):732–748, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Pinsonneault:2021:DOC

[PNT^{+}21] Andrew J. Pinsonneault, Patrick J. Neale, Maria Tzortziou, Elizabeth A. Canuel, Christina R. Pondell, Hannah Morrissette, Jonathan S. Lefcheck, and James Patrick Megonigal. Dissolved organic carbon sorption dynamics in tidal marsh soils. Limnology and Oceanography, 66(1):214–225, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Parent:2021:DWC

[PPB^{+}21] Geneviève J. Parent, Stéphane Plourde, Howard I. Browman, Julie Turgeon, and Adam Petrusek. Defining what constitutes a reliable dataset to test for hybridization and introgression in marine zooplankton: Comment on Choquet et al. 2020 “No evidence for hybridization between Calanus finmarchicus and C. glacialis in a subarctic area of sympatry”. Limnology and Oceanography, 66(10):3597–3602, October 2021. CODEN LIOCAH. ISSN 0024-3590. See reply [CSH21].

Pineda:2020:RSS

[PRS^{+}20] Jesús Pineda, Sally Rouse, Victoria Starczak, Karl Helfrich, and David Wiley. Response of small sharks to nonlinear internal waves. Limnology and Oceanography, 65(4):707–716, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Paillex:2020:HSI

[PSE^{+}20] Amael Paillex, Andre R. Siebers, Christian Ebi, Jorrit Mesman, and Christopher T. Robinson. High stream intermittency in an alpine fluvial network: Val Roseg, Switzerland. Limnology and Oceanography, 65(3):557–568, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Praetzel:2021:TSP

[PSK21] Leandra Stephanie Emilia Praetzel, Marcel Schmiedeskamp, and Klaus-Holger Knorr. Temperature and sediment properties drive spatiotemporal variability of methane ebullition in a small and shallow temperate lake. Limnology and Oceanography, 66(7):2598–2610, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Patriarca:2021:CEL

[PSNG^{+}21] Claudia Patriarca, Vicente T. Sedano-Núñez, Sarahi L. Garcia, Jonas Bergquist, Stefan Bertilsson, Per J. R. Sjöberg, Lars J. Tranvik, and Jeffrey A. Hawkes. Character and environmental lability of cyanobacteria-derived dissolved organic matter. Limnology and Oceanography, 66(2):496–509, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Prowe:2022:FWS

[PSNS22] A. E. Friederike Prowe, Bei Su, Jens C. Nejstgaard, and Markus Schartau. Food web structure and intraguild predation affect ecosystem functioning in an established plankton model. Limnology and Oceanography, 67(4):843–855, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Porskamp:2022:IME

[PSY^{+}22] Peter Porskamp, Alexandre C. G. Schimel, Mary Young, Alex Rattray, Yoann Ladroit, and Daniel Ierodiaconou. Integrating multibeam echosounder water-column data into benthic habitat mapping. Limnology and Oceanography, 67(8):1701– 1713, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Phillips:2023:OSS

[PUH^{+}23] Alexandra A. Phillips, Imanol Ulloa, Emily Hyde, Julia Agnich, Lewis Sharpnack, Katherine G. O’Malley, Samuel M. Webb, Kathryn M. Schreiner, Cody S. Sheik, Sergei Katsev, and Morgan Reed Raven. Organic sulfur from source to sink in low-sulfate Lake Superior. Limnology and Oceanography, 68(12):2716–2732, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Pinti:2022:FLP

[PVSP^{+}22] Jérôme Pinti, André W. Visser, Camila Serra-Pompei, Ken H. Andersen, Mark D. Ohman, and Thomas Kiørboe. Fear and loathing in the pelagic: How the seascape of fear impacts the biological carbon pump. Limnology and Oceanography, 67(6):1238–1256, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Pilla:2022:EIB

[PW22] Rachel M. Pilla and Craig E. Williamson. Earlier ice breakup induces changepoint responses in duration and variability of spring mixing and summer stratification in dimictic lakes. Limnology and Oceanography, 67(S1):S173–S183, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Peura:2020:OST

[PWS^{+}20] Sari Peura, Maxime Wauthy, Domenico Simone, Alexander Eiler, Karólína Einarsdóttir, Milla Rautio, and Stefan Bertilsson. Ontogenic succession of thermokarst thaw ponds is linked to dissolved organic matter quality and microbial degradation potential. Limnology and Oceanography, 66(6): S248–S263, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Pan:2020:IPD

[PYM^{+}20] Yongbo Pan, Jun Yang, George B. McManus, Senjie Lin, and Wenjing Zhang. Insights into protist diversity and biogeography in intertidal sediments sampled across a range of spatial scales. Limnology and Oceanography, 65(5):1103–1115, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Pei:2020:GRS

[PYY^{+}20] Lixin Pei, Siyuan Ye, Hongming Yuan, Shaofeng Pei, Shucheng Xie, and Jin Wang. Glomalin-related soil protein distributions in the wetlands of the Liaohe Delta, Northeast China: Implications for carbon sequestration and mineral weathering of coastal wetlands. Limnology and Oceanography, 65(5):979–991, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Qu:2021:EPE

[QBJG21] Liming Qu, John Beardall, Xiaowen Jiang, and Kunshan Gao. Elevated pCO_{2} enhances under light but reduces in darkness the growth rate of a diatom, with implications for the fate of phytoplankton below the photic zone. Limnology and Oceanography, 66(10):3630–3642, October 2021. CO-DEN LIOCAH. ISSN 0024-3590.

Quinlan:2021:RTP

[QFM^{+}21] Roberto Quinlan, Alessandro Filazzola, Octavia Mahdiyan, Arnab Shuvo, Kevin Blagrave, Carolyn Ewins, Luke Moslenko, Derek K. Gray, Catherine M. O’Reilly, and Sapna Sharma. Relationships of total phosphorus and chlorophyll in lakes worldwide. Limnology and Oceanography, 66(2):392– 404, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Qi:2021:BPE

[QMG21] Man Qi, Jessica MacGregor, and Keryn Gedan. Biogeomorphic patterns emerge with pond expansion in deteriorating marshes affected by relative sea level rise. Limnology and Oceanography, 66(4):1036–1049, April 2021. CODEN LI-OCAH. ISSN 0024-3590.

Quinones-Rivera:2022:DEB

[QRWT^{+}22] Zoraida J. Quiñones-Rivera, Björn Wissel, R. Eugene Turner, Nancy N. Rabalais, Dubravko Justić, Kerri P. Finlay, and Charles S. Milan. Divergent effects of biological and physical processes on dissolved oxygen and dissolved inorganic carbon dynamics on a eutrophied and hypoxic continental shelf. Limnology and Oceanography, 67(11):2603–2616, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Qin:2021:TRB

[QS21] Qubin Qin and Jian Shen. Typical relationships between phytoplankton biomass and transport time in river-dominated coastal aquatic systems. Limnology and Oceanography, 66 (8):3209–3220, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Risse-Buhl:2020:NSF

[RBAC^{+}20] Ute Risse-Buhl, Christine Anlanger, Antonis Chatzinotas, Christian Noss, Andreas Lorke, and Markus Weitere. Near streambed flow shapes microbial guilds within and across trophic levels in fluvial biofilms. Limnology and Oceanography, 65(10):2261–2277, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Rathbone:2022:THV

[RBD22] Michaela Rathbone, Kristen T. Brown, and Sophie Dove. Tolerance to a highly variable environment does not infer resilience to future ocean warming and acidification in a branching coral. Limnology and Oceanography, 67(1):272– 284, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Reyserhove:2020:FNA

[RBL^{+}20] Lien Reyserhove, Lore Bulteel, Jing Liu, Caroline Souffreau, Kristien I. Brans, Jessie M. T. Engelen, Luc De Meester, Frederik Hendrickx, Koenraad Muylaert, Steven A. J. Declerck, and Ellen Decaestecker. Food nutrient availability affects epibiont prevalence and richness in natural Daphnia populations. Limnology and Oceanography, 65(10):2529– 2540, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Rydin:2023:CDS

[RBR^{+}23] Emil Rydin, Elias Broman, Kasper Reitzel, Kaarina Lukkari, Martijn Hermans, Linda Kumblad, Magnus Karlsson, Anna Apler, Bo Ek, and Per J. R. Sjöberg. Contrasting distribution and speciation of sedimentary organic phosphorus among different basins of the Baltic Sea. Limnology and Oceanography, 68(4):767–779, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Roberts:2024:RHA

[RBR^{+}24] Megan E. Roberts, Maya P. Bhatia, Elden Rowland, Patrick L. White, Stephanie Waterman, Maria A. Cavaco, Patrick Williams, Jodi N. Young, Jenifer S. Spence, Jean-Éric Tremblay, Jean-Carlos Montero-Serrano, and Erin M. Bertrand. Rubisco in high Arctic tidewater glacier-marine systems: a new window into phytoplankton dynamics. Limnology and Oceanography, 69(4):802–817, April 2024. CO-DEN LIOCAH. ISSN 0024-3590.

Rulli:2022:SPN

[RBSB22] Mayra P. D. Rulli, Ann-Kristin Bergström, Ryan A. Sponseller, and Martin Berggren. Seasonal patterns in nutrient bioavailability in boreal headwater streams. Limnology and Oceanography, 67(5):1169–1183, May 2022. CODEN LI-OCAH. ISSN 0024-3590.

Ros:2020:UBB

[RCH^{+}20] Mickael Ros, Emma F. Camp, David J. Hughes, Joseph R. Crosswell, Mark E. Warner, William P. Leggat, and David J. Suggett. Unlocking the black-box of inorganic carbon-uptake and utilization strategies among coral endosymbionts (Symbiodiniaceae). Limnology and Oceanography, 65(8):1747– 1763, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Richardson:2021:MCA

[RCMA21] Murray Richardson, John Chételat, Gwyneth A. MacMillan, and Marc Amyot. Mercury concentrations and associations with dissolved organic matter are modified by water residence time in eastern Canadian lakes along a 30^{o} latitudinal gradient. Limnology and Oceanography, 66(S1):S64–S80, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Ruuskanen:2020:MGR

[RCS^{+}20] Matti O. Ruuskanen, Graham Colby, Kyra A. St. Pierre, Vincent L. St. Louis, Stéphane Aris-Brosou, and Alexandre J. Poulain. Microbial genomes retrieved from High Arctic lake sediments encode for adaptation to cold and oligotrophic environments. Limnology and Oceanography, 66(6):S233–S247, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Rock:2023:LPD

[RD23] Linnea A. Rock and Hilary A. Dugan. Lakes protect downstream riverine habitats from chloride toxicity. Limnology and Oceanography, 68(6):1216–1231, June 2023. CODEN LI-OCAH. ISSN 0024-3590.

Roberts:2021:SRS

[REF^{+}21] Derek C. Roberts, Galen C. Egan, Alexander L. Forrest, John L. Largier, Fabian A. Bombardelli, Bernard E. Laval, Stephen G. Monismith, and Geoffrey Schladow. The setup and relaxation of spring upwelling in a deep, rotationally influenced lake. Limnology and Oceanography, 66(4):1168– 1189, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Radice:2022:BNT

[RFB^{+}22] Veronica Z. Radice, Brian Fry, Kristen T. Brown, Sophie Dove, and Ove Hoegh-Guldberg. Biogeochemical niches and trophic plasticity of shallow and mesophotic corals recovering from mass bleaching. Limnology and Oceanography, 67(7): 1617–1630, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Richardson:2024:NRI

[RFW^{+}24] David C. Richardson, Alessandro Filazzola, R. Iestyn Woolway, M. Arshad Imrit, Damien Bouffard, Gesa A. Weyhenmeyer, John Magnuson, and Sapna Sharma. Nonlinear responses in interannual variability of lake ice to climate change. Limnology and Oceanography, 69(4):789–801, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Ribot:2022:CES

[RGP^{+}22] Miquel Ribot, Nancy B. Grimm, Lindsey D. Pollard, Daniel von Schiller, Amalia M. Handler, and Eugènia Martí. Consequences of an ecosystem state shift for nitrogen cycling in a desert stream. Limnology and Oceanography, 67(6):1274– 1286, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Rokitta:2023:FWS

[RGW^{+}23] Sebastian D. Rokitta, Christian H. Grossmann, Elisa Werner, Jannika Moye, Giulia Castellani, Eva-Maria Nöthig, and Björn Rost. Future warming stimulates growth and photosynthesis in an Arctic microalga more strongly than changes in light intensity or pCO_{2}. Limnology and Oceanography, 68 (12):2789–2799, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Ray:2023:STV

[RHA^{+}23] Nicholas E. Ray, Meredith A. Holgerson, Mikkel Rene Andersen, Jānis Bikše, Lauren E. Bortolotti, Martyn Futter, Ilga Kokorīte, Alan Law, Cory McDonald, Jorrit P. Mesman, Mike Peacock, David C. Richardson, Julien Arsenault, Sheel Bansal, Kaelin Cawley, McKenzie Kuhn, Amir Reza Shahabinia, and Facundo Smufer. Spatial and temporal variability in summertime dissolved carbon dioxide and methane in temperate ponds and shallow lakes. Limnology and Oceanography, 68(7):1530–1545, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Rothausler:2021:MWS

[RHM^{+}21] Eva Rothäusler, Iván A. Hinojosa, Julio Moraga, Matias Pizarro-Koch, Marcel Ramos, and Martin Thiel. At the mercy of the winds: The seasonal dynamics of floating and stranded seaweeds at mid-latitudes. Limnology and Oceanography, 66(12):4391–4402, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Roley:2023:CPP

[RHP^{+}23] Sarah S. Roley, Robert O. Hall, Jr., William Perkins, Vanessa A. Garayburu-Caruso, and James C. Stegen. Coupled primary production and respiration in a large river contrasts with smaller rivers and streams. Limnology and Oceanography, 68(11):2461–2475, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Rahav:2022:HBO

[RHS^{+}22] Eyal Rahav, Barak Herut, Dina Spungin, Adi Levi, Margaret R. Mulholland, and Ilana Berman-Frank. Heterotrophic bacteria outcompete diazotrophs for orthophosphate in the Mediterranean Sea. Limnology and Oceanography, 67(1):159– 171, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Ryan-Keogh:2023:STD

[RKBT^{+}23] Thomas James Ryan-Keogh, Emma Lewis Bone, Sandy J. Thomalla, Lisl Lain, Marie E. Smith, Stewart Bernard, and Marcello Vichi. Spatial and temporal drivers of fluorescence quantum yield variability in the Southern Ocean. Limnology and Oceanography, 68(3):569–582, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Ren:2021:WEI

[RLL^{+}21] Lijuan Ren, Yuanyuan Liu, Torben L. Lauridsen, Martin Søndergaard, Boping Han, Jianjun Wang, Erik Jeppesen, Jizhong Zhou, and Qinglong L. Wu. Warming exacerbates the impact of nutrient enrichment on microbial functional potentials important to the nutrient cycling in shallow lake mesocosms. Limnology and Oceanography, 66(6):2481–2495, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Ravaglioli:2020:OAA

[RLP^{+}20] Chiara Ravaglioli, Claudio Lardicci, Antonio Pusceddu, Eleonora Arpe, Silvia Bianchelli, Emanuela Buschi, and Fabio Bulleri. Ocean acidification alters meiobenthic assemblage composition and organic matter degradation rates in seagrass sediments. Limnology and Oceanography, 65(1):37– 50, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Rassweiler:2022:HDF

[RMH^{+}22] Andrew Rassweiler, Scott D. Miller, Sally J. Holbrook, Matthew Lauer, Mark A. Strother, Sarah E. Lester, Thomas C. Adam, Jean Wencélius, and Russell J. Schmitt. How do fisher responses to macroalgal overgrowth influence the resilience of coral reefs? Limnology and Oceanography, 67(S1):S365–S377, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Ray:2021:DPC

[RMW^{+}21] Raghab Ray, Toshihiro Miyajima, Atsushi Watanabe, Masaya Yoshikai, Charissa M. Ferrera, Iris Orizar, Takashi Nakamura, Maria Lourdes San Diego-McGlone, Eugene C. Herrera, and Kazuo Nadaoka. Dissolved and particulate carbon export from a tropical mangrove-dominated riverine system. Limnology and Oceanography, 66(11):3944–3962, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Ran:2023:SLL

[RNW^{+}23] Fengwei Ran, Xiaodong Nie, Shilan Wang, Tao Xiao, Changrong Yang, Yi Liu, Lingxia Wang, Yaojun Liu, Xin Chu, and Zhongwu Li. A significant loop of lake sedimentation state links catchment eco-surroundings and lake biogenic regime shifts. Limnology and Oceanography, 68(8): 1775–1790, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Royer:2020:SNP

[Roy20] Todd V. Royer. Stoichiometry of nitrogen, phosphorus, and silica loads in the Mississippi–Atchafalaya River basin reveals spatial and temporal patterns in risk for cyanobacterial blooms. Limnology and Oceanography, 65(2):325–335, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Rowe:2022:PLA

[RPA^{+}22] Mark D. Rowe, Sara E. Prendergast, Karen M. Alofs, David B. Bunnell, Edward S. Rutherford, and Eric J. Anderson. Predicting larval alewife transport in Lake Michigan using hydrodynamic and Lagrangian particle dispersion models. Limnology and Oceanography, 67(9):2042–2058, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Rii:2022:SEV

[RPKC22] Yoshimi M. Rii, Logan M. Peoples, David M. Karl, and Matthew J. Church. Seasonality and episodic variation in picoeukaryote diversity and structure reveal community resilience to disturbances in the North Pacific Subtropical Gyre. Limnology and Oceanography, 67(S1):S331–S351, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Roth-Rosenberg:2021:DMC

[RRAO^{+}21] Dalit Roth-Rosenberg, Dikla Aharonovich, Anne Willem Omta, Michael J. Follows, and Daniel Sher. Dynamic macromolecular composition and high exudation rates in Prochlorococcus. Limnology and Oceanography, 66(5):1759–1773, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Rocher-Ros:2021:MOP

[RRHS^{+}21] Gerard Rocher-Ros, Tamara K. Harms, Ryan A. Sponseller, Maria Väisänen, Carl-Magnus Mörth, and Reiner Giesler. Metabolism overrides photo-oxidation in CO_{2} dynamics of Arctic permafrost streams. Limnology and Oceanography, 66 (S1):S169–S181, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Reustle:2020:CCM

[RS20] Joseph W. Reustle and Delbert L. Smee. Cloudy with a chance of mesopredator release: Turbidity alleviates topdown control on intermediate predators through sensory disruption. Limnology and Oceanography, 65(10):2278–2290, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Robertson:2021:TCN

[RS21] Ben P. Robertson and Candida Savage. Thresholds in catchment nitrogen load for shifts from seagrass to nuisance macroalgae in shallow intertidal estuaries. Limnology and Oceanography, 65(3):1353–1366, April 2021. CODEN LI-OCAH. ISSN 0024-3590.

Rogener:2020:PMO

[RSH^{+}20] Mary Katherine Rogener, Rachel E. Sipler, Kimberley S. Hunter, Deborah A. Bronk, and Samantha B. Joye. Pelagic methane oxidation in the northern Chukchi Sea. Limnology and Oceanography, 65(1):96–110, January 2020. CODEN LI-OCAH. ISSN 0024-3590.

Rehn:2023:LTC

[RSLW23] Lukas Rehn, Ryan A. Sponseller, Hjalmar Laudon, and Marcus B. Wallin. Long-term changes in dissolved inorganic carbon across boreal streams caused by altered hydrology. Limnology and Oceanography, 68(2):409–423, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Rudberg:2024:CGC

[RSP^{+}24] David Rudberg, Jonathan Schenk, Gustav Pajala, Henrique Sawakuchi, Anna Sieczko, Ingrid Sundgren, Nguyen Thanh Duc, Jan Karlsson, Sally MacIntyre, John Melack, and David Bastviken. Contribution of gas concentration and transfer velocity to CO_{2} flux variability in northern lakes. Limnology and Oceanography, 69(4):818–833, April 2024. CODEN LI-OCAH. ISSN 0024-3590.

Rodellas:2021:CUG

[RST^{+}21] Valentí Rodellas, Thomas C. Stieglitz, Joseph J. Tamborski, Pieter van Beek, Aladin Andrisoa, and Peter G. Cook. Conceptual uncertainties in groundwater and porewater fluxes estimated by radon and radium mass balances. Limnology and Oceanography, 66(4):1237–1255, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Ryderheim:2023:SHT

[RTK23] Fredrik Ryderheim, Uffe H. Thygesen, and Thomas Kiørboe. Short handling times allow for active prey selection in suspension feeding copepods. Limnology and Oceanography, 68 (4):891–901, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Reading:2020:LUD

[RTM^{+}20] Michael J. Reading, Douglas R. Tait, Damien T. Maher, Luke C. Jeffrey, Arun Looman, Ceylena Holloway, Haile A.

Shishaye, Summer Barron, and Isaac R. Santos. Land use drives nitrous oxide dynamics in estuaries on regional and global scales. Limnology and Oceanography, 65(8):1903–1920, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Reading:2021:SGD

[RTM^{+}21] Michael J. Reading, Douglas R. Tait, Damien T. Maher, Luke C. Jeffrey, Rogger E. Correa, James P. Tucker, Haile A. Shishaye, and Isaac R. Santos. Submarine groundwater discharge drives nitrous oxide source/sink dynamics in a metropolitan estuary. Limnology and Oceanography, 66(5): 1665–1686, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Ramaekers:2023:EDC

[RVBP23] Lana Ramaekers, Bram Vanschoenwinkel, Luc Brendonck, and Tom Pinceel. Elevated dissolved carbon dioxide and associated acidification delays maturation and decreases calcification and survival in the freshwater crustacean Daphnia magna. Limnology and Oceanography, 68(7):1624–1635, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Rasigraf:2020:MCC

[RvHF^{+}20] Olivia Rasigraf, Niels A. G. M. van Helmond, Jeroen Frank, Wytze K. Lenstra, Matthias Egger, Caroline P. Slomp, and Mike S. M. Jetten. Microbial community composition and functional potential in Bothnian Sea sediments is linked to Fe and S dynamics and the quality of organic matter. Limnology and Oceanography, 65(S1):S113–S133, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Ruffing:2022:PSM

[RVS^{+}22] Claire M. Ruffing, Allison M. Veach, Anne Schechner, Janine Rüegg, Matt T. Trentman, and Walter K. Dodds. Prairie stream metabolism recovery varies based on antecedent hydrology across a stream network after a bank-full flood. Limnology and Oceanography, 67(9):1986–1999, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Rewrie:2023:SSI

[RVvB^{+}23] Louise C. V. Rewrie, Yoana G. Voynova, Justus E. E. van Beusekom, Tina Sanders, Arne Körtzinger, Holger Brix, Gregor Ollesch, and Burkard Baschek. Significant shifts in inorganic carbon and ecosystem state in a temperate estuary (1985–2018). Limnology and Oceanography, 68(8):1920–1935, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Ran:2024:DDB

[RWL^{+}24] Lihua Ran, Martin G. Wiesner, Yuzhao Liang, Wen Liang, Lanlan Zhang, Zhi Yang, Hongliang Li, and Jianfang Chen. Differential dissolution of biogenic silica significantly affects the utility of sediment diatoms as paleoceanographic proxies. Limnology and Oceanography, 69(3):467–481, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Regier:2023:SDD

[RWMP^{+}23] Peter J. Regier, Nicholas D. Ward, Allison N. Myers-Pigg, Jay Grate, Michael J. Freeman, and Ruby N. Ghosh. Seasonal drivers of dissolved oxygen across a tidal creek–marsh interface revealed by machine learning. Limnology and Oceanography, 68(10):2359–2374, October 2023. CODEN LI-OCAH. ISSN 0024-3590.

Roberts:2022:RDD

[RWS^{+}22] Keryn L. Roberts, Wei Wen Wong, Jeff Shimeta, Adam J. Kessler, and Perran L. M. Cook. Recovery of denitrification and dissimilatory nitrate reduction to ammonium following reoxygenation of sediments from a periodically hypoxic temperate lagoon. Limnology and Oceanography, 67(8):1879– 1890, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Robison:2021:STH

[RWT^{+}21] Andrew L. Robison, Wilfred M. Wollheim, Bonnie Turek, Cynthia Bova, Carter Snay, and Ruth K. Varner. Spatial and temporal heterogeneity of methane ebullition in lowland headwater streams and the impact on sampling design. Limnology and Oceanography, 66(12):4063–4076, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Subhas:2020:CAA

[SAD^{+}20] Adam V. Subhas, Jess F. Adkins, Sijia Dong, Nick E. Rollins, and William M. Berelson. The carbonic anhydrase activity of sinking and suspended particles in the North Pacific Ocean. Limnology and Oceanography, 65(3):637–651, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Shen:2020:MPP

[SB20] Yuan Shen and Ronald Benner. Molecular properties are a primary control on the microbial utilization of dissolved organic matter in the ocean. Limnology and Oceanography, 65(5):1061–1071, May 2020. CODEN LIOCAH. ISSN 0024-3590. See comment [LD22].

Shen:2022:RCC

[SB22] Yuan Shen and Ronald Benner. Reply to comment: Controls on turnover of marine dissolved organic matter — testing the null hypothesis of purely concentration-driven uptake. Limnology and Oceanography, 67(3):680–683, March 2022. CODEN LIOCAH. ISSN 0024-3590. See [LD22].

Segatto:2020:MCD

[SBB20] Pier Luigi Segatto, Tom J. Battin, and Enrico Bertuzzo. Modeling the coupled dynamics of stream metabolism and microbial biomass. Limnology and Oceanography, 65(7): 1573–1593, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Smith:2021:FIU

[SBB^{+}21] Jason M. Smith, Gordon Blasco, Mark A. Brzezinski, John M. Melack, Daniel C. Reed, and Robert J. Miller. Factors influencing urea use by giant kelp (Macrocystis pyrifera, Phaeophyceae). Limnology and Oceanography, 66(4):1190– 1200, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Saba:2021:TBU

[SBD^{+}21] Grace K. Saba, Adrian B. Burd, John P. Dunne, Santiago Hernández-León, Angela H. Martin, Kenneth A. Rose, Joseph Salisbury, Deborah K. Steinberg, Clive N. Trueman, Rod W. Wilson, and Stephanie E. Wilson. Toward a better understanding of fish-based contribution to ocean carbon flux. Limnology and Oceanography, 66(5):1639–1664, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Salis:2023:SIF

[SBH23] Romana K. Salis, Georgina L. Brennan, and Lars-Anders Hansson. Successful invasions to freshwater systems double with climate warming. Limnology and Oceanography, 68(4): 953–962, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Singh:2021:IIC

[SBL^{+}21] Arvind Singh, Lennart T. Bach, Carolin R. Löscher, Allanah J. Paul, Narendra Ojha, and Ulf Riebesell. Impact of increasing carbon dioxide on dinitrogen and carbon fixation rates under oligotrophic conditions and simulated upwelling. Limnology and Oceanography, 66(7):2855–2867, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Schanke:2021:BEV

[SBM^{+}21] Nicole L. Schanke, Francesco Bolinesi, Olga Mangoni, Christian Katlein, Philipp Anhaus, Mario Hoppmann, Peter A. Lee, and Giacomo R. DiTullio. Biogeochemical and ecological variability during the late summer–early autumn transition at an ice-floe drift station in the Central Arctic Ocean. Limnology and Oceanography, 66(S1):S363–S382, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Stirnimann:2021:PCC

[SBV^{+}21] Luca Stirnimann, Thomas G. Bornman, Hans M. Verheye, Marie-Lou Bachèlery, Janine van der Poel, and Sarah E. Fawcett. Plankton community composition and productivity near the Subantarctic Prince Edward Islands archipelago in autumn. Limnology and Oceanography, 66(12):4140–4158, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Sosa:2020:PCS

[SBW^{+}20] Oscar A. Sosa, Timothy J. Burrell, Samuel T. Wilson, Rhea K. Foreman, David M. Karl, and Daniel J. Repeta. Phosphonate cycling supports methane and ethylene supersaturation in the phosphate-depleted western North Atlantic Ocean. Limnology and Oceanography, 65(10):2443–2459, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Steiner:2021:PBI

[SBW21] Oscar Sepúlveda Steiner, Damien Bouffard, and Alfred Wüest. Persistence of bioconvection-induced mixed layers in a stratified lake. Limnology and Oceanography, 65(3):1531– 1547, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Su:2020:SPO

[SCB^{+}20] Jianzhong Su, Wei-Jun Cai, Jean Brodeur, Najid Hussain, Baoshan Chen, Jeremy M. Testa, K. Michael Scaboo, Deb P. Jaisi, Qiang Li, Minhan Dai, and Jeffrey Cornwell. Source partitioning of oxygen-consuming organic matter in the hypoxic zone of the Chesapeake Bay. Limnology and Oceanography, 65(8):1801–1817, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Selden:2021:CFH

[SCC^{+}21] Corday R. Selden, P. Dreux Chappell, Sophie Clayton, Alfonso Macías-Tapia, Peter W. Bernhardt, and Margaret R. Mulholland. A coastal N_{2} fixation hotspot at the Cape Hatteras front: Elucidating spatial heterogeneity in diazotroph activity via supervised machine learning. Limnology and Oceanography, 66(5):1832–1849, May 2021. CODEN LI-OCAH. ISSN 0024-3590.

Sprules:2022:MEW

[SCM22] William Gary Sprules, Hélène Cyr, and Charles W. Menza. Multiscale effects of wind-induced hydrodynamics on lake plankton distribution. Limnology and Oceanography, 67(7): 1631–1646, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Su:2021:SCC

[SCT^{+}21] Jianzhong Su, Wei-Jun Cai, Jeremy M. Testa, Jean R. Brodeur, Baoshan Chen, K. Michael Scaboo, Ming Li, Chunqi Shen, Margaret Dolan, Yuan-Yuan Xu, Yafeng Zhang, and Najid Hussain. Supply-controlled calcium carbonate dissolution decouples the seasonal dissolved oxygen and pH minima in Chesapeake Bay. Limnology and Oceanography, 66(10):3796–3810, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Solano:2023:SRE

[SDB^{+}23] Vanessa Solano, Clément Duvert, Christian Birkel, Damien T. Maher, Erica A. García, and Lindsay B. Hutley. Stream respiration exceeds CO_{2} evasion in a low-energy, oligotrophic tropical stream. Limnology and Oceanography, 68(5):1132– 1146, May 2023. CODEN LIOCAH. ISSN 0024-3590.

Shaked:2023:CAM

[SdBW^{+}23] Yeala Shaked, Dirk de Beer, Siyuan Wang, Futing Zhang, Anna-Neva Visser, Meri Eichner, and Subhajit Basu. Coacquisition of mineral-bound iron and phosphorus by natural Trichodesmium colonies. Limnology and Oceanography, 68 (5):1064–1077, May 2023. CODEN LIOCAH. ISSN 0024-

Shen:2022:SRB

[SDL22] Jian Shen, Jiabi Du, and Lisa V. Lucas. Simple relationships between residence time and annual nutrient retention, export, and loading for estuaries. Limnology and Oceanography, 67(4):918–933, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Suarez:2023:EDP

[SDS^{+}23] Ena Lucia Suarez, Lukas De Ventura, Arno Stöckli, César Ordóñez, Mridul K. Thomas, Bastiaan W. Ibelings, and Daniel F. McGinnis. The emergence and dominance of Planktothrix rubescens as an hypolimnetic cyanobacterium in response to re-oligotrophication of a deep peri-alpine lake. Limnology and Oceanography, 68(6):1346–1359, June 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Stukel:2021:SSG

[SDSGR21] Michael R. Stukel, Moira Décima, Karen E. Selph, and Andres Gutiérrez-Rodríguez. Size-specific grazing and competitive interactions between large salps and protistan grazers. Limnology and Oceanography, 65(6):2521–2534, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Stief:2021:RMS

[SEG21] Peter Stief, Marcus Elvert, and Ronnie N. Glud. Respiration by “marine snow” at high hydrostatic pressure: Insights from continuous oxygen measurements in a rotating pressure tank. Limnology and Oceanography, 66(7):2797–2809, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Stadler:2020:LTD

[SEK20] Masumi Stadler, Elisabet Ejarque, and Martin J. Kainz. Inlake transformations of dissolved organic matter composition in a subalpine lake do not change its biodegradability. Limnology and Oceanography, 65(7):1554–1572, July 2020. CO-DEN LIOCAH. ISSN 0024-3590.

Scholtysik:2022:GFB

[SGA^{+}22] Grzegorz Scholtysik, Tobias Goldhammer, Helge W. Arz, Matthias Moros, Ralf Littke, and Michael Hupfer. Geochemical focusing and burial of sedimentary iron, manganese, and phosphorus during lake eutrophication. Limnology and Oceanography, 67(4):768–783, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Schutze:2021:WPL

[SGGB21] Martin Schütze, Philipp Gatz, Benjamin-Silas Gilfedder, and Harald Biester. Why productive lakes are larger mercury sedimentary sinks than oligotrophic brown water lakes. Limnology and Oceanography, 66(4):1316–1332, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Sutherland:2021:RFH

[SGK^{+}21] Kevin M. Sutherland, Kalina C. Grabb, Jennifer S. Karolewski, Lina Taenzer, Colleen M. Hansel, and Scott D. Wankel. The redox fate of hydrogen peroxide in the marine water column. Limnology and Oceanography, 66(10):3828– 3841, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Socha:2023:UIP

[SGL^{+}23] Ellie Socha, Adrianna Gorsky, Noah R. Lottig, Gretchen Gerrish, Emily C. Whitaker, and Hilary A. Dugan. Under-ice plankton community response to snow removal experiment in bog lake. Limnology and Oceanography, 68(5):1001–1018, May 2023. CODEN LIOCAH. ISSN 0024-3590.

Stetler:2021:ASW

[SGM^{+}21] Jonathan T. Stetler, Scott Girdner, Jeremy Mack, Luke A. Winslow, Taylor H. Leach, and Kevin C. Rose. Atmospheric stilling and warming air temperatures drive long-term changes in lake stratification in a large oligotrophic lake. Limnology and Oceanography, 66(3):954–964, March 2021. CO-DEN LIOCAH. ISSN 0024-3590.

Sonnet:2022:LWS

[SGM^{+}22] Virginie Sonnet, Lionel Guidi, Colleen B. Mouw, Gavino Puggioni, and Sakina-Dorothée Ayata. Length, width, shape regularity, and chain structure: time series analysis of phytoplankton morphology from imagery. Limnology and Oceanography, 67(8):1850–1864, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Sullivan:2022:DOC

[SGS22] Kristy Lee Sullivan, Evelyn E. Gaiser, and Hilary M. Swain. Dissolved organic carbon as a driver of seasonal and multiyear phytoplankton assembly oscillations in a subtropical monomictic lake. Limnology and Oceanography, 67(S1):S416– S429, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Suominen:2022:IBM

[SGSD^{+}22] Saara Suominen, Gonzalo V. Gomez-Saez, Thorsten Dittmar, Jaap S. Sinninghe Damsté, and Laura Villanueva. Interplay between microbial community composition and chemodiversity of dissolved organic matter throughout the Black Sea water column redox gradient. Limnology and Oceanography, 67 (2):329–347, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Sherman:2020:PEC

[SGSF20] Jonathan Sherman, Maxim Y. Gorbunov, Oscar Schofield, and Paul G. Falkowski. Photosynthetic energy conversion efficiency in the West Antarctic Peninsula. Limnology and Oceanography, 66(6):2912–2925, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Sangil:2022:RLS

[SH22] Carlos Sangil and José Carlos Hernández. Recurrent largescale sea urchin mass mortality and the establishment of a long-lasting alternative macroalgae-dominated community state. Limnology and Oceanography, 67(S1):S430–S443, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Schmitt:2022:EPC

[SHBA22] Russell J. Schmitt, Sally J. Holbrook, Andrew J. Brooks, and Thomas C. Adam. Evaluating the precariousness of coral recovery when coral and macroalgae are alternative basins of attraction. Limnology and Oceanography, 67(S1):S285–S297, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Shiozaki:2021:AOU

[SHE^{+}21] Takuhei Shiozaki, Fuminori Hashihama, Hisashi Endo, Minoru Ijichi, Noriko Takeda, Akiko Makabe, Amane Fujiwara, Shigeto Nishino, and Naomi Harada. Assimilation and oxidation of urea-derived nitrogen in the summer Arctic Ocean. Limnology and Oceanography, 66(12):4159–4170, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Swanborn:2022:BSE

[SHPW22] Denise J. B. Swanborn, Veerle A. I. Huvenne, Simon J. Pittman, and Lucy C. Woodall. Bringing seascape ecology to the deep seabed: a review and framework for its application. Limnology and Oceanography, 67(1):66–88, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Santos:2022:CSA

[SHS^{+}22] Isaac R. Santos, Vanessa Hatje, Oscar Serrano, David Bastviken, and Dorte Krause-Jensen. Carbon sequestration in aquatic ecosystems: Recent advances and challenges. Limnology and Oceanography, 67(S2):S1–S5, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Smith:2021:CPT

[SHZ^{+}21] Alaina N. Smith, Gwenn M. M. Hennon, Erik R. Zinser, Benjamin C. Calfee, Jeremy W. Chandler, and Andrew D. Barton. Comparing Prochlorococcus temperature niches in the lab and across ocean basins. Limnology and Oceanography, 66(7):2632–2647, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Sawall:2021:CES

[SIP21] Yvonne Sawall, Maysa Ito, and Christian Pansch. Chronically elevated sea surface temperatures revealed high susceptibility of the eelgrass Zostera marina to winter and spring warming. Limnology and Oceanography, 66(12):4112–4124, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Saber:2020:LTF

[SJH20] Ali Saber, David E. James, and Donald F. Hayes. Longterm forecast of water temperature and dissolved oxygen profiles in deep lakes using artificial neural networks conjugated with wavelet transform. Limnology and Oceanography, 65(6): 1297–1317, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Song:2021:SRD

[SJJ^{+}21] Hongjun Song, Rubao Ji, Meibing Jin, Yun Li, Zhixuan Feng, Øystein Varpe, and Cabell S. Davis. Strong and regionally distinct links between ice-retreat timing and phytoplankton production in the Arctic Ocean. Limnology and Oceanography, 65(6):2498–2508, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Scholze:2021:PPS

[SJR21] Caroline Scholze, Bo Barker Jørgensen, and Hans Røy. Psychrophilic properties of sulfate-reducing bacteria in Arctic marine sediments. Limnology and Oceanography, 66(S1): S293–S302, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Strzalek:2024:RDI

[SKA^{+}24] Małgorzata Strzałek, Lech Kufel, Karina Apolinarska, Marcin Becher, Elżbieta Biardzka, Michał Brzozowski, Rafał Kiełczewski, Grzegorz Kowalewski, Andrzej Pukacz, Michał Woszczyk, and Mariusz Pełechaty. Recycling and deposition of inorganic carbon from calcium carbonate encrustations of charophytes. Limnology and Oceanography, 69(2):279–289, February 2024. CODEN LIOCAH. ISSN 0024-3590.

Sterling:2022:EHA

[SKB^{+}22] Alexa R. Sterling, Riley D. Kirk, Matthew J. Bertin, Tatiana A. Rynearson, David G. Borkman, Marissa C. Caponi, Jessica Carney, Katherine A. Hubbard, Meagan A. King, Lucie Maranda, Emily J. McDermith, Nina R. Santos, Jacob P. Strock, Erin M. Tully, Samantha B. Vaverka, Patrick D. Wilson, and Bethany D. Jenkins. Emerging harmful algal blooms caused by distinct seasonal assemblages of a toxic diatom. Limnology and Oceanography, 67(11):2341–2359, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Stoof-Leichsenring:2022:SDI

[SLHL^{+}22] Kathleen R. Stoof-Leichsenring, Sichao Huang, Sisi Liu, Weihan Jia, Kai Li, Xingqi Liu, Luidmila A. Pestryakova, and Ulrike Herzschuh. Sedimentary DNA identifies modern and past macrophyte diversity and its environmental drivers in high-latitude and high-elevation lakes in Siberia and China. Limnology and Oceanography, 67(5):1126–1141, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Schlosser:2022:LWE

[SLJ^{+}22] Tamara L. Schlosser, Andrew J. Lucas, Nicole L. Jones, Jonathan D. Nash, and Gregory N. Ivey. Local winds and encroaching currents drive summertime subsurface blooms over a narrow shelf. Limnology and Oceanography, 67(4):888–902, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Silverthorn:2024:RNS

[SLRS^{+}24] Teresa Silverthorn, Naiara López-Rojo, Romain Sarremejane, Arnaud Foulquier, Vincent Chanudet, Abdelkader Azougui, Rubén del Campo, Gabriel Singer, and Thibault Datry. River network-scale drying impacts the spatiotemporal dynamics of greenhouse gas fluxes. Limnology and Oceanography, 69(4):861–873, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Song:2021:RBN

[SLZ^{+}21a] Guodong Song, Sumei Liu, Jing Zhang, Zhuoyi Zhu, Guiling Zhang, Hannah K. Marchant, Marcel M. M. Kuypers, and Gaute Lavik. Response of benthic nitrogen cycling to estuarine hypoxia. Limnology and Oceanography, 66(3):652–666, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Soulignac:2021:RCR

[SLZ^{+}21b] Frédéric Soulignac, Ulrich Lemmin, Seyed Mahmood Hamze Ziabari, Htet Kyi Wynn, Benjamin Graf, and David Andrew Barry. Rapid changes in river plume dynamics caused by advected wind-driven coastal upwelling as observed in Lake Geneva. Limnology and Oceanography, 66(8):3116–3133, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Schwefel:2023:ODS

[SMCS23] Robert Schwefel, Sally MacIntyre, Alicia Cortés, and Steven Sadro. Oxygen depletion and sediment respiration in icecovered Arctic lakes. Limnology and Oceanography, 68(7): 1470–1489, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Strock:2021:TAA

[SMD21] Jacob P. Strock and Susanne Menden-Deuer. Temperature acclimation alters phytoplankton growth and production rates. Limnology and Oceanography, 66(3):740–752, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Shen:2020:HBM

[SMF20] Xinhui Shen, Marcos, and Henry C. Fu. How the bending mechanics of setae modulate hydrodynamic sensing in copepods. Limnology and Oceanography, 65(4):749–761, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Savoie:2022:ILR

[SMG^{+}22] Allison M. Savoie, Amy Moody, Melissa Gilbert, Kevin S. Dillon, Stephan D. Howden, Alan M. Shiller, and Christopher T. Hayes. Impact of local rivers on coastal acidification.

Limnology and Oceanography, 67(12):2779–2795, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Smith:2022:EIR

[SMKS22] Kathryn E. Smith, Pippa J. Moore, Nathan G. King, and Dan A. Smale. Examining the influence of regional-scale variability in temperature and light availability on the depth distribution of subtidal kelp forests. Limnology and Oceanography, 67(2):314–328, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Schauberger:2021:SVP

[SML^{+}21a] Clemens Schauberger, Mathias Middelboe, Morten Larsen, Logan M. Peoples, Douglas H. Bartlett, Finn Kirpekar, Ashley A. Rowden, Frank Wenzhöfer, Bo Thamdrup, and Ronnie N. Glud. Spatial variability of prokaryotic and viral abundances in the Kermadec and Atacama Trench regions. Limnology and Oceanography, 66(6):2095–2109, June 2021. CO-DEN LIOCAH. ISSN 0024-3590.

Shousha:2021:DFC

[SML21b] Stéphanie Shousha, Roxane Maranger, and Jean-François Lapierre. Different forms of carbon, nitrogen, and phosphorus influence ecosystem stoichiometry in a north temperate river across seasons and land uses. Limnology and Oceanography, 66(12):4285–4298, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Scully:2022:STV

[SMNT22] Malcolm E. Scully, Anna P. M. Michel, David P. Nicholson, and Shawnee Traylor. Spatial and temporal variations in atmospheric gas flux from the Hudson River: the estuarine gas exchange maximum. Limnology and Oceanography, 67(7): 1590–1603, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Seidov:2021:RWD

[SMP21] Dan Seidov, Alexey Mishonov, and Rost Parsons. Recent warming and decadal variability of Gulf of Maine and Slope Water. Limnology and Oceanography, 66(9):3472–3488, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Selden:2021:TRD

[SMW^{+}21] Corday R. Selden, Margaret R. Mulholland, Brittany Widner, Peter Bernhardt, and Amal Jayakumar. Toward resolving disparate accounts of the extent and magnitude of nitrogen fixation in the Eastern Tropical South Pacific oxygen deficient zone. Limnology and Oceanography, 66(5):1950–1960, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Schaefer:2022:WDS

[SN22] Rachel Beth Schaefer and Heidi Nepf. Wave damping by seagrass meadows in combined wave-current conditions. Limnology and Oceanography, 67(7):1554–1565, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Shchapov:2023:OBB

[SO23] Kirill Shchapov and Ted Ozersky. Opening the black box of winter: Full-year dynamics of crustacean zooplankton along a nearshore depth gradient in a large lake. Limnology and Oceanography, 68(7):1438–1451, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Schoenrock:2021:EBL

[SOO^{+}21] Kathryn M. Schoenrock, Rory O’Callaghan, Tony O’Callaghan, Aisha O’Connor, and Dagmar B. Stengel. An ecological baseline for Laminaria hyperborea forests in western Ireland. Limnology and Oceanography, 66(9):3439–3454, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Sebastian:2021:EGP

[SORGC^{+}21] Marta Sebastián, Eva Ortega-Retuerta, Laura Gómez-Consarnau, Marina Zamanillo, Marta Álvarez, Javier Arístegui, and Josep M. Gasol. Environmental gradients and physical barriers drive the basin-wide spatial structuring of Mediterranean Sea and adjacent eastern Atlantic Ocean prokaryotic communities. Limnology and Oceanography, 66 (12):4077–4095, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Satoh:2023:FRC

[SOSN23] Yuhi Satoh, Shigeyoshi Otosaka, Takashi Suzuki, and Takahiro Nakanishi. Factors regulating the concentration of particulate iodine in coastal seawater. Limnology and Oceanography, 68(7):1580–1594, July 2023. CODEN LI-OCAH. ISSN 0024-3590.

StPierre:2020:TED

[SOT^{+}20] Kyra A. St. Pierre, Allison A. Oliver, Suzanne E. Tank, Brian P. V. Hunt, Ian Giesbrecht, Colleen T. E. Kellogg, Jennifer M. Jackson, Ken P. Lertzman, William C. Floyd, and Maartje C. Korver. Terrestrial exports of dissolved and particulate organic carbon affect nearshore ecosystems of the Pacific coastal temperate rainforest. Limnology and Oceanography, 66(10):2657–2675, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Schmiedeskamp:2021:WLM

[SPBK21] Marcel Schmiedeskamp, Leandra Stephanie Emilia Praetzel, David Bastviken, and Klaus-Holger Knorr. Whole-lake methane emissions from two temperate shallow lakes with fluctuating water levels: Relevance of spatiotemporal patterns. Limnology and Oceanography, 66(6):2455–2469, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Seekell:2022:NDR

[SPHH22] David A. Seekell, Michael L. Pace, James B. Heffernan, and Sally J. Holbrook. Nonlinear dynamics, resilience, and regime shifts in aquatic communities and ecosystems: an overview. Limnology and Oceanography, 67(S1):S1–S4, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Smale:2022:EPE

[SPKM22] Dan A. Smale, Albert Pessarrodona, Nathan King, and Pippa J. Moore. Examining the production, export, and immediate fate of kelp detritus on open-coast subtidal reefs in the Northeast Atlantic. Limnology and Oceanography, 67 (S2):S36–S49, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Soria-Piriz:2020:WSD

[SPLJA^{+}20] Sara Soria-Píriz, Miguel Lara, Juan Luis Jiménez-Arias, Sokratis Papaspyrou, Bárbara Úbeda, Emilio García-Robledo, Julio Bohórquez, José Ángel Gálvez, Niels Peter Revsbech, and Alfonso Corzo. What supports the deep chlorophyll maximum in acidic lakes? The role of the bacterial CO_{2} production in the hypolimnion. Limnology and Oceanography, 65(6):1318–1335, June 2020. CODEN LI-OCAH. ISSN 0024-3590.

Siebers:2022:SFV

[SPR22] Andre R. Siebers, Amael Paillex, and Christopher T. Robinson. Seasonal and functional variation in the trophic base of intermittent alpine streams. Limnology and Oceanography, 67(5):1098–1110, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Sterner:2020:FAC

[SRL^{+}20] Robert W. Sterner, Kaitlin L. Reinl, Brenda Moraska Lafrancois, Sandra Brovold, and Todd R. Miller. A first assessment of cyanobacterial blooms in oligotrophic Lake Superior. Limnology and Oceanography, 66(6):2984–2998, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Stahl:2023:AAC

[SRM23] Angela R. Stahl, Tatiana A. Rynearson, and Kelton W. McMahon. Amino acid carbon isotope fingerprints are unique among eukaryotic microalgal taxonomic groups. Limnology and Oceanography, 68(6):1331–1345, June 2023. CODEN LI-OCAH. ISSN 0024-3590.

Schoutens:2021:STI

[SRN^{+}21] Ken Schoutens, Svenja Reents, Stefanie Nolte, Ben Evans, Maike Paul, Matthias Kudella, Tjeerd Bouma, Iris Möller, and Stijn Temmerman. Survival of the thickest? Impacts of extreme wave-forcing on marsh seedlings are mediated by species morphology. Limnology and Oceanography, 66(7): 2936–2951, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Shanks:2020:MHW

[SRV^{+}20] Alan L. Shanks, Leif K. Rasmuson, Jenna R. Valley, Marley A. Jarvis, Carly Salant, David A. Sutherland, Eleanor I. Lamont, MacKenna A. H. Hainey, and Richard B. Emlet. Marine heat waves, climate change, and failed spawning by coastal invertebrates. Limnology and Oceanography, 65(3): 627–636, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Suthers:2023:FEP

[SSA^{+}23] Iain M. Suthers, Amandine Schaeffer, Matthew Archer, Moninya Roughan, David A. Griffin, Christopher C. Chapman, Bernadette M. Sloyan, and Jason D. Everett. Frontal eddies provide an oceanographic triad for favorable larval fish habitat. Limnology and Oceanography, 68(5):1019–1036, May 2023. CODEN LIOCAH. ISSN 0024-3590.

Sun:2022:LAS

[SSB^{+}22] Xuerong Sun, Fang Shen, Robert J. W. Brewin, Mengyu Li, and Qing Zhu. Light absorption spectra of naturally mixed phytoplankton assemblages for retrieval of phytoplankton group composition in coastal oceans. Limnology and Oceanography, 67(4):946–961, April 2022. CODEN LI-OCAH. ISSN 0024-3590.

Stoltenberg:2020:SVC

[SSCE20] Laura Stoltenberg, Kai G. Schulz, Tyler Cyronak, and Bradley D. Eyre. Seasonal variability of calcium carbonate precipitation and dissolution in shallow coral reef sediments. Limnology and Oceanography, 65(4):876–891, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Simone:2021:OAM

[SSEO21] Michelle N. Simone, Kai G. Schulz, Bradley D. Eyre, and Joanne M. Oakes. Ocean acidification may mitigate negative effects of warming on carbon burial potential in subtidal unvegetated estuarine sediments. Limnology and Oceanography, 66(8):2953–2966, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Sato:2022:RDS

[SSH^{+}22] Mitsuhide Sato, Takuhei Shiozaki, Fuminori Hashihama, Taketoshi Kodama, Hiroshi Ogawa, Hiroaki Saito, Atsushi Tsuda, Shigenobu Takeda, and Ken Furuya. Relative depths of the subsurface peaks of phytoplankton abundance conserved over ocean provinces. Limnology and Oceanography, 67(11):2557–2571, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Szczepanek:2024:FSD

[SSK24] Marta Szczepanek, Marc J. Silberberger, and Monika Kedra. Functional structure and diversity of marine benthos reflect food availability and quality, as evidenced by isotopic diversity. Limnology and Oceanography, 69(4):757–771, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Schoepf:2022:CHT

[SSL22] Verena Schoepf, Hermione Sanderson, and Ellis Larcombe. Coral heat tolerance under variable temperatures: Effects of different variability regimes and past environmental history vs. current exposure. Limnology and Oceanography, 67(2): 404–418, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Saxton:2021:SRM

[SSM^{+}21a] Matthew A. Saxton, Vladimir A. Samarkin, Michael T. Madigan, Marshall W. Bowles, William Matthew Sattley, Charles A. Schutte, and Samantha B. Joye. Sulfate reduction and methanogenesis in the hypersaline deep waters and sediments of a perennially ice-covered lake. Limnology and Oceanography, 66(5):1804–1818, May 2021. CODEN LI-OCAH. ISSN 0024-3590.

Stamieszkin:2021:FPP

[SSM21b] Karen Stamieszkin, Deborah K. Steinberg, and Amy E. Maas. Fecal pellet production by mesozooplankton in the subarctic Northeast Pacific Ocean. Limnology and Oceanography, 66(7):2585–2597, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Soulie:2022:BRO

[SSM^{+}22a] Tanguy Soulié, Herwig Stibor, Sébastien Mas, Benjamin Braun, Johanna Knechtel, Jens C. Nejstgaard, Ulrich Sommer, Francesca Vidussi, and Behzad Mostajir. Brownification reduces oxygen gross primary production and community respiration and changes the phytoplankton community composition: an in situ mesocosm experiment with high-frequency sensor measurements in a North Atlantic bay. Limnology and Oceanography, 67(4):874–887, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Steinsdottir:2022:AAM

[SSM^{+}22b] Herdís G. R. Steinsdóttir, Clemens Schauberger, Snehit Mhatre, Bo Thamdrup, and Laura A. Bristow. Aerobic and anaerobic methane oxidation in a seasonally anoxic basin. Limnology and Oceanography, 67(6):1257–1273, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Schutte:2020:VSS

[SSP^{+}20] Charles A. Schutte, Vladimir A. Samarkin, Brian Peters, Michael T. Madigan, Marshall Bowles, Rachael Morgan-Kiss, Karen Casciotti, and Samantha Joye. Vertical stratification and stability of biogeochemical processes in the deep saline waters of Lake Vanda, Antarctica. Limnology and Oceanography, 65(3):569–581, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Stanley:2023:LSN

[SSR^{+}23] Brianna C. Stanley, Rachel E. Sipler, Quinn N. Roberts, Jenna L. Spackeen, E. Zane Norton, and Deborah A. Bronk. Late-season nitrogen uptake across the western coastal Alaskan Arctic. Limnology and Oceanography, 68(8): 1687–1703, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Sippo:2020:CCC

[SSS^{+}20] James Z. Sippo, Christian J. Sanders, Isaac R. Santos, Luke C. Jeffrey, Mitchell Call, Yota Harada, Kylie Maguire, Dylan Brown, Stephen R. Conrad, and Damien T. Maher. Coastal carbon cycle changes following mangrove loss. Limnology and Oceanography, 66(10):2642–2656, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Sebastian:2024:WAQ

[SSS^{+}24] Marta Sebastián, Pablo Sánchez, Guillem Salazar, Xosé A. Álvarez-Salgado, Isabel Reche, Xosé Anxelu G. Morán, Maria Montserrat Sala, Carlos M. Duarte, Silvia G. Acinas, and Josep M. Gasol. Water aging and the quality of organic carbon sources drive niche partitioning of the active bathypelagic prokaryotic microbiome. Limnology and Oceanography, 69(3):562–575, March 2024. CODEN LIOCAH. ISSN 0024-3590.

Steinsberger:2020:HOD

[SSWM20] Thomas Steinsberger, Robert Schwefel, Alfred Wüest, and Beat Müller. Hypolimnetic oxygen depletion rates in deep lakes: Effects of trophic state and organic matter accumulation. Limnology and Oceanography, 66(6):3128–3138, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Sutherland:2022:PTG

[ST22] Kelly R. Sutherland and Anne W. Thompson. Pelagic tunicate grazing on marine microbes revealed by integrative approaches. Limnology and Oceanography, 67(1):102–121, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Suzuki-Tellier:2022:MFD

[STAK22] Sei Suzuki-Tellier, Anders Andersen, and Thomas Kiørboe. Mechanisms and fluid dynamics of foraging in heterotrophic nanoflagellates. Limnology and Oceanography, 67(6):1287– 1298, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Schiffrine:2022:IET

[STB22] Nicolas Schiffrine, Jean-Éric Tremblay, and Marcel Babin. Interactive effects of temperature and nitrogen source on the elemental stoichiometry of a polar diatom. Limnology and Oceanography, 67(12):2750–2762, December 2022. CODEN LIOCAH. ISSN 0024-3590.

Schepers:2020:HPV

[SVBT20] Lennert Schepers, Alexander Van Braeckel, Tjeerd J. Bouma, and Stijn Temmerman. How progressive vegetation die-off in a tidal marsh would affect flow and sedimentation patterns: a field demonstration. Limnology and Oceanography, 65(2): 401–412, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Salk:2022:WCE

[SVC^{+}22] Kateri R. Salk, Jason J. Venkiteswaran, Raoul-Marie Couture, Scott N. Higgins, Michael J. Paterson, and Sherry L. Schiff. Warming combined with experimental eutrophication intensifies lake phytoplankton blooms. Limnology and Oceanography, 67(1):147–158, January 2022. CODEN LI-OCAH. ISSN 0024-3590.

Samper-Villarreal:2022:SOC

[SVMM^{+}22] Jimena Samper-Villarreal, Inés Mazarrasa, Pere Masqué, Oscar Serrano, and Jorge Cortés. Sediment organic carbon stocks were similar among four species compositions in a tropical seagrass meadow. Limnology and Oceanography, 67 (S2):S208–S225, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Stephens:2020:EZN

[SWB^{+}20] Brandon M. Stephens, Scott D. Wankel, J. Michael Beman, Ariel J. Rabines, Andrew E. Allen, and Lihini I. Aluwihare. Euphotic zone nitrification in the California Current ecosystem. Limnology and Oceanography, 65(4):790–806, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Shea:2023:SPH

[SWC^{+}23] Connor H. Shea, Paul K. Wojtal, Hilary G. Close, Amy E. Maas, Karen Stamieszkin, Joseph S. Cope, Deborah K. Steinberg, Natalie Wallsgrove, and Brian N. Popp. Small particles and heterotrophic protists support the mesopelagic zooplankton food web in the subarctic northeast Pacific Ocean. Limnology and Oceanography, 68(8):1949–1963, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Song:2023:HFV

[SWK^{+}23] Shuzhen Song, Zhaohui Aleck Wang, Kevin D. Kroeger, Meagan Eagle, Sophie N. Chu, and Jianzhong Ge. High-frequency variability of carbon dioxide fluxes in tidal water over a temperate salt marsh. Limnology and Oceanography, 68(9):2108– 2125, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Scofield:2020:DCM

[SWOR20] Anne E. Scofield, James M. Watkins, Eric Osantowski, and Lars G. Rudstam. Deep chlorophyll maxima across a trophic state gradient: a case study in the Laurentian Great Lakes. Limnology and Oceanography, 65(10):2460–2484, October 2020. CODEN LIOCAH. ISSN 0024-3590.

Strait:2022:NSI

[SWS^{+}22] Nicholas Strait, Taylor M. Williams, Alison R. Sherwood, Randall K. Kosaki, Louise Giuseffi, Celia M. Smith, and Heather L. Spalding. Nitrogen stable isotopes (δ ^{15}N) and tissue nitrogen in shallow-water and mesophotic macroalgae differ between the Main Hawaiian Islands and the Northwestern Hawaiian Islands. Limnology and Oceanography, 67(5): 1211–1226, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Sheng:2023:EDD

[SWZ^{+}23] Hua-Xia Sheng, Xianhui S. Wan, Bobo Zou, Youfang Sun, Buce Hanoch, Wenbin Zou, Weidi Yang, Siqi Wu, Hui Huang, Bess B. Ward, and Shuh-Ji Kao. An efficient diazotrophderived nitrogen transfer pathway in coral reef system. Limnology and Oceanography, 68(4):963–981, April 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Shen:2021:CMS

[SYMF21] Xinhui Shen, Xin Yao, Marcos, and Henry C. Fu. Can the mechanoreceptional setae of a feeding-current feeding copepod detect hydrodynamic disturbance induced by entrained free-floating prey? Limnology and Oceanography, 66(12): 4096–4111, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Shi:2021:ETI

[SYT^{+}21] Benwei Shi, Shi Lun Yang, Stijn Temmerman, Tjeerd Bouma, Tom Ysebaert, Sikai Wang, Yingxin Zhang, Jihua Wu, Haifei Yang, Longhui Zhang, Liqin Zuo, and Ya Ping Wang. Effect of typhoon-induced intertidal-flat erosion on dominant macrobenthic species (Meretrix meretrix). Limnology and Oceanography, 66(12):4197–4209, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Shogren:2021:ACD

[SZA^{+}21] Arial J. Shogren, Jay P. Zarnetske, Benjamin W. Abbott, Frances Iannucci, Alexander Medvedeff, Sam Cairns, Megan J. Duda, and William B. Bowden. Arctic concentration– discharge relationships for dissolved organic carbon and nitrate vary with landscape and season. Limnology and Oceanography, 66(S1):S197–S215, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Su:2020:MIS

[SZY^{+}20] Guangyi Su, Jakob Zopfi, Haoyi Yao, Lea Steinle, Helge Niemann, and Moritz F. Lehmann. Manganese/iron-supported sulfate-dependent anaerobic oxidation of methane by archaea in lake sediments. Limnology and Oceanography, 65(4):863– 875, April 2020. CODEN LIOCAH. ISSN 0024-3590.

To:2024:GSD

[TATC^{+}24] Sze-Wing To, Esteban Acevedo-Trejos, Subhendu Chakraborty, Francesco Pomati, and Agostino Merico. Grazing strategies determine the size composition of phytoplankton in eutrophic lakes. Limnology and Oceanography, 69(4):933–946, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Testa:2022:BSR

[TBH^{+}22] Jeremy M. Testa, Walter R. Boynton, Casey L. S. Hodgkins, Amanda L. Moore, Eva M. Bailey, and Johanna Rambo. Biogeochemical states, rates, and exchanges exhibit linear responses to large nutrient load reductions in a shallow, eutrophic urban estuary. Limnology and Oceanography, 67(4): 739–752, April 2022. CODEN LIOCAH. ISSN 0024-3590.

Traiger:2022:LBM

[TCP^{+}22] Sarah B. Traiger, Brian Cohn, Demetra Panos, Margaret Daly, Heidi K. Hirsh, Maria Martone, Isabella Gutierrez, David A. Mucciarone, Yuichiro Takeshita, Stephen G. Monismith, Robert B. Dunbar, and Kerry J. Nickols. Limited biogeochemical modification of surface waters by kelp forest canopies: Influence of kelp metabolism and site-specific hydrodynamics. Limnology and Oceanography, 67(2):392–403, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Traboni:2021:MUF

[TCS21] Claudia Traboni, Albert Calbet, and Enric Saiz. Mixotrophy upgrades food quality for marine calanoid copepods. Limnology and Oceanography, 66(12):4125–4139, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Tamborski:2021:PWE

[TEK^{+}21] Joseph J. Tamborski, Meagan Eagle, Barret L. Kurylyk, Kevin D. Kroeger, Zhaoihui Aleck Wang, Paul Henderson, and Matthew A. Charette. Pore water exchange-driven inorganic carbon export from intertidal salt marshes. Limnology and Oceanography, 66(5):1774–1792, May 2021. CODEN LI-OCAH. ISSN 0024-3590.

Tanaka:2021:PRT

[TGE^{+}21] Mamoru Tanaka, Amatzia Genin, Yoshinari Endo, Gregory N. Ivey, and Hidekatsu Yamazaki. The potential role of turbulence in modulating the migration of demersal zooplankton. Limnology and Oceanography, 66(3):855–864, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Treude:2021:RSC

[THL^{+}21] Tina Treude, Leila J. Hamdan, Sydnie Lemieux, Andrew W. Dale, and Stefan Sommer. Rapid sulfur cycling in sediments from the Peruvian oxygen minimum zone featuring simultaneous sulfate reduction and sulfide oxidation. Limnology and Oceanography, 66(7):2661–2671, July 2021. CODEN LI-OCAH. ISSN 0024-3590.

Trapp:2021:EPC

[THS21] Aubrey Trapp, Jan Heuschele, and Erik Selander. Eavesdropping on plankton — can zooplankton monitoring improve forecasting of biotoxins from harmful algae blooms? Limnology and Oceanography, 66(9):3455–3471, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Theroux:2020:SBA

[THT^{+}20] Susanna Theroux, Yongsong Huang, Jaime L. Toney, Robert Andersen, Paul Nyren, Rick Bohn, Jeffrey Salacup, Leslie Murphy, and Linda Amaral-Zettler. Successional blooms of alkenone-producing haptophytes in Lake George, North Dakota: Implications for continental paleoclimate reconstructions. Limnology and Oceanography, 65(2):413–425, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Thomsen:2023:MTS

[THVJ23] Esther Thomsen, Lucia S. Herbeck, Inés G. Viana, and Tim C. Jennerjahn. Meadow trophic status regulates the nitrogen filter function of tropical seagrasses in seasonally eutrophic coastal waters. Limnology and Oceanography, 68 (8):1906–1919, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Takahashi:2023:HIR

[TION23] Kazutaka Takahashi, Mutsuo Ichinomiya, Yuji Okazaki, and Yuichiro Nishibe. Higher ingestion rates and importance of ciliates in the diet of a large, subarctic copepod revealed by larger volume incubations. Limnology and Oceanography, 68 (4):790–802, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Thebault:2022:SSG

[TJW^{+}22] Julien Thébault, Aurélie Jolivet, Matthieu Waeles, Hélène Tabouret, Sophie Sabarot, Christophe Pécheyran, Aude Leynaert, Klaus Peter Jochum, Bernd R. Schöne, Lukas Fröhlich, Valentin Siebert, Erwan Amice, and Laurent Chauvaud. Scallop shells as geochemical archives of phytoplanktonrelated ecological processes in a temperate coastal ecosystem. Limnology and Oceanography, 67(1):187–202, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Thompson:2021:NPL

[TKA21] Anne W. Thompson, Kathleen Kouba, and Nathan A. Ahlgren. Niche partitioning of low-light adapted Prochlorococcus subecotypes across oceanographic gradients of the North Pacific subtropical front. Limnology and Oceanography, 65(3):1548–1562, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Tinta:2021:IJM

[TKH21] Tinkara Tinta, Katja Klun, and Gerhard J. Herndl. The importance of jellyfish–microbe interactions for biogeochemical cycles in the ocean. Limnology and Oceanography, 66(5): 2011–2032, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Thayne:2022:ALC

[TKM^{+}22] Michael W. Thayne, Benjamin M. Kraemer, Jorrit P. Mesman, Bastiaan W. Ibelings, and Rita Adrian. Antecedent lake conditions shape resistance and resilience of a shallow lake ecosystem following extreme wind storms. Limnology and Oceanography, 67(S1):S101–S120, February 2022. CO-DEN LIOCAH. ISSN 0024-3590.

Tsuchiya:2020:VBM

[TKS^{+}20] Kenji Tsuchiya, Kazuhiro Komatsu, Ryuichiro Shinohara, Akio Imai, Shin ichiro S. Matsuzaki, Ryuhei Ueno, Victor S. Kuwahara, and Ayato Kohzu. Variability of benthic methane-derived carbon along seasonal, biological, and sedimentary gradients in a polymictic lake. Limnology and Oceanography, 66(6):3017–3031, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Thompson:2023:CMC

[TKW^{+}23] Lauren M. Thompson, McKenzie A. Kuhn, Johanna C. Winder, Lucas P. P. Braga, Ryan H. S. Hutchins, Andrew J. Tanentzap, Vincent L. St. Louis, and David Olefeldt. Controls on methylmercury concentrations in lakes and streams of peatland-rich catchments along a 1700 km permafrost gradient. Limnology and Oceanography, 68(3):583–597, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Tomczak:2022:RSS

[TMKB^{+}22] Maciej T. Tomczak, Bärbel Müller-Karulis, Thorsten Blenckner, Eva Ehrnsten, Margit Eero, Bo Gustafsson, Alf Norkko, Saskia A. Otto, Karen Timmermann, and Christoph Humborg. Reference state, structure, regime shifts, and regulatory drivers in a coastal sea over the last century: The Central Baltic Sea case. Limnology and Oceanography, 67 (S1):S266–S284, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Tsunogai:2020:DSI

[TMM^{+}20] Urumu Tsunogai, Yuko Miyoshi, Toshiyuki Matsushita, Daisuke D. Komatsu, Masanori Ito, Chiho Sukigara, Fumiko Nakagawa, and Masahiro Maruo. Dual stable isotope characterization of excess methane in oxic waters of a mesotrophic lake. Limnology and Oceanography, 66(6):2937–2952, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Torstensson:2021:SIM

[TMS^{+}21] Anders Torstensson, Andrew R. Margolin, Gordon M. Showalter, Walker O. Smith, Jr., Elizabeth H. Shadwick, Shelly D. Carpenter, Francesco Bolinesi, and Jody W. Deming. Sea-ice microbial communities in the Central Arctic Ocean: Limited responses to short-term pCO_{2} perturbations. Limnology and Oceanography, 66(S1):S383–S400, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Turner:2022:PCC

[TMSL22] R. Eugene Turner, Charles S. Milan, Erick M. Swenson, and James M. Lee. Peak chlorophyll-a concentrations in the lower Mississippi River from 1997 to 2018. Limnology and Oceanography, 67(3):703–712, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Tanaka:2022:SFD

[TNK^{+}22] Kiyoshi Tanaka, Seiya Nagao, Yujiro Kitade, Masato Niki, Takaaki Katsumata, Toru Miyama, and Hiroshi Yoshinari. Spread of Fukushima-derived radiocesium over the coastal ocean in response to typhoon-induced flooding in September 2011. Limnology and Oceanography, 67(5):1184–1193, May 2022. CODEN LIOCAH. ISSN 0024-3590.

Thoral:2023:SLQ

[TPTS23] François Thoral, Matthew H. Pinkerton, Leigh W. Tait, and David R. Schiel. Spectral light quality on the seabed matters for macroalgal community composition at the extremities of light limitation. Limnology and Oceanography, 68(4):902– 916, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Tolar:2020:TSA

[TRS^{+}20] Bradley B. Tolar, Linta Reji, Jason M. Smith, Marguerite Blum, J. Timothy Pennington, Francisco P. Chavez, and Christopher A. Francis. Time series assessment of Thaumarchaeota ecotypes in Monterey Bay reveals the importance of water column position in predicting distribution– environment relationships. Limnology and Oceanography, 65 (9):2041–2055, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Tang:2023:RZI

[TSX^{+}23] Yali Tang, Ling Su, Ruohua Xu, Sirui Wang, Yaling Su, Zhengwen Liu, Jinlei Yu, Henri J. Dumont, and Erik Jeppesen. Response of zooplankton to inputs of terrestrial dissolved organic matter: Food quality constraints induced by microbes. Limnology and Oceanography, 68(3):709–722, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Tang:2022:NOP

[TTC^{+}22] Weiyi Tang, John C. Tracey, Julia Carroll, Elizabeth Wallace, Jenna A. Lee, Levy Nathan, Xin Sun, Amal Jayakumar, and Bess B. Ward. Nitrous oxide production in the Chesapeake Bay. Limnology and Oceanography, 67(9):2101–2116, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Tsui:2020:CIC

[TUR^{+}20] Martin Tsz-Ki Tsui, Habibullah Uzun, Alexander Ruecker, Hamed Majidzadeh, Yener Ulus, Hongyuan Zhang, Shaowu Bao, Joel D. Blum, Tanju Karanfil, and Alex T. Chow. Concentration and isotopic composition of mercury in a blackwater river affected by extreme flooding events. Limnology and Oceanography, 66(5):2158–2169, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Thingstad:2021:RVC

[TVB^{+}21] T. Frede Thingstad, Selina Våge, Gunnar Bratbak, Jorun Egge, Aud Larsen, Jens Christian Nejstgaard, and Ruth-Anne Sandaa. Reproducing the virus-to-copepod link in Arctic mesocosms using host fitness optimization. Limnology and Oceanography, 66(S1):S303–S313, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Tischer:2022:ISB

[TZF^{+}22] Jana Tischer, Jakob Zopfi, Claudia Frey, Paul M. Magyar, Andreas Brand, Kirsten Oswald, Corinne Jegge, Caitlin H. Frame, María R. Miracle, Xavier Sòria-Perpinyà, Eduardo Vicente, and Moritz F. Lehmann. Isotopic signatures of biotic and abiotic N_{2}O production and consumption in the water column of meromictic, ferruginous Lake La Cruz (Spain). Limnology and Oceanography, 67(8):1760–1775, August 2022. CODEN LIOCAH. ISSN 0024-3590.

VanMeter:2020:BAE

[VCBB20] Kim J. Van Meter, Shadman Chowdhury, Danyka K. Byrnes, and Nandita B. Basu. Biogeochemical asynchrony: Ecosystem drivers of seasonal concentration regimes across the Great Lakes Basin. Limnology and Oceanography, 65(4):848– 862, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Vane:2023:TBR

[VCT^{+}23] Kim Vane, Matthew R. D. Cobain, Clive N. Trueman, Tobias R. Vonnahme, Sebastian Rokitta, Nicholas V. C. Polunin, and Hauke Flores. Tracing basal resource use across sea-ice, pelagic, and benthic habitats in the early Arctic spring food web with essential amino acid carbon isotopes. Limnology and Oceanography, 68(4):862–877, April 2023. CODEN LIOCAH. ISSN 0024-3590.

vandePoll:2020:TSD

[vdPAV^{+}20] Willem H. van de Poll, Edwin Abdullah, Ronald J. W. Visser, Philipp Fischer, and Anita G. J. Buma. Taxon-specific dark survival of diatoms and flagellates affects Arctic phytoplankton composition during the polar night and early spring. Limnology and Oceanography, 65(5):903–914, May 2020. CODEN LIOCAH. ISSN 0024-3590.

vandePoll:2021:SRS

[vdPMF^{+}21] Willem H. van de Poll, Douwe S. Maat, Philipp Fischer, Ronald J. W. Visser, Corina P. D. Brussaard, and Anita G. J. Buma. Solar radiation and solar radiation driven cycles in warming and freshwater discharge control seasonal and inter-annual phytoplankton chlorophyll-a and taxonomic composition in a high Arctic fjord (Kongsfjorden, Spitsbergen). Limnology and Oceanography, 66(4):1221–1236, April 2021. CODEN LIOCAH. ISSN 0024-3590.

vanErk:2020:KDC

[vEMF^{+}20] Marit R. van Erk, Dimitri V. Meier, Timothy Ferdelman, Jens Harder, Ingeborg Bussmann, and Dirk de Beer. Kelp deposition changes mineralization pathways and microbial communities in a sandy beach. Limnology and Oceanography, 66(6):3066–3084, December 2020. CODEN LIOCAH. ISSN 0024-3590.

vanGrinsven:2021:NPT

[vGDH^{+}21] Sigrid van Grinsven, Jaap S. Sinninghe Damsté, John Harrison, Lubos Polerecky, and Laura Villanueva. Nitrate promotes the transfer of methane-derived carbon from the methanotroph Methylobacter sp. to the methylotroph Methylotenera sp. in eutrophic lake water. Limnology and Oceanography, 66(3):878–891, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Viana:2023:MTL

[VGSB23] Inés G. Viana, Rita García-Seoane, and Antonio Bode. A missing trophic link: Contribution of the microbial loop to the estimation of the trophic position of pelagic consumers. Limnology and Oceanography, 68(11):2587–2602, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Vanvelk:2021:IDP

[VGvdB^{+}21] Héléne Vanvelk, Lynn Govaert, Edwin M. van den Berg, Kristien I. Brans, and Luc De Meester. Interspecific differences, plastic, and evolutionary responses to a heat wave in three co-occurring Daphnia species. Limnology and Oceanography, 66(4):1201–1220, April 2021. CODEN LIOCAH. ISSN 0024-3590.

vanHespen:2021:ACS

[vHHP^{+}21] Rosanna van Hespen, Zhan Hu, Yisheng Peng, Bas W. Borsje, Maarten Kleinhans, Tom Ysebaert, and Tjeerd J. Bouma. Analysis of coastal storm damage resistance in successional mangrove species. Limnology and Oceanography, 66 (8):3221–3236, August 2021. CODEN LIOCAH. ISSN 0024-3590.

vanHespen:2022:ITB

[vHHP^{+}22] Rosanna van Hespen, Zhan Hu, Yisheng Peng, Zhenchang Zhu, Tom Ysebaert, and Tjeerd J. Bouma. Identifying traitbased tolerance to sediment dynamics during seedling establishment across eight mangrove species. Limnology and Oceanography, 67(10):2281–2295, October 2022. CODEN LI-OCAH. ISSN 0024-3590.

vanHelmond:2020:RMM

[vHLV^{+}20] Niels A. G. M. van Helmond, Bryan C. Lougheed, Annika Vollebregt, Francien Peterse, Guillaume Fontorbe, Daniel J. Conley, and Caroline P. Slomp. Recovery from multi-millennial natural coastal hypoxia in the Stockholm Archipelago, Baltic Sea, terminated by modern human activity. Limnology and Oceanography, 66(6):3085–3097, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Vizza:2022:PMD

[VJH^{+}22] Carmella Vizza, Stuart E. Jones, Julia A. Hart, William E. West, and Gary A. Lamberti. Pond methane dynamics, from microbial communities to ecosystem budget, during summer in Alaska. Limnology and Oceanography, 67(2):450–467, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Villa:2020:PMM

[VJS^{+}20] Jorge A. Villa, Yang Ju, Taylor Stephen, Camilo Rey-Sanchez, Kelly C. Wrighton, and Gil Bohrer. Plant-mediated methane transport in emergent and floating-leaved species of a temperate freshwater mineral-soil wetland. Limnology and Oceanography, 65(7):1635–1650, July 2020. CODEN LI-OCAH. ISSN 0024-3590.

Velthuis:2022:DEE

[VKB^{+}22] Mandy Velthuis, Joost. A. Keuskamp, Elisabeth S. Bakker, Maarten Boersma, Ulrich Sommer, Ellen van Donk, and Dedmer B. Van de Waal. Differential effects of elevated pCO_{2} and warming on marine phytoplankton stoichiometry. Limnology and Oceanography, 67(3):598–607, March 2022. CODEN LI-OCAH. ISSN 0024-3590.

Villalba:2022:FHC

[VKGW22] Luis Alberto Villalba, Rajat Karnatak, Hans-Peter Grossart, and Sabine Wollrab. Flexible habitat choice of pelagic bacteria increases system stability and energy flow through the microbial loop. Limnology and Oceanography, 67(6):1402– 1415, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Vesterinen:2021:NQL

[VKK^{+}21] Jussi Vesterinen, Ossi Keva, Kimmo K. Kahilainen, Ursula Strandberg, Minna Hiltunen, Paula Kankaala, and Sami J. Taipale. Nutritional quality of littoral macroinvertebrates and pelagic zooplankton in subarctic lakes. Limnology and Oceanography, 66(S1):S81–S97, February 2021. CODEN LI-OCAH. ISSN 0024-3590.

VanDam:2021:WTC

[VLP^{+}21] Bryce R. Van Dam, Christian C. Lopes, Pierre Polsenaere, René M. Price, Anna Rutgersson, and James W. Fourqurean. Water temperature control on CO_{2} flux and evaporation over a subtropical seagrass meadow revealed by atmospheric eddy covariance. Limnology and Oceanography, 66(2):510– 527, February 2021. CODEN LIOCAH. ISSN 0024-3590.

vanLeeuwe:2020:APP

[vLWV^{+}20] Maria A. van Leeuwe, Alison L. Webb, Hugh J. Venables, Ronald J. W. Visser, Mike P. Meredith, J. Theo M. Elzenga, and Jacqueline Stefels. Annual patterns in phytoplankton phenology in Antarctic coastal waters explained by environmental drivers. Limnology and Oceanography, 65(7):1651– 1668, July 2020. CODEN LIOCAH. ISSN 0024-3590.

Vilgrain:2021:TBA

[VMP^{+}21] Laure Vilgrain, Frédéric Maps, Marc Picheral, Marcel Babin, Cyril Aubry, Jean-Olivier Irisson, and Sakina-Dorothée Ayata. Trait-based approach using in situ copepod images reveals contrasting ecological patterns across an Arctic ice melt zone. Limnology and Oceanography, 66(4):1155–1167, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Vermaat:2020:NRL

[VMT^{+}20] Jan E. Vermaat, Andreas Matzinger, Sonja Trajanovska, Marina Talevska, and Susanne C. Schneider. Nutrient retention by the littoral vegetation of a large lake: Can Lake Ohrid cope with current and future loading? Limnology and Oceanography, 65(10):2390–2402, October 2020. CODEN LI-OCAH. ISSN 0024-3590.

Vanderwall:2024:MGI

[VMT^{+}24] Joseph W. Vanderwall, Clint C. Muhlfeld, Tyler H. Tappenbeck, Joseph Giersch, Ze Ren, and James J. Elser. Mountain glaciers influence biogeochemical and ecological characteristics of high-elevation lakes across the northern Rocky Mountains, USA. Limnology and Oceanography, 69(1):37–52, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Verhaegen:2023:DBD

[VSB^{+}23] Gerlien Verhaegen, Mehul Naresh Sangekar, Bastian Bentlage, Henk-Jan Hoving, Allen G. Collins, and Dhugal Lindsay. Drivers behind the diversity and distribution of a widespread midwater narcomedusa. Limnology and Oceanography, 68(9):2088–2107, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Vesterinen:2022:PKD

[VST^{+}22] Jussi Vesterinen, Ursula Strandberg, Sami J. Taipale, Martin J. Kainz, and Paula Kankaala. Periphyton as a key diet source of essential fatty acids for macroinvertebrates across a nutrient and dissolved organic carbon gradient in boreal lakes. Limnology and Oceanography, 67(7):1604–1616, July 2022. CODEN LIOCAH. ISSN 0024-3590.

Vogt-Vincent:2023:HFV

[VVMJ23] Noam S. Vogt-Vincent, Satoshi Mitarai, and Helen L. Johnson. High-frequency variability dominates potential connectivity between remote coral reefs. Limnology and Oceanography, 68(12):2733–2748, December 2023. CODEN LIOCAH. ISSN 0024-3590.

Volponi:2023:NFF

[VWR^{+}23] Sabrina N. Volponi, Heather L. Wander, David C. Richardson, Clayton J. Williams, Denise A. Bruesewitz, Shelley Arnott, Jennifer A. Brentrup, Hailee L. Edwards, Holly A. Ewing, Kristen Holeck, Lauren Johnson, Brian S. Kim, Ana M. Morales-Williams, Nisha Nadkarni, Beth C. Norman, Lianne Parmalee, Amy Shultis, Adrienne Tracy, Nicole K. Ward, Kathleen C. Weathers, Courtney R. Wigdahl-Perry, and Kiyoko Yokota. Nutrient function over form: Organic and inorganic nitrogen additions have similar effects on lake phytoplankton nutrient limitation. Limnology and Oceanography, 68(2):307–321, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Wong:2020:BAN

[WAPC20] Wei Wen Wong, Andrew Applegate, Seng Chee Poh, and Perran L. M. Cook. Biogeochemical attenuation of nitrate in a sandy subterranean estuary: Insights from two stable isotope approaches. Limnology and Oceanography, 66(6):3098–3113, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Wahl:2021:PPF

[WBB^{+}21] Martin Wahl, Francisco R. Barboza, Björn Buchholz, Sergey Dobretsov, Tamar Guy-Haim, Gil Rilov, Renate Schuett, Fabian Wolf, Jahangir Vajedsamiei, Maryam Yazdanpanah, and Christian Pansch. Pulsed pressure: Fluctuating impacts of multifactorial environmental change on a temperate macroalgal community. Limnology and Oceanography, 66 (12):4210–4226, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Wear:2020:BMP

[WCC20] Emma K. Wear, Craig A. Carlson, and Matthew J. Church. Bacterioplankton metabolism of phytoplankton lysates across a cyclone-anticyclone eddy dipole impacts the cycling of semi-labile organic matter in the photic zone. Limnology and Oceanography, 65(7):1608–1622, July 2020. CO-DEN LIOCAH. ISSN 0024-3590.

Williams:2021:ACC

[WCH^{+}21] Branwen Williams, Phoebe T. W. Chan, Jochen Halfar, Kathryn Hargan, and Walter Adey. Arctic crustose coralline alga resilient to recent environmental change. Limnology and Oceanography, 66(S1):S246–S258, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Wang:2020:DSB

[WCZ^{+}20] Yujing Wang, Xinyi Cao, Jin Zeng, Huabing Li, Dayong Zhao, and Qinglong L. Wu. Distinct shifts in bacterioplankton community composition and functional gene structure between macrophyte- and phytoplankton-dominated regimes in a large shallow lake. Limnology and Oceanography, 66(6): S208–S219, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Wu:2021:HFT

[WDG^{+}21] Yingxu Wu, Minhan Dai, Xianghui Guo, Jinshun Chen, Yi Xu, Xu Dong, Junwei Dai, and Zhirong Zhang. Highfrequency time-series autonomous observations of sea surface pCO_{2} and pH. Limnology and Oceanography, 66(3):588–606, March 2021. CODEN LIOCAH. ISSN 0024-3590.

Wojtal:2023:DMP

[WDS^{+}23] Paul K. Wojtal, Shannon C. Doherty, Connor H. Shea, Brian N. Popp, Claudia R. Benitez-Nelson, Ken O. Buesseler, Margaret L. Estapa, Montserrat Roca-Martí, and Hilary G. Close. Deconvolving mechanisms of particle flux attenuation using nitrogen isotope analyses of amino acids. Limnology and Oceanography, 68(9):1965–1981, September 2023. CO-DEN LIOCAH. ISSN 0024-3590.

Weisse:2024:PMP

[Wei24] Thomas Weisse. Physiological mortality of planktonic ciliates: Estimates, causes, and consequences. Limnology and Oceanography, 69(3):524–532, March 2024. CODEN LI-OCAH. ISSN 0024-3590.

Wang:2022:BAD

[WF22] Bin Wang and Katja Fennel. Biogeochemical-Argo data suggest significant contributions of small particles to the vertical carbon flux in the subpolar North Atlantic. Limnology and Oceanography, 67(11):2405–2417, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Waahlstrom:2020:CCC

[WHAR^{+}20] Iréne Wåhlström, Anders Höglund, Elin Almroth-Rosell, Brian R. MacKenzie, Matthias Gröger, Kari Eilola, Maris Plikshs, and Helén C. Andersson. Combined climate change and nutrient load impacts on future habitats and eutrophication indicators in a eutrophic coastal sea. Limnology and Oceanography, 66(5):2170–2187, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Wang:2021:HDI

[WHZ^{+}21] Sikai Wang, Qiang He, Youzheng Zhang, Qiang Sheng, Bo Li, and Jihua Wu. Habitat-dependent impacts of exotic plant invasions on benthic food webs in a coastal wetland. Limnology and Oceanography, 66(4):1256–1267, April 2021. CODEN LI-OCAH. ISSN 0024-3590.

Wang:2021:TLP

[WK21] Binbin Wang and Clifford E. Kraft. Thiamine limitation of periphyton in Adirondack Mountain streams. Limnology and Oceanography, 66(8):2988–2998, August 2021. CODEN LI-OCAH. ISSN 0024-3590.

Woodstock:2023:CMV

[WKRL^{+}23] Matthew S. Woodstock, Jeremy J. Kiszka, M. Rafael Ramírez-León, Tracey T. Sutton, Katja Fennel, Bin Wang, and Yuying Zhang. Cetacean-mediated vertical nitrogen transport in the oceanic realm. Limnology and Oceanography, 68(11):2445–2460, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Wong:2021:NOI

[WLK^{+}21] Wei Wen Wong, Moritz F. Lehmann, Thomas Kuhn, Caitlin Frame, Seng Chee Poh, Ian Cartwright, and Perran L. M. Cook. Nitrogen and oxygen isotopomeric constraints on the sources of nitrous oxide and the role of submarine groundwater discharge in a temperate eutrophic salt-wedge estuary. Limnology and Oceanography, 66(4):1068–1082, April 2021. CODEN LIOCAH. ISSN 0024-3590.

White:2021:SEK

[WLSD21] Lydia White, Stéphane Loisel, Laure Sevin, and Dominique Davoult. In situ estimates of kelp forest productivity in macro-tidal environments. Limnology and Oceanography, 66 (12):4227–4239, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Woosley:2020:FWA

[WM20] Ryan J. Woosley and Frank J. Millero. Freshening of the western Arctic negates anthropogenic carbon uptake potential. Limnology and Oceanography, 65(8):1834–1846, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Wurgaft:2022:WIF

[WMRL22] Eyal Wurgaft, Israela Musan, Tanya Rivlin, and Boaz Luz. Weak isotopic fractionation of dissolved O_{2} during community respiration. Limnology and Oceanography, 67(8):1794– 1804, August 2022. CODEN LIOCAH. ISSN 0024-3590.

Ward:2022:TME

[WMTJW22] Delphi F. L. Ward, Jessica Melbourne-Thomas, Craig R. Johnson, and Simon J. Wotherspoon. Trophic mediation and ecosystem stability: an assessment using qualitative network models. Limnology and Oceanography, 67(S1):S146–S162, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Wang:2021:LDG

[WNHY21] Jinhui Wang, Yijun Ni, Wei Hu, and Mingbo Yin. Lineage diversity and gene introgression in freshwater cladoceran crustaceans of the Chydorus sphaericus species complex. Limnology and Oceanography, 66(1):95–107, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Whitmore:2021:EZM

[WO21] Benjamin M. Whitmore and Mark D. Ohman. Zooglidermeasured association of zooplankton with the fine-scale vertical prey field. Limnology and Oceanography, 66(10):3811– 3827, October 2021. CODEN LIOCAH. ISSN 0024-3590.

Wong:2021:NIB

[WOK^{+}21] Kuo Hong Wong, Hajime Obata, Taejin Kim, Yoshiko Kondo, and Jun Nishioka. New insights into the biogeochemical cycling of copper in the subarctic Pacific: Distributions, size fractionation, and organic complexation. Limnology and Oceanography, 65(3):1424–1439, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Wing:2020:TMA

[WOMWR20] Stephen R. Wing, Sorrel A. O’Connell-Milne, Lucy C. Wing, and Malcolm R. Reid. Trace metals in Antarctic clam shells record the chemical dynamics of changing sea ice conditions. Limnology and Oceanography, 65(3):504–514, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Wagner:2023:DMC

[WOT^{+}23] Nicole D. Wagner, Felicia S. Osburn, Raegyn B. Taylor, Jeffrey A. Back, C. Kevin Chambliss, Bryan W. Brooks, and J. Thad Scott. Diazotrophy modulates cyanobacteria stoichiometry through functional traits that determine bloom magnitude and toxin production. Limnology and Oceanography, 68(2):348–360, February 2023. CODEN LIOCAH. ISSN 0024-3590.

Wu:2022:DCH

[WPO^{+}22] Fengrun Wu, Steven C. Pennings, Collin Ortals, Jennifer Ruiz, W. Reilly Farrell, Samuel M. McNichol, Christine Angelini, Amanda C. Spivak, Merryl Alber, and Chunfu Tong. Disturbance is complicated: Headward-eroding saltmarsh creeks produce multiple responses and recovery trajectories. Limnology and Oceanography, 67(S1):S86–S100, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Wang:2020:PPF

[WPY^{+}20] Yongming Wang, Jie Pan, Jun Yang, Zhichao Zhou, Yueping Pan, and Meng Li. Patterns and processes of free-living and particle-associated bacterioplankton and archaeaplankton communities in a subtropical river-bay system in South China. Limnology and Oceanography, 65(S1):S161–S179, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Wong:2023:DPR

[WRA^{+}23] Jane C. Y. Wong, John A. Raven, Montserrat Aldunate, Sebastián Silva, Juan Diego Gaitán-Espitia, Cristian A. Vargas, Osvaldo Ulloa, and Peter von Dassow. Do phytoplankton require oxygen to survive? A hypothesis and model synthesis from oxygen minimum zones. Limnology and Oceanography, 68(7):1417–1437, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Walters:2021:FWS

[WRC^{+}21] Andrea Walters, Marianne Robert, Pierre Cresson, Hervé Le Bris, and Dorothée Kopp. Food web structure in relation to environmental drivers across a continental shelf ecosystem. Limnology and Oceanography, 65(6):2563–2582, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Wolf:2022:PIA

[WRHR22] Klara K. E. Wolf, Sebastian D. Rokitta, Clara J. M. Hoppe, and Björn Rost. Pelagic and ice-associated microalgae under elevated light and pCO_{2}: Contrasting physiological strategies in two Arctic diatoms. Limnology and Oceanography, 67(9): 1895–1910, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Wasserman:2022:AED

[WRMP22] Ben A. Wasserman, Tanya L. Rogers, Stephan B. Munch, and Eric P. Palkovacs. Applying empirical dynamic modeling to distinguish abiotic and biotic drivers of population fluctuations in sympatric fishes. Limnology and Oceanography, 67(S1):S403–S415, February 2022. CODEN LIOCAH. ISSN 0024-3590.

White:2022:SIN

[WRS^{+}22] Margot E. White, Patrick A. Rafter, Brandon M. Stephens, Matthew R. Mazloff, Scott D. Wankel, and Lihini I. Aluwihare. Stable isotopes of nitrate record effects of the 2015–2016 El Niño and diatom iron limitation on nitrogen cycling in the eastern North Pacific Ocean. Limnology and Oceanography, 67(10):2140–2156, October 2022. CODEN LIOCAH. ISSN 0024-3590.

Wall:2019:SVB

[WRWPG19] Christopher B. Wall, Raphael Ritson-Williams, Brian N. Popp, and Ruth D. Gates. Spatial variation in the biochemical and isotopic composition of corals during bleaching and recovery. Limnology and Oceanography, 66(8):2011–2028, September 2019. CODEN LIOCAH. ISSN 0024-3590. See corrigendum [Ano21c].

Wang:2020:AFA

[WSB^{+}20] Linlin Wang, Junting Song, Hongsheng Bi, Matthew Gray, Chunlei Fan, Ming Liu, and Xian-Zhong Mao. Adaptive feeding in the American oyster Crassostrea virginica: Complex impacts of pulsatile flow during pseudofecal ejection events. Limnology and Oceanography, 65(9):2010–2023, September 2020. CODEN LIOCAH. ISSN 0024-3590.

Wei:2022:SSA

[WSLC22] Yuqiu Wei, Jun Sun, Liuyang Li, and Zhengguo Cui. Synechococcus silicon accumulation in oligotrophic oceans. Limnology and Oceanography, 67(3):552–566, March 2022. CO-DEN LIOCAH. ISSN 0024-3590.

Woodstock:2022:CTB

[WSZ22a] Matthew S. Woodstock, Tracey T. Sutton, and Yuying Zhang. Corrigendum: A trait-based carbon export model for mesopelagic fishes in the Gulf of Mexico with consideration of asynchronous vertical migration, flux boundaries, and feeding guilds. Limnology and Oceanography, 67(9):2117, September 2022. CODEN LIOCAH. ISSN 0024-3590. See [WSZ22b].

Woodstock:2022:TBC

[WSZ22b] Matthew S. Woodstock, Tracey T. Sutton, and Yuying Zhang. A trait-based carbon export model for mesopelagic fishes in the Gulf of Mexico with consideration of asynchronous vertical migration, flux boundaries, and feeding guilds. Limnology and Oceanography, 67(7):1443–1455, July 2022. CODEN LIOCAH. ISSN 0024-3590. See corrigenda [WSZ22a].

Wiesebron:2022:CSC

[WTvD^{+}22] Lauren Wiesebron, Lilian Teeuw, Jeroen van Dalen, Lennart van Ijzerloo, Karin Troost, Brenda Walles, Tom Ysebaert, and Tjeerd Bouma. Contrasting strategies to cope with storm-induced erosion events: a flume study comparing a native vs. introduced bivalve. Limnology and Oceanography, 67(11):2572–2585, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Wahl:2020:SAS

[WWB^{+}20] Martin Wahl, Franziska Julie Werner, Björn Buchholz, Stefanie Raddatz, Angelika Graiff, Birte Matthiessen, Ulf Karsten, Claas Hiebenthal, Jorin Hamer, Maysa Ito, Elisa Gülzow, Gil Rilov, and Tamar Guy-Haim. Season affects strength and direction of the interactive impacts of ocean warming and biotic stress in a coastal seaweed ecosystem. Limnology and Oceanography, 65(4):807–827, April 2020. CODEN LIOCAH. ISSN 0024-3590.

Wang:2021:DMP

[WWFS21] Jingyu Wang, Nicole D. Wagner, James M. Fulton, and J. Thad Scott. Diazotrophs modulate phycobiliproteins and nitrogen stoichiometry differently than other cyanobacteria in response to light and nitrogen availability. Limnology and Oceanography, 66(6):2333–2345, June 2021. CODEN LI-OCAH. ISSN 0024-3590.

Wall:2021:AAA

[WWGP21] Christopher B. Wall, Natalie J. Wallsgrove, Ruth D. Gates, and Brian N. Popp. Amino acid δ ^{13}C and δ ^{15}N analyses reveal distinct species-specific patterns of trophic plasticity in a marine symbiosis. Limnology and Oceanography, 66(5): 2033–2050, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Wong:2022:VSE

[WXK^{+}22] Kuo Hong Wong, Jiarui Xu, Yoshiko Kondo, Shigenobu Takeda, Asami S. Mashio, Hiroshi Hasegawa, and Hajime Obata. Very strong but exchangeable organic ligand of cobalt in the marginal sea. Limnology and Oceanography, 67(6): 1299–1312, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Wang:2020:STV

[WZCK20] Hualong Wang, Chuanlun Zhang, Feng Chen, and Jinjun Kan. Spatial and temporal variations of bacterioplankton in the Chesapeake Bay: a re-examination with high-throughput sequencing analysis. Limnology and Oceanography, 66(6): 3032–3045, December 2020. CODEN LIOCAH. ISSN 0024-3590.

Xiao:2022:ETD

[XDQ^{+}22] Qitao Xiao, Hongtao Duan, Boqiang Qin, Zhenghua Hu, Mi Zhang, Tianci Qi, and Xuhui Lee. Eutrophication and temperature drive large variability in carbon dioxide from China’s Lake Taihu. Limnology and Oceanography, 67(2): 379–391, February 2022. CODEN LIOCAH. ISSN 0024-3590.

Xia:2022:TAB

[XKP^{+}22] Jun Xia, Sohiko Kameyama, Florian Prodinger, Takashi Yoshida, Kyoung-Ho Cho, Jinyoung Jung, Sung-Ho Kang, Eun-Jin Yang, Hiroyuki Ogata, and Hisashi Endo. Tight association between microbial eukaryote and giant virus communities in the Arctic Ocean. Limnology and Oceanography, 67(6):1343–1356, June 2022. CODEN LIOCAH. ISSN 0024-3590.

Xu:2021:PFN

[XLN21] Zhaomeng Xu, Sumei Liu, and Xiaoyan Ning. Potential foraminiferal nitrate transport in sediments in contact with oxic overlying water. Limnology and Oceanography, 65(3): 1510–1530, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Xu:2021:CEN

[XMP^{+}21] Hai Xu, Mark J. McCarthy, Hans W. Paerl, Justin D. Brookes, Guangwei Zhu, Nathan S. Hall, Boqiang Qin, Yunlin Zhang, Mengyuan Zhu, Justyna J. Hampel, Silvia E. Newell, and Wayne S. Gardner. Contributions of external nutrient loading and internal cycling to cyanobacterial bloom dynamics in Lake Taihu, China: Implications for nutrient management. Limnology and Oceanography, 65(3): 1492–1509, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Xu:2021:DFD

[XSZ^{+}21] Huo Xu, Zhiyuan Shi, Xiaodong Zhang, Mengwen Pang, Ke Pan, and Hongbin Liu. Diatom frustules with different silica contents affect copepod grazing due to differences in the nanoscale mechanical properties. Limnology and Oceanography, 66(9):3408–3420, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Xing:2021:CMC

[XTJW21] Peng Xing, Ye Tao, Erik Jeppesen, and Qinglong L. Wu. Comparing microbial composition and diversity in freshwater lakes between Greenland and the Tibetan Plateau. Limnology and Oceanography, 66(S1):S142–S156, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Xing:2020:SMD

[XTL^{+}20] Peng Xing, Ye Tao, Jianhua Luo, Lina Wang, Biao Li, Huabing Li, and Qinglong L. Wu. Stratification of microbiomes during the holomictic period of Lake Fuxian, an alpine monomictic lake. Limnology and Oceanography, 65 (S1):S134–S148, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Xiao:2021:LCR

[XWL^{+}21] Kai Xiao, Alicia M. Wilson, Hailong Li, Isaac R. Santos, Joseph Tamborski, Erik Smith, Susan Q. Lang, Chunmiao Zheng, Xin Luo, Meiqing Lu, and Rogger E. Correa. Large CO_{2} release and tidal flushing in salt marsh crab burrows reduce the potential for blue carbon sequestration. Limnology and Oceanography, 66(1):14–29, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Xu:2023:DRT

[XYL^{+}23] Wei Xu, Change Yang, Yu Luo, Kai Zhang, Mingliang Chen, Shuangcheng Jiang, Hans-Peter Grossart, and Zhuhua Luo. Distinct response of total and active fungal communities and functions to seasonal changes in a semi-enclosed bay with mariculture (Dongshan Bay, Southern China). Limnology and Oceanography, 68(5):1048–1063, May 2023. CODEN LI-OCAH. ISSN 0024-3590.

Xu:2021:OCD

[XYZ^{+}21] Feng Xu, Shi-Bo Yan, Hong-Hai Zhang, Ying-Cui Wu, Qian-Yao Ma, Yu-Chen Song, Guang-Chao Zhuang, and Gui-Peng Yang. Occurrence and cycle of dimethyl sulfide in the western Pacific Ocean. Limnology and Oceanography, 66(7):2868– 2884, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Yuan:2023:PCR

[YAE^{+}23] Zhongwei Yuan, Eric P. Achterberg, Anja Engel, Zuozhu Wen, Linbin Zhou, Xunchi Zhu, Minhan Dai, and Thomas J. Browning. Phytoplankton community response to episodic wet and dry aerosol deposition in the subtropical North Atlantic. Limnology and Oceanography, 68(9):2126–2140, September 2023. CODEN LIOCAH. ISSN 0024-3590.

Yuan:2024:SBN

[YAE^{+}24] Zhongwei Yuan, Eric P. Achterberg, Anja Engel, Minhan Dai, and Thomas J. Browning. Switches between nitrogen limitation and nitrogen–phosphorus co-limitation in the subtropical North Atlantic Ocean. Limnology and Oceanography, 69(4):1005–1013, April 2024. CODEN LIOCAH. ISSN 0024-3590.

Yang:2021:CCD

[YAH^{+}21] Guang Yang, Angus Atkinson, Simeon L. Hill, Letterio Guglielmo, Antonia Granata, and Chaolun Li. Changing circumpolar distributions and isoscapes of Antarctic krill: Indo–Pacific habitat refuges counter long-term degradation of the Atlantic sector. Limnology and Oceanography, 66(1): 272–287, January 2021. CODEN LIOCAH. ISSN 0024-3590.

Yang:2022:MCB

[YAP^{+}22] Guang Yang, Angus Atkinson, Evgeny A. Pakhomov, Simeon L. Hill, and Marie-Fanny Racault. Massive circumpolar biomass of southern ocean zooplankton: Implications for food web structure, carbon export, and marine spatial planning. Limnology and Oceanography, 67(11):2516–2530, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Yi:2023:IUR

[YG23] Xiangqi Yi and Kunshan Gao. Impact of ultraviolet radiation nearly overrides the effects of elevated pCO_{2} on a prominent nitrogen-fixing cyanobacterium. Limnology and Oceanography, 68(3):557–568, March 2023. CODEN LIOCAH. ISSN 0024-3590.

Yu:2021:MRR

[YGD^{+}21] Liuqian Yu, Jianping Gan, Minhan Dai, Chiwing Rex Hui, Zhongming Lu, and Dou Li. Modeling the role of riverine organic matter in hypoxia formation within the coastal transition zone off the Pearl River Estuary. Limnology and Oceanography, 66(2):452–468, February 2021. CODEN LI-OCAH. ISSN 0024-3590.

Yuan:2020:RPC

[YJ20] Lester L. Yuan and John R. Jones. Rethinking phosphorus– chlorophyll relationships in lakes. Limnology and Oceanography, 65(8):1847–1857, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Yang:2020:RDS

[YJB^{+}20] Shu Yang, William P. Johnson, Frank J. Black, Ryan Rowland, Christine Rumsey, and Andrew Piskadlo. Response of density stratification, aquatic chemistry, and methylmercury to engineered and hydrologic forcings in an endorheic lake (Great Salt Lake, U.S.A.). Limnology and Oceanography, 65 (5):915–926, May 2020. CODEN LIOCAH. ISSN 0024-3590.

Yang:2020:RDF

[YLT^{+}20] Shi Lun Yang, Xiangxin Luo, Stijn Temmerman, Matthew Kirwan, Tjeerd Bouma, Kehui Xu, Saisai Zhang, Jiqing Fan, Benwei Shi, Haifei Yang, Ya Ping Wang, Xuefa Shi, and Shu Gao. Role of delta-front erosion in sustaining salt marshes under sea-level rise and fluvial sediment decline. Limnology and Oceanography, 65(9):1990–2009, September 2020. CO-DEN LIOCAH. ISSN 0024-3590.

Yang:2021:TIH

[YML^{+}21] Liqiang Yang, Shanli Mou, Hongmei Li, Zenghu Zhang, Nianzhi Jiao, and Yongyu Zhang. Terrestrial input of herbicides has significant impacts on phytoplankton and bacterioplankton communities in coastal waters. Limnology and Oceanography, 66(11):4028–4045, November 2021. CODEN LIOCAH. ISSN 0024-3590.

Yao:2020:HCC

[YMSH20] Hongming Yao, Melissa R. McCutcheon, Cory J. Staryk, and Xinping Hu. Hydrologic controls on CO_{2} chemistry and flux in subtropical lagoonal estuaries of the northwestern Gulf of Mexico. Limnology and Oceanography, 65(6):1380–1398, June 2020. CODEN LIOCAH. ISSN 0024-3590.

Yu:2023:TCM

[YOXC23] Hao Yu, Hannah Organ, Derry Xu, and Richard Coffin. Tidal control and mangrove dieback impact on methane emissions from a subtropical mangrove estuary. Limnology and Oceanography, 68(4):753–766, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Yamada:2023:BSR

[YPF^{+}23] Yosuke Yamada, Nirav Patel, Hideki Fukuda, Toshi Nagata, Satoshi Mitarai, and Farooq Azam. Bacterial surface roughness regulates nanoparticle scavenging in seawater. Limnology and Oceanography, 68(4):780–789, April 2023. CODEN LIOCAH. ISSN 0024-3590.

Ye:2022:EIC

[YSHS22] Youting Ye, William G. Sunda, Haizheng Hong, and Dalin Shi. Effect of increased CO_{2} on iron-light-CO_{2} co-limitation of growth in a marine diatom. Limnology and Oceanography, 67(1):172–186, January 2022. CODEN LIOCAH. ISSN 0024-3590.

Ye:2023:III

[YSHS23] Youting Ye, William G. Sunda, Haizheng Hong, and Dalin Shi. Interrelated influence of iron, light, and CO_{2} on carbon fixation in a Southern Ocean diatom. Limnology and Oceanography, 68(7):1504–1516, July 2023. CODEN LI-OCAH. ISSN 0024-3590.

Yan:2020:HTS

[YvdHC^{+}20] Jiaguo Yan, Tjisse van der Heide, Baoshan Cui, Junhong Bai, Tom Ysebaert, and Johan van de Koppel. A healthy trophic structure underlies the resistance of pristine seagrass beds to nutrient enrichment. Limnology and Oceanography, 66(10):2748–2756, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Yang:2020:MSP

[YWLY20] Bernard Yang, Mathew G. Wells, Jingzhi Li, and Joelle Young. Mixing, stratification, and plankton under lake-ice during winter in a large lake: Implications for spring dissolved oxygen levels. Limnology and Oceanography, 66(10):

2713–2729, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Ye:2021:RMI

[YWZ^{+}21] Zhiqiang Ye, Emily Williams, Chaoxian Zhao, Carolyn W. Burns, and Michael Lynch. The rapid, mass invasion of New Zealand by North American Daphnia “pulex”. Limnology and Oceanography, 66(7):2672–2683, July 2021. CODEN LI-OCAH. ISSN 0024-3590.

Ye:2020:MPO

[YWZZ20] Wang-Wang Ye, Xiao-Lei Wang, Xiao-Hua Zhang, and Gui-Ling Zhang. Methane production in oxic seawater of the western North Pacific and its marginal seas. Limnology and Oceanography, 65(10):2352–2365, October 2020. CODEN LI-OCAH. ISSN 0024-3590.

Yau:2022:AEO

[YXC^{+}22] Yvonne Y. Y. Yau, Pei Xin, Xiaogang Chen, Lucheng Zhan, Mitchell Call, Stephen R. Conrad, Christian J. Sanders, Linwei Li, Jinzhou Du, and Isaac R. Santos. Alkalinity export to the ocean is a major carbon sequestration mechanism in a macrotidal saltmarsh. Limnology and Oceanography, 67 (S2):S158–S170, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Zhou:2021:CEM

[ZBNH^{+}21] Kuanbo Zhou, Claudia R. Benitez-Nelson, Jie Huang, Peng Xiu, Zhenyu Sun, and Minhan Dai. Cyclonic eddies modulate temporal and spatial decoupling of particulate carbon, nitrogen, and biogenic silica export in the North Pacific Subtropical Gyre. Limnology and Oceanography, 66(9):3508–3522, September 2021. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2023:MDF

[ZCZ^{+}23] Lianbao Zhang, Mingming Chen, Yue Zheng, Jianning Wang, Xilin Xiao, Xiaowei Chen, Chen Hu, Jiaming Shen, Jihua Liu, Kai Tang, Dapeng Xu, Qiang Shi, Xiaoyan Ning, Helmuth Thomas, Wei Qin, Meixun Zhao, Nianzhi Jiao, and Yao Zhang. Microbially driven fate of terrigenous particulate organic matter in oceans. Limnology and Oceanography, 68(1): 148–164, January 2023. CODEN LIOCAH. ISSN 0024-3590.

Zhu:2020:IBU

[ZFQ^{+}20] Zhu Zhu, Feixue Fu, Pingping Qu, Esther Wing Kwan Mak, Haibo Jiang, Ruifeng Zhang, Zhuoyi Zhu, Kunshan Gao, and David A. Hutchins. Interactions between ultraviolet radiation exposure and phosphorus limitation in the marine nitrogen-fixing cyanobacteria Trichodesmium and Crocosphaera. Limnology and Oceanography, 65(2):363–376, February 2020. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2022:FRE

[ZJ22a] Haoran Zhang and Xiaodong Jiang. Forward and reverse evolution of multivariate responses to cyanobacteria in experimental waterflea populations. Limnology and Oceanography, 67(11):2418–2430, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2022:RDZ

[ZJ22b] Haoran Zhang and Xiaodong Jiang. Resurrection of dormant zooplankton grazers reveals multiple evolutionary responses to toxic cyanobacteria. Limnology and Oceanography, 67(9): 2000–2011, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Zhao:2022:PSO

[ZJW^{+}22] Bo Zhao, Yunlong Jia, Shengjie Wu, Lili Wei, Jian Li, Hualong Hong, Chongling Yan, Mark A. Williams, and Qiang Wang. Preservation of soil organic carbon in coastal wetlands promoted by glomalin–iron–organic carbon ternary system. Limnology and Oceanography, 67(S2):S180–S192, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Zhou:2023:REC

[ZLG^{+}23] Qiming Zhou, Qi Liu, Lei Gu, Yunfei Sun, and Zhou Yang. Resting eggs of Ceriodaphnia cornuta can sense predation risk and population crowding signals but cannot distinguish the types of predation risk. Limnology and Oceanography, 68 (8):1791–1801, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Zetsche:2020:FDA

[ZLIP20] Eva-Maria Zetsche, Ann I. Larsson, Morten H. Iversen, and Helle Ploug. Flow and diffusion around and within diatom aggregates: Effects of aggregate composition and shape. Limnology and Oceanography, 65(8):1818–1833, August 2020. CODEN LIOCAH. ISSN 0024-3590.

Zang:2024:SKF

[ZLJ^{+}24] Lin Zang, Yongqin Liu, Nianzhi Jiao, Kevin Xu Zhong, Xuanying Song, Yunlan Yang, Lanlan Cai, Keshao Liu, Guannan Mao, Mukan Ji, and Rui Zhang. Salinity as a key factor affecting viral activity and life strategies in alpine lakes. Limnology and Oceanography, 69(4):961–975, April 2024. CO-DEN LIOCAH. ISSN 0024-3590.

Zhou:2021:AIN

[ZLL^{+}21] Linbin Zhou, Fengjie Liu, Qingxia Liu, Claude Fortin, Yehui Tan, Liangmin Huang, and Peter G. C. Campbell. Aluminum increases net carbon fixation by marine diatoms and decreases their decomposition: Evidence for the iron–aluminum hypothesis. Limnology and Oceanography, 66(7):2712–2727, July 2021. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2024:BGT

[ZLL^{+}24] Zhe-Xuan Zhang, Jiwei Li, Hongxuan Lu, Huan Yang, Yige Zhang, Yongjie Tang, Meiyan Fu, and Xiaotong Peng. Bacterial glycerol tetraethers as a potential tool to trace marine methane cycling. Limnology and Oceanography, 69(1):104– 120, January 2024. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2021:SWD

[ZLN21] Xiaoxia Zhang, Pengzhi Lin, and Heidi Nepf. A simple-wave damping model for flexible marsh plants. Limnology and Oceanography, 66(12):4182–4196, December 2021. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2021:GDC

[ZLS^{+}21] Yong Zhang, Zhengke Li, Kai G. Schulz, Yingyu Hu, Andrew J. Irwin, and Zoe V. Finkel. Growth-dependent changes in elemental stoichiometry and macromolecular allocation in the coccolithophore Emiliania huxleyi under different environmental conditions. Limnology and Oceanography, 66(8): 2999–3009, August 2021. CODEN LIOCAH. ISSN 0024-

Zhao:2020:DIC

[ZLU^{+}20] Yangyang Zhao, Jing Liu, Khanittha Uthaipan, Xue Song, Yi Xu, Biyan He, Hongbin Liu, Jianping Gan, and Minhan Dai. Dynamics of inorganic carbon and pH in a large subtropical continental shelf system: Interaction between eutrophication, hypoxia, and ocean acidification. Limnology and Oceanography, 65(6):1359–1379, June 2020. CODEN LI-OCAH. ISSN 0024-3590.

Zheng:2021:HEC

[ZLW^{+}21a] Qiang Zheng, Wenxin Lin, Yu Wang, Yunyun Li, Chen He, Yuan Shen, Weidong Guo, Quan Shi, and Nianzhi Jiao. Highly enriched n-containing organic molecules of Synechococcus lysates and their rapid transformation by heterotrophic bacteria. Limnology and Oceanography, 66(2): 335–348, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Zhou:2021:CPR

[ZLW21b] Weiwen Zhou, Qian P. Li, and Zhengchao Wu. Coastal phytoplankton responses to atmospheric deposition during summer. Limnology and Oceanography, 66(4):1298–1315, April 2021. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2020:FIR

[ZN20] Xiaoxia Zhang and Heidi Nepf. Flow-induced reconfiguration of aquatic plants, including the impact of leaf sheltering. Limnology and Oceanography, 66(10):2697–2712, November 2020. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2022:RCC

[ZPK^{+}22] Shuang Zhang, Noah J. Planavsky, Joachim Katchinoff, Peter A. Raymond, Yoshiki Kanzaki, Tom Reershemius, and Christopher T. Reinhard. River chemistry constraints on the carbon capture potential of surficial enhanced rock weathering. Limnology and Oceanography, 67(S2):S148–S157, November 2022. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2021:LBA

[ZQB^{+}21] Sibo Zhang, Wei Qin, Yubei Bai, Zhenrui Zhang, Junfeng Wang, Hui Gao, Ji-Dong Gu, and Xinghui Xia. Linkages between anammox and denitrifying bacterial communities and nitrogen loss rates in high-elevation rivers. Limnology and Oceanography, 66(3):765–778, March 2021. CODEN LI-OCAH. ISSN 0024-3590.

Zehr:2023:QGC

[ZR23] Jonathan P. Zehr and Lasse Riemann. Quantification of gene copy numbers is valuable in marine microbial ecology: a comment to Meiler et al. (2022). Limnology and Oceanography, 68(6):1406–1412, June 2023. CODEN LIOCAH. ISSN 0024-3590. See [MBD^{+}22].

Zhang:2020:IHD

[ZSD^{+}20] Zhouling Zhang, Xiaole Sun, Minhan Dai, Zhimian Cao, Guillaume Fontorbe, and Daniel J. Conley. Impact of human disturbance on the biogeochemical silicon cycle in a coastal sea revealed by silicon isotopes. Limnology and Oceanography, 65(3):515–528, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Zhou:2023:TDH

[ZSF^{+}23] Zhengquan Zhou, Natalie Steiner, Gregory S. Fivash, Francesco Cozzoli, Daniel B. Blok, Lennart van Ijzerloo, Jeroen van Dalen, Tom Ysebaert, Brenda Walles, and Tjeerd J. Bouma. Temporal dynamics of heatwaves are key drivers of sediment mixing by bioturbators. Limnology and Oceanography, 68(5):1105–1116, May 2023. CODEN LIOCAH. ISSN 0024-3590.

Zabelina:2021:CET

[ZSK^{+}21] Svetlana A. Zabelina, Liudmila S. Shirokova, Sergey I. Klimov, Artem V. Chupakov, Artem G. Lim, Yuri M. Polishchuk, Vladimir Y. Polishchuk, Alexander N. Bogdanov, Ildar N. Muratov, Frederic Guerin, Jan Karlsson, and Oleg S. Pokrovsky. Carbon emission from thermokarst lakes in NE European tundra. Limnology and Oceanography, 66(S1): S216–S230, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Zhu:2021:MOR

[ZSZ^{+}21] Xuexia Zhu, Yunfei Sun, Lu Zhang, Jun Wang, Lei Gu, Yuan Huang, and Zhou Yang. Multi-omics reveal the pathways involved in induced defensive colony formation of Tetradesmus obliquus in response to Daphnia grazing cues. Limnology and Oceanography, 66(5):1819–1831, May 2021. CODEN LI-OCAH. ISSN 0024-3590.

Zhu:2022:RTW

[ZSZ^{+}22] Zhenchang Zhu, Aimee Slangen, Qin Zhu, Theo Gerkema, Tjeerd J. Bouma, and Zhifeng Yang. The role of tides and winds in shaping seed dispersal in coastal wetlands. Limnology and Oceanography, 67(3):646–659, March 2022. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2023:EDM

[ZTT23] Naiyu Zhang, Charlotte E. L. Thompson, and Ian H. Townend. The effects of disturbance on the microbial mediation of sediment stability. Limnology and Oceanography, 68(7): 1567–1579, July 2023. CODEN LIOCAH. ISSN 0024-3590.

Zhu:2020:VRN

[ZvBZ^{+}20] Zhenchang Zhu, Jim van Belzen, Qin Zhu, Johan van de Koppel, and Tjeerd J. Bouma. Vegetation recovery on neighboring tidal flats forms an Achilles’ heel of saltmarsh resilience to sea level rise. Limnology and Oceanography, 65(1):51–62, January 2020. CODEN LIOCAH. ISSN 0024-3590.

Zhu:2022:SGS

[ZWM22] Qingguang Zhu, Patricia L. Wiberg, and Karen J. McGlathery. Seasonal growth and senescence of seagrass alters sediment accumulation rates and carbon burial in a coastal lagoon. Limnology and Oceanography, 67(9):1931–1942, September 2022. CODEN LIOCAH. ISSN 0024-3590.

Zhang:2023:SCS

[ZXE^{+}23] Yu Zhang, Zhantang Xu, Jens K. Ehn, Aura Diaz, and Yuezhong Yang. A significant change in sea ice diffuse attenuation coefficient with temperature and its implications for the Arctic Ocean. Limnology and Oceanography, 68(11): 2431–2444, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Zou:2023:WDA

[ZXZ^{+}23] Wei Zou, Hai Xu, Guangwei Zhu, Mengyuan Zhu, Chaoxuan Guo, Man Xiao, Yunlin Zhang, and Boqiang Qin. Why do algal blooms intensify under reduced nitrogen and fluctuating phosphorus conditions: The underappreciated role of nonalgal light attenuation. Limnology and Oceanography, 68(10): 2274–2287, October 2023. CODEN LIOCAH. ISSN 0024-

Zhou:2021:HHA

[ZYZ^{+}21] Yongqiang Zhou, Xiaolong Yao, Lei Zhou, Zhonghua Zhao, Xiaolong Wang, Kyoung-Soon Jang, Wei Tian, Yunlin Zhang, David C. Podgorski, Robert G. M. Spencer, Dolly N. Kothawala, Erik Jeppesen, and Fengchang Wu. How hydrology and anthropogenic activity influence the molecular composition and export of dissolved organic matter: Observations along a large river continuum. Limnology and Oceanography, 66(5):1730–1742, May 2021. CODEN LIOCAH. ISSN 0024-3590.

Zhu:2020:NIV

[ZZB^{+}20] Longji Zhu, Yue Zhao, Sicong Bai, Haixuan Zhou, Xiaomeng Chen, and Zimin Wei. New insights into the variation of dissolved organic matter components in different latitudinal lakes of northeast China. Limnology and Oceanography, 65 (3):471–481, March 2020. CODEN LIOCAH. ISSN 0024-3590.

Zhao:2023:REZ

[ZZCD23] Shuai-Ying Zhao, Libin Zhou, Guangjie Chen, and Steven A. J. Declerck. Rapidly evolving zooplankton in a salinizing world: To what extent does microevolutionary adaptation to one salt increase tolerance to another one? Limnology and Oceanography, 68(11):2576–2586, November 2023. CODEN LIOCAH. ISSN 0024-3590.

Zhan:2023:SRP

[ZZHZ23] Weikang Zhan, Ying Zhang, Qingyou He, and Haigang Zhan. Shifting responses of phytoplankton to atmospheric and oceanic forcing in a prolonged marine heatwave. Limnology and Oceanography, 68(8):1821–1834, August 2023. CODEN LIOCAH. ISSN 0024-3590.

Zhao:2021:OSS

[ZZL^{+}21] Zhiyuan Zhao, Liquan Zhang, Xiuzhen Li, Lin Yuan, and Tjeerd J. Bouma. The onset of secondary seed dispersal is controlled by germination-features: a neglected process in sudden saltmarsh establishment. Limnology and Oceanography, 66(8):3070–3084, August 2021. CODEN LIOCAH. ISSN 0024-3590.

Zhan:2021:HRD

[ZZO^{+}21] Liyang Zhan, Jiexia Zhang, Zhangxian Ouyang, Ruibo Lei, Suqing Xu, Di Qi, Zhongyong Gao, Heng Sun, Yuhong Li, Man Wu, Jian Liu, and Liqi Chen. High-resolution distribution pattern of surface water nitrous oxide along a cruise track from the Okhotsk Sea to the western Arctic Ocean. Limnology and Oceanography, 66(S1):S401–S410, February 2021. CODEN LIOCAH. ISSN 0024-3590.

Zhou:2021:EAB

[ZZT^{+}21] Lei Zhou, Yongqiang Zhou, Xiangming Tang, Yunlin Zhang, Guangwei Zhu, Anna J. Székely, and Erik Jeppesen. Eutrophication alters bacterial co-occurrence networks and increases the importance of chromophoric dissolved organic matter composition. Limnology and Oceanography, 66(6): 2319–2332, June 2021. CODEN LIOCAH. ISSN 0024-3590.

Zheng:2022:HIA

[ZZWL22] Wenxiu Zheng, Enlou Zhang, Rong Wang, and Peter Guy Langdon. Human impacts alter driver–response relationships in lakes of southwest China. Limnology and Oceanography, 67(S1):S390–S402, February 2022. CODEN LIOCAH. ISSN 0024-3590.